### v1

# Safe MPC Alignment with Human Directional Feedback

Abstract
In safety-critical robot planning or control, manually specifying safety constraints or learning them from demonstrations can be challenging. In this paper, we propose a certifiable alignment method for a robot to learn a safety constraint in its model predictive control (MPC) policy with human online directional feedback. To our knowledge, it is the first method to learn safety constraints from human feedback. The proposed method is based on an empirical observation: human directional feedback, when available, tends to guide the robot toward safer regions. The method only requires the direction of human feedback to update the learning hypothesis space. It is certifiable, providing an upper bound on the total number of human feedback in the case of successful learning of safety constraints, or declaring the misspecification of the hypothesis space, i.e., the true implicit safety constraint cannot be found within the specified hypothesis space. We evaluated the proposed method using numerical examples and user studies in two developed simulation games. Additionally, we implemented and tested the proposed method on a real-world Franka robot arm performing mobile water-pouring tasks in a user study. The simulation and experimental results demonstrate the efficacy and efficiency of our method, showing that it enables a robot to successfully learn safety constraints with a small handful (tens) of human directional corrections.
Index Terms:
Learning from Human Feedback, Constraint Inference, Model Predictive Control, Human-Robot Interaction, Cutting-Plane Method.
I Introduction
In developing robot control policies, ensuring safety is always the top priority. This is especially crucial for safety-critical applications, such as autonomous vehicles or collaborative robots operating near human users, where violating safety constraints can lead to irreversible loss or damage.
While a significant line of work in safe control [1] and safe reinforcement learning [2] focuses on the safe policy design given robot behavior constraints, less attention has been paid to how to specify the safety constraints in the first place. In fact, for many daily tasks and applications, specifying proper safety constraints can be challenging due to two main reasons. First, explicit safety boundary for robot behavior can be hard to model or define. For example, in the task of a household robot feeding an elderly person, the complex nature of the task makes it very difficult to define a precise constraint for safe feeding behavior. Second, safety specifications can vary among users and require personalization. An example is lane-switching of an autonomous vehicle: aggressive drivers may prefer frequent lane-switching behavior, while conservative users may consider frequent lane-switching unsafe.
The challenges in safety specifications lead to two key questions, which we seek to answer in this paper. Q1: can a robot learn and align its safety constraints with human intent in an intuitive and online manner? Q2: can such learning process to be user-data efficient such that a robot is able to quickly adapt to different users?
Recent work [3, 4, 5] has proposed learning safety constraints from human demonstrations. However, these approaches face limitations that prevent them from fully addressing the two key questions. First, obtaining valid user demonstrations can be costly and time-consuming, especially with high-dimensional robot systems. Second, demonstration-based learning typically follows an offline or episodic learning procedure, which does not allow for on-the-fly interaction where human feedback and robot updates occur in real time.
Figure 1: Overall illustration of the proposed method. The proposed safe MPC alignment online updates the safety constraint in the robot safe MPC policy with human directional feedback, such that the robot eventually aligns the safe MPC policy with human intent for accomplishing safety-critical tasks.
This paper aims to answer the previous two questions by proposing a new paradigm for a robot to online and efficiently learn safety constraints in its model predictive control (MPC) policy with human feedback. We name the method safe MPC alignment. As shown in Fig. 1, a human user, while observing the robot’s potentially unsafe behavior, provides directional feedback for the robot’s motion. Once receiving the human feedback, the robot immediately updates the safety constraints in its MPC policy and continues its motion with the updated policy. Human feedback is in the form of physical corrections [6, 7] and sparse in time. The safe MPC alignment only utilizes the direction of human correction, regardless of the magnitude, for the safety constraints update, Importantly, the procedure of the proposed alignment is able to guarantee the exponential convergence with finite rounds of human corrections, or output failure with identified hypothesis misspecification. Specifically, we highlight our main contributions below.
(I) We propose the method of safe MPC alignment with human directional feedback. The method enables a robot to online learn a safe MPC policy from human directional corrections. To our knowledge, it is the first work for interactive learning of safety constraints from human feedback.
(II) We theoretically show the human-data efficiency of the proposed safe MPC alignment method: the proposed method is guaranteed to have exponential convergence with finite human interactions/corrections. Furthermore, we also prove that the proposed algorithm enables the identification of hypothesis space misspecification failure.
(III) We conduct extensive user studies to validate the effectiveness of the proposed method, including user studies on two MuJoCo-based simulation games and one real-world experiment of physically teaching a Franka robot arm for mobile liquid pouring. The results demonstrate the effectiveness and efficiency of the proposed method and show that a robot can successfully learn a safe MPC policy with only tens of human corrections.
The remaining paper is organized as follows. Section II reviews the related work, and Section III formulates the safe MPC alignment problem. The main algorithm is presented in Section IV. Section V focuses on the implementation details of the algorithm and presents the main results on convergence and misspecification. Sections VI and VII provide the results of numerical experiments and user studies. Conclusions are drawn in Section VIII.
II Related Work
II-A Learning from Demonstrations
Under the general theme of learning from human data, the proposed work is loosely connected to the topic of learning from demonstrations, where objective functions are learned from expert (human) data using techniques of inverse reinforcement learning (IRL) and inverse optimal control (IOC) [8, 9]. Since the demonstration data typically comes as time sequences of data, IRL/IOC methods typically run in an offline or episodic manner: a robot first obtains human demonstrated trajectories, and then performs reward/cost learning [9, 10, 11, 12, 13, 14]. As pointed in [15], acquiring optimal demonstration data can be difficult in practice. The challenge can be either attributed to higher degrees of freedom of the robot, or the lack of prior knowledge about the desired robot behavior [16].
In this work, we focus on online and interactive learning of safety constraints. Different from learning from demonstrations that needs the exhibition of desired robot behavior, the proposed learning paradigm only requires a human user to incrementally improve the current robot motion by providing directional feedback, thus more accessible to human users.
II-B Learning from Human Interaction
Orthogonal to learning from demonstrations is robot learning from human interaction. The line of this work typically focus on learning a reward/cost function from human feedback, which can be in the form of preference and corrections/feedback. The learned reward/cost function is then used to derive human-preferred robot policy. In learning from human preferences, human preference typically given as user’s ranking of two or multiple robot candidate trajectories [17, 18, 19, 20, 21, 22, 23], and Bradley-Terry model [24] is used to update the reward/cost function. The recently variant in this line of work seeks to directly learn a policy from human preference [25].
Human interaction can also come as corrections, where human users may intervene and correct the robot at runtime. An intended robot reward/cost function can be inferred from human corrections. Specific human corrections include language [26, 27, 28, 29, 30] and physical corrections in trajectory/action space [31, 32, 6]. Since human corrections can be sparse in time, many methods [31, 33, 32, 34, 35, 6, 36] use trajectory deformation to convert sparse corrections to a continuous human-intended trajectory using a temporal propagation kernel. The reward/cost function is updated to explain the deformed trajectories. Some strategies are further proposed to deal with the uncertainty and improve the robustness of the learning process, like updating one parameter dimension at a time [32] or formulating the reward/cost learning in the framework of Kalman Filter [33].
Our work also belongs to the category of learning from human physical corrections. But instead of learning reward/cost functions, we focus on learning safety constraints from human corrections, which has not been addressed before. Methodologically, we propose a certifiable learning method, which guarantees the exponential learning convergence with few human corrections or identifies the hypothesis misspecification when the learning hypothesis space does not contain the true human intent. Those capabilities are missing in existing methods. For example, the state-of-the-art methods [37] for learning from human preference requires hundreds or thousands of human feedback, while the proposed method can successfully learn a complex safe policy only with tens of human corrections.
II-C Learning Constraints from Demonstrations
A increasing number of work focuses on learning safety constraints from demonstrations. One vanilla idea is to formulate constraint learning as a binary classification. However, different from learning reward/cost, since a robot usually runs within a safe region, negative demonstration samples that violate safety constraints are challenge to collect. To collect negative samples, different strategies have been proposed. [3, 38] consider the trajectories of lower cost as the unsafe samples. [39] proposes adjusting the constraints toward decreasing the safety margin of feasible policies. Safety constraints can also be learned via mixed integer programming [3, 38], pessimistic adjustment [39], and no-regret algorithm [40].
Another line of work attempts to learn safety constraints only from safe demonstration data. For example, some work exploits the optimality of demonstration with unknown constraints [41, 42, 43]. Different probabilistic formulations are also proposed to learn constraints from demonstration data, such as maximum likelihood [5, 44, 45], maximum causal entropy [46, 47], and Bayesian inference [48, 49]. In the constraint learning, safety constraints can be represented as a set of safe states [41], the belief of potential constraints [42], Gaussian process [43], or directly as a convex hull in policy space [4].
The vast majority of existing approaches focus on learning safety constraints from human demonstrations. They are mostly developed in the setting of reinforcement learning, needing to repetitively solve policy optimization in each learning update. Those methods cannot be used in an online, interactive setting, where a robot is expected to immediately update its safe policy and respond to feedback given by a human user. To achieve so, we will develop the first method, to our best knowledge, to enable learning the safe constraint from human feedback.
III Preliminary and Problem Formulation
III-A Safe Model Predictive Control
We consider the following safe model predictive control (Safe MPC) for a robotic system, where the safety constraint is learnable and parameterized by $\boldsymbol{\theta}\in\mathbb{R}^{r}$.
$\min_{\boldsymbol{u}_{0:T-1}}&\quad J(\boldsymbol{u}_{0:T-1};\boldsymbol{x}_{0})=\sum_{t=0}^{T-1}c(\boldsymbol{x}_{t},\boldsymbol{u}_{t})+h(\boldsymbol{x}_{T})\\ \text{subject to}&\quad\boldsymbol{x}_{t+1}=\boldsymbol{f}(\boldsymbol{x}_{t},\boldsymbol{u}_{t})\\ &\quad{g}_{\boldsymbol{\theta}}(\boldsymbol{x}_{0:T},\boldsymbol{u}_{0:T-1})\leq 0\end{aligned}.$ (1)
Here, $\boldsymbol{x}_{t}\in\mathbb{R}^{n}$ and $\boldsymbol{u}_{t}\in\mathbb{R}^{m}$ are robot state and action, respectively, at time step $t=0,1,2,...,T$ with $T$ being the prediction horizon; $\boldsymbol{x}_{0:T}=(\boldsymbol{x}_{0},\boldsymbol{x}_{1},...,\boldsymbol{x}_{T})$ and $\boldsymbol{u}_{0:T-1}=(\boldsymbol{u}_{0},\boldsymbol{u}_{1},...,\boldsymbol{u}_{T-1})$; $J(\cdot)$ is the given cost function including the running cost term ${c}(\cdot)$ and final cost term $h(\cdot)$, representing the given task; $\boldsymbol{f}(\cdot,\cdot)$ is the system dynamics; ${g}_{\boldsymbol{\theta}}(\cdot)$ is the safety constraint, which is parameterized by $\boldsymbol{\theta}$. Given starting state $\boldsymbol{x}_{0}$ and constraint parameter $\boldsymbol{\theta}$, the trajectory solution to (1), denoted as $\boldsymbol{\xi}_{\boldsymbol{\theta}}=\{\boldsymbol{u}^{\boldsymbol{\theta}}_{0:T-1},\boldsymbol{x}^{\boldsymbol{\theta}}_{0:T}\}$, is an implicit function of $\boldsymbol{\theta}$.
In a receding horizon implementation, $\boldsymbol{x}_{0}$ is set as the robot’s current state (estimation) $\boldsymbol{{x}}_{k}$, where $k=0,1,2,...,$ is the policy rollout time step; and only the first action $\boldsymbol{u}^{\boldsymbol{\theta}}_{0}$ in $\boldsymbol{{u}}_{0:T}^{\boldsymbol{\theta}}$ is applied to the robot to move the robot to the next state $\boldsymbol{x}_{k+1}$. This leads to a closed-loop Safe MPC policy for the robotic system, denoted as $\boldsymbol{\pi}_{\boldsymbol{\theta}}$, mapping from robot current state $\boldsymbol{x}_{k}$ to its action $\boldsymbol{u}^{\boldsymbol{\theta}}_{0}$.
III-B Safe Asymptotic Approximation of Safe MPC
We present a penalty-based approximation to the original Safe MPC problem (1). Define variable $\boldsymbol{\xi}=\{\boldsymbol{u}_{0:T-1},\boldsymbol{x}_{0:T}\}$ and a penalty objective for (1) that incorporates the cost $J(\boldsymbol{\xi})$ and safety constraint $g_{\boldsymbol{\theta}}(\boldsymbol{\xi})$:
$B(\boldsymbol{\xi},\boldsymbol{\theta})=J(\boldsymbol{\xi})-\gamma\ln(-g_{\boldsymbol{\theta}}(\boldsymbol{\xi})).$ (2)
We have the following result from our prior work [50], stating an asymptotic and safe approximation to the solution $\boldsymbol{\xi}_{\boldsymbol{\theta}}$ to (1) by the following optimization
$\min_{\boldsymbol{u}_{0:T-1}}\quad B(\boldsymbol{\xi},\boldsymbol{\theta})\qquad\text{subject to}\quad\boldsymbol{x}_{t+1}=\boldsymbol{f}(\boldsymbol{x}_{t},\boldsymbol{u}_{t}).$ (3)
Lemma 1.
(Theorem 2 in [50]) If the original Safe MPC problem (1) satisfies certain mild conditions (i.e., local second-order condition and strict complementarity), then for any small $\gamma>0$, there exists a (local) solution $\boldsymbol{\xi}_{\boldsymbol{\theta}}^{\gamma}=\{\boldsymbol{u}_{0:T-1}^{\boldsymbol{\theta},\gamma},\boldsymbol{x}_{0:T-1}^{\boldsymbol{\theta},\gamma}\}$ to (3), and $\boldsymbol{\xi}_{\boldsymbol{\theta}}^{\gamma}$ is safe with respect to the original constraint, i.e., $g_{\boldsymbol{\theta}}(\boldsymbol{\xi}_{\boldsymbol{\theta}}^{\gamma})<0$. Furthermore, $\boldsymbol{\xi}_{\boldsymbol{\theta}}^{\gamma}$ is asymptotically approximate the original solution to (1): $\boldsymbol{\xi}_{\boldsymbol{\theta}}^{\gamma}\rightarrow\boldsymbol{\xi}_{\boldsymbol{\theta}}$ as $\gamma\rightarrow 0$.
The above lemma asserts that by solving (3) one can safely use the approximate solution $\boldsymbol{\xi}_{\boldsymbol{\theta}}^{\gamma}$ to replace the original solution with arbitrary controlled accuracy by $\gamma$. As shown in [50], $\gamma$ usually takes a small value for good approximation accuracy. In the development below, we fix the parameter $\gamma$ in (3). We denote the approximate Safe MPC in (3) as $\boldsymbol{\pi}^{\gamma}_{\theta}$.
III-C Problem Formulation
III-C1 Representation of Human Intended Safety
We consider the scenario of human-robot coexistence, where a human is supervising the robot for a given task. The user has his/her intent of robot safety. In Safe MPC (1), we hypothesize that the user intended safety can be represented by the set
$\begin{split}\mathcal{S}(\boldsymbol{\theta}_{H})=\{(\boldsymbol{x}_{0:T},\boldsymbol{u}_{0:T-1})\,\,|\,\,\ {g}_{\boldsymbol{\theta}_{H}}(\boldsymbol{x}_{0:T},\boldsymbol{u}_{0:T-1})\leq 0\}.\end{split}$ (4)
Here, $\boldsymbol{\theta}_{H}$ is user implicit intent parameter picked from an implicit set $\Theta_{H}\subset\mathbb{R}^{r}$, i.e., $\boldsymbol{\theta}_{H}\in\Theta_{H}$. $\Theta_{H}$ is implicit to the human user and unknown to the robot. For the robot policy $\boldsymbol{\pi}_{\boldsymbol{\theta}}^{\gamma}$, if $\boldsymbol{\theta}\in\Theta_{H}$, the user considers the robot behavior to be safe; otherwise if $\boldsymbol{\theta}\notin\Theta_{H}$, the robot motion by $\boldsymbol{\pi}_{\boldsymbol{\theta}}^{\gamma}$ might not be safe at certain time in the view of the user, and thus the user provides corrections to improve the safety of the robot.
We use an implicit set $\Theta_{H}\in\mathbb{R}^{r}$, instead of single point $\boldsymbol{\theta}_{H}$, to capture the potential variation of user intents. We suppose that a volume measure of $\Theta_{H}$, $\operatorname{{Vol}}(\Theta_{H})$, is non-zero, and a user intent $\boldsymbol{\theta}_{H}\in\Theta_{H}$ can vary at different steps and context, but always within $\Theta_{H}$.
III-C2 Human Online Directional Correction
We consider a human user who gives correction to the robot in an online fashion while the robot is running with the most updated $\boldsymbol{\pi}_{\boldsymbol{\theta}}^{\gamma}$. Specifically, denote $i$ to be the index of learning update, and suppose the robot policy currently is $\boldsymbol{\pi}_{\boldsymbol{\theta}_{i}}^{\gamma}$. At the time steps $k_{i}$ of rolling out $\boldsymbol{\pi}_{\boldsymbol{\theta}_{i}}^{\gamma}$, a human user gives directional correction $\boldsymbol{a}_{k_{i}}$ in robot action space at its state $\boldsymbol{x}_{k_{i}}$. After receiving $\boldsymbol{a}_{k_{i}}$, the robot updates the safety parameter $\boldsymbol{\theta}_{i}\rightarrow\boldsymbol{\theta}_{i+1}$ in (1) and obtain the updated policy $\boldsymbol{\pi}_{\boldsymbol{\theta}_{i+1}}^{\gamma}$. Then, the robot continues with the updated policy $\boldsymbol{\pi}_{\boldsymbol{\theta}_{i}}^{\gamma}$ until the next user possible correction. This process of robot MPC policy update and user correction repeats until the convergence of the policy update or the user is satisfied with the robot motion. An illustration of one update is shown in Fig. 2.
Figure 2: Update of safe MPC policy with human directional corrections.
In the above-stated human feedback process, we consider that a human user only provides directional correction $\boldsymbol{a}$, i.e., the magnitude $\left\lVert\boldsymbol{a}\right\rVert$ does not matter (theoretical justification will be given later in Section IV-A). For example, when correcting a mobile robot, the user’s correction can be "left" or "right", without specifying how much the robot should go. Also, we assume that the user feedback is applied in the action space. For some robotic systems, this assumption can be readily met, such as turning the driving wheels for autonomous vehicles. In other cases, the correction can be applied through a dedicated (hardware or algorithmic) interface to the action space; such as contact forces from end-effector to joint torque, and teleoperation device for remote control. In those cases, $\boldsymbol{a}$ is the correction signal after the interface. The reason why we only consider the correction in action space is that otherwise the corrections in robot state space can lead to dynamics unfeasibility or uncontrollability for under-actuated robotic systems [7, 51].
III-C3 Robot Operation Safety and Emergency Stop
In the aforementioned policy update procedure, the robot will respond to human corrections by executing the updated safe MPC policy $\pi_{\boldsymbol{\theta}}$. However, it is possible that the robot may still exit the user-intended safety zone during the correction process. In such cases, a human user may trigger an emergency stop to stop the robot motion and potentially restart it from a desired safety region. This emergency stop mechanism is common in real-world robotic applications. Until an emergency stop is triggered, human corrections typically occur while the robot is still within the intended safety zone; otherwise, the corrections would be useless. For instance, correcting an autonomous vehicle after it has been crashed is futile.
III-C4 Parametric Safety Constraints
In the safe MPC (1), the safety set is defined as a zero sublevel set of a safety value function ${g}_{\boldsymbol{\theta}}$. We note that if ${g}_{\boldsymbol{\theta}}$ has sufficient representation power, the safety set can represent complex safety region, which can be alternatively represented by the intersection of multiple constraints (e.g., a polytope safety set defined by multiple hyperplanes). In the following, we mainly focus on a weighted-features parameterization of ${g}_{\theta}$, which has been commonly used by the extensive prior work on learning cost/reward [4, 7, 52, 53, 54]. Specifically, we parameterize $\boldsymbol{g}_{\theta}$
${g}_{\boldsymbol{\theta}}(\boldsymbol{x}_{0:T},\boldsymbol{u}_{0:T-1})=\phi_{0}(\boldsymbol{x}_{0:T},\boldsymbol{u}_{0:T-1})+\theta_{1}\phi_{1}(\boldsymbol{x}_{0:T},\boldsymbol{u}_{0:T-1})\\ +\theta_{2}\phi_{2}(\boldsymbol{x}_{0:T},\boldsymbol{u}_{0:T-1})+\cdots+\theta_{r}\phi_{r}(\boldsymbol{x}_{0:T},\boldsymbol{u}_{0:T-1}),$ (5)
with $\boldsymbol{\theta}=[\theta_{1},\theta_{2},...\theta_{r}]^{\mkern-1.5mu\mathsf{T}}\in\mathbb{R}^{r}$ and $\boldsymbol{\phi}=[\phi_{1},\phi_{2},\dots,\phi_{r}]^{\mkern-1.5mu\mathsf{T}}$ the weight and feature vectors, respectively. We fix the weight of $\phi_{0}$ as unit, because otherwise, the $\boldsymbol{\theta}$ is scaling ambiguous, i.e., any scaled $c\cdot\boldsymbol{\theta}$ with $c>0$ lead to the same MPC solutions. We do not impose any constraints on the choice of feature vector as long as it is differentiable. For example, $\boldsymbol{\phi}$ can be neural features. Later, we will discuss our proposed method for generic parameterization $\boldsymbol{g}_{\boldsymbol{\theta}}$.
III-C5 Problem of Safe MPC Alignment
Following the procedure of online human correction and robot policy update described in Subsection III-C and given a weighted-feature safety constraint (5), the interest of this work is to develop a policy update rule: $\boldsymbol{\theta}_{i}\rightarrow\boldsymbol{\theta}_{i+1}$ with each human directional correction, such that $\boldsymbol{\theta}_{i}\rightarrow\boldsymbol{\theta}_{H}\in\Theta_{H}$, as $i=1,2,3,\dots,$. i.e., the safety constraint in the robot safe MPC policy will be aligned with human intent.
IV Algorithm Overview for Safe MPC Alignment
IV-A Hypothesis on User Directional Feedback
In this session, we will introduce a hypothesis on user directional correction, which lays the foundation for our alignment framework proposed subsequently. Informally, we have the following intuitions on user feedback. First, user directional corrections, on average, improve the robot safety. The similar assumption is the basis for reinforcement learning with human feedback [55]. Second, user correction on robot safety is typically proactive and typically happens before the robot enters into the true unsafe region. In other words, human correction is made when the robot operates within $\mathcal{S}(\boldsymbol{\theta}_{H})$ in (4). For example, it is useless to correct a ground vehicle after its collision with obstacles. Third, with the second intuition, user feedback is typically safety-margin aware: the closer a robot gets to the boundary of the unsafe region, the more significant user safety correction will be. This is observed in our prior user study [7]. In the following, we will formalize those intuitions.
Suppose that robot motion plan $\boldsymbol{\xi}_{\boldsymbol{\theta}}^{\gamma}$ is solved from (3) with parameter $\boldsymbol{\theta}$, and a human user applies directional correction $\boldsymbol{a}$ at first time step 00. Define
$\mathbf{a}=(\boldsymbol{a}^{\mkern-1.5mu\mathsf{T}},0,0,...)\in\mathbb{R}^{mT}.$ (6)
Formally, we assume there is a nonzero-volume set $\bar{\Theta}_{H}$ (containing a ball) in user implicit intent set $\Theta_{H}$, i.e., $\bar{\Theta}_{H}\subseteq\Theta_{H}$, such that for any $\boldsymbol{\theta}_{H}\in\bar{\Theta}_{H}$, the user directional correction $\boldsymbol{a}$, on average, indicates a direction of lowering the penalty value $B(\boldsymbol{\xi}_{\boldsymbol{\theta}}^{\gamma},\boldsymbol{\theta}_{H})$ for the robot motion plan $\boldsymbol{\xi}_{\boldsymbol{\theta}}^{\gamma}$. Concretely, it means $\mathbf{a}$ on averages points in a direction of the negative gradient of $B(\boldsymbol{\xi},\boldsymbol{\theta}_{H})$ evaluated at $\boldsymbol{\xi}_{\boldsymbol{\theta}}^{\gamma}$, i.e.,
$\forall\boldsymbol{\theta}_{H}\in\bar{\Theta}_{H},\ \operatorname{\mathbb{E}}_{\mathbf{a}}\left\langle\mathbf{a},-\nabla B(\boldsymbol{\xi}_{\boldsymbol{\theta}}^{\gamma},\boldsymbol{\theta}_{H})\right\rangle=\\ \left\langle\mathbf{\bar{a}},-\nabla B(\boldsymbol{\xi}_{\boldsymbol{\theta}}^{\gamma},\boldsymbol{\theta}_{H})\right\rangle\geq 0.$ (7)
Here, ${\operatorname{\mathbb{E}}}_{\mathbf{a}}(\mathbf{a})=\mathbf{\bar{a}}$ and the equality is due to the second term $\nabla B(\boldsymbol{\xi},\boldsymbol{\theta})$ is not dependent on $\mathbf{a}$. $\nabla B(\boldsymbol{\xi},\boldsymbol{\theta})$ denotes the full gradient of $B$ with respect to $\boldsymbol{u}_{0:T-1}$ evaluated at trajectory $\boldsymbol{\xi}=\{\boldsymbol{u}_{0:T-1},\boldsymbol{x}_{0:T}\}$. Similar notations apply to $J$ and $g$ below. Since state sequence $\boldsymbol{x}_{0:T}$ is the rollout of $\boldsymbol{u}_{0:T-1}$ on dynamics, the differentiation goes through the dynamics equation.
One of the important observations to (7) is that the magnitude of the user correction $\mathbf{a}$ does not affect the inequality. This brings one advantage of directional correction compared to the magnitude-based correction: regardless of robot safeMPC plan $\boldsymbol{\xi}_{\boldsymbol{\theta}}^{\gamma}$, the user directional correction always accounts for half of the action space, as shown in Figure 3. Extending (7) leads to
$\forall\boldsymbol{\theta}_{H}\in\bar{\Theta}_{H},\left<\mathbf{\bar{a}},-\nabla J(\boldsymbol{\xi}_{\boldsymbol{\theta}}^{\gamma})\right>+\\ \gamma\frac{1}{-g_{\boldsymbol{\theta}_{H}}(\boldsymbol{\xi}_{\boldsymbol{\theta}}^{\gamma})}\left<\mathbf{\bar{a}},-\nabla g_{\boldsymbol{\theta}_{H}}(\boldsymbol{\xi}_{\boldsymbol{\theta}}^{\gamma})\right>\geq 0.$ (8)
Figure 3: Illustration of directional corrections in the space of robot motion plan $\boldsymbol{\xi}$s. The grey circles represent the level contours of $B(\boldsymbol{\xi},\boldsymbol{\theta}_{H})$. The black dot stands for current motion plan $\boldsymbol{\xi}_{\boldsymbol{\theta}}^{\gamma}$ and the red point stands for the plan $\boldsymbol{\xi}_{\boldsymbol{\theta}_{H}}^{\gamma}$ with human intended safety. The blue region stands for the half-space of all possible human corrections under our assumption. Some samples of human corrections are shown by blue arrows.
For the above equation, we have the following interpretations.
(I) The first and second terms on the left side of (8) correspond to the user’s correction toward lowering the cost and improving the safety value of robot plan $\boldsymbol{\xi}_{\boldsymbol{\theta}}^{\gamma}$. $\gamma$ serves as a weight between two improvements.
(II) If we interpret the value of $-{g_{\boldsymbol{\theta}_{H}}(\boldsymbol{\xi}_{\boldsymbol{\theta}}^{\gamma})}$ as the safety margin of robot plan $\boldsymbol{\xi}_{\boldsymbol{\theta}}^{\gamma}$, (8) indicates that user’s correction is safety-margin aware: the closer the robot motion get to the unsafe region, the more significant a human user correction will be to drive the system back to safety interior. This property is consistent with our previous intuition.
(III) The above hypothesis (7) requires that $B(\boldsymbol{\xi}_{\boldsymbol{\theta}}^{\gamma},\boldsymbol{\theta}_{H})$ is well-defined in the first place, i.e.,
$\forall\boldsymbol{\theta}_{H}\in\bar{\Theta}_{H},\ \quad g_{\boldsymbol{\theta}_{H}}(\boldsymbol{\xi}_{\boldsymbol{\theta}}^{\gamma})<0.$ (9)
The requirement corresponds to the third intuition — a user typically makes correction only when the robot is within user intended safety region, i.e., $\boldsymbol{\xi}_{\boldsymbol{\theta}}^{\gamma}\in\mathcal{S}(\boldsymbol{\theta}_{H})$.
In sum, we have used (7) and (9) to formalize the intuitions we have observed at the beginning of this session. Those two constraints on $\boldsymbol{\theta}_{H}$ will be used in the robot learning pipeline developed in the next session.
IV-B Alignment Algorithm: Learning by Hypothesis Space Cutting
To find an implicit $\boldsymbol{\theta}_{H}\in\bar{\Theta}_{H}\subseteq\Theta$ for the safe MPC (1), our idea is that at any learning iteration $i=0,1,2,....$, the robot monitors and updates a hypothesis space $\Theta_{i}\in\mathbb{R}^{r}$, a set containing all possible $\bar{\Theta}_{H}$. At initial, the robot may choose a hypothesis space $\Theta_{0}$ that is large enough to contain $\bar{\Theta}_{H}$. Specifically, one can choose
$\Theta_{0}=\{\boldsymbol{\theta}\ |\ \boldsymbol{c}_{l}\leq\boldsymbol{\theta}\leq\boldsymbol{c}_{h}\}\subset\mathbb{R}^{r},$ (10)
in which $\boldsymbol{c}_{l},\boldsymbol{c}_{h}$ are the given bounding vectors. By specifying a $\Theta_{0}$ large enough, we assume
$\bar{\Theta}_{H}\subset\Theta_{0}.$ (11)
Later, we will that our alignment algorithm can self-access the failure of learning in case of misspecification, where (11) is not met.
Recall the procedure of human correction in Section III-C and Fig. 2. At $i$-th learning update, $i=1,2,3...$, the robot safe MPC policy $\boldsymbol{\pi}_{\boldsymbol{\theta}_{i}}^{\gamma}$ with $\boldsymbol{\theta}_{i}\in\Theta_{i-1}$. The user gives a directional correction $\boldsymbol{a}_{k_{i}}$ at rollout step ${k_{i}}$. By letting $\mathbf{\bar{a}}_{i}=(\boldsymbol{a}_{k_{i}}^{T},0,0,...)$ as in (6), we can establish the following two constraints on $\bar{\Theta}_{H}$ from the previous hypothesis (7) and (9)
$\bar{\Theta}_{H}\subset\mathcal{C}_{i}=\left\{\boldsymbol{\theta}|\left<\mathbf{\bar{a}}_{i},\nabla B(\boldsymbol{\xi}_{\boldsymbol{\theta}_{i}}^{\gamma},\boldsymbol{\theta})\right>\leq 0,\,\,\,g_{\boldsymbol{\theta}}(\boldsymbol{\xi}_{\boldsymbol{\theta}_{i}}^{\gamma})<0\right\},$ (12)
which we refer to as the cutting set. Using $\mathcal{C}_{i}$, one can update the hypothesis space
$\Theta_{i}=\Theta_{i-1}\cap\mathcal{C}_{i}.$ (13)
The above update can be viewed as a step of using the cutting set (12) to cut the current hypothesis space $\Theta_{i-1}$ to remove the portion where $\bar{\Theta}_{H}$ is impossible. The remaining hypothesis space $\Theta_{i}$ will be new hypothesis used in the next iteration. With the above update, we have the following result.
Lemma 2.
Given $\bar{\Theta}_{H}\subset\Theta_{0}$, if one follows the update (13), then $\bar{\Theta}_{H}\subset\Theta_{i}$ for all $i=1,2,3,...$.
Proof.
See the proof in Appendix A-A. ∎
As shown by Fig. 4, the procedure of robot learning is to cut and move hypothesis space based on human directional correction. Therefore, we name our alignment method Hypothesis Space Cutting. As a summary, we present the three main steps of the hypothesis space cutting updates below.
Figure 4: The update of constraint parameter (2D examples). Two cutting lines are used to cut $\Theta_{i-1}$ to $\Theta_{i}$. The user implicit intent set $\bar{\Theta}_{H}$ is contained in $\Theta_{i}$. New $\boldsymbol{\theta}_{i+1}$ is chosen as the parameter for the next safe MPC policy.
Safe MPC alignment by Hypothesis Space Cutting Initialize $\Theta_{0}$ that contains $\bar{\Theta}_{H}$. Repeat the following three steps for $i=1,2,...,K$ until $\boldsymbol{\theta}_{i}\in\bar{\Theta}_{H}$. Step 1: Robot chooses parameter $\boldsymbol{\theta}_{i}\in\Theta_{i-1}$ for its current safe MPC policy $\boldsymbol{\pi}_{\boldsymbol{\theta}_{i}}^{\gamma}$. Step 2: Human user applies a directional correction $\boldsymbol{a}_{k_{i}}$ while robot executes $\boldsymbol{\pi}_{\boldsymbol{\theta}_{i}}^{\gamma}$ Step 3: Perform hypothesis space cut using (13) to get new hypothesis space $\Theta_{i}$.
At $i$th iteration of the Hypothesis Space Cutting, the robot chooses $\boldsymbol{\theta}_{i}\in\Theta_{i-1}$ from the most recent hypothesis space $\Theta_{i-1}$ for its current safe MPC $\boldsymbol{\pi}_{\boldsymbol{\theta}_{i}}^{\gamma}$. While executing $\boldsymbol{\pi}_{\boldsymbol{\theta}_{i}}^{\gamma}$, the robot receives human directional correction $\boldsymbol{a}_{k_{i}}$ at the rollout time step $k_{i}$. This correction is used to induce the cutting set $\mathcal{C}_{i}$ (12), which will be used to update $\Theta_{i-1}$ to $\Theta_{i}$ in (13). The above iteration repeats $i=1,2,3,...$.
From (13) and by following the above hypothesis space cutting method, one can easily see that $\Theta_{0}\supseteq\Theta_{1}\supseteq\Theta_{2}\supseteq...\supseteq\bar{\Theta}_{H}$. Given a volume measure, it follows $\operatorname{{Vol}}\Theta_{0}\geq\operatorname{{Vol}}\Theta_{1}\geq\operatorname{{Vol}}\Theta_{2}\geq...,\geq\operatorname{{Vol}}{\bar{\Theta}_{H}}$. Thus, the volume of the hypothesis space is at least not increasing along with user corrections. It is hoped that the volume of search space is strictly decreasing, such that new $\boldsymbol{\theta}_{i+1}$ has a very good chance of falling inside $\bar{\Theta}_{H}$, which means a successful safety alignment between the robot and human. We will later (in Section V) theoretically show that the above Safe MPC Alignment method, within finite interactions $K$, will either
  • (I)
    surely have $\boldsymbol{\theta}_{i}\in\bar{\Theta}_{H}\subset\Theta_{H}$, $i<K$, or
  • (II)
    declare the failure due to misspecification ($\bar{\Theta}_{H}\cap\Theta_{0}=\emptyset$)
Also in Section V, we will provide implementation details for the above hypothesis space cutting procedure.
IV-C Properties and Geometric Interpretations
In this session, we present some properties and the geometric interpretations for the proposed safe MPC alignment method. The properties will be used to establish the convergence of the algorithm in the next section.
IV-C1 Piecewise Linearity of Cuts
Recall the weight-feature parameterization of the safety constraint (5). From (12), we can examine $g_{\boldsymbol{\theta}}(\boldsymbol{\xi}_{\boldsymbol{\theta}_{i}}^{\gamma})<0$ is a linear inequality in $\boldsymbol{\theta}$.
Based on the definition of the penalty objective function (2), $\mathbf{\bar{a}}_{i},\nabla B(\boldsymbol{\xi}_{\boldsymbol{\theta}_{i}}^{\gamma},\boldsymbol{\theta})\right>\leq 0$ is a linear fractional inequality
$\mathbf{\bar{a}}_{i},\nabla J(\boldsymbol{\xi}_{\boldsymbol{\theta}_{i}}^{\gamma})\right>+\gamma\frac{1}{-g_{\boldsymbol{\theta}}(\boldsymbol{\xi}_{\boldsymbol{\theta}_{i}}^{\gamma})}\left<\mathbf{\bar{a}}_{i},\nabla g_{\boldsymbol{\theta}}(\boldsymbol{\xi}_{\boldsymbol{\theta}_{i}}^{\gamma})\right>\leq 0,$ (14)
with a positive denominator $-g_{\boldsymbol{\theta}}(\boldsymbol{\xi}_{\boldsymbol{\theta}_{i}}^{\gamma})>0$ due to (12). Multiplying the denominator on both side yields another linear inequality. Formally, we give a lemma here about the piecewise linearity of the cut $\mathcal{C}_{i}$ in (12).
Lemma 3.
$\forall i$, $\mathcal{C}_{i}$ in (12) can be expressed as an intersection of two half-spaces in $\mathbb{R}^{r}$:
$\mathcal{C}_{i}=\{\boldsymbol{\theta}\ |\ \boldsymbol{\theta}^{\mkern-1.5mu\mathsf{T}}\boldsymbol{h}_{i}\leq b_{i}\}\cap\{\boldsymbol{\theta}\ |\ {\boldsymbol{\theta}}^{\mkern-1.5mu\mathsf{T}}\boldsymbol{\phi}(\boldsymbol{\xi}_{\boldsymbol{\theta}_{i}}^{\gamma})<-\phi_{0}(\boldsymbol{\xi}_{\boldsymbol{\theta}_{i}}^{\gamma})\},$ (15)
with
$\small\boldsymbol{h}_{i}=-\left<\mathbf{\bar{a}}_{i},\nabla J(\boldsymbol{\xi}_{\boldsymbol{\theta}_{i}}^{\gamma})\right>\boldsymbol{\phi}(\boldsymbol{\xi}_{\boldsymbol{\theta}_{i}}^{\gamma})+\gamma\frac{\partial\boldsymbol{\phi}}{\partial\boldsymbol{u}}(\boldsymbol{\xi}_{\boldsymbol{\theta}_{i}}^{\gamma})\mathbf{\bar{a}}_{i},$ (16)
[…]
Proof.
Equation (15) can be directly obtained from (12), (2) and (5). This completes the proof. ∎
With the lemma above, the intersection $\Theta_{i-1}\cap\mathcal{C}_{i}$ can be viewed as cutting $\Theta_{i-1}$ with the boundary of $\mathcal{C}_{i}$ which consists of two piecewise hyperplanes:
$\begin{split}\mathbf{bd}\ \mathcal{C}_{i}=\{\boldsymbol{\theta}\ |\ \boldsymbol{\theta}^{\mkern-1.5mu\mathsf{T}}\boldsymbol{h}_{i}=b_{i},{\boldsymbol{\theta}}^{\mkern-1.5mu\mathsf{T}}\boldsymbol{\phi}(\boldsymbol{\xi}_{\boldsymbol{\theta}_{i}}^{\gamma})\leq-\phi_{0}(\boldsymbol{\xi}_{\boldsymbol{\theta}_{i}}^{\gamma})\}\ \cup\\ \{\boldsymbol{\theta}\ |\ \boldsymbol{\theta}^{\mkern-1.5mu\mathsf{T}}\boldsymbol{h}_{i}\leq b_{i},{\boldsymbol{\theta}}^{\mkern-1.5mu\mathsf{T}}\boldsymbol{\phi}(\boldsymbol{\xi}_{\boldsymbol{\theta}_{i}}^{\gamma})=-\phi_{0}(\boldsymbol{\xi}_{\boldsymbol{\theta}_{i}}^{\gamma})\},\end{split}$ (18)
which is shown in Fig. 4
As the initial hypothesis space $\Theta_{0}\in\mathbb{R}^{r}$ in (10) is a polytope. Iteratively cutting with piecewise hyperplanes yields a series of convex polytopes in $\mathbb{R}^{r}$. Consequently, all $\Theta_{i}$s, $i=0,1,2,...$, are convex polytopes.
IV-C2 Relationship between $\boldsymbol{\theta}_{i}$ and $\mathbf{bd}\ \mathcal{C}_{i}$
According to (15), the cut of $\mathcal{C}_{i}$ depends on robot plan $\boldsymbol{\xi}_{\boldsymbol{\theta}_{i}}^{\gamma}$ solved by robot MPC $\boldsymbol{\pi}_{\boldsymbol{\theta}_{i}}^{\gamma}$ in (3) with $\boldsymbol{\theta}_{i}\in\Theta_{i-1}$. This suggests some relationship between $\boldsymbol{\theta}_{i}$ and the boundary $\mathcal{C}_{i}$ in (18). In fact, the relationship between $\boldsymbol{\theta}_{i}$ and $\mathbf{bd}\ \mathcal{C}_{i}$ is given by following lemma.
Lemma 4.
$\forall i,\ \boldsymbol{\theta}_{i}$ is located on $\mathbf{bd}\ \mathcal{C}_{i}$ in (18). Specifically, $\boldsymbol{\theta}_{i}$ satisfies
$\boldsymbol{\theta}_{i}^{\mkern-1.5mu\mathsf{T}}\boldsymbol{h}_{i}=b_{i}\quad\text{and}\quad\boldsymbol{\theta}_{i}^{\mkern-1.5mu\mathsf{T}}\boldsymbol{\phi}(\boldsymbol{\xi}_{\boldsymbol{\theta}_{i}}^{\gamma})<-\phi_{0}(\boldsymbol{\xi}_{\boldsymbol{\theta}_{i}}^{\gamma})$ (19)
for all $i=1,2,...$.
Proof.
See Appendix A-B. ∎
The above lemma states $\boldsymbol{\theta}_{i}\in\mathbf{bd}\ \mathcal{C}_{i}$. Its geometric interpretation is in Fig. 5: the piecewise cut of $\mathcal{C}_{i}$ always passes $\boldsymbol{\theta}_{i}$. This suggests that the choice of $\boldsymbol{\theta}_{i}\in\Theta_{i-1}$ is critical in the learning process: it directly affects the position of the cut $\mathbf{bd}\ \mathcal{C}_{i}$ and consequently the volume to be cut out. Thus, it is important to choose a good $\boldsymbol{\theta}_{i}\in\Theta_{i-1}$ such that a large hypothesis space is removed after each cut.
We here briefly discuss the implication of different choices of $\boldsymbol{\theta}_{i}\in\Theta_{i-1}$. When choosing $\boldsymbol{\theta}_{i}$ near the boundary of $\Theta_{i-1}$, as shown in Fig. 5 and Fig. 5, the removed volume of the hypothesis space depends on the specific direction of cut $\mathbf{bd}\,\mathcal{C}_{i}$. Since the direction of the cut $\mathbf{bd}\,\mathcal{C}_{i}$ will be eventually dependent upon the user’s correction direction according to (18), the removed volume of the hypothesis space can be relatively small, as in Fig. 5, or large, as in Fig. 5, depending on the specific direction of the correction applied by the user.
As the specific direction of user correction is hard to assume, one can intuitively see that one good choice of $\boldsymbol{\theta}_{i}\in\Theta_{i-1}$ will be near the center of $\Theta_{i-1}$. As such, we can ensure an averaged large cut of the hypothesis space regardless of the direction of $\mathbf{bd}\ \mathcal{C}_{i}$, as shown in Fig. 5. We will rigorously discuss the centering choice of $\boldsymbol{\theta}_{i}$ and the guarantee for volume reduction rate of the hypothesis space in the next section.
Figure 5: Illustration of the different choices of $\boldsymbol{\theta}_{i}$. (a) and (b): When $\boldsymbol{\theta}_{i}$ is chosen near the set boundary, the cut volume (i.e., the volume of the removed hypothesis space) depends on the specific direction of $\mathbf{bd}\ \mathcal{C}_{i}$ (i.e., the user-specific directional correction). (c): When $\boldsymbol{\theta}_{i}$ is near the center, the cut volume is large on average regardless of the specific user directional correction.
V Implementation Details and Main Results
In this section, we will present the details for how to select $\boldsymbol{\theta}_{i}\in\Theta_{i-1}$ at each iteration $i$ of the hypothesis space cutting in our proposed Safe MPC Alignment algorithm. Based on the preceding geometric intuition, our choice of $\boldsymbol{\theta}_{i}$ is to achieve a large averaged reduction rate for hypothesis space volume given human directional corrections. The reduction rate enables us to establish an exponential convergence of the learning process. As a result, we can find an upper bound of total user corrections to guarantee the success of learning, i.e., $\boldsymbol{\theta}_{i}\in\bar{\Theta}_{H}$, or identify the misspecification of hypothesis space.
V-A The Choice of $\boldsymbol{\theta}_{i}$
As discussed previously, $\boldsymbol{\theta}_{i}$ needs to be picked from the "center" of $\Theta_{i-1}$ in the proposed algorithm. An apparent option is to choose the mass center of $\Theta_{i-1}$. However, computing the centroid of a polytope can be #P-Hard [56]. The methods, such as Hit-And-Run algorithm [57], which estimate the mass center by sampling, also suffer from the high sample complxity, thus are impractical in our online real-time setting. In [58], the authors provide an another method to choose the center of a polytope using a set of inscribed balls, but it lacks of the guarantee of the volume reduction rate. For computational efficiency and guaranteed convergence, we will use the center of the Maximum Volume Ellipsoid (MVE) inscribed in the convex set $\Theta_{i-1}$ as $\boldsymbol{\theta}_{i}$, shown in Fig. 6.
Figure 6: $\boldsymbol{\theta}_{i}$ is chosen as the center of MVE of $\Theta_{i-1}$ (2D example).
Definition 1 (Maximum Volume inscribed Ellipsoid (MVE)[59]).
The MVE of a given convex set $\Theta\subset\mathbb{R}^{r}$ is defined as the image of the unit ball under an affine transformation
=$\mathcal{E}=\{\bar{H}\boldsymbol{\theta}+\bar{\boldsymbol{d}}\ |\ \|\boldsymbol{\theta}\|_{2}\leq 1\}.$ (20)
The affine transformation parameter $\bar{H}\in\mathbb{S}^{r}_{++}$ (a $r\times r$ positive definite matrix), and $\bar{\boldsymbol{d}}\in\mathbb{R}^{r}$ is called the center of $\mathcal{E}$. $\bar{H},\bar{\boldsymbol{d}}$ are determined via the following optimization problem for the volume of the ellipsoid $\det H$:
$\begin{split}\bar{H},\bar{\boldsymbol{d}}=\underset{H\in\mathbb{S}^{r}_{++},\boldsymbol{d}}{\arg\max}\quad&\log\det H\\ \rm{s.t.}\quad&\sup\nolimits_{\|\boldsymbol{\theta}\|_{2}\leq 1}\mathbbm{1}_{\Theta}(H\boldsymbol{\theta}+\boldsymbol{d})\leq 0\end{split},$ (21)
where the indicator function $\mathbbm{1}_{\Theta}(\boldsymbol{\theta})=0$ for $\boldsymbol{\theta}\in\Theta$ and otherwise $\mathbbm{1}_{\Theta}(\boldsymbol{\theta})=+\infty$. Specifically, when $\mathbbm{1}_{\Theta}$ is a polytope $\{\boldsymbol{\theta}^{\mkern-1.5mu\mathsf{T}}\boldsymbol{h}_{i}\leq b_{i},\ i=1,\dots,m\}$, the problem can be converted to a convex optimization problem with $m$ constraints:
$\begin{split}\bar{H},\bar{\boldsymbol{d}}=\underset{H\in\mathbb{S}^{r}_{++},\boldsymbol{d}}{\arg\min}\quad&\log\det H^{-1}\\ \rm{s.t.}\quad&\left\lVert H\boldsymbol{h}_{i}\right\rVert_{2}+\boldsymbol{d}^{\mkern-1.5mu\mathsf{T}}\boldsymbol{h}_{i}\leq b_{i},\ i=1,\dots,m\end{split}.$ (22)
In the Safe MPC Alignment algorithm, by the above definition, let $\mathcal{E}_{i-1}$ denote the MVE of $\Theta_{i-1}$ and $\bar{\boldsymbol{d}}_{i-1}$ denote the center of $\mathcal{E}_{i-1}$. Then, the choice of $\boldsymbol{\theta}_{i}$ is
$\boldsymbol{\theta}_{i}=\bar{\boldsymbol{d}}_{i-1}.$ (23)
Given $\Theta_{i-1}$ a convex polytope, the optimization problem (21) can be solved efficiently by convex programming (22).
V-B Result 1: Finite Iteration Convergence
In this section, we give the main result that, when using the choice of $\boldsymbol{\theta}_{i}=\bar{\boldsymbol{d}}_{i-1}$ mentioned above, the Safe MPC Alignment algorithm can converge in a finite number of iterations to find $\boldsymbol{\theta}_{H}\in\bar{\Theta}_{H}\subset\Theta_{H}$. This result comes from the fact that the volume reduction rate after each hypothesis space cut is upper bounded, as formally stated below.
Lemma 5.
In the Safe MPC Alignment algorithm, with $\boldsymbol{\theta}_{i}$ as the center $\boldsymbol{\bar{d}}_{i-1}$ of MVE of $\Theta_{i-1}$, one has
$\frac{\mathbf{Vol}(\Theta_{i})}{\mathbf{Vol}(\Theta_{i-1})}\leq 1-\frac{1}{r}.$ (24)
for all $i=1,2,3,...$ and any enclosure volume measure $\bf{Vol}$.
Proof.
See Appendix A-C. ∎
The lemma above asserts an upper bound of the volume of the updated hypothesis space $\mathbf{Vol}(\Theta_{i})$ relative to the previous volume. With the above result, we can now establish the finite iteration convergence of the proposed Safe MPC Alignment, stated below.
Theorem 1 (Successful learning in finite iterations).
In the Safe MPC Alignment algorithm, $\boldsymbol{\theta}_{i}$ is selected as the center $\boldsymbol{\bar{d}}_{i-1}$ of MVE of $\Theta_{i-1}$, $i=1,2,3,...$. If $\bar{\Theta}_{H}\subset\Theta_{0}$, the algorithm will terminate successfully within
$K=\left\lceil\frac{\ln\frac{\tau_{r}\rho_{H}^{r}}{\mathbf{Vol}(\Theta_{0})}}{\ln(1-\frac{1}{r})}\right\rceil.$ (25)
iterations, where $\rho_{H}$ is the radius of a ball contained within $\bar{\Theta}_{H}$ and $\tau_{r}$ is the volume of a unit ball in $\mathbb{R}^{r}$ space. That is, there exists $i<K$, such that $\boldsymbol{\theta}_{i}\in\bar{\Theta}_{H}\subset\Theta_{H}$.
Proof.
See Appendix A-D. ∎
The above main result asserts a maximum iteration count for the proposed algorithm to find $\boldsymbol{\theta}_{i}\in\Theta_{H}$. In other words, within $K$ rounds of human corrections, the robot can successfully learn a safety set intended by the user. The above result requires a pre-set of $\rho_{H}^{r}$, which is the radius of a ball contained in $\bar{\Theta}_{H}$. Since the size of $\bar{\Theta}_{H}$ is unknown in practice, $\rho_{H}>0$ can be interpreted as an underestimated size of the unknown user intent set $\bar{\Theta}_{H}$. It also serves as a termination threshold for learning convergence: a smaller $\rho_{H}$ can lead to a more concentrated learned $\boldsymbol{\theta}$.
However, the finite iteration convergence certificate in Theorem 1 comes with a prerequisite: $\bar{\Theta}_{H}\in\Theta_{0}$ has to be satisfied in the first place in the Safe MPC Alignment algorithm. Although such a condition can be achieved by empirically choosing $\Theta_{0}$ large enough to contain all possible $\bar{\Theta}_{H}$ (10), it is also desired to have a systematic way to detect the hypothesis misspecification. In the next session, we will discuss the identification of hypothesis misspecification, where such condition is not met.
V-C Result 2: Hypothesis Misspecification Declaration
We consider the case of hypothesis space misspecification, where $\bar{\Theta}_{H}\cap\Theta_{0}=\emptyset$. In this case, no $\boldsymbol{\theta}\in\Theta_{0}$ can account for the user intended safety (4). As mentioned in [54, 60], this kind of gap between hypothesis space and true human model can lead to undesired behavior of the robot. The following result states the proposed Safe MPC Alignment algorithm can automatically certify the misspecification of the initial hypothesis space.
Theorem 2 (Identifying misspecification).
In the Safe MPC Alignment algorithm, $\boldsymbol{\theta}_{i}$ is selected as the center $\boldsymbol{\bar{d}}_{i-1}$ of MVE of $\Theta_{i-1}$, $i=1,2,3,...$. If $\bar{\Theta}_{H}\cap\Theta_{0}=\emptyset$, the algorithm converges to a set $\Theta^{*}$ as $i\rightarrow\infty$, such that $\mathbf{Vol}(\Theta^{*})=0$ and $\Theta^{*}\cap\mathbf{int}\ \Theta_{0}=\emptyset$.
Proof.
See Appendix A-E. ∎
The above result states that, when the initial hypothesis space is misspecified, $\bar{\Theta}_{H}\cap\Theta_{0}=\emptyset$, the Safe MPC Alignment algorithm will finally converge to a set of zero volume on the boundary of the initial hypothesis space $\Theta_{0}$.
The above theorem inspires us to empirically identify the misspecification by examining the position of MVE centers during the converging process: if after the maximum iteration step $K-1$ in (25), the MVE center $\boldsymbol{\theta}_{K-1}$ is very close to the boundary of initial hypothesis set (10), there is likely hypothesis space misspecification and further adjustment on initial hypothesis space should be made. See Fig. 7.
Figure 7: When there is misspecification, the Safe MPC Alignment algorithm will converge to some point on the boundary of $\Theta_{0}$. So we can use the distance from $\boldsymbol{\theta}_{K-1}$ to the boundary as a signal for misspecification.
Specifically, recall (10) that $\Theta_{0}$ is a box with bounding vector $\boldsymbol{c}_{l},\boldsymbol{c}_{h}$, we can pre-define a small threshold $\epsilon$ to examine every entry of $\boldsymbol{\theta}_{K-1}$ and identify possible misspecification. If $\boldsymbol{\theta}_{K-1}$ satisfies the criterion
$\min_{j}\ \lvert\boldsymbol{c}_{l}[j]-\boldsymbol{\theta}_{K-1}[j]\rvert\leq\epsilon\ \text{or}\ \min_{j}\ \lvert\boldsymbol{c}_{h}[j]-\boldsymbol{\theta}_{K-1}[j]\rvert\leq\epsilon,$ (26)
it means $\boldsymbol{\theta}_{K-1}$ is very close to the boundary of $\Theta_{0}$. We consider there is likely misspecification in this case.
V-D Detailed Algorithm of Safe MPC Alignment
With the detailed centering choice method of $\boldsymbol{\theta}_{i}$ as well as the main results on convergence and misspecification, we can establish the full safe MPC alignment. With initializing $\Theta_{0}$ and pre-define a termination threshold $\rho_{H}$, three steps in the hypothesis space cutting are executed until user satisfies robot motion or the iteration count reaches the maximum count $K$ in (25). Note that during the execution of the MPC policy, $\boldsymbol{\pi}_{\boldsymbol{\theta}_{i}}^{\gamma}$, the human can perform emergency stopping and reset the robot’s state, which corresponds to Section III-C3.
When reaching and exist after $K-1$ iterations, the algorithm checks the misspecification by (26). The Safe MPC Alignment algorithm with misspecification identification is presented in the Algorithm 1.
Initialization: Parameterized constraint $\boldsymbol{g}_{\theta}$, initial hypothesis space $\Theta_{0}$, termination threshold $\rho_{H}$ (i.e., approximate radius of $\bar{\Theta}_{H}$), $i=1$
Compute maximum correction count $K$ from $\rho_{H}$ in (25);
while $i<K$ do
       Choose $\boldsymbol{\theta}_{i}\in\Theta_{i-1}$ using (23);
      Robot executes $\boldsymbol{\pi}_{\boldsymbol{\theta}_{i}}^{\gamma}$;
      if Human provides directional correction $\boldsymbol{a}_{k_{i}}$ then
             Generate $\bar{\boldsymbol{\mathbf{a}}}_{i}$ using (6);
             Calculate the cutting set $\mathcal{C}_{i}$ using (15);
             Update hypothesis space $\Theta_{i}=\Theta_{i-1}\cap\mathcal{C}_{i}$;
             $i\leftarrow i+1$;
      else if  Human triggers emergency stop  then
             Human restarts robot from safe region.;
      else if  Human user is satisfied  then
             return $\boldsymbol{\theta}_{i}$;
end while
if $\boldsymbol{\theta}_{K-1}$ satisfies (26) then
       return misspecification failure;
else
       return $\boldsymbol{\theta}_{K-1}$
end if
Algorithm 1 Safe MPC Alignment Algorithm
VI Numerical Example and User Study in Simulation Games
In this section, we will first evaluate the proposed Safe MPC Alignment method using a numerical example with simulated human corrections knowing ground-truth safety constraints. Then, we will test the performance of the methods in two user studies using two simulation games developed based on MuJoCo [61].We consider two typical robot tasks. One is a drone navigation task and the other a 7-DOF robot arm reaching task. In both user studies, the robot is controlled by an MPC controller and human participants provide simulation robots with real-time correction instructions through a keyboard. The objective of each game is to teach a robot to successfully learn a safe MPC to move through an unknown safety-critical environment.
The organization in this section is as follows: Subsection VI-A presents the numerical evaluation and its analysis. Subsections VI-B and VI-C describe the setting of the two developed simulation games used in our user study. Subsection VI-D details the study procedure and the recruitment of human users. The results of the user studies and some analysis are given in VI-E. Please access the accompanying codes at https://github.com/asu-iris/Safe-MPC-Alignment.
VI-A Numerical Example
VI-A1 Environment Setup
We use the safe control of an inverted pendulum as the numerical example. The dynamics equation of an inverted pendulum is
$\ddot{\alpha}=\frac{3}{ml^{2}}(-\frac{1}{2}m\mathrm{g}l\sin{\alpha}+u-d\dot{\alpha}),$ (27)
with $\alpha$ being the angle between the pendulum and the direction of gravity, $\mathrm{g}=10\mathrm{m/s^{2}}$ the gravity constant, $u$ the pivot torque, $m=1\mathrm{kg}$, $l=1\mathrm{m}$ and $d=0.4$ being the length, inertia, and dampening of the pendulum. We set the system state as $\boldsymbol{x}=[\alpha,\dot{\alpha}]^{\mkern-1.5mu\mathsf{T}}$ and discretize (27) by the Euler method with $\triangle t=0.02\mathrm{s}$. A zero-mean Gaussian noise with covariance $\Sigma=\mathbf{diag}(1,4)\times 10^{-5}$ is added to $\boldsymbol{x}$ at each control step to simulate the state noise. The task is to control the pendulum to the target $\boldsymbol{x}^{*}=[\pi,0]^{\mkern-1.5mu\mathsf{T}}$. We additionally pose the following unknown state affine constraint to the motion of the pendulum
$0.6\alpha+\dot{\alpha}=\boldsymbol{\theta}^{H}\boldsymbol{x}\leq 3,$ (28)
where $\boldsymbol{\theta}^{H}=[0.6,1]^{\mkern-1.5mu\mathsf{T}}$ is to be learned.
In Safe MPC formulation (1), we set $T=40$, the step cost $c(\boldsymbol{x}_{t},u_{t})$ and terminal cost $h(\boldsymbol{x}_{T})$
$\begin{gathered}c(\boldsymbol{x}_{t},u_{t})=Qu_{t}^{2}\\ h(\boldsymbol{x}_{T})=(\boldsymbol{x}_{T}-\boldsymbol{x}^{*})^{T}R(\boldsymbol{x}_{T}-\boldsymbol{x}^{*}),\end{gathered}$ (29)
respectively, with $Q=0.1$ and $R=\mathrm{diag}(25,10)$.
To ensure the safe MPC rollout trajectory to satisfy a global constraint, it is sufficient to learn a safety constraint in safe MPC defined on the closest controllable state. Thus, in (1), we aim to define and learn a constraint defined on the first controlled state:
${g}_{\boldsymbol{\theta}}(\boldsymbol{\xi})=-3+\boldsymbol{\theta}^{\mkern-1.5mu\mathsf{T}}\boldsymbol{x}_{1}.$ (30)
In asymptotic MPC approximation, we set $\gamma=0.1$.
Suppose the user intent set $\bar{\Theta}_{H}=\{\boldsymbol{\theta}|\left\lVert\boldsymbol{\theta}-\boldsymbol{\theta}_{H}\right\rVert\leq 0.02\}$. Then, simulated human corrections $\mathbf{a}$ can be generated as follows: when the system is running near the human safety boundary, i.e., ${g}_{\boldsymbol{\theta}_{H}}(\boldsymbol{\xi}_{\boldsymbol{\theta}}^{\gamma})>-\epsilon_{g}$ ($\epsilon_{g}=0.25$ here), a correction $\mathbf{a}$ can be generated in probability $p=0.3$ using the first element the signed gradient vector $\mathrm{sign}(-\nabla B(\boldsymbol{\xi}_{\boldsymbol{\theta}}^{\gamma},\boldsymbol{\theta}_{H}))$. Recall that $B(\boldsymbol{\xi},\boldsymbol{\theta})$ is the barrier function in (2). It is trivial to verify that this correction $\mathbf{a}$ has positive product with $-\nabla B(\boldsymbol{\xi}_{\boldsymbol{\theta}}^{\gamma},\boldsymbol{\theta}_{H})$, and by definition it is applied to a safe trajectory $\boldsymbol{\xi}_{\boldsymbol{\theta}}^{\gamma}$. The both hypotheses in section IV-A is then satisfied, and $\mathbf{a}$ can be used as a valid human correction.
When the system state violates ${g}_{\boldsymbol{\theta}_{H}}(\boldsymbol{\xi}_{\boldsymbol{\theta}}^{\gamma})\leq 0$ or it is near the target, the pendulum state is uniformly randomly reset between $[0,0]^{\mkern-1.5mu\mathsf{T}}$ and $[\frac{2\pi}{3},3]^{\mkern-1.5mu\mathsf{T}}$. Simulated human correction will continue to be applied until the learned parameters $\boldsymbol{\theta}_{i}$ fall in $\bar{\Theta}_{H}$, which terminates one trial.
VI-A2 Convergence Analysis
In the first session, we evaluate the convergence of the learning algorithm without hypothesis misspecification. In Algorithm 1, we set the initial hypothesis $\Theta_{0}$ with box boundaries $\boldsymbol{c}_{l}=[-6,-6]^{\mkern-1.5mu\mathsf{T}}$ and $\boldsymbol{c}_{h}=[2,2]^{\mkern-1.5mu\mathsf{T}}$; thus $\Theta_{0}$ includes $\bar{\Theta}_{H}$. The initial value of $\boldsymbol{\theta}_{1}$ is set as $[-2,-2]^{\mkern-1.5mu\mathsf{T}}$ (i.e., the center of $\Theta_{0}$).
With simulated human corrections, we run Algorithm 1 and track the update of $\boldsymbol{\theta}_{i}$ after every correction. $\boldsymbol{\theta}_{i}$ falls in the intention set $\bar{\Theta}_{H}$ after 14 corrections, triggering the termination of learning. We then respectively evaluate each history $\boldsymbol{\theta}_{i}$ by using it as the MPC constraint parameter to control the system starting from $\boldsymbol{x}=[0,0]^{\mkern-1.5mu\mathsf{T}}$. We plot the corresponding MPC rollout trajectory in $(\alpha,\dot{\alpha})$ phase plot in Fig. 8.The affine ground-truth constraint (28) and the corresponding learned constraints $\boldsymbol{g}_{\boldsymbol{\theta}_{i}}$ are also shown in Fig. 8.
(a) Initial trajectory (with $\boldsymbol{\theta}_{1}$)
(b) After 3 corrections (with $\boldsymbol{\theta}_{4}$)
(c) After 6 corrections (with $\boldsymbol{\theta}_{7}$)
(d) After 14 corrections (with $\boldsymbol{\theta}_{15}$)
Figure 8: The pendulum rollout trajectory at different correction interactions: (a) trajectory with $\boldsymbol{\theta}_{1}$, (b) with $\boldsymbol{\theta}_{4}$, (c) with $\boldsymbol{\theta}_{7}$, (d) with $\boldsymbol{\theta}_{15}$. The learned constraint with $\boldsymbol{\theta}_{1}$ is omitted since it covers the whole first quadrant. As the learning proceeds, the learned constraint boundary (orange line) converges to the ground-truth constraint boundary (red line). When the learning process terminates, the learned and ground-truth constraint boundaries almost overlap, suggesting the success of constraint learning from corrections.
From Fig. 8, we can conclude that the ground-truth constraint can be successfully recovered from the simulated corrections. The precision of this recovery is very high and the learned and ground-truth safety boundary almost overlap.
To further analyze the convergence of the algorithm, we run the algorithm for 10 times, all of which terminate after 12-17 corrections. In each run, we keep track of the following two metrics over correction count, shown in Fig. 9.
M1 (left panel of Fig. 9): The logarithm volume measure of the solved MVE $\log\det\bar{H}$ in (21). This provides a lower bound of $\Theta_{i}$ to show the exponential convergence property.
M2 (right panel of Fig. 9): The euclidean distance between learned parameter $\boldsymbol{\theta}_{i}$ and $\boldsymbol{\theta}_{H}$, i.e $\left\lVert\boldsymbol{\theta}_{i}-\boldsymbol{\theta}_{H}\right\rVert_{2}$. This suggests the similarity between the learned constraint and the ground-truth constraint, see Fig. 9.
From Fig. 9, we can conclude the proposed algorithm is
  • (I)
    Fast-converging. During learning, the logarithm volume $\log\det\bar{H}$ has near-linear decreasing property. This observation is consistent with the exponential convergence in theory stated in Lemma 5.
  • (II)
    Correction-efficient. During learning, $\left\lVert\boldsymbol{\theta}_{i}-\boldsymbol{\theta}_{H}\right\rVert_{2}$ converges to the value lower than 0.02 within 17 corrections. By using 2D area as a volume measure, we can compute the upper bound of correction count by Theorem 1 :
    ${\ln\frac{\tau_{r}\rho_{H}^{r}}{\mathbf{Vol}(\Theta_{0})}}/{\ln(1-\frac{1}{r})}=17.28.$ (31)
    This suggests that the algorithm is guaranteed to converge within finite $K=17$ corrections, this is consistent with the observations in the experiment.
(a) Logarithm MVE Volume
(b) Loss to Ground-truth
Figure 9: (a) The plot of $\log\det\bar{H}$ with respect to correction numbers. In all algorithm runs, the plot shows the near-linear decrease of the logarithm volume, suggesting the exponential decrease of the volume of MVE of the hypothesis spaces, as asserted in Theorem 1. (b) The plot of the parameter estimation error $\left\lVert\boldsymbol{\theta}_{i}-\boldsymbol{\theta}_{H}\right\rVert_{2}$ versus correction numbers. In all runs, the parameter estimation error can converge to a value lower than 0.02 within 17 corrections.
VI-A3 Result Analysis for Misspecification Case
To analyze the performance of the proposed algorithm in misspecification cases, we set the initial hypothesis $\Theta_{0}$ as $\hat{\boldsymbol{c}_{l}}=[-1,-1]^{\mkern-1.5mu\mathsf{T}}$ and $\hat{\boldsymbol{c}_{h}}=[0.8,0.8]^{\mkern-1.5mu\mathsf{T}}$ which does not contain $\bar{\Theta}_{H}$. In this setting, the algorithm automatically stops after 7th correction. The path of $\boldsymbol{\theta}_{i},\ i=1,\dots,8$ is shown in Fig. 10. In Fig. 10, $\boldsymbol{\theta}_{i}$ is continuously moving towards the boundary of the initial hypothesis space $\Theta_{0}$. This fact verifies Theorem 2 and justifies the proposed method is capable of misspecification detection.
Figure 10: Illustration of misspecification case. The blue dashed lines are the boundary of $\Theta_{0}$. $\boldsymbol{\theta}_{i},\ i=1,\dots,7$ is shown in he orange dots in one run. Black arrows show the update directions of $\boldsymbol{\theta}_{i}$s. The center of the ground-truth intent set is marked as the red star, and the $\boldsymbol{\theta}_{8}$ at the termination of learning is marked as the green triangle. When misspecification happens, $\boldsymbol{\theta}$ will converge toward the boundary of $\Theta_{0}$, as stated by Theorem 2.
VI-B Drone Navigation Game
VI-B1 Environment Setup
As shown in Fig. 11, the goal of this drone navigation game is let a human user to teach a done to successfully navigate through a narrow gate (rectangle frame in yellow) whose position is unknown, given different flying target positions (shown in yellow points behind the gate). The drone starts from $[x_{\text{init}},y_{\text{init}},z_{\text{init}}]^{\mkern-1.5mu\mathsf{T}}$ with $x_{\text{init}}=0.1$, $z_{\text{init}}=0.5$, and $y_{\text{init}}$ uniformly randomly in an interval $[4,6]$. The starting quaternion $\boldsymbol{q}_{\text{init}}=[1,0,0,0]^{\mkern-1.5mu\mathsf{T}}$. We set three different target positions at $[19,1,9]^{\mkern-1.5mu\mathsf{T}}$, $[19,5,9]^{\mkern-1.5mu\mathsf{T}}$ and $[19,9,9]^{\mkern-1.5mu\mathsf{T}}$, with target quaternion $[1,0,0,0]^{\mkern-1.5mu\mathsf{T}}$. For the narrow gate (yellow rectangle), both its size and position are unknown, and one only roughly know its x position $x_{\text{gate}}\approx 10.0$. The drone navigation game is developed using MuJoCo physics engine [61].
(a) Tracking View
(b) Fixed View
Figure 11: Drone Navigation Game. (a) The tracking view seen by the user in the drone user study. (b) The fixed view seen by the user in the drone user study. Users can precept the motion of the drone in 3D space with the help of these two views. The $xyz$-axis (red, green, and blue lines), the gate (yellow bars for the edge, orange bars for the outer frame), the drone (white robot model), the local plan outputted from MPC (red curve) and current motion target (yellow balls) are visible to users. in runtime, the user will see one target at a time.
Since the location of the gate is unknown, the drone needs to learn a safety constraint in its safe MPC policy with humans guidance in order to successfully fly through the gate. In the drone MPC policy (1), the dynamics and its parameters are set following ([7]); and we generically set the constraint function ${g}_{\boldsymbol{\theta}}(\boldsymbol{\xi})$ using a set of $N$ Gaussian radial basis functions (RBFs)
${g}_{\boldsymbol{\theta}}(\boldsymbol{\xi})=\phi_{0}+\sum\nolimits_{i=1}^{N}\theta_{i}\phi_{i}(\boldsymbol{\xi}).$ (32)
As discussed in Section III-C4, we fix $\phi_{0}=-1$ to avoid learning ambiguity, and use $N=20$ temporally accumulated RBFs defined as
$\phi_{i}(\boldsymbol{\xi})=\sum\nolimits_{t=0}^{T}\beta_{t}\mathrm{RBF}_{i}(\boldsymbol{x}_{t}),$ (33)
where $\beta_{t}$ is a weighting factor to control or balance the immediate and future constraint satisfaction of robot MPC control. If one emphasizes instant safety, more weights will be given to the initial time step. In our implementation, inspired by the practice in RL, we choose an exponentially decaying $\beta_{t}$ starting at $t=5$, i.e. $\beta_{t}=0,\ t<5$ and $\beta_{t}=0.9^{t-5},\ t\geq 5$ to achieve better predictive safety performance. In (33), each Guassian RBF is defined as
$\mathrm{RBF}_{i}(\boldsymbol{x}_{t})=\begin{cases}\exp(-\epsilon^{2}\left\lVert r_{xy,t}-c_{xy,i}\right\rVert^{2}),\,\,i=1,...,10\\ \exp(-\epsilon^{2}\left\lVert r_{xz,t}-c_{xz,i}\right\rVert^{2}),\,\,i=11,...,20\end{cases}.$ (34)
Here, the variables $r_{xy,t}=[r_{x,t},r_{y,t}]^{\mkern-1.5mu\mathsf{T}}$ and $r_{xz,t}=[r_{x,t},r_{z,t}]^{\mkern-1.5mu\mathsf{T}}$ with $r_{x},r_{y},r_{z}$ from $\boldsymbol{x}_{t}$. The $i$th center $c_{xy,i}=[c_{x,i},c_{y,i}]^{\mkern-1.5mu\mathsf{T}}$ or $c_{xz,i}=[c_{x,i},c_{z,i}]^{\mkern-1.5mu\mathsf{T}}$ is defined with $c_{x,i}={x_{\text{gate}}}=9.8$, $c_{y,i}$ and $c_{z,i}$ are evenly chosen on interval $[0,10]$. We set $\epsilon=0.45$. The initial hypothesis space $\Theta_{0}$ for possible $\boldsymbol{\theta}$ in (32) is set as $\boldsymbol{c}_{l}=[-80,\dots,-80]^{\mkern-1.5mu\mathsf{T}}$ and $\boldsymbol{c}_{h}=[100,\dots,100]^{\mkern-1.5mu\mathsf{T}}$. $\gamma$ is 50. This initialization blocks the drone from advancing to the gate at the beginning of the game, leading to conservative behavior in the learning process. The drone will proceed toward the gate after few corrections.
VI-B2 Human Correction Interface
The motion of the drone is rendered in two camera views in MuJoCo visualization (Fig. 11) for better human perception: one is the first-person view tracking the drone, the other a third-person fixed view. The gate, the current MPC planning solution, and the target position are visible to the users.
Human users apply corrections to the drone via a keyboard during the drone flying. The user is informed of the mappings between the key press and the corresponding correction in the game, listed in Table I. The safety constraint (32) in the drone’s MPC policy is immediately updated once a user correction is detected. When the user presses "ENTER", the drone is reset to the initial position and the target will also change.
TABLE I: keyboard commands in uav user study
Key Correction Interpretation
"up" $[-1,0,1,0]^{\mkern-1.5mu\mathsf{T}}$ Go forward (adding positive y-axis torque)
"down" $[1,0,-1,0]^{\mkern-1.5mu\mathsf{T}}$ Go backward (adding negative y-axis torque)
"left" $[0,1,0,-1]^{\mkern-1.5mu\mathsf{T}}$ Go left (adding negative x-axis torque)
"right" $[0,-1,0,1]^{\mkern-1.5mu\mathsf{T}}$ Go right (adding positive x-axis torque)
"shift" $[1,1,1,1]^{\mkern-1.5mu\mathsf{T}}$ Go up (increasing all thrusts)
"ctrl" $-[1,1,1,1]^{\mkern-1.5mu\mathsf{T}}$ Go down (decreasing all thrusts)
"enter" N/A Reset robot position and switch target.
VI-C Franka Constrained Reaching Game
VI-C1 Environment Setup
As shown in Fig. 12, the goal of the Franka constrained reaching game is to teach a simulated Franka robot arm such that its gripper hand successfully reaches different target positions (shown in cyan ball) through a narrow slot (shown by two yellow bars) with unknown position or geometry. Specifically, the arm starts at a configuration in which the end-effector position $\boldsymbol{r}_{\text{init}}=[0.22,-0.5,0.48]^{\mkern-1.5mu\mathsf{T}}$ and altitude $\boldsymbol{q}_{\text{init}}=[0,-0.707,0,0.707]^{\mkern-1.5mu\mathsf{T}}$. The reaching target position $\boldsymbol{r}^{*}$ will be dynamically chosen from three options $[-0.6,-0.5,0.4]^{\mkern-1.5mu\mathsf{T}}$, $[-0.6,-0.5,0.5]^{\mkern-1.5mu\mathsf{T}}$, $[-0.6,-0.5,0.6]^{\mkern-1.5mu\mathsf{T}}$. The reaching target attitude is $\boldsymbol{q}^{*}=[0,-0.707,0,0.707]^{\mkern-1.5mu\mathsf{T}}$. The slot size and position are unknown except its rough x-position which is $x_{\text{slot}}\approx-0.15$.
(a) Starting Configuration
(b) Motion Illustration
Figure 12: Franka Constrained Reaching Game. (a) The starting configuration of the arm in the user study. RGB arrows denote the world $xyz$-frame. (b) Illustration of one successful reach by adjusting the position and altitude of the gripper, long green arrow denotes the motion direction. Users can precept the motion of the robot arm in 3D space by adjusting the view. The bars, the robot arm, and current motion target (cyan ball) are all visible to users.
We adopt a two-level control scheme, where a safe MPC policy (high-level) is defined for the end effector, which generates the desired velocity for a given low-level operational space controller to track. In the safe MPC policy, we model the end-effector dynamics as
$\begin{gathered}\dot{\boldsymbol{r}}=\boldsymbol{v}^{d},\\ \dot{\boldsymbol{q}}=\frac{1}{2}\Omega(R^{\mkern-1.5mu\mathsf{T}}(\boldsymbol{q})\boldsymbol{w}^{d})\boldsymbol{q},\end{gathered}$ (35)
where $\boldsymbol{r}$ and $\boldsymbol{q}$ is the position and the quaternion of the arm end-effector in the world frame. $R^{\mkern-1.5mu\mathsf{T}}(\boldsymbol{q})$ is the rotation matrix from the end-effector frame to the body frame. $\Omega(\cdot)$ is the matrix operator of the angular velocity for quaternion multiplication. Define the robot state and control input
$\boldsymbol{x}_{t}=[\boldsymbol{r}_{t},\ \boldsymbol{q}_{t}]\in\mathbb{R}^{7},\quad\boldsymbol{u}_{t}=[\boldsymbol{v}^{d}_{t},\ \boldsymbol{w}^{d}_{t}]\in\mathbb{R}^{6},$ (36)
respectively.
The above end-effector dynamics is discretized with $\triangle t=0.1\mathrm{s}$. In (1), the control horizon is $T=20$. Let $\boldsymbol{x}^{*}=[\boldsymbol{r}^{*}\ \boldsymbol{q}^{*}]$ denote the target end-effector pose. The step cost $c(\boldsymbol{x}_{t},\boldsymbol{u}_{t})$ and terminal cost $h(\boldsymbol{x}_{T})$ in (1) is set to
$c(\boldsymbol{x}_{t},\boldsymbol{u}_{t})=$ $(\boldsymbol{r}_{t}-\boldsymbol{r}^{*})^{\mkern-1.5mu\mathsf{T}}Q_{r}(\boldsymbol{r}_{t}-\boldsymbol{r}^{*})+k_{q}\mathbf{diff}(\boldsymbol{q}_{t},\boldsymbol{q}^{*})$ (37)
$\quad\quad+\boldsymbol{v}_{I,t}^{\mkern-1.5mu\mathsf{T}}R_{v}\boldsymbol{v}_{I,t}+\boldsymbol{w}_{I,t}^{\mkern-1.5mu\mathsf{T}}R_{w}\boldsymbol{w}_{I,t},$
$h(\boldsymbol{x}_{T})=$ $(\boldsymbol{r}_{T}-\boldsymbol{r}^{*})^{\mkern-1.5mu\mathsf{T}}S_{r}(\boldsymbol{r}_{T}-\boldsymbol{r}^{*})+\mu_{q}\mathbf{diff}(\boldsymbol{q}_{T},\boldsymbol{q}^{*}).$
In this game, the cost parameters are set as $Q=\mathrm{diag}(2,2,2)$, $k_{q}=1$, $R_{v}=\mathrm{diag}(30,30,1)$, $R_{w}=\mathrm{diag}(0.85,0.85,0.85)$, $S_{r}=\mathrm{diag}(80,80,80)$, and $\mu_{q}=60$. The output of the MPC controller is sent to the low-level controller to track.
Similar to Section VI-B, to constrain $r_{z}$ and $\boldsymbol{q}$ for proper height and tilt in order to pass through the slot. We use a set of $N=20$ Gaussian radial basis function (RBFs) to parameter the safety constraint:
${g}_{\boldsymbol{\theta}}(\boldsymbol{\xi})=\phi_{0}+\sum\nolimits_{i=1}^{N}\theta_{i}\phi_{i}(\boldsymbol{\xi}),$ (38)
with
$\phi_{i}(\boldsymbol{\xi})=\sum\nolimits_{t=0}^{T}\beta_{t}\mathrm{RBF}_{i}(\boldsymbol{x}_{t}),\quad\text{and}$ (39)
$\mathrm{RBF}_{i}(\boldsymbol{x}_{t}){=}\begin{cases}\sigma(r_{x,t})k_{z}\exp{(-\epsilon_{z}^{2}(r_{z,t}-c_{z,i})^{2})},\,\,i{=}1,...,10\\ \sigma(r_{x,t})\exp{(-\epsilon_{q}^{2}\mathbf{diff}(\boldsymbol{q}_{t},\boldsymbol{q}_{c,i}))},\,\,i{=}11,...,20\end{cases}.$ (40)
Here, $k_{z}$ is used to balance the scaling between position and orientation $\mathrm{RBF}$s, $\sigma(r_{x,t})$ is defined as
$\sigma(r_{x,t})=\mathrm{sigmoid}(50\big{(}(r_{x,t}-x_{\text{slot}})^{2}-(\frac{w}{2})^{2}\big{)}),$ (41)
which is used to constrain the $\mathrm{RBF}_{i}$ only taking effect around $x$ range $[x_{\text{slot}}-\frac{w}{2},x_{\text{slot}}+\frac{w}{2}]$ (recall slot’s $x$ location, slot, is a approximate value). In the above definition, $w=0.3$, $k_{z}=0.05$, $\epsilon_{z}=12$ and $\epsilon_{q}=1.8$. The center of positional $\mathrm{RBF}$, ${c_{z,i}}$, is evenly chosen on interval $[0,1]$, and the center of orientation $\mathrm{RBF}$, ${\boldsymbol{q}_{c,i}}$, is generated by $(\cos{\frac{\alpha}{2}},\sin{\frac{\alpha}{2}},0,0)*\boldsymbol{q}_{\text{init}}$ with $\alpha$ evenly choosen from $[-\pi/4,3\pi/4]$ to make sure ${\boldsymbol{q}_{c,i}}$ covers all possible orientations. $\beta_{t}=0.9^{t}$ if $t<5$ and 00 otherwise. The bounds of the initial hypothesis space $\Theta_{0}$ is set to $\boldsymbol{c}_{l}=[-3,\dots,-3]^{\mkern-1.5mu\mathsf{T}}$ and $\boldsymbol{c}_{h}=[5,\dots,5]^{\mkern-1.5mu\mathsf{T}}$. $\gamma$ is set to 1.5. This initialization leads to a conservative MPC policy, preventing the arm from approaching the target at the beginning.
VI-C2 Human Correction Interface
The motion of the robot arm is rendered in the free-view mode in MuJoCo visualization (Fig. 12), and users can drag to select his/her comfortable view. The slot and the random reaching target are visible to the users.
Human users apply their corrections to the robot arm via a keyboard. Users are informed with the following key mapping between the key press and the corresponding correction in the game. The safety constraint (38) in Franka arm’s safe MPC policy is updated immediately once a user correction is detected. When user press "ENTER", the arm is reset to the initial position and the reaching target will also change.
TABLE II: keyboard commands in robot arm user study
Key Correction Interpretation
"left" $[0,0,0,0,0,1]^{\mkern-1.5mu\mathsf{T}}$ Clock-wise rotation of the gripper
"right" $[0,0,0,0,0,-1]^{\mkern-1.5mu\mathsf{T}}$ Anti-clock-wise rotation of the gripper
"shift" $[0,0,0,-1,0,0]^{\mkern-1.5mu\mathsf{T}}$ Go up in world frame
"ctrl" $[0,0,0,1,0,0]^{\mkern-1.5mu\mathsf{T}}$ Go down in world frame
"Enter" N/A Reset robot position and switch target.
(a) After 2 corrections
(b) After 6 corrections
(c) After 12 corrections
Figure 13: Illustrations of the drone MPC rollout trajectory after $k=2,6,12$ corrections. The red stars denote the target positions and the blue curves denote the trajectories. The gate is drew in black lines and collision is denoted by green crosses. User’s corrections typically lower the flying drone to make it pass through the gate.
VI-D Participants, Procedures, and Metrics
VI-D1 Participants
We recruited 15 participants from the student population of Arizona State University, covering undergraduate, master’s, and PhD students with different backgrounds. Among all participants, 13 are male and 2 female. All participants are novices to the user study. This user study had been reviewed and approved by the Institutional Review Board (IRB) of Arizona State University, and all participants had signed the consent forms.
VI-D2 Procedure
Each participant was instructed to play the previously developed two simulation games. In each game, each participant plays 10 trials.
In the drone navigation game, a trial is deemed successful when the drone navigates through the gate without collision and successfully reaches different targets 6 consecutive times with the learned safe MPC policy. A trial is deemed failed when there are too many corrections, causing the crash of the MVE solver, or the user believes the robot cannot be corrected anymore and gives up.
In the Franka reaching game, a trial is deemed successful when the robot arm reaches different targets without collision with the slot frame with the learned safe MPC controller. The failure of one trial is determined when there are too many corrections or the user believes the robot cannot be corrected anymore and gives up.
The result of every trial is recorded. Before formally starting the study, every user is instructed about the settings of the game and is given 2 chances for hands-on trying of the environments to get familiar with the keys and commands.
VI-D3 Metrics for Evaluation
Two major metrics are used to evaluate the performance of the proposed safe MPC alignment algorithm in our user study:
  • (I)
    Success rate: The overall and individual success rate suggest the effectiveness of the proposed method.
  • (II)
    Numbers of corrections needed for a successful trial: The averaged number of corrections performed by the users in successful trials is a measure for the efficiency of the proposed method. Fewer corrections in successful trials suggest better efficiency in teaching the robot.
VI-E Result and Analysis
VI-E1 User Study Metrics Report
TABLE III: Result of Simulated User Study
Environment Success Rate Correction count
Drone navigation $89.3\%\pm 7.7\%$ $15.7\pm 10.6$
Franka arm reaching $89.3\%\pm 11.2\%$ $16.5\pm 6.1$
For the user study, we report the success rate and correction count in a successful game trial in Table III. Both metrics are averaged across individual users. The success rate is $89.3\%\pm 7.7\%$ for the drone navigation game and $89.3\%\pm 11.2\%$ for the Franka arm constrained reaching game. The number of corrections for a successful trial is $15.7\pm 10.6$ for the drone navigation game and $16.5\pm 6.1$ Franka arm reaching games.
The above result illustrates that the proposed method is effective (i.e., with a high success rate) in enabling a robot to learn a safety constraint for varying tasks (i.e. reaching targets are changing). It also shows the efficiency of the proposed method — successful learning of a safety constraint only requires tens of human corrections. Since this method is online and the user can perform multiple corrections in one robot rollout trajectory, the constraint can be learned to guide the robot to complete safety-critical tasks in a few runs. In Table III, one can note that the drone navigation task has a higher variance in the correction count than the Franka arm reaching task. This could be attributed to the larger motion range and longer task rollout horizon of drone navigation. As a result, in the drone game, the human correction strategies, including the correction position and correction timing, can vary significantly. Such diverse correction strategies can lead to a larger variance of total correction count.
VI-E2 Analysis for Learned Behavior and Learned Constraint
To demonstrate the learning progress, in Fig. 13 and Fig. 14, we illustratively show the intermediate robot motion at different numbers of human corrections. Fig. 13 shows the flight trajectory of the drone after different numbers of human corrections. Fig. 14 shows the position and altitude of the robot arm’s end effector when passing the slot after the different number of human corrections. From the illustration examples, we can conclude that the learned constraint generally guides the robot to complete the task.
Figure 14: Illustrations of the robot arm pose when passing the $yz$-plane of the bars after $k=4,8,17$ corrections. (a): The motion after 4 corrections with three different targets (cyan ball). (b): The motion after 8 corrections. (c): The motion after 17 corrections. The ends of the gripper are denoted with red balls. At the beginning, the gripper cannot pass the slot. But after several corrections, it succeeds in passing the slot for all target positions.
To better understand the learned safety constraint function ${g}_{\boldsymbol{\theta}}(\boldsymbol{\xi})$, in Fig. 15, we draw the heat map of the value of ${g}_{\boldsymbol{\theta}_{H}}(\boldsymbol{\xi})$ with $\boldsymbol{\theta}_{H}$ the averaged value of weights learned in all successful trials. In drawing such heat map, since ${g}_{\boldsymbol{\theta}_{H}}$ is defined by RBFs spreading only on $yz$-plane, we fix $r_{x,t}=10$ (recall $x_{\text{gate}}=9.8$) and $r_{y,t},r_{z,t}$ are generated by the grid in $[0,10]\times[0,10]$ for all states $\boldsymbol{x}_{t}$ in $\boldsymbol{\xi}$. In the heat map, we draw the 00-level contour of ${g}_{\boldsymbol{\theta}_{H}}$ in the dashed line, which is the learned safety boundary. Also, we plot the $yz$-plane projection of the gate, which is the black box.
In Fig. 15, we can note the learned safety boundary (corresponding to ${g}_{\boldsymbol{\theta}_{H}}(\boldsymbol{\xi})\leq 0$) is only similar to the top edge of the gate. This is explained by the following reasons. As shown in Fig. 11, the target positions are set higher than the gate, and thus during the learning process, the motion of the drone is always going up, as shown in Fig. 13. As a result, the majority of human directional corrections happen when the drone approaches the top edge of the gate, leaving the learned safety boundary accurate near the top edge of the gate and bottom boundary untouched. It is expected that in order to learn a safety constraint that fully represents the gate frame, lower target positions should be added so that the drone will receive human corrections near the bottom of the gate.
Figure 15: Heatmap for the learned constraint function in the drone user study. The black solid line denotes the actual position of the gate in $yz$-frame. The black dashed curve denotes the boundary of the safety zone. Since the target positions are higher than the gate (see Fig. 11), thus during the learning process, the motion of the drone is always going up, as shown in Fig. 13. As a result, the majority of human directional corrections happen when the drone approaches the top edge of the gate, leaving the learned safety boundary accurate near the top edge of the gate and bottom boundary untouched.
VII User Study in Real-world Robot Experiment
In this section, we conduct a user study to evaluate the performance of our safe MPC alignment algorithm using a hardware Franka Research 3 robot experiment. We focus on the task of mobile liquid pouring, i.e., the robot performs liquid pouring while moving from one location to another. In certain scenarios, this kind of task can be safety-critical: for example, if the liquid is corrosive/chemical/toxic, inappropriate mobile pouring will cause safety concerns. The task is challenging because it is hard to manually define a safety constraint for the robot to accomplish liquid pouring without spilling. The Franka robot arm will learn such a safe MPC controller from human physical corrections.
VII-A Hardware Experiment Setup
[…]
We specifically focus on mobile water pouring by the Franka robot arm. A task setup is shown in Fig. 16. A water pot is gripped by the robot end-effector. The robot needs to move from the starting point $[0.22,-0.5,0.48]^{\mkern-1.5mu\mathsf{T}}$ (with orientation quaternion $[0,-0.707,0,0.707]^{\mkern-1.5mu\mathsf{T}}$) to target point $[-0.65,-0.5,0.5]^{\mkern-1.5mu\mathsf{T}}$ (with orientation quaternion $[0,-0.707,0,0.707]^{\mkern-1.5mu\mathsf{T}}$). During the motion, the robot needs to pour more than 50 grams of water into a bar sink. The bar sink is approximately placed in the middle of the robot’s moving path, but the precise size and location of the bar sink are unknown.
Figure 17: Components of the hardware Franka arm learning platform
As shown Fig. 17, the hardware Franka arm platform has three main components: the parameterized safe MPC controller (high-level controller), the low-level controller, and a learning module. The high-level safe MPC controller follows the safe MPC formulation of Franka arm reaching game in Section VI-C (recall that the unknown constraint is parameterized by RBFs). The other control parameters of this Safe MPC controller are listed in Table IV. The high-level safe MPC controller receives the end-effector’s current pose $\boldsymbol{x}$ from the low-level controller and solves for the current desired velocity $\boldsymbol{u}$ (defined in (36)) of the end-effector. The low-level controller implements the operation space velocity control to track the desired velocity $\boldsymbol{u}$ given by the high-level controller.
The learning module updates the parameterize $\boldsymbol{\theta}$ of the high-level safe MPC controller. The update only happens when the learning module detects a human correction $\boldsymbol{a}$, which is always accompanied by an end-effector pose $\boldsymbol{x}_{\boldsymbol{a}}$ where such human correction $\boldsymbol{a}$ is applied to. The learning module implements the proposed safe MPC alignment algorithm.
TABLE IV: parameter list in real-world user study
Parameter Value
$Q_{r}$ $\mathrm{diag}(0.8,0.8,0.8)$
$k_{q}$ 0
$R_{v}$ $\mathrm{diag}(30,30,1)$
$R_{w}$ $\mathrm{diag}(0.85,0.85,0.85)$
$S_{r}$ $\mathrm{diag}(80,80,80)$
$\mu_{q}$ 80
In the high-level MPC policy, the cost of the MPC controller is set to penalize the current and target gripper position in Cartesian space, as in Section VI-C. Since it is difficult to explicitly define a parameterized constraint using the state of pouring water, we instead define the constraints of safe water pouring using the robot state. Specifically, we define the same RBF (40) parameterized constraint in Section VI-C, constraining the tilt angle and height of the robot gripper for pouring the water into the sink.
Figure 18: Illustrations of human correction in the real-world user study. (a) A human user applies force to the robot. (b) The human moves the robot to the desired configuration to pour the water into the bar sink. The delta position of the end effector is used as human directional correction, as in (42).
As shown in Fig. 18, to give a physical correction to the Franka robot arm, a human user can physically touch/hold any link of the robot arm, and the applied force will immediately activate the zero-torque control mode. Suppose the joint positions of the robot arm before and after human physical correction are $\mathbf{q}_{\text{before}}$ and ${\mathbf{q}_{\text{after}}}$, respectively, we calculate the corresponding raw human correction $\boldsymbol{a}_{r}$ using
$\boldsymbol{a}_{r}=\mathrm{J}(\mathbf{q}_{\text{before}})({\mathbf{q}}_{\text{after}}-\mathbf{q}_{\text{before}}),$ (42)
with $\mathrm{J}$ the Jacobian of the end-effector. The above calculated $\boldsymbol{a}_{r}\in\mathbb{R}^{6}$ is considered as the directional displacement of the end-effector pose, intended by the human user. In the computed $\boldsymbol{a}_{r}\in\mathbb{R}^{6}$, we mask the entries in the other dimensions with 00s and only keep the entries corresponding to the tilt angle along $x$-axis and the height in $z$-axis, since they are the only dimensions relevant to the given task. The masked vector are is as the actual human correction $\boldsymbol{a}$ fed to the learning module.
(a) After 2 Corrections
(b) After 3 Corrections
(c) After 5 Corrections
Figure 19: Illustrations of one trial of the real-world study. The trajectories of the watering pot is denoted by blue dashed lines. (a) After 2 corrections, the arm begins tilting the pot but spills out the water. (b) After 3 corrections, the arm can pour water into the sink but the human is not satisfied with its height and altitude of pouring. (c) After 5 corrections, the arm learns to tilt only along the direction of the $x-$axis and successfully aligns with human intent.
VII-B Participants, Procedures, and Metrics
All participants recruited for the two simulation games in Section VI-D also take part in the hardware experiment user study. They are told of the objective before starting trials. Similar to part VI-C, one episode is defined as one robot motion from the starting gripper position to the target position. One trial means one run of the control with multiple episodes (with a maximum of 15 corrections). During corrections, the nozzle of the water pot is sealed, and we switch to the unsealed nozzle and measure the amount of water poured into the flower pot after the user is satisfied with the motion of the robot motion. The success trial is defined when the amount of poured water is over 50g. Each user performs 5 trials, and the same metrics in Section VI-D are recorded.
VII-C Result and Analysis
In real-world robot experiments, the success rate is $84.0\%\pm 16.7\%$ and the average number of corrections in successful trials is $6.88\pm 3.26$. The results show the efficacy and efficiency of the proposed method in real-world robot learning applications. To illustrate the robot learning progress, in Fig. 19, we use one trial of the user study to show the robot behavior at different numbers of human corrections.
Since the task of this real-world Franka arm experiment is similar to the simulated Franka arm reaching game in Section VI-C, we compare the user study results for both and have the following conclusion and analysis.
(I) The real-world experiment has a slightly lower success rate: While the difference between simulations and real-world experiments is small in success rate, we have identified some factors during our experiment that may contribute to such lower success rate, which includes both the human factors and hardware factors. On human factors, most of the participants have no experience or knowledge in operating robot hardware, and thus when directly interacting with hardware, their corrections can be conservative or ineffective due to safety considerations. On the hardware side, some error sources, such as noisy correction signals, and the possible timing lag between the recorded correction-happening robot state and the human intended correction-happening state, can contribute to the failure of the learning.
(II) The real-world experiment has a less number of human corrections: This could be attributed to two reasons. First, the constraint needed for accomplishing the water pouring task in the real-world experiment could be more relaxed than the constraint for accomplishing the narrow slot passing task in the simulation game in Section VI-C. This makes the safety constraints easy to learn for the real-world task. Second, different from simulation games, where human users can only perceive the motion of the robot arm from the screen, a human user has a better perception of the robot’s motion in a real-world experiment. A good perception of the robot’s motion could help humans give more effective feedback each time, reducing the total number of human interactions for successful robot learning.
VII-D Discussion
Based on the observations in both the hardware and simulation user studies, we discuss some design choices and potential limitations of the proposed safe MPC alignment algorithm.
VII-D1 Learning Task-Independent Safety Constraints
In the real-world and simulation experiments, due to the insufficiency of task sampling, the learned safety constraint can be task-dependent, instead of task-independent. For instance, in the drone navigation game in Section VI-B, since the three task targets are placed above the gate, the learned constraint shown in Fig. 15 can only reflect the upper edge of the gate constraints. One limitation of this is that the learned task constraints have poor generalization to new tasks, which are far from previously encountered. For example, in drone navigation, if a new target is placed below the gate, the drone with the learned safety constraint Fig. 15 will fail.
To improve the generalization of the learned safety constraints or learning task-independent safety constraints, one solution is to increase the diversity of task samples. By doing so, it is expected that the robot can receive human correction on the entire boundary of the safety region, making the learned safety constraints task irrelevant.
VII-D2 Robustness Against Human Irrationality
Our proposed safe MPC alignment method is based on the cutting hypothesis. In our formulation, we define a hypothesis cut (7) based on an average of multiple human corrections, such an averaging treatment is to mitigate human mistakes for one cut. But still, in practice, human irrationality can lead to the wrong cut of the hypothesis space, which constitutes the main reason for failure in user studies. For example, in the drone navigation game, we found that failure game trials usually happen when a human gives too many wrong directional corrections at the beginning of the trial. A failure case is shown in Fig. 20.
Figure 20: One illustration of the failure case in the drone user study. A human user makes a false "up" correction at the beginning and fails to correct the motion of the drone.
To further improve the robustness of the method against human mistakes, two extensions of the current method can be used on top of our proposed algorithm. The first is to add a relaxation factor $\epsilon\geq 0$ in our cutting hypothesis. For example, we can re-write (7) as
$\forall\boldsymbol{\theta}_{H}\in\bar{\Theta}_{H},\ \operatorname{\mathbb{E}}_{\mathbf{a}}\left\langle\mathbf{a},-\nabla B(\boldsymbol{\xi}_{\boldsymbol{\theta}}^{\gamma},\boldsymbol{\theta}_{H})\right\rangle+\epsilon\geq 0.$ (43)
Another extension is to add a selection mechanism on human correction. For example, one can use some statistical method to distinguish and rule out mistake corrections before performing a hypothesis cut. We will leave those possible extensions in our future work.
VIII Conclusion
This paper proposes safe MPC alignment method to online learn a safety constraint in robot safe MPC controllers from human directional feedback. The method, named hypothesis space cutting, has geometric interpretation: it exponentially cuts the hypothesis space based on each received human directional feedback. The method has a provable guarantee for the upper bound of the total number of human feedback, and also can declare hypothesis misspecification. A numeric example, two user studies on simulation games, and a user study on the real-world robot arm, covering varying robot tasks, were used to validate the effectiveness and efficiency of the proposed approach. The results show that the proposed method can learn safety constraints effectively and efficiently. To our knowledge, this is the first certifiable method to enable online learning of safety constraints with human feedback.
Appendix A Proof of lemmas and theorems
A-A Proof of Lemma 2
Proof.
We prove this lemma by induction. First, we have $\bar{\Theta}_{H}\subset\Theta_{0}$. Suppose $\bar{\Theta}_{H}\subset\Theta_{i-1}$, $i=1,2,3,...$. Due to the condition $\bar{\Theta}_{H}\subset\mathcal{C}_{i}$ in (12), following the update (13), we have $\bar{\Theta}_{H}\subset\Theta_{i}=\Theta_{i-1}\cap\mathcal{C}_{i}$, which completes the proof. ∎
A-B Proof of Lemma 4
Proof.
For any $i=1,2,3,...$, the plan $\boldsymbol{\xi}_{\boldsymbol{\theta}_{i}}^{\gamma}=\{\boldsymbol{u}^{\boldsymbol{\theta}_{i},\gamma}_{0:T-1},\boldsymbol{x}^{\boldsymbol{\theta}_{i},\gamma}_{0:T}\}$ is a solution to the optimization (3), with $\boldsymbol{\theta}=\boldsymbol{\theta}_{i}$. Recall Lemma 1, $g_{\boldsymbol{\theta}_{i}}(\boldsymbol{\xi}_{\boldsymbol{\theta}_{i}}^{\gamma})<0$. Extending the feature-weight expression of $g_{\boldsymbol{\theta}_{i}}(\boldsymbol{\xi}_{\boldsymbol{\theta}_{i}}^{\gamma})$, one can get $\boldsymbol{\theta}_{i}^{\mkern-1.5mu\mathsf{T}}\boldsymbol{\phi}(\boldsymbol{\xi}_{\boldsymbol{\theta}_{i}}^{\gamma})<-\phi_{0}(\boldsymbol{\xi}_{\boldsymbol{\theta}_{i}}^{\gamma})$. Since the optimal solution $\boldsymbol{u}^{\boldsymbol{\theta}_{i},\gamma}_{0:T-1}$ is a minimizer to (3), it means
$\nabla B(\boldsymbol{\xi}_{\boldsymbol{\theta}_{i}}^{\gamma},\boldsymbol{\theta}_{i})=0.$ (44)
Note that with $g_{\boldsymbol{\theta}_{i}}(\boldsymbol{\xi}_{\boldsymbol{\theta}_{i}}^{\gamma})<0$, the above $\nabla B(\boldsymbol{\xi}_{\boldsymbol{\theta}_{i}}^{\gamma},\boldsymbol{\theta}_{i})$ is well defined. With (44),
$\mathbf{a},\nabla B(\boldsymbol{\xi}_{\boldsymbol{\theta}_{i}}^{\gamma},\boldsymbol{\theta}_{i})\right>=0,$ (45)
always holds. Extending the above (45) leads to
$\mathbf{a},\nabla J(\boldsymbol{\xi}_{\boldsymbol{\theta}_{i}}^{\gamma})\right>-\gamma\frac{1}{g_{\boldsymbol{\theta}_{i}}(\boldsymbol{\xi}_{\boldsymbol{\theta}_{i}}^{\gamma})}\left<\mathbf{a},\nabla g_{\boldsymbol{\theta}_{i}}(\boldsymbol{\xi}_{\boldsymbol{\theta}_{i}}^{\gamma})\right>=0.$ (46)
Due to $-g_{\boldsymbol{\theta}_{i}}(\boldsymbol{\xi}_{\boldsymbol{\theta}_{i}}^{\gamma})>0$, one can multiple $-g_{\boldsymbol{\theta}_{i}}(\boldsymbol{\xi}_{\boldsymbol{\theta}_{i}}^{\gamma})$ on both sides of the above equality, which leads to $\boldsymbol{\theta}_{i}^{\mkern-1.5mu\mathsf{T}}\boldsymbol{h}_{i}=b_{i}$ based on the definition (16). The second term in (19) directly follows $g_{\boldsymbol{\theta}_{i}}(\boldsymbol{\xi}_{\boldsymbol{\theta}_{i}}^{\gamma})<0$. This concludes the proof. ∎
A-C Proof of Lemma 5
Proof.
By Lemma 4, the MVE center $\boldsymbol{\theta}_{i}=\bar{\boldsymbol{d}}_{i-1}$ is on the cutting hyperplane $\boldsymbol{\theta}^{\mkern-1.5mu\mathsf{T}}\boldsymbol{h}_{i}=b_{i}$. Directly from [62], one has
$\frac{\mathbf{Vol}(\Theta_{i-1}\cap\{\boldsymbol{\theta}|\boldsymbol{\theta}^{\mkern-1.5mu\mathsf{T}}\boldsymbol{h}_{i}\leq b_{i}\})}{\mathbf{Vol}(\Theta_{i-1})}\leq 1-\frac{1}{r}.$ (47)
Since
$\mathbf{Vol}(\Theta_{i-1}\cap\{\boldsymbol{\theta}|\boldsymbol{\theta}^{\mkern-1.5mu\mathsf{T}}\boldsymbol{h}_{i}\leq b_{i}\})\geq\mathbf{Vol}(\Theta_{i-1}\cap\mathcal{C}_{i})=\mathbf{Vol}(\Theta_{i}),$ (48)
based on (12), (13), one can get
$\frac{\mathbf{Vol}(\Theta_{i})}{\mathbf{Vol}(\Theta_{i-1})}\leq 1-\frac{1}{r}.$ (49)
Which concludes the proof. ∎
A-D Proof of Theorem 1
Proof.
We complete the proof in two parts. First, we prove the number of learning iterations is less than $K$, then we use this to prove there will be a $\boldsymbol{\theta}_{k}\in\bar{\Theta}_{H}$, $k=0,1,2...,K-1.$
We first prove the total number of iterations must be less than $K$ by contradiction. Suppose the learning iteration number $i\geq K$. With recursively applying lemma 5, one can obtain
$\mathbf{Vol}(\Theta_{i})\leq(1-\frac{1}{r})^{i}\mathbf{Vol}(\Theta_{0}).$ (50)
When $i\geq K$,
$\mathbf{Vol}(\Theta_{i})<\frac{\tau_{r}\rho_{H}^{r}}{\mathbf{Vol}(\Theta_{0})}\mathbf{Vol}(\Theta_{0})=\tau_{r}\rho_{H}^{r}\leq\mathbf{Vol}(\bar{\Theta}_{H})$ (51)
However, from Lemma 2, one has $\bar{\Theta}_{H}\subset\Theta_{i}$, meaning $\mathbf{Vol}(\bar{\Theta}_{H})\leq\mathbf{Vol}(\Theta_{i})$. This leads to a contradiction with (51). Therefore, the total number of iterations must be less than $K$.
Next, we prove $\exists k<K$ such that $\boldsymbol{\theta}_{k}\in\bar{\Theta}_{H}$. In contradiction, suppose $\forall k<K$, $\boldsymbol{\theta}_{k}\notin\bar{\Theta}_{H}$. By Lemma 2, this means that for all $i<K$, $\bar{\Theta}_{H}\subset\mathcal{C}_{i}$. Then, the algorithm will continue for $i=K$. Consequently, the total number of iterations can reach $K$, which contradicts that the algorithm will terminate within $K$ iterations. This concludes the proof. ∎
A-E Proof of Theorem 2
Proof.
We complete the proof in two parts. First we prove the algorithm converges to $\Theta^{*}$ which has $\mathbf{Vol}(\Theta^{*})=0$. Based on this, we next prove $\Theta^{*}\cap\mathbf{int}\ \Theta_{0}=\emptyset$.
We first prove the algorithm converges to $\Theta^{*}$ which has $\mathbf{Vol}(\Theta^{*})=0$. Given the condition $\Theta_{0}\cap\bar{\Theta}_{H}=\emptyset$, the algorithm will never terminate because $\forall i=1,2,3,...\infty$, $\mathcal{C}_{i}\supset\bar{\Theta}_{H}$ due to Lemma 2. Since the volume of the initial hypothesis space is finite, by applying Lemma 5 we have
$\mathbf{Vol}(\Theta^{*})=\lim_{i\rightarrow\infty}(1-\frac{1}{r})^{i}\mathbf{Vol}(\Theta_{0})=0.$ (52)
Next, we prove $\Theta^{*}\cap\mathbf{int}\ \Theta_{0}=\emptyset$ by contradiction. Suppose there exists $\widehat{\boldsymbol{\theta}}\in\Theta^{*}\cap\mathbf{int}\ \Theta_{0}$. Define the ball contained in $\bar{\Theta}_{H}$ to be $\mathcal{B}({\boldsymbol{\theta}^{c}_{H}},{\rho_{H}})=\{\boldsymbol{\theta}\ |\ \left\lVert\boldsymbol{\theta}-\boldsymbol{\theta}_{H}^{c}\right\rVert\leq\rho_{H}\}\subseteq\bar{\Theta}_{H}$ with $\boldsymbol{\theta}_{H}^{c}$ as the center. By the equation (13), we can note
$\Theta^{*}=\Theta_{0}\cap{\mathcal{C}},$ (53)
with ${\mathcal{C}}=\bigcap_{i=1}^{\infty}\mathcal{C}_{i}$ a convex set. So we have
$\Theta^{*}\subseteq\mathcal{C}.$ (54)
By equation (12),
$\mathcal{B}_{H}\subseteq\bar{\Theta}_{H}\subseteq\mathcal{C}.$ (55)
From the relationship between sets above, $\boldsymbol{\theta}_{H}^{c}\in\mathcal{C},\widehat{\boldsymbol{\theta}}\in\mathcal{C}$.
Since $\boldsymbol{\theta}_{H}^{c}\notin\Theta_{0}$, there exists $\mu>0$ such that
$\widetilde{\boldsymbol{\theta}}=\widehat{\boldsymbol{\theta}}+\mu(\boldsymbol{\theta}_{H}^{c}-\widehat{\boldsymbol{\theta}})\in\mathbf{int}\ \Theta_{0}.$ (56)
Then, one can define a ball $\mathcal{B}(\widetilde{\boldsymbol{\theta}},d)$ centered at $\widetilde{\boldsymbol{\theta}}$ with radius
$d=\min_{\boldsymbol{\theta}\in\mathbf{bd}\ \Theta_{0}}\left\lVert\widetilde{\boldsymbol{\theta}}-\boldsymbol{\theta}\right\rVert.$ (57)
Thus, $\mathcal{B}(\widetilde{\boldsymbol{\theta}},d)\subset\mathbf{int}\ \Theta_{0}$.
Next, we will show $\mathcal{B}(\widetilde{\boldsymbol{\theta}},\mu\rho_{H})\subset\mathcal{C}$. In fact, for any $\boldsymbol{\theta}$ satisfying $\left\lVert\boldsymbol{\theta}-\widetilde{\boldsymbol{\theta}}\right\rVert\leq\mu\rho_{H}$, one has
$\begin{split}\boldsymbol{\theta}&=\mu(\frac{1}{\mu}(\boldsymbol{\theta}-\widetilde{\boldsymbol{\theta}})+\boldsymbol{\theta}_{H}^{c})+\widetilde{\boldsymbol{\theta}}-\mu\boldsymbol{\theta}_{H}^{c}\\ &=\mu\underbrace{(\frac{1}{\mu}(\boldsymbol{\theta}-\widetilde{\boldsymbol{\theta}})+\boldsymbol{\theta}_{H}^{c})}_{\in\mathcal{B}_{H}\subseteq\mathcal{C}}+(1-\mu)\underbrace{\widehat{\boldsymbol{\theta}}}_{\in\Theta^{*}\subseteq\mathcal{C}}\end{split}.$ (58)
The convexity of $\mathcal{C}$ leads to $\boldsymbol{\theta}\in\mathcal{C}$, which means the ball $\mathcal{B}_{\mu}\triangleq\{\boldsymbol{\theta}\ |\ \left\lVert\widetilde{\boldsymbol{\theta}}-\boldsymbol{\theta}\right\rVert\leq\mu\rho_{H}\}\subset\mathcal{C}$.
Let a new ball be $\mathcal{B}(\widetilde{\boldsymbol{\theta}},\min{(\mu\rho_{H},d)})$. Based on the previous two results of $\mathcal{B}(\widetilde{\boldsymbol{\theta}},d)\subset\mathbf{int}\Theta_{0}$ and $\mathcal{B}(\widetilde{\boldsymbol{\theta}},\mu\rho_{H})\subset\mathcal{C}$, one has $\mathcal{B}(\widetilde{\boldsymbol{\theta}},\min{(\mu\rho_{H},d)})\subset(\Theta_{0}\cap\widetilde{\mathcal{C}})=\Theta^{*}$. As a result,
$\mathbf{Vol}(\Theta^{*})\geq\mathbf{Vol}\,\,\mathcal{B}(\widetilde{\boldsymbol{\theta}},\min{(\mu\rho_{H},d)})>0$ (59)
which contradicts $\mathbf{Vol}(\Theta^{*})=0$ from (52). Thus, $\Theta^{*}\cap\mathbf{int}\ \Theta_{0}=\emptyset$ must hold. This completes the proof. ∎

### v2

@@
 In safety-critical robot planning or control, manually specifying safety constraints or learning them from demonstrations can be challenging.
-In this paper, we propose a certifiable alignment method for a robot to learn a safety constraint in its model predictive control (MPC) policy with human online directional feedback.
+In this article, we propose a certifiable alignment method for a robot to learn a safety constraint in its model predictive control (MPC) policy with human online directional feedback.
 To our knowledge, it is the first method to learn safety constraints from human feedback.
@@
 The method only requires the direction of human feedback to update the learning hypothesis space.
-It is certifiable, providing an upper bound on the total number of human feedback in the case of successful learning of safety constraints, or declaring the misspecification of the hypothesis space, i.e., the true implicit safety constraint cannot be found within the specified hypothesis space.
-We evaluated the proposed method using numerical examples and user studies in two developed simulation games.
-Additionally, we implemented and tested the proposed method on a real-world Franka robot arm performing mobile water-pouring tasks in a user study.
-The simulation and experimental results demonstrate the efficacy and efficiency of our method, showing that it enables a robot to successfully learn safety constraints with a small handful (tens) of human directional corrections.
+It is certifiable, providing an upper bound on the total number of human feedback in the case of successful learning, or declaring the hypothesis misspecification, i.e., the true implicit safety constraint cannot be found within the specified hypothesis space.
+We evaluated the proposed method using numerical examples and user studies in two simulation games.
+Additionally, we implemented and tested the proposed method on a real-world Franka robot arm performing mobile water-pouring tasks.
+The results demonstrate the efficacy and efficiency of our method, showing that it enables a robot to successfully learn safety constraints with a small handful (tens) of human directional corrections.
 Index Terms:
@@
 I Introduction
-In developing robot control policies, ensuring safety is always the top priority.
-This is especially crucial for safety-critical applications, such as autonomous vehicles or collaborative robots operating near human users, where violating safety constraints can lead to irreversible loss or damage.
-While a significant line of work in safe control [1] and safe reinforcement learning [2] focuses on the safe policy design given robot behavior constraints, less attention has been paid to how to specify the safety constraints in the first place.
-In fact, for many daily tasks and applications, specifying proper safety constraints can be challenging due to two main reasons.
-First, explicit safety boundary for robot behavior can be hard to model or define.
-For example, in the task of a household robot feeding an elderly person, the complex nature of the task makes it very difficult to define a precise constraint for safe feeding behavior.
-Second, safety specifications can vary among users and require personalization.
-An example is lane-switching of an autonomous vehicle: aggressive drivers may prefer frequent lane-switching behavior, while conservative users may consider frequent lane-switching unsafe.
-The challenges in safety specifications lead to two key questions, which we seek to answer in this paper.
-Q1: can a robot learn and align its safety constraints with human intent in an intuitive and online manner?
-Q2: can such learning process to be user-data efficient such that a robot is able to quickly adapt to different users?
-Recent work [3, 4, 5] has proposed learning safety constraints from human demonstrations.
-However, these approaches face limitations that prevent them from fully addressing the two key questions.
-First, obtaining valid user demonstrations can be costly and time-consuming, especially with high-dimensional robot systems.
-Second, demonstration-based learning typically follows an offline or episodic learning procedure, which does not allow for on-the-fly interaction where human feedback and robot updates occur in real time.
-Figure 1: Overall illustration of the proposed method.
-The proposed safe MPC alignment online updates the safety constraint in the robot safe MPC policy with human directional feedback, such that the robot eventually aligns the safe MPC policy with human intent for accomplishing safety-critical tasks.
-This paper aims to answer the previous two questions by proposing a new paradigm for a robot to online and efficiently learn safety constraints in its model predictive control (MPC) policy with human feedback.
+In robot policy design, safety is always the top priority, especially in safety-critical applications, such as autonomous vehicles or robots operating near humans, where safety violation can lead to irreversible loss or damage.
+While extensive work in safe control [1] and reinforcement learning [2] focuses on designing safe policy given robot motion constraints, less attention has been paid to how to specify the safety constraints in the first place.
+In many applications, specifying proper safety constraints can be challenging.
+First, explicit safety boundary for robot behavior can be hard to model.
+For example, when delivering food in a crowded restaurant, a service robot must balance the tray to avoid spills; this requires constraints on robot speed, food weight distributions, and interaction with its surroundings, and all those constraint aspects can be complex to model.
+Second, safety perception can vary among users.
+In the same example of the delivery robot: some diner may feel secure with the robot staying within a foot or two when delivering items, while others may feel uncomfortable with close proximity and prefer the robot to stay at a slight distance.
+Figure 1: Overall illustration of the proposed safe MPC alignment.
+Our method online updates the safety constraint in robot safe MPC policy based on human directional feedback, such that the robot eventually aligns the safe MPC policy with human intent for accomplishing safety-critical tasks.
+The challenges in safety specifications lead to two questions, which this paper seeks to answer.
+Q1: can a robot learn and align its safety constraints with human intent in an interactive manner?
+Q2: can such learning process to be user-data efficient such that a robot can quickly adapt to different users?
+Recent work [3, 4, 5] has explored learning safety constraints from human demonstrations.
+However, acquiring safety-informative demonstrations could be challenging.
+An alternative is learning from human preference feedback [6, 7, 8, 9, 10, 11, 12], where users express preferences between trajectories to guide robot learning.
+Despite the success in reward alignment [6], the query trajectories can be difficult for humans to judge when they are equally suboptimal, especially in early learning stages.
+To address such issue, [13, 14, 15, 16, 17, 18, 19] proposed learning from human physical corrections.
+While promising, the recent work [20] suggests those methods are potentially sensitive to correction magnitude: improper human corrections can lead to overshooting and inefficiency of learning.
+Importantly, all those feedback learning methods primarily focus on learning the reward function and do not address the challenges of learning safety constraints from human feedback.
+This paper aims to answer the previous two questions by proposing a new paradigm for a robot to efficiently learn safety constraints in its model predictive control (MPC) policy with human feedback.
 We name the method safe MPC alignment.
 As shown in Fig.
-1, a human user, while observing the robot’s potentially unsafe behavior, provides directional feedback for the robot’s motion.
-Once receiving the human feedback, the robot immediately updates the safety constraints in its MPC policy and continues its motion with the updated policy.
-Human feedback is in the form of physical corrections [6, 7] and sparse in time.
-The safe MPC alignment only utilizes the direction of human correction, regardless of the magnitude, for the safety constraints update, Importantly, the procedure of the proposed alignment is able to guarantee the exponential convergence with finite rounds of human corrections, or output failure with identified hypothesis misspecification.
+1, a human user, observing the robot’s potentially unsafe behavior, provides feedback to the robot.
+Upon receiving human feedback, the robot immediately updates the safety constraints in its MPC policy and continues with the updated policy.
+Different from existing learning from feedback methods, our method only requires human directional corrections [20], which can be sparse in time.
+For example, when a human user corrects an autonomous car, he/she only needs to give the command of ‘turn left‘ or ‘turn right‘, without dictating how much the car needs to turn.
+The safe MPC alignment only utilizes the direction of human correction, regardless of the magnitude, for the safety constraints update.
+As shown in our previous work [20] or later, the type of directional feedback provides a consistently large and flexible correction space, effectively overcoming the over-shooting correction, compared to existing magnitude feedback.
+Importantly, the proposed alignment is able to guarantee the exponential convergence with finite rounds of human corrections, or output failure with certified hypothesis misspecification.
 Specifically, we highlight our main contributions below.
@@
 To our knowledge, it is the first work for interactive learning of safety constraints from human feedback.
-(II) We theoretically show the human-data efficiency of the proposed safe MPC alignment method: the proposed method is guaranteed to have exponential convergence with finite human interactions/corrections.
-Furthermore, we also prove that the proposed algorithm enables the identification of hypothesis space misspecification failure.
-(III) We conduct extensive user studies to validate the effectiveness of the proposed method, including user studies on two MuJoCo-based simulation games and one real-world experiment of physically teaching a Franka robot arm for mobile liquid pouring.
-The results demonstrate the effectiveness and efficiency of the proposed method and show that a robot can successfully learn a safe MPC policy with only tens of human corrections.
+(II) We theoretically show the human-data efficiency of the safe MPC alignment: the proposed method is guaranteed to have exponential convergence with finite human corrections.
+Furthermore, we prove that the proposed method enables the certification of hypothesis space misspecification failure.
+(III) We conduct extensive user studies to validate the proposed method, including two MuJoCo-based simulation games and a real-world experiment of teaching a Franka robot arm for mobile liquid pouring.
+The results show the effectiveness and efficiency of our method.
 The remaining paper is organized as follows.
 Section II reviews the related work, and Section III formulates the safe MPC alignment problem.
-The main algorithm is presented in Section IV.
-Section V focuses on the implementation details of the algorithm and presents the main results on convergence and misspecification.
-Sections VI and VII provide the results of numerical experiments and user studies.
-Conclusions are drawn in Section VIII.
+The main method is presented in Section IV.
+Section V gives the implementation details and presents the main results.
+Sections VI and VII provide the experiments and user studies.
+Conclusions are in Section VIII.
 II Related Work
 II-A Learning from Demonstrations
-Under the general theme of learning from human data, the proposed work is loosely connected to the topic of learning from demonstrations, where objective functions are learned from expert (human) data using techniques of inverse reinforcement learning (IRL) and inverse optimal control (IOC) [8, 9].
-Since the demonstration data typically comes as time sequences of data, IRL/IOC methods typically run in an offline or episodic manner: a robot first obtains human demonstrated trajectories, and then performs reward/cost learning [9, 10, 11, 12, 13, 14].
+Under the general theme of learning from human data, the proposed method is loosely connected to the topic of learning from demonstrations, where cost/reward functions are learned from expert (human) data using inverse reinforcement learning (IRL) and inverse optimal control (IOC) methods [21, 22].
+IRL/IOC methods typically run offline: human demonstration data is first obtained, and then reward/cost learning is performed [22, 23, 24, 25, 26, 27].
 As pointed in [28], acquiring optimal demonstration data can be difficult in practice.
-The challenge can be either attributed to higher degrees of freedom of the robot, or the lack of prior knowledge about the desired robot behavior [16].
-In this work, we focus on online and interactive learning of safety constraints.
-Different from learning from demonstrations that needs the exhibition of desired robot behavior, the proposed learning paradigm only requires a human user to incrementally improve the current robot motion by providing directional feedback, thus more accessible to human users.
-II-B Learning from Human Interaction
-Orthogonal to learning from demonstrations is robot learning from human interaction.
-The line of this work typically focus on learning a reward/cost function from human feedback, which can be in the form of preference and corrections/feedback.
-The learned reward/cost function is then used to derive human-preferred robot policy.
-In learning from human preferences, human preference typically given as user’s ranking of two or multiple robot candidate trajectories [17, 18, 19, 20, 21, 22, 23], and Bradley-Terry model [24] is used to update the reward/cost function.
-The recently variant in this line of work seeks to directly learn a policy from human preference [25].
-Human interaction can also come as corrections, where human users may intervene and correct the robot at runtime.
+The challenge can be either attributed to the higher degrees of freedom of the robot, or the lack of prior knowledge about the desired robot behaviour to perform demonstrations with high-quality [29].
+In this work, we focus on interactive learning of safety constraints.
+Different from learning from demonstrations that needs the exhibition of desired robot motion, the proposed method only requires a human user to incrementally improve the current robot motion by providing directional feedback, thus more accessible to human users.
+II-B Learning from Human Feedback
+The line of this work typically focuses on learning rewards/costs from human feedback, which can be given in preference and corrections.
+In learning from human preferences, human preference is typically given as user’s ranking of two or multiple robot candidate trajectories [6, 7, 8, 9, 10, 11, 12], and Bradley-Terry model [30] is used to update the rewards/costs.
+The recent variant seeks to directly learn a policy from human preference [31].
+Human feedback can also come as corrections, where human users may intervene and correct the robot at runtime.
 An intended robot reward/cost function can be inferred from human corrections.
-Specific human corrections include language [26, 27, 28, 29, 30] and physical corrections in trajectory/action space [31, 32, 6].
-Since human corrections can be sparse in time, many methods [31, 33, 32, 34, 35, 6, 36] use trajectory deformation to convert sparse corrections to a continuous human-intended trajectory using a temporal propagation kernel.
-The reward/cost function is updated to explain the deformed trajectories.
+Specific human corrections include language [32, 33, 34, 35, 36] and physical corrections [13, 15, 18].
+Since human corrections can be sparse in time, many methods [13, 14, 15, 16, 17, 18, 19] use trajectory deformation methods to interpolate sparse corrections as a continuous trajectory.
+The rewards/costs are updated to explain the deformed trajectories.
 Some strategies are further proposed to deal with the uncertainty and improve the robustness of the learning process, like updating one parameter dimension at a time [15] or formulating the reward/cost learning in the framework of Kalman Filter [14].
-Our work also belongs to the category of learning from human physical corrections.
-But instead of learning reward/cost functions, we focus on learning safety constraints from human corrections, which has not been addressed before.
-Methodologically, we propose a certifiable learning method, which guarantees the exponential learning convergence with few human corrections or identifies the hypothesis misspecification when the learning hypothesis space does not contain the true human intent.
-Those capabilities are missing in existing methods.
-For example, the state-of-the-art methods [37] for learning from human preference requires hundreds or thousands of human feedback, while the proposed method can successfully learn a complex safe policy only with tens of human corrections.
+Our work belongs to the category of learning from human corrections.
+But instead of learning rewards/costs, we focus on learning safety constraints from human corrections, which has not been addressed before.
+Methodologically, we propose a certifiable method, which guarantees the exponential learning convergence with finite human corrections, and certifies the hypothesis misspecification if the learning hypothesis space does not contain the true human intent.
+Those theoretical results are missing in existing methods.
 II-C Learning Constraints from Demonstrations
-A increasing number of work focuses on learning safety constraints from demonstrations.
+An increasing number of work focuses on learning safety constraints from demonstrations.
 One vanilla idea is to formulate constraint learning as a binary classification.
-However, different from learning reward/cost, since a robot usually runs within a safe region, negative demonstration samples that violate safety constraints are challenge to collect.
-To collect negative samples, different strategies have been proposed.
-[3, 38] consider the trajectories of lower cost as the unsafe samples.
+However, since a robot usually runs within a safe region, negative demonstration samples that violate safety are challenging to collect.
+To address such, [3, 37] consider the trajectories of lower cost as the unsafe samples.
 [38] proposes adjusting the constraints toward decreasing the safety margin of feasible policies.
@@
 Different probabilistic formulations are also proposed to learn constraints from demonstration data, such as maximum likelihood [5, 43, 44], maximum causal entropy [45, 46], and Bayesian inference [47, 48].
-In the constraint learning, safety constraints can be represented as a set of safe states [41], the belief of potential constraints [42], Gaussian process [43], or directly as a convex hull in policy space [4].
-The vast majority of existing approaches focus on learning safety constraints from human demonstrations.
-They are mostly developed in the setting of reinforcement learning, needing to repetitively solve policy optimization in each learning update.
-Those methods cannot be used in an online, interactive setting, where a robot is expected to immediately update its safe policy and respond to feedback given by a human user.
-To achieve so, we will develop the first method, to our best knowledge, to enable learning the safe constraint from human feedback.
+In the above constraint learning, safety constraints can be represented as a set of safe states [40], the belief of potential constraints [41], Gaussian process [42], or directly as a convex hull in policy space [4].
+The vast majority of the above methods formulate the constraint learning problem in the setting of reinforcement learning, and thus the learning process usually involves repetitive policy optimization.
+Those methods cannot be deployed in an online setting, where a robot is expected to on-the-fly update its safe policy and respond to human feedback.
+To fill this gap, we will develop the first method, to our best knowledge, to enable online learning of safe constraint from human feedback.
 III Preliminary and Problem Formulation
@@
 $\min_{\boldsymbol{u}_{0:T-1}}&\quad J(\boldsymbol{u}_{0:T-1};\boldsymbol{x}_{0})=\sum_{t=0}^{T-1}c(\boldsymbol{x}_{t},\boldsymbol{u}_{t})+h(\boldsymbol{x}_{T})\\ \text{subject to}&\quad\boldsymbol{x}_{t+1}=\boldsymbol{f}(\boldsymbol{x}_{t},\boldsymbol{u}_{t})\\ &\quad{g}_{\boldsymbol{\theta}}(\boldsymbol{x}_{0:T},\boldsymbol{u}_{0:T-1})\leq 0\end{aligned}.$ (1)
-Here, $\boldsymbol{x}_{t}\in\mathbb{R}^{n}$ and $\boldsymbol{u}_{t}\in\mathbb{R}^{m}$ are robot state and action, respectively, at time step $t=0,1,2,...,T$ with $T$ being the prediction horizon; $\boldsymbol{x}_{0:T}=(\boldsymbol{x}_{0},\boldsymbol{x}_{1},...,\boldsymbol{x}_{T})$ and $\boldsymbol{u}_{0:T-1}=(\boldsymbol{u}_{0},\boldsymbol{u}_{1},...,\boldsymbol{u}_{T-1})$; $J(\cdot)$ is the given cost function including the running cost term ${c}(\cdot)$ and final cost term $h(\cdot)$, representing the given task; $\boldsymbol{f}(\cdot,\cdot)$ is the system dynamics; ${g}_{\boldsymbol{\theta}}(\cdot)$ is the safety constraint, which is parameterized by $\boldsymbol{\theta}$.
+Here, $\boldsymbol{x}_{t}\in\mathbb{R}^{n}$ and $\boldsymbol{u}_{t}\in\mathbb{R}^{m}$ are robot state and action, respectively; $t=0,1,2,...,T$ is the time step with prediction horizon $T$; $\boldsymbol{x}_{0:T}=(\boldsymbol{x}_{0},\boldsymbol{x}_{1},...,\boldsymbol{x}_{T})$ and $\boldsymbol{u}_{0:T-1}=(\boldsymbol{u}_{0},\boldsymbol{u}_{1},...,\boldsymbol{u}_{T-1})$; $J(\cdot)$ is the given cost function including the running cost ${c}(\cdot)$ and final cost $h(\cdot)$, representing the given task; $\boldsymbol{f}(\cdot,\cdot)$ is the system dynamics; ${g}_{\boldsymbol{\theta}}(\cdot)$ is the safety constraint, which is parameterized by $\boldsymbol{\theta}\in\mathbb{R}^{r}$.
 Given starting state $\boldsymbol{x}_{0}$ and constraint parameter $\boldsymbol{\theta}$, the trajectory solution to (1), denoted as $\boldsymbol{\xi}_{\boldsymbol{\theta}}=\{\boldsymbol{u}^{\boldsymbol{\theta}}_{0:T-1},\boldsymbol{x}^{\boldsymbol{\theta}}_{0:T}\}$, is an implicit function of $\boldsymbol{\theta}$.
 In a receding horizon implementation, $\boldsymbol{x}_{0}$ is set as the robot’s current state (estimation) $\boldsymbol{{x}}_{k}$, where $k=0,1,2,...,$ is the policy rollout time step; and only the first action $\boldsymbol{u}^{\boldsymbol{\theta}}_{0}$ in $\boldsymbol{{u}}_{0:T}^{\boldsymbol{\theta}}$ is applied to the robot to move the robot to the next state $\boldsymbol{x}_{k+1}$.
-This leads to a closed-loop Safe MPC policy for the robotic system, denoted as $\boldsymbol{\pi}_{\boldsymbol{\theta}}$, mapping from robot current state $\boldsymbol{x}_{k}$ to its action $\boldsymbol{u}^{\boldsymbol{\theta}}_{0}$.
+This leads to a closed-loop Safe MPC policy for the robot, denoted as $\boldsymbol{\pi}_{\boldsymbol{\theta}}$, mapping from robot current state $\boldsymbol{x}_{k}$ to its action $\boldsymbol{u}^{\boldsymbol{\theta}}_{0}$.
 III-B Safe Asymptotic Approximation of Safe MPC
-We present a penalty-based approximation to the original Safe MPC problem (1).
-Define variable $\boldsymbol{\xi}=\{\boldsymbol{u}_{0:T-1},\boldsymbol{x}_{0:T}\}$ and a penalty objective for (1) that incorporates the cost $J(\boldsymbol{\xi})$ and safety constraint $g_{\boldsymbol{\theta}}(\boldsymbol{\xi})$:
+We present a penalty-based approximation to the Safe MPC in (1). Define trajectory variable $\boldsymbol{\xi}=\{\boldsymbol{u}_{0:T-1},\boldsymbol{x}_{0:T}\}$ and the following penalty objective:
 $B(\boldsymbol{\xi},\boldsymbol{\theta})=J(\boldsymbol{\xi})-\gamma\ln(-g_{\boldsymbol{\theta}}(\boldsymbol{\xi})).$ (2)
-We have the following result from our prior work [50], stating an asymptotic and safe approximation to the solution $\boldsymbol{\xi}_{\boldsymbol{\theta}}$ to (1) by the following optimization
+Here, the above defined penalty $B(\boldsymbol{\xi},\boldsymbol{\theta})$ includes the original control cost in (1) and safety constraint $g_{\boldsymbol{\theta}}(\boldsymbol{\xi})$ via logarithmic barrier transformation, $\gamma>0$ is named the barrier parameter, controlling the weight of safety penalty $-\ln(-g_{\boldsymbol{\theta}}(\boldsymbol{\xi}))$ relative the control cost.
+With $B(\boldsymbol{\xi},\boldsymbol{\theta})$, one can establish the following penalty-based MPC formulation
 $\min_{\boldsymbol{u}_{0:T-1}}\quad B(\boldsymbol{\xi},\boldsymbol{\theta})\qquad\text{subject to}\quad\boldsymbol{x}_{t+1}=\boldsymbol{f}(\boldsymbol{x}_{t},\boldsymbol{u}_{t}).$ (3)
+We denote the solution to the above penalty-based MPC (3) as $\boldsymbol{\xi}_{\boldsymbol{\theta}}^{\gamma}=\{\boldsymbol{u}_{0:T-1}^{\boldsymbol{\theta},\gamma},\boldsymbol{x}_{0:T-1}^{\boldsymbol{\theta},\gamma}\}$.
+The following result is from our prior work [49], stating that the penalty-based MPC (3) is an asymptotic and safe approximation to the Safe MPC (1).
 Lemma 1.
-(Theorem 2 in [50]) If the original Safe MPC problem (1) satisfies certain mild conditions (i.e., local second-order condition and strict complementarity), then for any small $\gamma>0$, there exists a (local) solution $\boldsymbol{\xi}_{\boldsymbol{\theta}}^{\gamma}=\{\boldsymbol{u}_{0:T-1}^{\boldsymbol{\theta},\gamma},\boldsymbol{x}_{0:T-1}^{\boldsymbol{\theta},\gamma}\}$ to (3), and $\boldsymbol{\xi}_{\boldsymbol{\theta}}^{\gamma}$ is safe with respect to the original constraint, i.e., $g_{\boldsymbol{\theta}}(\boldsymbol{\xi}_{\boldsymbol{\theta}}^{\gamma})<0$.
-Furthermore, $\boldsymbol{\xi}_{\boldsymbol{\theta}}^{\gamma}$ is asymptotically approximate the original solution to (1): $\boldsymbol{\xi}_{\boldsymbol{\theta}}^{\gamma}\rightarrow\boldsymbol{\xi}_{\boldsymbol{\theta}}$ as $\gamma\rightarrow 0$.
-The above lemma asserts that by solving (3) one can safely use the approximate solution $\boldsymbol{\xi}_{\boldsymbol{\theta}}^{\gamma}$ to replace the original solution with arbitrary controlled accuracy by $\gamma$.
-As shown in [50], $\gamma$ usually takes a small value for good approximation accuracy.
+(Theorem 2 in [49]) If the original Safe MPC (1) satisfies certain mild conditions (i.e., local second-order condition and strict complementarity), then for any small $\gamma>0$, the penalty-based MPC (3) has a (local) solution $\boldsymbol{\xi}_{\boldsymbol{\theta}}^{\gamma}$, and $\boldsymbol{\xi}_{\boldsymbol{\theta}}^{\gamma}$ is safe with respect to the original constraint, i.e., $g_{\boldsymbol{\theta}}(\boldsymbol{\xi}_{\boldsymbol{\theta}}^{\gamma})<0$, and $\boldsymbol{\xi}_{\boldsymbol{\theta}}^{\gamma}$ satisifies the first-order condition of (3).
+Furthermore, $\boldsymbol{\xi}_{\boldsymbol{\theta}}^{\gamma}$ is asymptotically approximate $\xi_{\boldsymbol{\theta}}$: $\boldsymbol{\xi}_{\boldsymbol{\theta}}^{\gamma}\rightarrow\boldsymbol{\xi}_{\boldsymbol{\theta}}$ as $\gamma\rightarrow 0$.
+The above lemma asserts that solving the Safe MPC (1) can be replaced by solving penalty-based MPC (3).
+The solution approximation is controlled by $\gamma$ for arbitrary accuracy, as $\boldsymbol{\xi}_{\boldsymbol{\theta}}^{\gamma}\rightarrow\boldsymbol{\xi}_{\boldsymbol{\theta}}$ as $\gamma\rightarrow 0$.
+A good approximation is achieved by smaller $\gamma$.
+on the contrary, if $\gamma$ is larger, penalty-based MPC (3) will lead to more safety-conservative solution $\boldsymbol{\xi}_{\boldsymbol{\theta}}^{\gamma}$.
 In the development below, we fix the parameter $\gamma$ in (3).
-We denote the approximate Safe MPC in (3) as $\boldsymbol{\pi}^{\gamma}_{\theta}$.
+We denote the penalty-based MPC policy in (3) as $\boldsymbol{\pi}^{\gamma}_{\theta}$.
 III-C Problem Formulation
 III-C1 Representation of Human Intended Safety
-We consider the scenario of human-robot coexistence, where a human is supervising the robot for a given task.
-The user has his/her intent of robot safety.
-In Safe MPC (1), we hypothesize that the user intended safety can be represented by the set
+Consider a human-robot coexistence task, where a human is supervising the robot motion.
+The user has his/her intent of robot motion safety.
+In Safe MPC (1), we hypothesize user-intended safety can be represented by a subset of the trajectory space:
 $\begin{split}\mathcal{S}(\boldsymbol{\theta}_{H})=\{(\boldsymbol{x}_{0:T},\boldsymbol{u}_{0:T-1})\,\,|\,\,\ {g}_{\boldsymbol{\theta}_{H}}(\boldsymbol{x}_{0:T},\boldsymbol{u}_{0:T-1})\leq 0\}.\end{split}$ (4)
-Here, $\boldsymbol{\theta}_{H}$ is user implicit intent parameter picked from an implicit set $\Theta_{H}\subset\mathbb{R}^{r}$, i.e., $\boldsymbol{\theta}_{H}\in\Theta_{H}$.
-$\Theta_{H}$ is implicit to the human user and unknown to the robot.
-For the robot policy $\boldsymbol{\pi}_{\boldsymbol{\theta}}^{\gamma}$, if $\boldsymbol{\theta}\in\Theta_{H}$, the user considers the robot behavior to be safe; otherwise if $\boldsymbol{\theta}\notin\Theta_{H}$, the robot motion by $\boldsymbol{\pi}_{\boldsymbol{\theta}}^{\gamma}$ might not be safe at certain time in the view of the user, and thus the user provides corrections to improve the safety of the robot.
+Here, $\boldsymbol{\theta}_{H}$ is intent parameter from an implicit set $\boldsymbol{\theta}_{H}\in\Theta_{H}\subset\mathbb{R}^{r}$.
+$\Theta_{H}$ is implicit to human and unknown to the robot.
+For the robot policy $\boldsymbol{\pi}_{\boldsymbol{\theta}}^{\gamma}$, if $\boldsymbol{\theta}\in\Theta_{H}$, the user considers the robot motion to be safe; otherwise if $\boldsymbol{\theta}\notin\Theta_{H}$, the robot motion by $\boldsymbol{\pi}_{\boldsymbol{\theta}}^{\gamma}$ might not be safe at certain time in the view of the user, and thus the user provides corrections to improve its safety.
 We use an implicit set $\Theta_{H}\in\mathbb{R}^{r}$, instead of single point $\boldsymbol{\theta}_{H}$, to capture the potential variation of user intents.
 We suppose that a volume measure of $\Theta_{H}$, $\operatorname{{Vol}}(\Theta_{H})$, is non-zero, and a user intent $\boldsymbol{\theta}_{H}\in\Theta_{H}$ can vary at different steps and context, but always within $\Theta_{H}$.
-III-C2 Human Online Directional Correction
-We consider a human user who gives correction to the robot in an online fashion while the robot is running with the most updated $\boldsymbol{\pi}_{\boldsymbol{\theta}}^{\gamma}$.
-Specifically, denote $i$ to be the index of learning update, and suppose the robot policy currently is $\boldsymbol{\pi}_{\boldsymbol{\theta}_{i}}^{\gamma}$.
-At the time steps $k_{i}$ of rolling out $\boldsymbol{\pi}_{\boldsymbol{\theta}_{i}}^{\gamma}$, a human user gives directional correction $\boldsymbol{a}_{k_{i}}$ in robot action space at its state $\boldsymbol{x}_{k_{i}}$.
-After receiving $\boldsymbol{a}_{k_{i}}$, the robot updates the safety parameter $\boldsymbol{\theta}_{i}\rightarrow\boldsymbol{\theta}_{i+1}$ in (1) and obtain the updated policy $\boldsymbol{\pi}_{\boldsymbol{\theta}_{i+1}}^{\gamma}$.
-Then, the robot continues with the updated policy $\boldsymbol{\pi}_{\boldsymbol{\theta}_{i}}^{\gamma}$ until the next user possible correction.
-This process of robot MPC policy update and user correction repeats until the convergence of the policy update or the user is satisfied with the robot motion.
+III-C2 Human Directional Correction
+We consider human gives correction to a robot in an online fashion while the robot is executing its most updated $\boldsymbol{\pi}_{\boldsymbol{\theta}}^{\gamma}$.
+Specifically, let $i$ be the index of learning update, and suppose robot current policy is $\boldsymbol{\pi}_{\boldsymbol{\theta}_{i}}^{\gamma}$.
+During rolling out $\boldsymbol{\pi}_{\boldsymbol{\theta}_{i}}^{\gamma}$, the robot receives a sparse human directional correction $\boldsymbol{a}_{k_{i}}\in\mathbb{R}^{m}$ at the rollout time step $k_{i}$.
+Here, we do not assume precise timing $k_{i}$ or duration of the human correction due to the sparsity or intermittency of human feedback.
+In other words, human can apply one or multiple corrections at any time in robot rollout time steps for any duration as he/she finds necessary.
+But we here assume human directional corrections $\boldsymbol{a}_{k_{i}}$ are in robot action space.
+For example, in autonomous driving where the robot action space is steering angle and acceleration, human can directly apply feedback with steering wheel and pedals.
+In some other applications, the action space correction can be achieved via specific mapping or interface.
+For example, to correct a torque-controlled robot arm, human’s correction on robot end-effector can be mapped into joint torque via Jacobian transpose matrices.
+The reason why we only consider the action space correction is because first, corrections in input spaces may be easier to implement, as shown in our later experiments, and second, corrections in state spaces can be infeasible for the under-actuated robot system [20].
+After receiving each $\boldsymbol{a}_{k_{i}}$, the robot updates the safety parameter $\boldsymbol{\theta}_{i}\rightarrow\boldsymbol{\theta}_{i+1}$ in (1) for its MPC policy.
+Then, the robot continues with $\boldsymbol{\pi}_{\boldsymbol{\theta}_{i+1}}^{\gamma}$ until the next possible user corrections.
+This process repeats until the convergence of the policy update or the user is satisfied with the robot motion.
 An illustration of one update is shown in Fig.
@@
 Figure 2: Update of safe MPC policy with human directional corrections.
-In the above-stated human feedback process, we consider that a human user only provides directional correction $\boldsymbol{a}$, i.e., the magnitude $\left\lVert\boldsymbol{a}\right\rVert$ does not matter (theoretical justification will be given later in Section IV-A).
+In the above learning process, we consider that a human user only provides directional correction $\boldsymbol{a}$, i.e., the magnitude $\left\lVert\boldsymbol{a}\right\rVert$ does not matter (theoretical justification will be given later).
 For example, when correcting a mobile robot, the user’s correction can be "left" or "right", without specifying how much the robot should go.
-Also, we assume that the user feedback is applied in the action space.
-For some robotic systems, this assumption can be readily met, such as turning the driving wheels for autonomous vehicles.
-In other cases, the correction can be applied through a dedicated (hardware or algorithmic) interface to the action space; such as contact forces from end-effector to joint torque, and teleoperation device for remote control.
-In those cases, $\boldsymbol{a}$ is the correction signal after the interface.
-The reason why we only consider the correction in action space is that otherwise the corrections in robot state space can lead to dynamics unfeasibility or uncontrollability for under-actuated robotic systems [7, 51].
+As shown in our previous work [20] or later, the type of directional feedback provides a consistently large and flexible correction space, effectively overcoming the over-shooting correction, compared to existing magnitude feedback.
 III-C3 Robot Operation Safety and Emergency Stop
-In the aforementioned policy update procedure, the robot will respond to human corrections by executing the updated safe MPC policy $\pi_{\boldsymbol{\theta}}$.
-However, it is possible that the robot may still exit the user-intended safety zone during the correction process.
-In such cases, a human user may trigger an emergency stop to stop the robot motion and potentially restart it from a desired safety region.
-This emergency stop mechanism is common in real-world robotic applications.
-Until an emergency stop is triggered, human corrections typically occur while the robot is still within the intended safety zone; otherwise, the corrections would be useless.
-For instance, correcting an autonomous vehicle after it has been crashed is futile.
+In the aforementioned learning procedure, the robot will respond to human corrections by executing the updated safe MPC policy $\pi_{\boldsymbol{\theta}}^{\gamma}$.
+However, it is possible that the robot may still exit the user-intended safety region.
+In such cases, a human user may trigger an emergency stop mechanism to stop the robot motion and potentially restart it from a desired safety region.
+In this sense, all trajectories before the human triggers the emergency stop are considered safe in human intent.
+In other words, until an emergency stop is triggered, human corrections typically occur while the robot is still within the intended safety zone; otherwise, the corrections would be useless.
+For instance, correcting a vehicle after it is crashed is futile.
 III-C4 Parametric Safety Constraints
 In the safe MPC (4), the safety set is defined as a zero sublevel set of a safety value function ${g}_{\boldsymbol{\theta}}$.
-We note that if ${g}_{\boldsymbol{\theta}}$ has sufficient representation power, the safety set can represent complex safety region, which can be alternatively represented by the intersection of multiple constraints (e.g., a polytope safety set defined by multiple hyperplanes).
-In the following, we mainly focus on a weighted-features parameterization of ${g}_{\theta}$, which has been commonly used by the extensive prior work on learning cost/reward [4, 7, 52, 53, 54].
+We note that if ${g}_{\boldsymbol{\theta}}$ has sufficient representation power, the safety set can represent complex (non-convex) safety region, which can be alternatively represented by the intersection of multiple constraints.
+In the following, we mainly focus on a weighted-features parameterization of ${g}_{\theta}$, which has been commonly used by the prior work on learning cost/reward [4, 20, 50, 51, 52].
 Specifically, we parameterize $\boldsymbol{g}_{\theta}$
-${g}_{\boldsymbol{\theta}}(\boldsymbol{x}_{0:T},\boldsymbol{u}_{0:T-1})=\phi_{0}(\boldsymbol{x}_{0:T},\boldsymbol{u}_{0:T-1})+\theta_{1}\phi_{1}(\boldsymbol{x}_{0:T},\boldsymbol{u}_{0:T-1})\\ +\theta_{2}\phi_{2}(\boldsymbol{x}_{0:T},\boldsymbol{u}_{0:T-1})+\cdots+\theta_{r}\phi_{r}(\boldsymbol{x}_{0:T},\boldsymbol{u}_{0:T-1}),$ (5)
+${g}_{\boldsymbol{\theta}}(\boldsymbol{\xi})=\phi_{0}(\boldsymbol{\xi})+\theta_{1}\phi_{1}(\boldsymbol{\xi})+\theta_{2}\phi_{2}(\boldsymbol{\xi})+\cdots+\theta_{r}\phi_{r}(\boldsymbol{\xi}),$ (5)
 with $\boldsymbol{\theta}=[\theta_{1},\theta_{2},...\theta_{r}]^{\mkern-1.5mu\mathsf{T}}\in\mathbb{R}^{r}$ and $\boldsymbol{\phi}=[\phi_{1},\phi_{2},\dots,\phi_{r}]^{\mkern-1.5mu\mathsf{T}}$ the weight and feature vectors, respectively.
-We fix the weight of $\phi_{0}$ as unit, because otherwise, the $\boldsymbol{\theta}$ is scaling ambiguous, i.e., any scaled $c\cdot\boldsymbol{\theta}$ with $c>0$ lead to the same MPC solutions.
-We do not impose any constraints on the choice of feature vector as long as it is differentiable.
-For example, $\boldsymbol{\phi}$ can be neural features.
+We fix the weight of $\phi_{0}$ as unit, because otherwise, $\boldsymbol{\theta}$ has scaling ambiguity, i.e., any scaled $c\cdot\boldsymbol{\theta}$ with $c>0$ lead to the same MPC solutions.
+The dimension (number) $r$ of features can be chosen such that $\boldsymbol{\phi}$ has enough expressiveness to represent the human-intended safety constraint.
+We do not impose any constraints on the choice of feature vector as long as it is differentiable, for example, $\boldsymbol{\phi}$ can be neural features.
+In practice the feature vector can be chosen by trial and error, i.e., starting from a few features and gradually adding more until the learned constraint is satisfactory to the human user.
 Later, we will discuss our proposed method for generic parameterization $\boldsymbol{g}_{\boldsymbol{\theta}}$.
 III-C5 Problem of Safe MPC Alignment
-Following the procedure of online human correction and robot policy update described in Subsection III-C and given a weighted-feature safety constraint (5), the interest of this work is to develop a policy update rule: $\boldsymbol{\theta}_{i}\rightarrow\boldsymbol{\theta}_{i+1}$ with each human directional correction, such that $\boldsymbol{\theta}_{i}\rightarrow\boldsymbol{\theta}_{H}\in\Theta_{H}$, as $i=1,2,3,\dots,$.
-i.e., the safety constraint in the robot safe MPC policy will be aligned with human intent.
-IV Algorithm Overview for Safe MPC Alignment
-IV-A Hypothesis on User Directional Feedback
-In this session, we will introduce a hypothesis on user directional correction, which lays the foundation for our alignment framework proposed subsequently.
-Informally, we have the following intuitions on user feedback.
-First, user directional corrections, on average, improve the robot safety.
-The similar assumption is the basis for reinforcement learning with human feedback [55].
-Second, user correction on robot safety is typically proactive and typically happens before the robot enters into the true unsafe region.
-In other words, human correction is made when the robot operates within $\mathcal{S}(\boldsymbol{\theta}_{H})$ in (4).
-For example, it is useless to correct a ground vehicle after its collision with obstacles.
-Third, with the second intuition, user feedback is typically safety-margin aware: the closer a robot gets to the boundary of the unsafe region, the more significant user safety correction will be.
-This is observed in our prior user study [7].
+Following the procedure of human correction and robot policy update in Subsection III-C and given a weighted-feature safety constraint (5), we aim to develop a policy update rule: $\boldsymbol{\theta}_{i}\rightarrow\boldsymbol{\theta}_{i+1}$ for each received human directional correction, such that $\boldsymbol{\theta}_{i}\rightarrow\boldsymbol{\theta}_{H}\in\Theta_{H}$, as $i=1,2,3,\dots,$.
+i.e., safety constraints in the robot’s safe MPC policy will be aligned with human intent.
+IV Overview of Safe MPC Alignment Algorithm
+IV-A Hypothesis on Human Directional Corrections
+Informally, we have the following intuitions on human feedback.
+First, human corrections, on average, improve the robot safety.
+Second, human correction on robot safety is typically proactive and happens before the robot leaves the human-intended safety set.
+In other words, human correction is made when the robot is still in $\mathcal{S}(\boldsymbol{\theta}_{H})$ in (4).
+For example, it is useless to correct a vehicle after its crash.
+Third, human feedback is typically safety-margin aware: as a robot approaches the safety boundary, human users tend to prioritize safety over the task objective.
+For instance, when a vehicle moves closer to obstacles, humans may instinctively shift their focus to avoiding collisions rather than pursuing the original target.
 In the following, we will formalize those intuitions.
-Suppose that robot motion plan $\boldsymbol{\xi}_{\boldsymbol{\theta}}^{\gamma}$ is solved from (3) with parameter $\boldsymbol{\theta}$, and a human user applies directional correction $\boldsymbol{a}$ at first time step 00.
+Without loss of generality, suppose the motion plan $\boldsymbol{\xi}_{\boldsymbol{\theta}}^{\gamma}$ is a local solution to the penalty MPC (3), and when the robot execute the first action $\boldsymbol{u}_{0}^{\boldsymbol{\theta},\gamma}$ (receding horizon implementation), a human applies a directional correction $\boldsymbol{a}$ in robot action space.
 Define
 $\mathbf{a}=(\boldsymbol{a}^{\mkern-1.5mu\mathsf{T}},0,0,...)\in\mathbb{R}^{mT}.$ (6)
-Formally, we assume there is a nonzero-volume set $\bar{\Theta}_{H}$ (containing a ball) in user implicit intent set $\Theta_{H}$, i.e., $\bar{\Theta}_{H}\subseteq\Theta_{H}$, such that for any $\boldsymbol{\theta}_{H}\in\bar{\Theta}_{H}$, the user directional correction $\boldsymbol{a}$, on average, indicates a direction of lowering the penalty value $B(\boldsymbol{\xi}_{\boldsymbol{\theta}}^{\gamma},\boldsymbol{\theta}_{H})$ for the robot motion plan $\boldsymbol{\xi}_{\boldsymbol{\theta}}^{\gamma}$.
-Concretely, it means $\mathbf{a}$ on averages points in a direction of the negative gradient of $B(\boldsymbol{\xi},\boldsymbol{\theta}_{H})$ evaluated at $\boldsymbol{\xi}_{\boldsymbol{\theta}}^{\gamma}$, i.e.,
+We assume there is a nonzero volume set $\bar{\Theta}_{H}$ (containing a ball) in user implicit set $\Theta_{H}$, i.e., $\bar{\Theta}_{H}\subseteq\Theta_{H}$, such that for any $\boldsymbol{\theta}_{H}\in\bar{\Theta}_{H}$, the human correction $\boldsymbol{a}$, on average, indicates a direction of lowering the penalty value $B(\boldsymbol{\xi}_{\boldsymbol{\theta}}^{\gamma},\boldsymbol{\theta}_{H})$ for the robot motion plan $\boldsymbol{\xi}_{\boldsymbol{\theta}}^{\gamma}$.
+Concretely, it means $\mathbf{a}$ on averages points in the direction of the negative gradient of $B(\boldsymbol{\xi},\boldsymbol{\theta}_{H})$ at $\boldsymbol{\xi}_{\boldsymbol{\theta}}^{\gamma}$, i.e.,
 $\forall\boldsymbol{\theta}_{H}\in\bar{\Theta}_{H},\ \operatorname{\mathbb{E}}_{\mathbf{a}}\left\langle\mathbf{a},-\nabla B(\boldsymbol{\xi}_{\boldsymbol{\theta}}^{\gamma},\boldsymbol{\theta}_{H})\right\rangle=\\ \left\langle\mathbf{\bar{a}},-\nabla B(\boldsymbol{\xi}_{\boldsymbol{\theta}}^{\gamma},\boldsymbol{\theta}_{H})\right\rangle\geq 0.$ (7)
-Here, ${\operatorname{\mathbb{E}}}_{\mathbf{a}}(\mathbf{a})=\mathbf{\bar{a}}$ and the equality is due to the second term $\nabla B(\boldsymbol{\xi},\boldsymbol{\theta})$ is not dependent on $\mathbf{a}$.
-$\nabla B(\boldsymbol{\xi},\boldsymbol{\theta})$ denotes the full gradient of $B$ with respect to $\boldsymbol{u}_{0:T-1}$ evaluated at trajectory $\boldsymbol{\xi}=\{\boldsymbol{u}_{0:T-1},\boldsymbol{x}_{0:T}\}$.
-Similar notations apply to $J$ and $g$ below.
-Since state sequence $\boldsymbol{x}_{0:T}$ is the rollout of $\boldsymbol{u}_{0:T-1}$ on dynamics, the differentiation goes through the dynamics equation.
-One of the important observations to (7) is that the magnitude of the user correction $\mathbf{a}$ does not affect the inequality.
-This brings one advantage of directional correction compared to the magnitude-based correction: regardless of robot safeMPC plan $\boldsymbol{\xi}_{\boldsymbol{\theta}}^{\gamma}$, the user directional correction always accounts for half of the action space, as shown in Figure 3.
-Extending (7) leads to
+[…]
+(7) indicates that the magnitude of $\mathbf{a}$ does not affect the inequality.
+This brings one advantage of directional correction compared to the magnitude-based correction: regardless of robot plan $\boldsymbol{\xi}_{\boldsymbol{\theta}}^{\gamma}$, the allowable human corrections satisfying (7) always account for half of the action space, making directional corrections more accessible for human users, as shown in Fig.
+3.
+Recall that $B(\boldsymbol{\xi},\boldsymbol{\theta})=J(\boldsymbol{\xi})-\gamma\ln(-g_{\boldsymbol{\theta}}(\boldsymbol{\xi}))$, by extension of the gradient, (7) becomes
 $\forall\boldsymbol{\theta}_{H}\in\bar{\Theta}_{H},\left<\mathbf{\bar{a}},-\nabla J(\boldsymbol{\xi}_{\boldsymbol{\theta}}^{\gamma})\right>+\\ \gamma\frac{1}{-g_{\boldsymbol{\theta}_{H}}(\boldsymbol{\xi}_{\boldsymbol{\theta}}^{\gamma})}\left<\mathbf{\bar{a}},-\nabla g_{\boldsymbol{\theta}_{H}}(\boldsymbol{\xi}_{\boldsymbol{\theta}}^{\gamma})\right>\geq 0.$ (8)
@@
 For the above equation, we have the following interpretations.
-(I) The first and second terms on the left side of (8) correspond to the user’s correction toward lowering the cost and improving the safety value of robot plan $\boldsymbol{\xi}_{\boldsymbol{\theta}}^{\gamma}$.
+(I) The first and second terms on the left side of (8) correspond to human correction toward lowering the cost and improving the safety of $\boldsymbol{\xi}_{\boldsymbol{\theta}}^{\gamma}$, respectively.
 $\gamma$ serves as a weight between two improvements.
-(II) If we interpret the value of $-{g_{\boldsymbol{\theta}_{H}}(\boldsymbol{\xi}_{\boldsymbol{\theta}}^{\gamma})}$ as the safety margin of robot plan $\boldsymbol{\xi}_{\boldsymbol{\theta}}^{\gamma}$, (8) indicates that user’s correction is safety-margin aware: the closer the robot motion get to the unsafe region, the more significant a human user correction will be to drive the system back to safety interior.
+(II) If we interpret the value of $-{g_{\boldsymbol{\theta}_{H}}(\boldsymbol{\xi}_{\boldsymbol{\theta}}^{\gamma})}$ as the safety margin of $\boldsymbol{\xi}_{\boldsymbol{\theta}}^{\gamma}$, (8) suggests that the human correction is safety-margin aware: the closer the robot motion get to the unsafe region, the more will human user emphasize on improving the robot’s safety than minimizing the objective.
 This property is consistent with our previous intuition.
-(III) The above hypothesis (7) requires that $B(\boldsymbol{\xi}_{\boldsymbol{\theta}}^{\gamma},\boldsymbol{\theta}_{H})$ is well-defined in the first place, i.e.,
+(III) (7) implies $B(\boldsymbol{\xi}_{\boldsymbol{\theta}}^{\gamma},\boldsymbol{\theta}_{H})$ is well-defined in the first place:
 $\forall\boldsymbol{\theta}_{H}\in\bar{\Theta}_{H},\ \quad g_{\boldsymbol{\theta}_{H}}(\boldsymbol{\xi}_{\boldsymbol{\theta}}^{\gamma})<0.$ (9)
-The requirement corresponds to the third intuition — a user typically makes correction only when the robot is within user intended safety region, i.e., $\boldsymbol{\xi}_{\boldsymbol{\theta}}^{\gamma}\in\mathcal{S}(\boldsymbol{\theta}_{H})$.
+This corresponds to the third intuition: a user typically gives correction only when the robot is in user intended safety region, i.e., $\boldsymbol{\xi}_{\boldsymbol{\theta}}^{\gamma}\in\mathcal{S}(\boldsymbol{\theta}_{H})$.
 In sum, we have used (7) and (9) to formalize the intuitions we have observed at the beginning of this session.
-Those two constraints on $\boldsymbol{\theta}_{H}$ will be used in the robot learning pipeline developed in the next session.
+Those two constraints on $\boldsymbol{\theta}_{H}$ will be used in the robot learning next.
 IV-B Alignment Algorithm: Learning by Hypothesis Space Cutting
-To find an implicit $\boldsymbol{\theta}_{H}\in\bar{\Theta}_{H}\subseteq\Theta$ for the safe MPC (1), our idea is that at any learning iteration $i=0,1,2,....$, the robot monitors and updates a hypothesis space $\Theta_{i}\in\mathbb{R}^{r}$, a set containing all possible $\bar{\Theta}_{H}$.
-At initial, the robot may choose a hypothesis space $\Theta_{0}$ that is large enough to contain $\bar{\Theta}_{H}$.
-Specifically, one can choose
-$\Theta_{0}=\{\boldsymbol{\theta}\ |\ \boldsymbol{c}_{l}\leq\boldsymbol{\theta}\leq\boldsymbol{c}_{h}\}\subset\mathbb{R}^{r},$ (10)
-in which $\boldsymbol{c}_{l},\boldsymbol{c}_{h}$ are the given bounding vectors. By specifying a $\Theta_{0}$ large enough, we assume
+To find $\boldsymbol{\theta}_{H}\in\bar{\Theta}_{H}\subseteq\Theta$ for the safe MPC (1), our idea is that for each human directional correction received $i=0,1,2,....$, the robot performs one learning iteration, by updating a hypothesis space $\Theta_{i}\in\mathbb{R}^{r}$, a set containing all possible $\bar{\Theta}_{H}$.
+At initial, the robot may choose a large enough hypothesis space $\Theta_{0}$.
+Specifically, one can choose $\Theta_{0}$ as a box space
+$\Theta_{0}=\{\boldsymbol{\theta}\ |\ \boldsymbol{c}_{l}\leq\boldsymbol{\theta}\leq\boldsymbol{c}_{u}\}\subset\mathbb{R}^{r},$ (10)
+in which $\boldsymbol{c}_{l},\boldsymbol{c}_{u}$ are the vectors of lower and upper bounds of the box space, resepctively, and the inequalities applied element-wise.
+By saying a $\Theta_{0}$ large enough, we assume
 $\bar{\Theta}_{H}\subset\Theta_{0}.$ (11)
-Later, we will that our alignment algorithm can self-access the failure of learning in case of misspecification, where (11) is not met.
+Later, we will show that our method can self-access the learning failure in case of misspecification, where (11) is not met.
 Recall the procedure of human correction in Section III-C and Fig.
 2.
-At $i$-th learning update, $i=1,2,3...$, the robot safe MPC policy $\boldsymbol{\pi}_{\boldsymbol{\theta}_{i}}^{\gamma}$ with $\boldsymbol{\theta}_{i}\in\Theta_{i-1}$.
+At $i$-th learning iteration, $i=1,2,3...$, the robot executes $\boldsymbol{\pi}_{\boldsymbol{\theta}_{i}}^{\gamma}$ with $\boldsymbol{\theta}_{i}\in\Theta_{i-1}$.
 The user gives a directional correction $\boldsymbol{a}_{k_{i}}$ at rollout step ${k_{i}}$.
-By letting $\mathbf{\bar{a}}_{i}=(\boldsymbol{a}_{k_{i}}^{T},0,0,...)$ as in (6), we can establish the following two constraints on $\bar{\Theta}_{H}$ from the previous hypothesis (7) and (9)
+By letting $\mathbf{\bar{a}}_{i}=(\boldsymbol{a}_{k_{i}}^{T},0,0,...)$ as in (6), we can establish the following two constraints on $\bar{\Theta}_{H}$ from the previous hypotheses (7) and (9):
 $\bar{\Theta}_{H}\subset\mathcal{C}_{i}=\left\{\boldsymbol{\theta}|\left<\mathbf{\bar{a}}_{i},\nabla B(\boldsymbol{\xi}_{\boldsymbol{\theta}_{i}}^{\gamma},\boldsymbol{\theta})\right>\leq 0,\,\,\,g_{\boldsymbol{\theta}}(\boldsymbol{\xi}_{\boldsymbol{\theta}_{i}}^{\gamma})<0\right\},$ (12)
-which we refer to as the cutting set. Using $\mathcal{C}_{i}$, one can update the hypothesis space
+which we refer to as the cutting set.
+Note that the first inequality $\mathbf{\bar{a}}_{i},\nabla B(\boldsymbol{\xi}_{\boldsymbol{\theta}_{i}}^{\gamma},\boldsymbol{\theta})\right>\leq 0$ is deduced by multiplying -1 on both sides of (7).
+Using $\mathcal{C}_{i}$, one can update the hypothesis space
 $\Theta_{i}=\Theta_{i-1}\cap\mathcal{C}_{i}.$ (13)
-The above update can be viewed as a step of using the cutting set (12) to cut the current hypothesis space $\Theta_{i-1}$ to remove the portion where $\bar{\Theta}_{H}$ is impossible.
-The remaining hypothesis space $\Theta_{i}$ will be new hypothesis used in the next iteration.
+The above update can be viewed as a step of using (12) to cut (intersect with) the current hypothesis space $\Theta_{i-1}$ to remove the portion where $\bar{\Theta}_{H}$ is impossible.
+The remaining hypothesis space $\Theta_{i}$ will be new hypothesis used for the next iteration.
 With the above update, we have the following result.
@@
 See the proof in Appendix A-A. ∎
-As shown by Fig.
+As shown in Fig.
 4, the procedure of robot learning is to cut and move hypothesis space based on human directional correction.
 Therefore, we name our alignment method Hypothesis Space Cutting.
-As a summary, we present the three main steps of the hypothesis space cutting updates below.
+As a summary, we present the three main steps of our algorithm below.
 Figure 4: The update of constraint parameter (2D examples).
@@
 Step 2: Human user applies a directional correction $\boldsymbol{a}_{k_{i}}$ while robot executes $\boldsymbol{\pi}_{\boldsymbol{\theta}_{i}}^{\gamma}$ Step 3: Perform hypothesis space cut using (13) to get new hypothesis space $\Theta_{i}$.
-At $i$th iteration of the Hypothesis Space Cutting, the robot chooses $\boldsymbol{\theta}_{i}\in\Theta_{i-1}$ from the most recent hypothesis space $\Theta_{i-1}$ for its current safe MPC $\boldsymbol{\pi}_{\boldsymbol{\theta}_{i}}^{\gamma}$.
+At $i$th learning iteration, the robot chooses $\boldsymbol{\theta}_{i}\in\Theta_{i-1}$ from the most recent hypothesis space $\Theta_{i-1}$ for its current MPC $\boldsymbol{\pi}_{\boldsymbol{\theta}_{i}}^{\gamma}$.
 While executing $\boldsymbol{\pi}_{\boldsymbol{\theta}_{i}}^{\gamma}$, the robot receives human directional correction $\boldsymbol{a}_{k_{i}}$ at the rollout time step $k_{i}$.
-This correction is used to induce the cutting set $\mathcal{C}_{i}$ (12), which will be used to update $\Theta_{i-1}$ to $\Theta_{i}$ in (13).
+This correction leads to a cutting set $\mathcal{C}_{i}$ (12), which will be used to update $\Theta_{i-1}$ to $\Theta_{i}$ in (13).
 The above iteration repeats $i=1,2,3,...$.
-From (13) and by following the above hypothesis space cutting method, one can easily see that $\Theta_{0}\supseteq\Theta_{1}\supseteq\Theta_{2}\supseteq...\supseteq\bar{\Theta}_{H}$.
-Given a volume measure, it follows $\operatorname{{Vol}}\Theta_{0}\geq\operatorname{{Vol}}\Theta_{1}\geq\operatorname{{Vol}}\Theta_{2}\geq...,\geq\operatorname{{Vol}}{\bar{\Theta}_{H}}$.
-Thus, the volume of the hypothesis space is at least not increasing along with user corrections.
-It is hoped that the volume of search space is strictly decreasing, such that new $\boldsymbol{\theta}_{i+1}$ has a very good chance of falling inside $\bar{\Theta}_{H}$, which means a successful safety alignment between the robot and human.
-We will later (in Section V) theoretically show that the above Safe MPC Alignment method, within finite interactions $K$, will either
+From (13), one can easily see that $\Theta_{0}\supseteq\Theta_{1}\supseteq...\supseteq\bar{\Theta}_{H}$.
+Given a volume measure, it follows $\operatorname{{Vol}}\Theta_{0}\geq\operatorname{{Vol}}\Theta_{1}\geq...,\geq\operatorname{{Vol}}{\bar{\Theta}_{H}}$.
+Thus, the volume of the hypothesis space is at least not increasing after each user correction.
+It is hoped that the volume of the hypothesis space decreases quickly, such that new $\boldsymbol{\theta}_{i+1}$ increases its chance of falling inside $\bar{\Theta}_{H}$ — a successful safety alignment between the robot and human.
+For the above Safe MPC Alignment method, we will later (in Section V) show that within finite number $K$ of corrections, the method will either
   • (I)
@@
 IV-C Properties and Geometric Interpretations
-In this session, we present some properties and the geometric interpretations for the proposed safe MPC alignment method.
-The properties will be used to establish the convergence of the algorithm in the next section.
+In this session, we present some properties and the geometric interpretations for the proposed safe MPC alignment.
+Those properties will be used to establish the convergence of the algorithm in the next section.
 IV-C1 Piecewise Linearity of Cuts
-Recall the weight-feature parameterization of the safety constraint (5).
-From (12), we can examine $g_{\boldsymbol{\theta}}(\boldsymbol{\xi}_{\boldsymbol{\theta}_{i}}^{\gamma})<0$ is a linear inequality in $\boldsymbol{\theta}$.
-Based on the definition of the penalty objective function (2), $\mathbf{\bar{a}}_{i},\nabla B(\boldsymbol{\xi}_{\boldsymbol{\theta}_{i}}^{\gamma},\boldsymbol{\theta})\right>\leq 0$ is a linear fractional inequality
-$\mathbf{\bar{a}}_{i},\nabla J(\boldsymbol{\xi}_{\boldsymbol{\theta}_{i}}^{\gamma})\right>+\gamma\frac{1}{-g_{\boldsymbol{\theta}}(\boldsymbol{\xi}_{\boldsymbol{\theta}_{i}}^{\gamma})}\left<\mathbf{\bar{a}}_{i},\nabla g_{\boldsymbol{\theta}}(\boldsymbol{\xi}_{\boldsymbol{\theta}_{i}}^{\gamma})\right>\leq 0,$ (14)
-with a positive denominator $-g_{\boldsymbol{\theta}}(\boldsymbol{\xi}_{\boldsymbol{\theta}_{i}}^{\gamma})>0$ due to (12).
-Multiplying the denominator on both side yields another linear inequality.
-Formally, we give a lemma here about the piecewise linearity of the cut $\mathcal{C}_{i}$ in (12).
+By recalling the weight-feature parameterization of the safety constraint (5), one can examine that the inequities in (12) are both linear in $\theta$.
+This means that the cutting set $\mathcal{C}_{i}$ in (12) is a piecewise linear cut, as stated formally in the following lemma.
 Lemma 3.
@@
 Proof.
-Equation (15) can be directly obtained from (12), (2) and (5). This completes the proof. ∎
-With the lemma above, the intersection $\Theta_{i-1}\cap\mathcal{C}_{i}$ can be viewed as cutting $\Theta_{i-1}$ with the boundary of $\mathcal{C}_{i}$ which consists of two piecewise hyperplanes:
+See Appendix A-B. ∎
+According to the above lemma, $\Theta_{i-1}\cap\mathcal{C}_{i}$ can be viewed as cutting $\Theta_{i-1}$ with the boundary of $\mathcal{C}_{i}$ which is two hyperplanes:
 $\begin{split}\mathbf{bd}\ \mathcal{C}_{i}=\{\boldsymbol{\theta}\ |\ \boldsymbol{\theta}^{\mkern-1.5mu\mathsf{T}}\boldsymbol{h}_{i}=b_{i},{\boldsymbol{\theta}}^{\mkern-1.5mu\mathsf{T}}\boldsymbol{\phi}(\boldsymbol{\xi}_{\boldsymbol{\theta}_{i}}^{\gamma})\leq-\phi_{0}(\boldsymbol{\xi}_{\boldsymbol{\theta}_{i}}^{\gamma})\}\ \cup\\ \{\boldsymbol{\theta}\ |\ \boldsymbol{\theta}^{\mkern-1.5mu\mathsf{T}}\boldsymbol{h}_{i}\leq b_{i},{\boldsymbol{\theta}}^{\mkern-1.5mu\mathsf{T}}\boldsymbol{\phi}(\boldsymbol{\xi}_{\boldsymbol{\theta}_{i}}^{\gamma})=-\phi_{0}(\boldsymbol{\xi}_{\boldsymbol{\theta}_{i}}^{\gamma})\},\end{split}$ (17)
-which is shown in Fig. 4
-As the initial hypothesis space $\Theta_{0}\in\mathbb{R}^{r}$ in (10) is a polytope.
-Iteratively cutting with piecewise hyperplanes yields a series of convex polytopes in $\mathbb{R}^{r}$.
+shown in Fig. 4.
+As the initial hypothesis space $\Theta_{0}\in\mathbb{R}^{r}$ in (10) is a convex polytope, iteratively cutting with hyperplane pairs yields a series of convex polytopes in $\mathbb{R}^{r}$.
 Consequently, all $\Theta_{i}$s, $i=0,1,2,...$, are convex polytopes.
 IV-C2 Relationship between $\boldsymbol{\theta}_{i}$ and $\mathbf{bd}\ \mathcal{C}_{i}$
-According to (15), the cut of $\mathcal{C}_{i}$ depends on robot plan $\boldsymbol{\xi}_{\boldsymbol{\theta}_{i}}^{\gamma}$ solved by robot MPC $\boldsymbol{\pi}_{\boldsymbol{\theta}_{i}}^{\gamma}$ in (3) with $\boldsymbol{\theta}_{i}\in\Theta_{i-1}$.
-This suggests some relationship between $\boldsymbol{\theta}_{i}$ and the boundary $\mathcal{C}_{i}$ in (18).
-In fact, the relationship between $\boldsymbol{\theta}_{i}$ and $\mathbf{bd}\ \mathcal{C}_{i}$ is given by following lemma.
+According to (14), the cutting set $\mathcal{C}_{i}$ depends on plan $\boldsymbol{\xi}_{\boldsymbol{\theta}_{i}}^{\gamma}$ which is a solution to penalty-based MPC (3) with $\boldsymbol{\theta}_{i}\in\Theta_{i-1}$.
+The following assertion states a relationship between $\boldsymbol{\theta}_{i}$ and $\mathbf{bd}\ \mathcal{C}_{i}$.
 Lemma 4.
-$\forall i,\ \boldsymbol{\theta}_{i}$ is located on $\mathbf{bd}\ \mathcal{C}_{i}$ in (18). Specifically, $\boldsymbol{\theta}_{i}$ satisfies
+$\forall i,\ \boldsymbol{\theta}_{i}$ is on $\mathbf{bd}\ \mathcal{C}_{i}$ in (17). Specifically, $\boldsymbol{\theta}_{i}$ satisfies
 $\boldsymbol{\theta}_{i}^{\mkern-1.5mu\mathsf{T}}\boldsymbol{h}_{i}=b_{i}\quad\text{and}\quad\boldsymbol{\theta}_{i}^{\mkern-1.5mu\mathsf{T}}\boldsymbol{\phi}(\boldsymbol{\xi}_{\boldsymbol{\theta}_{i}}^{\gamma})<-\phi_{0}(\boldsymbol{\xi}_{\boldsymbol{\theta}_{i}}^{\gamma})$ (18)
@@
 Proof.
-See Appendix A-B. ∎
-The above lemma states $\boldsymbol{\theta}_{i}\in\mathbf{bd}\ \mathcal{C}_{i}$.
+See Appendix A-C. ∎
+Lemma 4 states $\boldsymbol{\theta}_{i}\in\mathbf{bd}\ \mathcal{C}_{i}$.
 Its geometric interpretation is in Fig.
-5: the piecewise cut of $\mathcal{C}_{i}$ always passes $\boldsymbol{\theta}_{i}$.
-This suggests that the choice of $\boldsymbol{\theta}_{i}\in\Theta_{i-1}$ is critical in the learning process: it directly affects the position of the cut $\mathbf{bd}\ \mathcal{C}_{i}$ and consequently the volume to be cut out.
-Thus, it is important to choose a good $\boldsymbol{\theta}_{i}\in\Theta_{i-1}$ such that a large hypothesis space is removed after each cut.
-We here briefly discuss the implication of different choices of $\boldsymbol{\theta}_{i}\in\Theta_{i-1}$.
+5: the cut $\mathbf{bd}\ \mathcal{C}_{i}$ always passes $\boldsymbol{\theta}_{i}$.
+This suggests that the choice of $\boldsymbol{\theta}_{i}\in\Theta_{i-1}$ is critical in the learning process: it directly affects the position of the cut $\mathbf{bd}\ \mathcal{C}_{i}$ and consequently the removal volume.
+It is important to choose a good $\boldsymbol{\theta}_{i}\in\Theta_{i-1}$ such that a large hypothesis space is removed after each cut.
+We briefly discuss the implication of different choices of $\boldsymbol{\theta}_{i}\in\Theta_{i-1}$.
 When choosing $\boldsymbol{\theta}_{i}$ near the boundary of $\Theta_{i-1}$, as shown in Fig.
 5 and Fig.
-5, the removed volume of the hypothesis space depends on the specific direction of cut $\mathbf{bd}\,\mathcal{C}_{i}$.
-Since the direction of the cut $\mathbf{bd}\,\mathcal{C}_{i}$ will be eventually dependent upon the user’s correction direction according to (18), the removed volume of the hypothesis space can be relatively small, as in Fig.
+5, the removal of the hypothesis space depends on the specific direction of cut $\mathbf{bd}\,\mathcal{C}_{i}$.
+Since the direction of the cut $\mathbf{bd}\,\mathcal{C}_{i}$ will be eventually determined by the direction of user correction according to (17), the removal volume of the hypothesis space can be relatively small, as in Fig.
 5, or large, as in Fig.
-5, depending on the specific direction of the correction applied by the user.
-As the specific direction of user correction is hard to assume, one can intuitively see that one good choice of $\boldsymbol{\theta}_{i}\in\Theta_{i-1}$ will be near the center of $\Theta_{i-1}$.
-As such, we can ensure an averaged large cut of the hypothesis space regardless of the direction of $\mathbf{bd}\ \mathcal{C}_{i}$, as shown in Fig.
+5, depending on the specific direction of the correction.
+As the specific direction of user correction is hard to assume, one can intuitively see that a good choice of $\boldsymbol{\theta}_{i}\in\Theta_{i-1}$ will be near the center of $\Theta_{i-1}$.
+As such, it can lead to an averaged large cut of the hypothesis space regardless of the direction of $\mathbf{bd}\ \mathcal{C}_{i}$, as shown in Fig.
 5.
-We will rigorously discuss the centering choice of $\boldsymbol{\theta}_{i}$ and the guarantee for volume reduction rate of the hypothesis space in the next section.
+We will rigorously discuss the centering choice of $\boldsymbol{\theta}_{i}$ and the guarantee of volume reduction rate of the hypothesis space in the next section.
 Figure 5: Illustration of the different choices of $\boldsymbol{\theta}_{i}$.
-(a) and (b): When $\boldsymbol{\theta}_{i}$ is chosen near the set boundary, the cut volume (i.e., the volume of the removed hypothesis space) depends on the specific direction of $\mathbf{bd}\ \mathcal{C}_{i}$ (i.e., the user-specific directional correction).
+(a) and (b): When $\boldsymbol{\theta}_{i}$ is chosen near the set boundary, the removal volume depends on the specific direction of $\mathbf{bd}\ \mathcal{C}_{i}$.
 (c): When $\boldsymbol{\theta}_{i}$ is near the center, the cut volume is large on average regardless of the specific user directional correction.
-V Implementation Details and Main Results
-In this section, we will present the details for how to select $\boldsymbol{\theta}_{i}\in\Theta_{i-1}$ at each iteration $i$ of the hypothesis space cutting in our proposed Safe MPC Alignment algorithm.
-Based on the preceding geometric intuition, our choice of $\boldsymbol{\theta}_{i}$ is to achieve a large averaged reduction rate for hypothesis space volume given human directional corrections.
+V Main Results and Implementation Details
+In this section, we will present the details for how to select $\boldsymbol{\theta}_{i}\in\Theta_{i-1}$ at each learning iteration $i$ in our Safe MPC Alignment algorithm.
+Based on the previous geometric intuition, our choice of $\boldsymbol{\theta}_{i}$ is to achieve a large averaged reduction rate in the hypothesis space volume for all possible human directional corrections.
 The reduction rate enables us to establish an exponential convergence of the learning process.
-As a result, we can find an upper bound of total user corrections to guarantee the success of learning, i.e., $\boldsymbol{\theta}_{i}\in\bar{\Theta}_{H}$, or identify the misspecification of hypothesis space.
+As a result, we can find an upper bound of total human corrections for successful learning, i.e., $\boldsymbol{\theta}_{i}\in\bar{\Theta}_{H}$, or identify the misspecification of hypothesis space.
 V-A The Choice of $\boldsymbol{\theta}_{i}$
-As discussed previously, $\boldsymbol{\theta}_{i}$ needs to be picked from the "center" of $\Theta_{i-1}$ in the proposed algorithm.
-An apparent option is to choose the mass center of $\Theta_{i-1}$.
+As discussed previously, $\boldsymbol{\theta}_{i}$ needs to be picked from the "center" of $\Theta_{i-1}$ in each learning iteration.
+One option is to choose the mass center of $\Theta_{i-1}$.
 However, computing the centroid of a polytope can be #P-Hard [53].
-The methods, such as Hit-And-Run algorithm [57], which estimate the mass center by sampling, also suffer from the high sample complxity, thus are impractical in our online real-time setting.
-In [58], the authors provide an another method to choose the center of a polytope using a set of inscribed balls, but it lacks of the guarantee of the volume reduction rate.
-For computational efficiency and guaranteed convergence, we will use the center of the Maximum Volume Ellipsoid (MVE) inscribed in the convex set $\Theta_{i-1}$ as $\boldsymbol{\theta}_{i}$, shown in Fig.
+While methods like the Hit-And-Run algorithm [54] estimate the mass center by sampling, they can suffer from high sample complexity.
+Another possible choice is to use the center of the maximum inscribed ball of a polytope [55], but it lacks of the guarantee in volume reduction rate.
+For computational efficiency and guaranteed convergence, we will choose $\boldsymbol{\theta}_{i}$ as the center of the Maximum Volume Ellipsoid (MVE) inscribed in the convex set $\Theta_{i-1}$, shown in Fig.
 6.
@@
 The MVE of a given convex set $\Theta\subset\mathbb{R}^{r}$ is defined as the image of the unit ball under an affine transformation
-=$\mathcal{E}=\{\bar{H}\boldsymbol{\theta}+\bar{\boldsymbol{d}}\ |\ \|\boldsymbol{\theta}\|_{2}\leq 1\}.$ (20)
-The affine transformation parameter $\bar{H}\in\mathbb{S}^{r}_{++}$ (a $r\times r$ positive definite matrix), and $\bar{\boldsymbol{d}}\in\mathbb{R}^{r}$ is called the center of $\mathcal{E}$.
-$\bar{H},\bar{\boldsymbol{d}}$ are determined via the following optimization problem for the volume of the ellipsoid $\det H$:
-$\begin{split}\bar{H},\bar{\boldsymbol{d}}=\underset{H\in\mathbb{S}^{r}_{++},\boldsymbol{d}}{\arg\max}\quad&\log\det H\\ \rm{s.t.}\quad&\sup\nolimits_{\|\boldsymbol{\theta}\|_{2}\leq 1}\mathbbm{1}_{\Theta}(H\boldsymbol{\theta}+\boldsymbol{d})\leq 0\end{split},$ (21)
+=$\mathcal{E}=\{\bar{H}\boldsymbol{\theta}+\bar{\boldsymbol{d}}\ |\ \|\boldsymbol{\theta}\|_{2}\leq 1\},$ (19)
+where the affine transformation parameter $\bar{H}\in\mathbb{S}^{r}_{++}$ (a $r\times r$ symmetric positive definite matrix), and $\bar{\boldsymbol{d}}\in\mathbb{R}^{r}$ is called the center of $\mathcal{E}$.
+$\bar{H},\bar{\boldsymbol{d}}$ are determined via the following optimization problem to maximize the volume of the ellipsoid $\det H$:
+$\begin{split}\bar{H},\bar{\boldsymbol{d}}=\underset{H\in\mathbb{S}^{r}_{++},\boldsymbol{d}}{\arg\max}\quad&\log\det H\\ \rm{s.t.}\quad&\sup\nolimits_{\|\boldsymbol{\theta}\|_{2}\leq 1}\mathbbm{1}_{\Theta}(H\boldsymbol{\theta}+\boldsymbol{d})\leq 0,\end{split}$ (20)
 where the indicator function $\mathbbm{1}_{\Theta}(\boldsymbol{\theta})=0$ for $\boldsymbol{\theta}\in\Theta$ and otherwise $\mathbbm{1}_{\Theta}(\boldsymbol{\theta})=+\infty$.
-Specifically, when $\mathbbm{1}_{\Theta}$ is a polytope $\{\boldsymbol{\theta}^{\mkern-1.5mu\mathsf{T}}\boldsymbol{h}_{i}\leq b_{i},\ i=1,\dots,m\}$, the problem can be converted to a convex optimization problem with $m$ constraints:
-$\begin{split}\bar{H},\bar{\boldsymbol{d}}=\underset{H\in\mathbb{S}^{r}_{++},\boldsymbol{d}}{\arg\min}\quad&\log\det H^{-1}\\ \rm{s.t.}\quad&\left\lVert H\boldsymbol{h}_{i}\right\rVert_{2}+\boldsymbol{d}^{\mkern-1.5mu\mathsf{T}}\boldsymbol{h}_{i}\leq b_{i},\ i=1,\dots,m\end{split}.$ (22)
-In the Safe MPC Alignment algorithm, by the above definition, let $\mathcal{E}_{i-1}$ denote the MVE of $\Theta_{i-1}$ and $\bar{\boldsymbol{d}}_{i-1}$ denote the center of $\mathcal{E}_{i-1}$.
+Specifically, when $\Theta$ is a polytope $\{\boldsymbol{\theta}^{\mkern-1.5mu\mathsf{T}}\boldsymbol{n}_{i}\leq c_{i},\ i=1,\dots,m\}$, the problem can be converted to a convex optimization with $m$ constraints, see chapter 8.4.2 in [56]:
+$\begin{split}\bar{H},\bar{\boldsymbol{d}}=\underset{H\in\mathbb{S}^{r}_{++},\boldsymbol{d}}{\arg\min}\quad&\log\det H^{-1}\\ \rm{s.t.}\quad&\left\lVert H\boldsymbol{n}_{i}\right\rVert_{2}+\boldsymbol{d}^{\mkern-1.5mu\mathsf{T}}\boldsymbol{n}_{i}\leq c_{i},\ i=1,...,m.\end{split}$ (21)
+Recall that in our Safe MPC Alignment the learning update (13) is to iteratively cut the hypothesis space by a pair of hyperplanes (17), and that the initial hypothesis space (11) is a box.
+By induction, it leads to that any updated hypothesis space $\Theta_{i}$ is a convex polytope, $i=0,1,2,...$.
+Thus, for any $\Theta_{i-1}$, by the above definition, let $\mathcal{E}_{i-1}$ denote the MVE of $\Theta_{i-1}$ and $\bar{\boldsymbol{d}}_{i-1}$ denote the center of $\mathcal{E}_{i-1}$.
 Then, the choice of $\boldsymbol{\theta}_{i}$ is
 $\boldsymbol{\theta}_{i}=\bar{\boldsymbol{d}}_{i-1}.$ (22)
-Given $\Theta_{i-1}$ a convex polytope, the optimization problem (21) can be solved efficiently by convex programming (22).
+The convex optimization (20) can be solved efficiently [57].
 V-B Result 1: Finite Iteration Convergence
-In this section, we give the main result that, when using the choice of $\boldsymbol{\theta}_{i}=\bar{\boldsymbol{d}}_{i-1}$ mentioned above, the Safe MPC Alignment algorithm can converge in a finite number of iterations to find $\boldsymbol{\theta}_{H}\in\bar{\Theta}_{H}\subset\Theta_{H}$.
-This result comes from the fact that the volume reduction rate after each hypothesis space cut is upper bounded, as formally stated below.
+When choosing $\boldsymbol{\theta}_{i}$ as the center of MVE of $\Theta_{i-1}$, the following lemma states that the volume ratio ${\mathbf{Vol}(\Theta_{i})}/{\mathbf{Vol}(\Theta_{i-1})}$ after each hypothesis space cut can be upper bounded.
 Lemma 5.
-In the Safe MPC Alignment algorithm, with $\boldsymbol{\theta}_{i}$ as the center $\boldsymbol{\bar{d}}_{i-1}$ of MVE of $\Theta_{i-1}$, one has
-$\frac{\mathbf{Vol}(\Theta_{i})}{\mathbf{Vol}(\Theta_{i-1})}\leq 1-\frac{1}{r}.$ (24)
+In the Safe MPC Alignment algorithm, when choosing $\boldsymbol{\theta}_{i}\in\mathbb{R}^{r}$ as the center of MVE of $\Theta_{i-1}$, one has
+$\frac{\mathbf{Vol}(\Theta_{i})}{\mathbf{Vol}(\Theta_{i-1})}\leq 1-\frac{1}{r},$ (23)
 for all $i=1,2,3,...$ and any enclosure volume measure $\bf{Vol}$.
 Proof.
-See Appendix A-C. ∎
+See Appendix A-D. ∎
 The lemma above asserts an upper bound of the volume of the updated hypothesis space $\mathbf{Vol}(\Theta_{i})$ relative to the previous volume.
-With the above result, we can now establish the finite iteration convergence of the proposed Safe MPC Alignment, stated below.
+With this, we can now establish the finite iteration convergence of the Safe MPC Alignment, stated below.
 Theorem 1 (Successful learning in finite iterations).
-In the Safe MPC Alignment algorithm, $\boldsymbol{\theta}_{i}$ is selected as the center $\boldsymbol{\bar{d}}_{i-1}$ of MVE of $\Theta_{i-1}$, $i=1,2,3,...$.
+In the Safe MPC Alignment algorithm, $\boldsymbol{\theta}_{i}$ is selected as the center of MVE of $\Theta_{i-1}$, $i=1,2,3,...$.
 If $\bar{\Theta}_{H}\subset\Theta_{0}$, the algorithm will terminate successfully within
@@
 Proof.
-See Appendix A-D. ∎
+See Appendix A-E. ∎
 The above main result asserts a maximum iteration count for the proposed algorithm to find $\boldsymbol{\theta}_{i}\in\Theta_{H}$.
-In other words, within $K$ rounds of human corrections, the robot can successfully learn a safety set intended by the user.
+In other words, within $K$ rounds of human corrections, the robot can successfully learn a safety constraint intended by the user.
 The above result requires a pre-set of $\rho_{H}^{r}$, which is the radius of a ball contained in $\bar{\Theta}_{H}$.
 Since the size of $\bar{\Theta}_{H}$ is unknown in practice, $\rho_{H}>0$ can be interpreted as an underestimated size of the unknown user intent set $\bar{\Theta}_{H}$.
-It also serves as a termination threshold for learning convergence: a smaller $\rho_{H}$ can lead to a more concentrated learned $\boldsymbol{\theta}$.
-However, the finite iteration convergence certificate in Theorem 1 comes with a prerequisite: $\bar{\Theta}_{H}\in\Theta_{0}$ has to be satisfied in the first place in the Safe MPC Alignment algorithm.
-Although such a condition can be achieved by empirically choosing $\Theta_{0}$ large enough to contain all possible $\bar{\Theta}_{H}$ (10), it is also desired to have a systematic way to detect the hypothesis misspecification.
-In the next session, we will discuss the identification of hypothesis misspecification, where such condition is not met.
+It also serves as a termination threshold for learning convergence: a smaller $\rho_{H}$ can lead to a more concentrated of $\boldsymbol{\theta}_{H}$ learned at convergence.
+However, the finite iteration convergence in Theorem 1 comes with a prerequisite: $\bar{\Theta}_{H}\in\Theta_{0}$ has to be satisfied in the first place in the Safe MPC Alignment algorithm.
+Although such a condition can be achieved by empirically choosing $\Theta_{0}$ large enough to contain all possible $\bar{\Theta}_{H}$ (10), it is also desired to have a guaranteed method to detect the hypothesis misspecification, where such condition is not met.
+Next, we will discuss the declaration of hypothesis misspecification.
 V-C Result 2: Hypothesis Misspecification Declaration
@@
 In this case, no $\boldsymbol{\theta}\in\Theta_{0}$ can account for the user intended safety (4).
-As mentioned in [54, 60], this kind of gap between hypothesis space and true human model can lead to undesired behavior of the robot.
+As shown in [52, 58], the hypothesis misspecification can lead to the undesired behavior of the robot.
 The following result states the proposed Safe MPC Alignment algorithm can automatically certify the misspecification of the initial hypothesis space.
-Theorem 2 (Identifying misspecification).
-In the Safe MPC Alignment algorithm, $\boldsymbol{\theta}_{i}$ is selected as the center $\boldsymbol{\bar{d}}_{i-1}$ of MVE of $\Theta_{i-1}$, $i=1,2,3,...$.
+Theorem 2 (Certifying misspecification).
+In the Safe MPC Alignment algorithm, $\boldsymbol{\theta}_{i}$ is selected as the center of MVE of $\Theta_{i-1}$, $i=1,2,3,...$.
 If $\bar{\Theta}_{H}\cap\Theta_{0}=\emptyset$, the algorithm converges to a set $\Theta^{*}$ as $i\rightarrow\infty$, such that $\mathbf{Vol}(\Theta^{*})=0$ and $\Theta^{*}\cap\mathbf{int}\ \Theta_{0}=\emptyset$.
 Proof.
-See Appendix A-E. ∎
-The above result states that, when the initial hypothesis space is misspecified, $\bar{\Theta}_{H}\cap\Theta_{0}=\emptyset$, the Safe MPC Alignment algorithm will finally converge to a set of zero volume on the boundary of the initial hypothesis space $\Theta_{0}$.
-The above theorem inspires us to empirically identify the misspecification by examining the position of MVE centers during the converging process: if after the maximum iteration step $K-1$ in (25), the MVE center $\boldsymbol{\theta}_{K-1}$ is very close to the boundary of initial hypothesis set (10), there is likely hypothesis space misspecification and further adjustment on initial hypothesis space should be made.
-See Fig.
-7.
+See Appendix A-F. ∎
+The above result states that when the initial hypothesis space is misspecified, $\bar{\Theta}_{H}\cap\Theta_{0}=\emptyset$, the Safe MPC Alignment algorithm will finally converge to a set of zero volume on the boundary of the initial hypothesis space $\Theta_{0}$.
+The above theorem inspires us to empirically identify the misspecification by monitoring the position of the MVE center of $\Theta_{i}$ during the learning update.
+As shown in Fig.
+7, if after the maximum iteration $K-1$ in (24), the MVE center $\boldsymbol{\theta}_{K-1}$ is very close to the boundary of the initial hypothesis set $\Theta_{0}$, there is likely hypothesis space misspecification, and the algorithm terminates with misspecification failure.
 Figure 7: When there is misspecification, the Safe MPC Alignment algorithm will converge to some point on the boundary of $\Theta_{0}$.
-So we can use the distance from $\boldsymbol{\theta}_{K-1}$ to the boundary as a signal for misspecification.
-Specifically, recall (10) that $\Theta_{0}$ is a box with bounding vector $\boldsymbol{c}_{l},\boldsymbol{c}_{h}$, we can pre-define a small threshold $\epsilon$ to examine every entry of $\boldsymbol{\theta}_{K-1}$ and identify possible misspecification.
-If $\boldsymbol{\theta}_{K-1}$ satisfies the criterion
-$\min_{j}\ \lvert\boldsymbol{c}_{l}[j]-\boldsymbol{\theta}_{K-1}[j]\rvert\leq\epsilon\ \text{or}\ \min_{j}\ \lvert\boldsymbol{c}_{h}[j]-\boldsymbol{\theta}_{K-1}[j]\rvert\leq\epsilon,$ (26)
+So we can use the distance of $\boldsymbol{\theta}_{K-1}$ to the boundary as an indicator for misspecification.
+Specifically, recall that $\Theta_{0}$ in (10) is a box with bounds $\boldsymbol{c}_{l},\boldsymbol{c}_{h}$, we can pre-define a small threshold $\epsilon$ to examine if $\boldsymbol{\theta}_{K-1}$ is on the boundary of $\Theta_{0}$.
+If $\boldsymbol{\theta}_{K-1}$ satisfies
+$\min_{j{=}1,...,r}\ \min\big{(}\lvert\boldsymbol{c}_{l}[j]-\boldsymbol{\theta}_{K-1}[j]\rvert,\lvert\boldsymbol{c}_{u}[j]-\boldsymbol{\theta}_{K-1}[j]\rvert\big{)}\leq\epsilon$ (25)
 it means $\boldsymbol{\theta}_{K-1}$ is very close to the boundary of $\Theta_{0}$. We consider there is likely misspecification in this case.
 V-D Detailed Algorithm of Safe MPC Alignment
-With the detailed centering choice method of $\boldsymbol{\theta}_{i}$ as well as the main results on convergence and misspecification, we can establish the full safe MPC alignment.
-With initializing $\Theta_{0}$ and pre-define a termination threshold $\rho_{H}$, three steps in the hypothesis space cutting are executed until user satisfies robot motion or the iteration count reaches the maximum count $K$ in (25).
+With the main results on convergence and misspecification, we can establish the full safe MPC alignment algorithm in Algorithm 1.
+With initializing $\Theta_{0}$ and pre-define a termination threshold $\rho_{H}$, three steps in the hypothesis space cutting are executed until the user satisfies robot motion or the iteration count reaches the maximum count $K$ in (24).
 Note that during the execution of the MPC policy, $\boldsymbol{\pi}_{\boldsymbol{\theta}_{i}}^{\gamma}$, the human can perform emergency stopping and reset the robot’s state, which corresponds to Section III-C3.
-When reaching and exist after $K-1$ iterations, the algorithm checks the misspecification by (26).
-The Safe MPC Alignment algorithm with misspecification identification is presented in the Algorithm 1.
+After $K-1$ iterations, the algorithm checks the misspecification by (25).
 Initialization: Parameterized constraint $\boldsymbol{g}_{\theta}$, initial hypothesis space $\Theta_{0}$, termination threshold $\rho_{H}$ (i.e., approximate radius of $\bar{\Theta}_{H}$), $i=1$
@@
 Algorithm 1 Safe MPC Alignment Algorithm
-VI Numerical Example and User Study in Simulation Games
-In this section, we will first evaluate the proposed Safe MPC Alignment method using a numerical example with simulated human corrections knowing ground-truth safety constraints.
-Then, we will test the performance of the methods in two user studies using two simulation games developed based on MuJoCo [61].We consider two typical robot tasks.
-One is a drone navigation task and the other a 7-DOF robot arm reaching task.
-In both user studies, the robot is controlled by an MPC controller and human participants provide simulation robots with real-time correction instructions through a keyboard.
-The objective of each game is to teach a robot to successfully learn a safe MPC to move through an unknown safety-critical environment.
-The organization in this section is as follows: Subsection VI-A presents the numerical evaluation and its analysis.
-Subsections VI-B and VI-C describe the setting of the two developed simulation games used in our user study.
-Subsection VI-D details the study procedure and the recruitment of human users.
-The results of the user studies and some analysis are given in VI-E.
-Please access the accompanying codes at https://github.com/asu-iris/Safe-MPC-Alignment.
-VI-A Numerical Example
-VI-A1 Environment Setup
-We use the safe control of an inverted pendulum as the numerical example. The dynamics equation of an inverted pendulum is
+VI Numerical Examples and Analysis
+In this section, we will evaluate the proposed Safe MPC Alignment using two numerical examples: inverted pendulum and two-link torque-controlled robot arm moving in constrained environment.
+Simulated human corrections knowing ground-truth safety constraints will be used.
+We will compare the proposed method with two baselines, and conduct ablation studies on the hyperparameters of the algorithm.
+All numerical experiments are implemented in Python, running on a Dell desktop with Intel i9-13900K CPU.
+All our experiment codes are open-sourced at at https://github.com/asu-iris/Safe-MPC-Alignment.
+VI-A Inverted Pendulum
+VI-A1 Setup
+The dynamics of an inverted pendulum is
 $\ddot{\alpha}=\frac{3}{ml^{2}}(-\frac{1}{2}m\mathrm{g}l\sin{\alpha}+u-d\dot{\alpha}),$ (26)
-with $\alpha$ being the angle between the pendulum and the direction of gravity, $\mathrm{g}=10\mathrm{m/s^{2}}$ the gravity constant, $u$ the pivot torque, $m=1\mathrm{kg}$, $l=1\mathrm{m}$ and $d=0.4$ being the length, inertia, and dampening of the pendulum.
+with $\alpha$ being the angle between the pendulum and gravity, $\mathrm{g}$ the gravity constant, $u$ input torque, $m$, $l$ and $d$ the length, inertia, and damping ratio.
 We set the system state as $\boldsymbol{x}=[\alpha,\dot{\alpha}]^{\mkern-1.5mu\mathsf{T}}$ and discretize (26) by the Euler method with $\triangle t=0.02\mathrm{s}$.
@@
 The task is to control the pendulum to the target $\boldsymbol{x}^{*}=[\pi,0]^{\mkern-1.5mu\mathsf{T}}$.
-We additionally pose the following unknown state affine constraint to the motion of the pendulum
+We additionally pose an unknown global state constraint on pendulum motion:
 $0.6\alpha+\dot{\alpha}=\boldsymbol{\theta}^{H}\boldsymbol{x}\leq 3,$ (27)
-where $\boldsymbol{\theta}^{H}=[0.6,1]^{\mkern-1.5mu\mathsf{T}}$ is to be learned.
-In Safe MPC formulation (1), we set $T=40$, the step cost $c(\boldsymbol{x}_{t},u_{t})$ and terminal cost $h(\boldsymbol{x}_{T})$
-$\begin{gathered}c(\boldsymbol{x}_{t},u_{t})=Qu_{t}^{2}\\ h(\boldsymbol{x}_{T})=(\boldsymbol{x}_{T}-\boldsymbol{x}^{*})^{T}R(\boldsymbol{x}_{T}-\boldsymbol{x}^{*}),\end{gathered}$ (29)
-respectively, with $Q=0.1$ and $R=\mathrm{diag}(25,10)$.
-To ensure the safe MPC rollout trajectory to satisfy a global constraint, it is sufficient to learn a safety constraint in safe MPC defined on the closest controllable state.
-Thus, in (1), we aim to define and learn a constraint defined on the first controlled state:
+where $\boldsymbol{\theta}^{H}=[0.6,1]^{\mkern-1.5mu\mathsf{T}}$ is to be learned. All other parameter values are given in Appendix.
+In Safe MPC (1), we set $T=40$ and the cost functions as
+$\begin{gathered}c(\boldsymbol{x}_{t},u_{t})=Ru_{t}^{2}\\ h(\boldsymbol{x}_{T})=(\boldsymbol{x}_{T}-\boldsymbol{x}^{*})^{T}Q(\boldsymbol{x}_{T}-\boldsymbol{x}^{*}),\end{gathered}$ (28)
+respectively, with $R=0.1$ and $Q=\mathrm{diag}(25,10)$.
+To ensure the safe MPC policy to satisfy a global constraint, it is sufficient to learn a safety constraint defined on the closest controllable state.
+Thus, in (1), we define
 ${g}_{\boldsymbol{\theta}}(\boldsymbol{\xi})=-3+\boldsymbol{\theta}^{\mkern-1.5mu\mathsf{T}}\boldsymbol{x}_{1}.$ (29)
-In asymptotic MPC approximation, we set $\gamma=0.1$.
+In the penalty-based MPC (3), we set $\gamma=0.1$.
 Suppose the user intent set $\bar{\Theta}_{H}=\{\boldsymbol{\theta}|\left\lVert\boldsymbol{\theta}-\boldsymbol{\theta}_{H}\right\rVert\leq 0.02\}$.
-Then, simulated human corrections $\mathbf{a}$ can be generated as follows: when the system is running near the human safety boundary, i.e., ${g}_{\boldsymbol{\theta}_{H}}(\boldsymbol{\xi}_{\boldsymbol{\theta}}^{\gamma})>-\epsilon_{g}$ ($\epsilon_{g}=0.25$ here), a correction $\mathbf{a}$ can be generated in probability $p=0.3$ using the first element the signed gradient vector $\mathrm{sign}(-\nabla B(\boldsymbol{\xi}_{\boldsymbol{\theta}}^{\gamma},\boldsymbol{\theta}_{H}))$.
-Recall that $B(\boldsymbol{\xi},\boldsymbol{\theta})$ is the barrier function in (2).
-It is trivial to verify that this correction $\mathbf{a}$ has positive product with $-\nabla B(\boldsymbol{\xi}_{\boldsymbol{\theta}}^{\gamma},\boldsymbol{\theta}_{H})$, and by definition it is applied to a safe trajectory $\boldsymbol{\xi}_{\boldsymbol{\theta}}^{\gamma}$.
-The both hypotheses in section IV-A is then satisfied, and $\mathbf{a}$ can be used as a valid human correction.
-When the system state violates ${g}_{\boldsymbol{\theta}_{H}}(\boldsymbol{\xi}_{\boldsymbol{\theta}}^{\gamma})\leq 0$ or it is near the target, the pendulum state is uniformly randomly reset between $[0,0]^{\mkern-1.5mu\mathsf{T}}$ and $[\frac{2\pi}{3},3]^{\mkern-1.5mu\mathsf{T}}$.
-Simulated human correction will continue to be applied until the learned parameters $\boldsymbol{\theta}_{i}$ fall in $\bar{\Theta}_{H}$, which terminates one trial.
+The simulated human corrections $\mathbf{a}$ are generated as follows: when the system state is near the safety boundary, i.e., ${g}_{\boldsymbol{\theta}_{H}}(\boldsymbol{\xi}_{\boldsymbol{\theta}}^{\gamma})>-\epsilon_{g}$ ($\epsilon_{g}=0.25$ here), a directional correction $\mathbf{a}$ is be generated in probability $p=0.3$ using $\mathrm{sign}(-\nabla B(\boldsymbol{\xi}_{\boldsymbol{\theta}}^{\gamma},\boldsymbol{\theta}_{H}))$ (from which only the first entry is taken considering (6)).
+The correction is probabilistic to simulate the randomness of human providing corrections.
+It is easy to verify that the simulated $\mathbf{a}$ is valid in term of satisfying the hypotheses in section IV-A.
 VI-A2 Convergence Analysis
-In the first session, we evaluate the convergence of the learning algorithm without hypothesis misspecification.
-In Algorithm 1, we set the initial hypothesis $\Theta_{0}$ with box boundaries $\boldsymbol{c}_{l}=[-6,-6]^{\mkern-1.5mu\mathsf{T}}$ and $\boldsymbol{c}_{h}=[2,2]^{\mkern-1.5mu\mathsf{T}}$; thus $\Theta_{0}$ includes $\bar{\Theta}_{H}$.
-The initial value of $\boldsymbol{\theta}_{1}$ is set as $[-2,-2]^{\mkern-1.5mu\mathsf{T}}$ (i.e., the center of $\Theta_{0}$).
-With simulated human corrections, we run Algorithm 1 and track the update of $\boldsymbol{\theta}_{i}$ after every correction.
-$\boldsymbol{\theta}_{i}$ falls in the intention set $\bar{\Theta}_{H}$ after 14 corrections, triggering the termination of learning.
-We then respectively evaluate each history $\boldsymbol{\theta}_{i}$ by using it as the MPC constraint parameter to control the system starting from $\boldsymbol{x}=[0,0]^{\mkern-1.5mu\mathsf{T}}$.
-We plot the corresponding MPC rollout trajectory in $(\alpha,\dot{\alpha})$ phase plot in Fig.
-8.The affine ground-truth constraint (28) and the corresponding learned constraints $\boldsymbol{g}_{\boldsymbol{\theta}_{i}}$ are also shown in Fig.
+We first evaluate the convergence of our algorithm without hypothesis misspecification.
+In Algorithm 1, we set the initial hypothesis $\Theta_{0}$ with box boundaries $\boldsymbol{c}_{l}=[-6,-6]^{\mkern-1.5mu\mathsf{T}}$ and $\boldsymbol{c}_{u}=[2,2]^{\mkern-1.5mu\mathsf{T}}$; thus $\Theta_{0}$ includes $\bar{\Theta}_{H}$.
+The initial $\boldsymbol{\theta}_{1}=[-2,-2]^{\mkern-1.5mu\mathsf{T}}$ (i.e., the center of $\Theta_{0}$).
+With each simulated human correction, we track $\boldsymbol{\theta}_{i}$ in each learning update.
+Each update of ${\Theta}_{i}$ takes less than 30ms.
+$\boldsymbol{\theta}_{i}$ falls in the ground-truth set $\bar{\Theta}_{H}$ after 14 corrections.
+The MPC rollout trajectory (phase plot) with each $\boldsymbol{\theta}_{i}$ is shown in Fig.
 8.
+The ground-truth constraint (27) and the learned constraints $\boldsymbol{g}_{\boldsymbol{\theta}_{i}}$ are also plotted in Fig.
+8.
+From Fig.
+8, we can conclude that the ground-truth constraint can be successfully learned from simulated corrections.
 (a) Initial trajectory (with $\boldsymbol{\theta}_{1}$)
@@
 When the learning process terminates, the learned and ground-truth constraint boundaries almost overlap, suggesting the success of constraint learning from corrections.
+To further analyze the learning convergence, we run the algorithm for 10 times.
+In each run, we keep track of two metrics over correction (iteration) count and show them in Fig.
+9.
+The first metric (left panel) is the logarithm volume measure of the solved MVE $\log\det\bar{H}$ in (20); this provides a lower bound of $\Theta_{i}$ to show the exponential convergence property.
+The second (right panel) is the parameter error $\left\lVert\boldsymbol{\theta}_{i}-\boldsymbol{\theta}_{H}\right\rVert_{2}$.
 From Fig.
-8, we can conclude that the ground-truth constraint can be successfully recovered from the simulated corrections.
-The precision of this recovery is very high and the learned and ground-truth safety boundary almost overlap.
-To further analyze the convergence of the algorithm, we run the algorithm for 10 times, all of which terminate after 12-17 corrections.
-In each run, we keep track of the following two metrics over correction count, shown in Fig.
-9.
-M1 (left panel of Fig. 9): The logarithm volume measure of the solved MVE $\log\det\bar{H}$ in (21). This provides a lower bound of $\Theta_{i}$ to show the exponential convergence property.
-M2 (right panel of Fig.
-9): The euclidean distance between learned parameter $\boldsymbol{\theta}_{i}$ and $\boldsymbol{\theta}_{H}$, i.e $\left\lVert\boldsymbol{\theta}_{i}-\boldsymbol{\theta}_{H}\right\rVert_{2}$.
-This suggests the similarity between the learned constraint and the ground-truth constraint, see Fig.
-9.
-From Fig. 9, we can conclude the proposed algorithm is
+9, we can conclude the proposed method is
   • (I)
-    Fast-converging.
-During learning, the logarithm volume $\log\det\bar{H}$ has near-linear decreasing property.
-This observation is consistent with the exponential convergence in theory stated in Lemma 5.
+    Fast-converging. The logarithm volume $\log\det\bar{H}$ has near-linear decreasing speed. This is consistent with the exponential convergence in theory stated in Lemma 5.
   • (II)
     Correction-efficient.
-During learning, $\left\lVert\boldsymbol{\theta}_{i}-\boldsymbol{\theta}_{H}\right\rVert_{2}$ converges to the value lower than 0.02 within 17 corrections.
-By using 2D area as a volume measure, we can compute the upper bound of correction count by Theorem 1 :
+$\left\lVert\boldsymbol{\theta}_{i}-\boldsymbol{\theta}_{H}\right\rVert_{2}$ converges to the value lower than 0.02 within 17 corrections.
+This is consistent with the theoretical upper bound of correction count from given by Theorem 1 :
     ${\ln\frac{\tau_{r}\rho_{H}^{r}}{\mathbf{Vol}(\Theta_{0})}}/{\ln(1-\frac{1}{r})}=17.28.$ (30)
-    This suggests that the algorithm is guaranteed to converge within finite $K=17$ corrections, this is consistent with the observations in the experiment.
 (a) Logarithm MVE Volume
-(b) Loss to Ground-truth
+(b) Parameter Estimation Error
 Figure 9: (a) The plot of $\log\det\bar{H}$ with respect to correction numbers.
@@
 In all runs, the parameter estimation error can converge to a value lower than 0.02 within 17 corrections.
-VI-A3 Result Analysis for Misspecification Case
-To analyze the performance of the proposed algorithm in misspecification cases, we set the initial hypothesis $\Theta_{0}$ as $\hat{\boldsymbol{c}_{l}}=[-1,-1]^{\mkern-1.5mu\mathsf{T}}$ and $\hat{\boldsymbol{c}_{h}}=[0.8,0.8]^{\mkern-1.5mu\mathsf{T}}$ which does not contain $\bar{\Theta}_{H}$.
+VI-A3 Analysis for Misspecification Case
+To analyze the performance of our algorithm to certify misspecification cases, we set the initial hypothesis space $\Theta_{0}$ as ${\boldsymbol{c}_{l}}=[-1,-1]^{\mkern-1.5mu\mathsf{T}}$ and ${\boldsymbol{c}_{u}}=[0.8,0.8]^{\mkern-1.5mu\mathsf{T}}$ which does not contain $\bar{\Theta}_{H}$.
 In this setting, the algorithm automatically stops after 7th correction.
@@
 In Fig.
-10, $\boldsymbol{\theta}_{i}$ is continuously moving towards the boundary of the initial hypothesis space $\Theta_{0}$.
+10, $\boldsymbol{\theta}_{i}$ is continuously moving towards the boundary of $\Theta_{0}$.
 This fact verifies Theorem 2 and justifies the proposed method is capable of misspecification detection.
@@
 When misspecification happens, $\boldsymbol{\theta}$ will converge toward the boundary of $\Theta_{0}$, as stated by Theorem 2.
-VI-B Drone Navigation Game
-VI-B1 Environment Setup
+VI-B Two-link Robot Arm Reaching with Complex Constraint
+VI-B1 Setup
+The dynamics if torque controlled two-link robot arm moving in the horizontal plane is $M(\boldsymbol{q})\ddot{\boldsymbol{q}}+\boldsymbol{c}(\boldsymbol{q},\dot{\boldsymbol{q}})=\boldsymbol{\tau}$, where $M(\boldsymbol{q})$ is the inertia matrix; $\boldsymbol{c}(\boldsymbol{q},\dot{\boldsymbol{q}})$ is the Coriolis term.
+$\boldsymbol{q}=[q_{1},q_{2}]^{\mkern-1.5mu\mathsf{T}}$ defines the two joint angles of the robot.
+The system state is defined as $\boldsymbol{x}=[\boldsymbol{q},\dot{\boldsymbol{q}}]$.
+The control input $\boldsymbol{u}$ is the joint torque $\boldsymbol{\tau}=[\tau_{1},\tau_{2}]^{\mkern-1.5mu\mathsf{T}}$, i.e.
+$\boldsymbol{u}=\boldsymbol{\tau}$.
+The other detailed physical parameters are given in Appendix.
 As shown in Fig.
-11, the goal of this drone navigation game is let a human user to teach a done to successfully navigate through a narrow gate (rectangle frame in yellow) whose position is unknown, given different flying target positions (shown in yellow points behind the gate).
-The drone starts from $[x_{\text{init}},y_{\text{init}},z_{\text{init}}]^{\mkern-1.5mu\mathsf{T}}$ with $x_{\text{init}}=0.1$, $z_{\text{init}}=0.5$, and $y_{\text{init}}$ uniformly randomly in an interval $[4,6]$.
-The starting quaternion $\boldsymbol{q}_{\text{init}}=[1,0,0,0]^{\mkern-1.5mu\mathsf{T}}$.
-We set three different target positions at $[19,1,9]^{\mkern-1.5mu\mathsf{T}}$, $[19,5,9]^{\mkern-1.5mu\mathsf{T}}$ and $[19,9,9]^{\mkern-1.5mu\mathsf{T}}$, with target quaternion $[1,0,0,0]^{\mkern-1.5mu\mathsf{T}}$.
-For the narrow gate (yellow rectangle), both its size and position are unknown, and one only roughly know its x position $x_{\text{gate}}\approx 10.0$.
-The drone navigation game is developed using MuJoCo physics engine [61].
+12, the robot starts at $\boldsymbol{x}_{\mathrm{init}}=[-\frac{\pi}{2},0,0,0]^{\mkern-1.5mu\mathsf{T}}$ and the task is to reach a target $\boldsymbol{x}^{*}=[\frac{\pi}{2},0,0,0]^{\mkern-1.5mu\mathsf{T}}$.
+In Safe MPC (1), we set $T=30$, the step cost functions
+$\begin{gathered}c(\boldsymbol{x}_{t},\boldsymbol{u}_{t})=\boldsymbol{x}_{t}^{\mkern-1.5mu\mathsf{T}}P\boldsymbol{x}_{t}+\boldsymbol{u}_{t}^{\mkern-1.5mu\mathsf{T}}R\boldsymbol{u}_{t},\\ h(\boldsymbol{x}_{T})=(\boldsymbol{x}_{T}-\boldsymbol{x}^{*})^{T}Q(\boldsymbol{x}_{T}-\boldsymbol{x}^{*}),\end{gathered}$ (31)
+with $P=\mathrm{diag}(1,1,2,2)$, $R=\mathrm{diag}(0.1,0.1)$ and $Q=\mathrm{diag}(5,5,10,10)$.
+As shown by the black lines (obstacles) in Fig.
+12, we define a complex reaching environment with non-convex obstacles and narrow corridor.
+The ground-truth global safety constraint corresponding to Fig.
+12 is given
+${g}_{\boldsymbol{\theta}_{H}}(\boldsymbol{\xi})=-1+\boldsymbol{\theta}_{H}^{\mkern-1.5mu\mathsf{T}}\boldsymbol{\phi}$ (32)
+with $\boldsymbol{\phi}{=}[\mathrm{tanh}(x_{e}y_{e}),\mathrm{tanh}(x_{e}^{2},\mathrm{tanh}(y_{e}^{2}),\mathrm{tanh}(x_{e}),\mathrm{tanh}(y_{e})]^{\mkern-1.5mu\mathsf{T}}$, $\boldsymbol{\theta}_{H}=[0.78,0.91,-0.52,0.52,-1.04]^{\mkern-1.5mu\mathsf{T}}$, and $[x_{e},y_{e}]^{\mkern-1.5mu\mathsf{T}}$ is the end-effector position of the robot arm.
+The human intent set is $\bar{\Theta}_{H}{=}\{\boldsymbol{\theta}|\left\lVert\boldsymbol{\theta}-\boldsymbol{\theta}_{H}\right\rVert\leq 0.2\}$.
+We set the initial hypothesis $\Theta_{0}$ with box boundaries $\boldsymbol{c}_{l}=[-3,\dots,-3]^{\mkern-1.5mu\mathsf{T}}$ and $\boldsymbol{c}_{u}=[3,\dots,3]^{\mkern-1.5mu\mathsf{T}}$.
+The simulated human corrections are generated in the similar way to that in Section VI-A.
+Figure 11: The heatmap of a learned constraint $g$. The learned and ground-truth constraints are almost identical, showing the success of our method for learning complex constraint from corrections.
+VI-B2 Result
+Our method can successfully learn the safety constraint within 25-30 simulated corrections.
+One instance of the learned constraint is shown in Fig.
+11.
+The comparison of the robot reaching motion without and with the learned safety constraint is shown in Fig.
+12.
+The robot is capable of operating in the intended safe region after the algorithm converges.
+Those results show that the efficacy of the proposed method for solving system with complex safety constraints.
+(a) Trajectory Before Corrections
+(b) Trajectory After Corrections
+Figure 12: The trajectory of the two-link robot arm (a) without and (b) with learned safety constraint. The robot successfully achieves the task with the learned safety constraint.
+VI-C Baseline Comparison
+In this session, we compare our method with two baseline methods using magnitude corrections in the pendulum environment (the task and model settings are identical to Section VI-A).
+As reviewed previously, the direct literature learning safety constraints from human feedback is sparse to the authors’ best knowledge.
+Thus, the chosen baselines are an adaptation from the most relevant methods.
+Baseline 1: gradient matching.
+This baseline is inspired by [59], where raw human correction (with magnitude) is directly used as the gradient direction.
+In a similar spirit, when a human feedback $\mathbf{a}$ is made on robot trajectory $\xi_{\boldsymbol{\theta}}^{\gamma}$, we directly use the human magnitude correction as the true gradient label to learn the constraint $\boldsymbol{\theta}$.
+This leads to a learning loss
+$\mathcal{L}=\left\lVert\mathbf{a}+\nabla B(\boldsymbol{\xi}_{\boldsymbol{\theta}}^{\gamma})\right\rVert^{2}$ (33)
+Baseline 2: maximum likelihood learning.
+This baseline follows [60], where it views $e^{{g}_{\boldsymbol{\theta}}(\boldsymbol{\xi})}$ as a probability of human applying the "stop-feedback" to robot motion $\boldsymbol{\xi}$, when the robot reaches in a $\epsilon_{g}$-proximity of the safety boundary, i.e., when ${g}_{\boldsymbol{\theta}_{H}}(\boldsymbol{\xi}_{\boldsymbol{\theta}}^{\gamma})>-\epsilon_{g}$ Thus, a dataset of trajectory-correction paris, i.e., $\mathcal{D}=\{\mathbf{a}_{i},\boldsymbol{\xi}_{i}\}_{i=0}^{N}$, can be used as the training data to learn $\boldsymbol{\theta}$ using maximum likelihood, while also penalizing positivity of ${g}_{\boldsymbol{\theta}}(\boldsymbol{\xi}_{i})$ as a regulizer term.
+This leads to the following loss
+$\mathcal{L}=\sum\nolimits_{i=1}^{N}\big{(}-{g}_{\boldsymbol{\theta}}(\boldsymbol{\xi}_{i})+\alpha\cdot\mathrm{softplus}(g_{\boldsymbol{\theta}}(\boldsymbol{\xi}_{i}))\big{)}$ (34)
+In our comparison, we choose $N=15$ (similar to correction number needed for our method to converge) and $\alpha=5$ for its best performance.
+The comparison between our method and Baseline 1 focuses on data efficiency.
+The convergence criteria of $\left\lVert\boldsymbol{\theta}-\boldsymbol{\theta}_{H}\right\rVert\leq 0.02$ is set for both methods.
+50 independent runs are performed and the number of correction needed for convergence are counted.
+We use Adam optimizer [61] with learning rate 0.02.
+We summarize the results in Table I.
+The parameter estimation error $\left\lVert\boldsymbol{\theta}-\boldsymbol{\theta}_{H}\right\rVert$ versus learning iteration (correction count) for a best run (i.e., best performance for the baseline) is plotted in Fig.
+13(a).
+The comparison results demonstrate the significant data efficiency of the proposed method compared to gradient matching baseline, even though the latter uses both magnitude and direction information of corrections to learn.
+TABLE I: Comparison of Safe MPC Alignment and Baseline I
+Method Correction count for convergence
+Safe MPC Alignment $11.76\pm 2.14$
+Gradient Matching $240.38\pm 120.70$
+(a) Comparison with Baseline 1
+(b) Comparison with Baseline 2
+Figure 13: (a) Comparison with gradient matching baseline: the parameter estimation error $\left\lVert\boldsymbol{\theta}-\boldsymbol{\theta}_{H}\right\rVert$ versus learning iteration (correction count).
+This is best run (among 50 runs) with the least corrections for the baseline.
+Our method converges after 12 corrections, while the baseline needs more than 160 corrections to converge.
+(b) Comparison with the maximum likelihood baseline under the different $\epsilon_{g}$.
+In the plot, the y-axis is $\left\lVert\boldsymbol{\theta}-\boldsymbol{\theta}_{H}\right\rVert$ at convergence and the error bar denotes the standard deviation.
+The maximum likelihood baseline fails to learn the correct weights when $\epsilon_{g}$ is large, while our method consistently shows high learning accuracy.
+The comparison between our method and maximum likelihood baseline focuses on the varying positions of simulated correction, i.e., varying $\epsilon_{g}$.
+Recall that the simulated corrections will be generated only when ${g}_{\boldsymbol{\theta}_{H}}(\boldsymbol{\xi}_{\boldsymbol{\theta}}^{\gamma})>-\epsilon_{g}$.
+The value of $\epsilon_{g}$ determines the position of the correction applied.
+In our comparison, we vary $\epsilon_{g}$ from 0.1 to 0.9 and track the converged parameter error for both methods.
+As shown in Fig.
+13(b), safe MPC alignment consistently converges to the ground-truth $\boldsymbol{\theta}_{H}$ regardless of $\epsilon_{g}$, while the maximum likelihood baseline fails when $|\epsilon_{g}|$ is large.
+The results suggest the robustness of the proposed method against different correcting conditions compared to maximum likelihood baseline method.
+VI-D Ablation Study
+In this session, we conduct two ablation studies to show the performance of our algorithm given different design parameters/aspects.
+The task is still based on pendulum environment with settings identical to Section VI-A).
+VI-D1 Choosing $\boldsymbol{\theta}_{i}$: MVE centering versus random sampling
+In the first study, we compare two strategies in selecting $\boldsymbol{\theta}_{i}$ from the hypothesis space $\Theta_{i-1}$ in our algorithm: MVE centering (used in our method) and random sampling.
+For the latter, $\boldsymbol{\theta}_{i}$ is chosen by sampling from the uniform distribution supported by $\Theta_{i-1}$.
+We compare the total number of corrections needed at learning convergence under the two strategies.
+With 50 independent runs, we observe that the MVE centering requires $11.76\pm 2.14$ corrections, while random sampling requires $11.88\pm 3.50$ corrections.
+While the mean value of the correction count looks similar, the random sampling shows a slightly larger variance.
+This is understandable given the sampling randomness.
+We also observed that sampling process introduced large computational overhead, which is around 25% time slower than MVE centering.
+In contrast, the MVE centering used in our algorithm shows low variance and computational efficiency in finding $\boldsymbol{\theta}_{i}$.
+Because selecting $\boldsymbol{\theta}_{i}$ is casted as a convex programming for MVE centering, such advantage would be more significant for high-dimensional parameter space.
+VI-D2 Choosing $\gamma$
+In the second study, we investigate how the setting of $\gamma$ affects the learning performance.
+We vary $\gamma$ from 0.01 to 10.
+Fig.
+14 shows the total number of corrections needed at convergence versus different $\gamma$s.
+The result in Fig.
+14 shows that the learning performance remains the same given different settings of $\gamma$.
+Since $\gamma$ is weights parameter to control the safety conservativeness in the penalty-based MPC policy in (3), it means our method can successfully learn a safety constraint when the robot executes with safe MPC policy of different conservativeness levels.
+Figure 14: The total number of corrections needed for convergence under different $\gamma$, the x-axis is of log scale.
+Safe MPC alignment can converge with a similar number of corrections (iterations) under different levels of $\gamma$.
+VII User Study in Simulation Games
+In this section, we test the Safe MPC Alignment method in two real-human user studies using two simulation games developed based on MuJoCo [62].
+One is the drone navigation game and the other the Franka robot arm reaching game.
+In both games, the robot is controlled by an MPC controller and human participants provide real-time corrections through a keyboard.
+The objective of each game is to teach a robot to successfully learn a safe MPC to move through environments with unknown obstacles.
+In following, Subsection VII-A describes the basic setting of the two games.
+Subsection VII-B details the user study procedure and the recruitment of human participants.
+The results and analysis are given in VII-C.
+VII-A Setting of Two Simulation Games
+In both user studies, human participants are asked to give correction to the following two physics-based simulation games developed based on MuJoCo.
+(1) Drone Navigation Game.
+As shown in Fig.
+15, the goal of this game is to let a drone to successfully navigate through a narrow gate (yellow rectangle) whose position is unknown, reaching different targets (shown in yellow points behind the gate).
+The safety constraint $\boldsymbol{\theta}^{\mkern-1.5mu\mathsf{T}}\boldsymbol{\phi}$ is parameterized using a linear combination of spacial radial basis functions (RBFs), defined on y-z plane at $x=10$.
+The drone’s control inputs are the thrusts of four rotors and human correction is applied via keyboard, with specific keys mapping to specific correction commands to the drone’s control space.
+See the detailed task setup and keyboard interface in Appendix B-C.
+(2) Franka arm reaching game.
+As in Fig.
+16, the goal of the Franka arm reaching game is to teach a Franka arm such that its gripper hand successfully reaches different target positions (cyan ball) through an unknown narrow slot (two yellow bars).
+Note that the safety constraint is only applied on the end effector and the collision between the link and the slot is ignored.
+The safety constraint function $\boldsymbol{\theta}^{\mkern-1.5mu\mathsf{T}}\boldsymbol{\phi}$ is parameterized as a weighted combination of RBFs defined on the $z$ axis and rotations around the $x$ axis.
+Human corrections are applied via keyboard, with specific keys mapping to specific correction commands applied to the robot control space $\mathbb{R}^{6}$ (we use operational space control).
+See the detailed task setup and keyboard interface in Appendix B-D.
+In both games, the robots are controlled with MPC controller (3) but fail to maintain safety with the initial safety constraints.
+To simulate an emergency stop, the “Enter” key is used as a reset command, allowing users to reset the environment to its initial state whenever the participant deems the robot violating the safety constraint.
 (a) Tracking View
@@
 Figure 15: Drone navigation game.
-(a) The tracking view seen by the user in the drone user study.
-(b) The fixed view seen by the user in the drone user study.
-Users can precept the motion of the drone in 3D space with the help of these two views.
-The $xyz$-axis (red, green, and blue lines), the gate (yellow bars for the edge, orange bars for the outer frame), the drone (white robot model), the local plan outputted from MPC (red curve) and current motion target (yellow balls) are visible to users.
-in runtime, the user will see one target at a time.
-Since the location of the gate is unknown, the drone needs to learn a safety constraint in its safe MPC policy with humans guidance in order to successfully fly through the gate.
-In the drone MPC policy (1), the dynamics and its parameters are set following ([7]); and we generically set the constraint function ${g}_{\boldsymbol{\theta}}(\boldsymbol{\xi})$ using a set of $N$ Gaussian radial basis functions (RBFs)
-${g}_{\boldsymbol{\theta}}(\boldsymbol{\xi})=\phi_{0}+\sum\nolimits_{i=1}^{N}\theta_{i}\phi_{i}(\boldsymbol{\xi}).$ (32)
-As discussed in Section III-C4, we fix $\phi_{0}=-1$ to avoid learning ambiguity, and use $N=20$ temporally accumulated RBFs defined as
-$\phi_{i}(\boldsymbol{\xi})=\sum\nolimits_{t=0}^{T}\beta_{t}\mathrm{RBF}_{i}(\boldsymbol{x}_{t}),$ (33)
-where $\beta_{t}$ is a weighting factor to control or balance the immediate and future constraint satisfaction of robot MPC control.
-If one emphasizes instant safety, more weights will be given to the initial time step.
-In our implementation, inspired by the practice in RL, we choose an exponentially decaying $\beta_{t}$ starting at $t=5$, i.e.
-$\beta_{t}=0,\ t<5$ and $\beta_{t}=0.9^{t-5},\ t\geq 5$ to achieve better predictive safety performance.
-In (33), each Guassian RBF is defined as
-$\mathrm{RBF}_{i}(\boldsymbol{x}_{t})=\begin{cases}\exp(-\epsilon^{2}\left\lVert r_{xy,t}-c_{xy,i}\right\rVert^{2}),\,\,i=1,...,10\\ \exp(-\epsilon^{2}\left\lVert r_{xz,t}-c_{xz,i}\right\rVert^{2}),\,\,i=11,...,20\end{cases}.$ (34)
-Here, the variables $r_{xy,t}=[r_{x,t},r_{y,t}]^{\mkern-1.5mu\mathsf{T}}$ and $r_{xz,t}=[r_{x,t},r_{z,t}]^{\mkern-1.5mu\mathsf{T}}$ with $r_{x},r_{y},r_{z}$ from $\boldsymbol{x}_{t}$.
-The $i$th center $c_{xy,i}=[c_{x,i},c_{y,i}]^{\mkern-1.5mu\mathsf{T}}$ or $c_{xz,i}=[c_{x,i},c_{z,i}]^{\mkern-1.5mu\mathsf{T}}$ is defined with $c_{x,i}={x_{\text{gate}}}=9.8$, $c_{y,i}$ and $c_{z,i}$ are evenly chosen on interval $[0,10]$.
-We set $\epsilon=0.45$.
-The initial hypothesis space $\Theta_{0}$ for possible $\boldsymbol{\theta}$ in (32) is set as $\boldsymbol{c}_{l}=[-80,\dots,-80]^{\mkern-1.5mu\mathsf{T}}$ and $\boldsymbol{c}_{h}=[100,\dots,100]^{\mkern-1.5mu\mathsf{T}}$.
-$\gamma$ is 50.
-This initialization blocks the drone from advancing to the gate at the beginning of the game, leading to conservative behavior in the learning process.
-The drone will proceed toward the gate after few corrections.
-VI-B2 Human Correction Interface
-The motion of the drone is rendered in two camera views in MuJoCo visualization (Fig.
-11) for better human perception: one is the first-person view tracking the drone, the other a third-person fixed view.
-The gate, the current MPC planning solution, and the target position are visible to the users.
-Human users apply corrections to the drone via a keyboard during the drone flying.
-The user is informed of the mappings between the key press and the corresponding correction in the game, listed in Table I.
-The safety constraint (32) in the drone’s MPC policy is immediately updated once a user correction is detected.
-When the user presses "ENTER", the drone is reset to the initial position and the target will also change.
-TABLE I: keyboard commands in uav user study
-Key Correction Interpretation
-"up" $[-1,0,1,0]^{\mkern-1.5mu\mathsf{T}}$ Go forward (adding positive y-axis torque)
-"down" $[1,0,-1,0]^{\mkern-1.5mu\mathsf{T}}$ Go backward (adding negative y-axis torque)
-"left" $[0,1,0,-1]^{\mkern-1.5mu\mathsf{T}}$ Go left (adding negative x-axis torque)
-"right" $[0,-1,0,1]^{\mkern-1.5mu\mathsf{T}}$ Go right (adding positive x-axis torque)
-"shift" $[1,1,1,1]^{\mkern-1.5mu\mathsf{T}}$ Go up (increasing all thrusts)
-"ctrl" $-[1,1,1,1]^{\mkern-1.5mu\mathsf{T}}$ Go down (decreasing all thrusts)
-"enter" N/A Reset robot position and switch target.
-VI-C Franka Constrained Reaching Game
-VI-C1 Environment Setup
-As shown in Fig.
-12, the goal of the Franka constrained reaching game is to teach a simulated Franka robot arm such that its gripper hand successfully reaches different target positions (shown in cyan ball) through a narrow slot (shown by two yellow bars) with unknown position or geometry.
-Specifically, the arm starts at a configuration in which the end-effector position $\boldsymbol{r}_{\text{init}}=[0.22,-0.5,0.48]^{\mkern-1.5mu\mathsf{T}}$ and altitude $\boldsymbol{q}_{\text{init}}=[0,-0.707,0,0.707]^{\mkern-1.5mu\mathsf{T}}$.
-The reaching target position $\boldsymbol{r}^{*}$ will be dynamically chosen from three options $[-0.6,-0.5,0.4]^{\mkern-1.5mu\mathsf{T}}$, $[-0.6,-0.5,0.5]^{\mkern-1.5mu\mathsf{T}}$, $[-0.6,-0.5,0.6]^{\mkern-1.5mu\mathsf{T}}$.
-The reaching target attitude is $\boldsymbol{q}^{*}=[0,-0.707,0,0.707]^{\mkern-1.5mu\mathsf{T}}$.
-The slot size and position are unknown except its rough x-position which is $x_{\text{slot}}\approx-0.15$.
+(a) Tracking view.
+(b) Fixed view.
+Users can see the motion of the drone with the help of these two views.
+The $xyz$-axis (red, green, and blue axes), the gate (yellow bars), the drone, the local plan solved from the current MPC (red curve) and the motion target (yellow balls) are visible to users.
+In each game reset, the user will see one target randomly.
 (a) Starting Configuration
 (b) Motion Illustration
-Figure 12: Franka Constrained Reaching Game.
-(a) The starting configuration of the arm in the user study.
+Figure 16: Franka arm reaching game.
+(a) The starting pose of the arm.
 RGB arrows denote the world $xyz$-frame.
-(b) Illustration of one successful reach by adjusting the position and altitude of the gripper, long green arrow denotes the motion direction.
-Users can precept the motion of the robot arm in 3D space by adjusting the view.
-The bars, the robot arm, and current motion target (cyan ball) are all visible to users.
-We adopt a two-level control scheme, where a safe MPC policy (high-level) is defined for the end effector, which generates the desired velocity for a given low-level operational space controller to track.
-In the safe MPC policy, we model the end-effector dynamics as
-$\begin{gathered}\dot{\boldsymbol{r}}=\boldsymbol{v}^{d},\\ \dot{\boldsymbol{q}}=\frac{1}{2}\Omega(R^{\mkern-1.5mu\mathsf{T}}(\boldsymbol{q})\boldsymbol{w}^{d})\boldsymbol{q},\end{gathered}$ (35)
-where $\boldsymbol{r}$ and $\boldsymbol{q}$ is the position and the quaternion of the arm end-effector in the world frame.
-$R^{\mkern-1.5mu\mathsf{T}}(\boldsymbol{q})$ is the rotation matrix from the end-effector frame to the body frame.
-$\Omega(\cdot)$ is the matrix operator of the angular velocity for quaternion multiplication.
-Define the robot state and control input
-$\boldsymbol{x}_{t}=[\boldsymbol{r}_{t},\ \boldsymbol{q}_{t}]\in\mathbb{R}^{7},\quad\boldsymbol{u}_{t}=[\boldsymbol{v}^{d}_{t},\ \boldsymbol{w}^{d}_{t}]\in\mathbb{R}^{6},$ (36)
-respectively.
-The above end-effector dynamics is discretized with $\triangle t=0.1\mathrm{s}$.
-In (1), the control horizon is $T=20$.
-Let $\boldsymbol{x}^{*}=[\boldsymbol{r}^{*}\ \boldsymbol{q}^{*}]$ denote the target end-effector pose.
-The step cost $c(\boldsymbol{x}_{t},\boldsymbol{u}_{t})$ and terminal cost $h(\boldsymbol{x}_{T})$ in (1) is set to
-$c(\boldsymbol{x}_{t},\boldsymbol{u}_{t})=$ $(\boldsymbol{r}_{t}-\boldsymbol{r}^{*})^{\mkern-1.5mu\mathsf{T}}Q_{r}(\boldsymbol{r}_{t}-\boldsymbol{r}^{*})+k_{q}\mathbf{diff}(\boldsymbol{q}_{t},\boldsymbol{q}^{*})$ (37)
-$\quad\quad+\boldsymbol{v}_{I,t}^{\mkern-1.5mu\mathsf{T}}R_{v}\boldsymbol{v}_{I,t}+\boldsymbol{w}_{I,t}^{\mkern-1.5mu\mathsf{T}}R_{w}\boldsymbol{w}_{I,t},$
-$h(\boldsymbol{x}_{T})=$ $(\boldsymbol{r}_{T}-\boldsymbol{r}^{*})^{\mkern-1.5mu\mathsf{T}}S_{r}(\boldsymbol{r}_{T}-\boldsymbol{r}^{*})+\mu_{q}\mathbf{diff}(\boldsymbol{q}_{T},\boldsymbol{q}^{*}).$
-In this game, the cost parameters are set as $Q=\mathrm{diag}(2,2,2)$, $k_{q}=1$, $R_{v}=\mathrm{diag}(30,30,1)$, $R_{w}=\mathrm{diag}(0.85,0.85,0.85)$, $S_{r}=\mathrm{diag}(80,80,80)$, and $\mu_{q}=60$.
-The output of the MPC controller is sent to the low-level controller to track.
-Similar to Section VI-B, to constrain $r_{z}$ and $\boldsymbol{q}$ for proper height and tilt in order to pass through the slot.
-We use a set of $N=20$ Gaussian radial basis function (RBFs) to parameter the safety constraint:
-${g}_{\boldsymbol{\theta}}(\boldsymbol{\xi})=\phi_{0}+\sum\nolimits_{i=1}^{N}\theta_{i}\phi_{i}(\boldsymbol{\xi}),$ (38)
-with
-$\phi_{i}(\boldsymbol{\xi})=\sum\nolimits_{t=0}^{T}\beta_{t}\mathrm{RBF}_{i}(\boldsymbol{x}_{t}),\quad\text{and}$ (39)
-$\mathrm{RBF}_{i}(\boldsymbol{x}_{t}){=}\begin{cases}\sigma(r_{x,t})k_{z}\exp{(-\epsilon_{z}^{2}(r_{z,t}-c_{z,i})^{2})},\,\,i{=}1,...,10\\ \sigma(r_{x,t})\exp{(-\epsilon_{q}^{2}\mathbf{diff}(\boldsymbol{q}_{t},\boldsymbol{q}_{c,i}))},\,\,i{=}11,...,20\end{cases}.$ (40)
-Here, $k_{z}$ is used to balance the scaling between position and orientation $\mathrm{RBF}$s, $\sigma(r_{x,t})$ is defined as
-$\sigma(r_{x,t})=\mathrm{sigmoid}(50\big{(}(r_{x,t}-x_{\text{slot}})^{2}-(\frac{w}{2})^{2}\big{)}),$ (41)
-which is used to constrain the $\mathrm{RBF}_{i}$ only taking effect around $x$ range $[x_{\text{slot}}-\frac{w}{2},x_{\text{slot}}+\frac{w}{2}]$ (recall slot’s $x$ location, slot, is a approximate value).
-In the above definition, $w=0.3$, $k_{z}=0.05$, $\epsilon_{z}=12$ and $\epsilon_{q}=1.8$.
-The center of positional $\mathrm{RBF}$, ${c_{z,i}}$, is evenly chosen on interval $[0,1]$, and the center of orientation $\mathrm{RBF}$, ${\boldsymbol{q}_{c,i}}$, is generated by $(\cos{\frac{\alpha}{2}},\sin{\frac{\alpha}{2}},0,0)*\boldsymbol{q}_{\text{init}}$ with $\alpha$ evenly choosen from $[-\pi/4,3\pi/4]$ to make sure ${\boldsymbol{q}_{c,i}}$ covers all possible orientations.
-$\beta_{t}=0.9^{t}$ if $t<5$ and 00 otherwise.
-The bounds of the initial hypothesis space $\Theta_{0}$ is set to $\boldsymbol{c}_{l}=[-3,\dots,-3]^{\mkern-1.5mu\mathsf{T}}$ and $\boldsymbol{c}_{h}=[5,\dots,5]^{\mkern-1.5mu\mathsf{T}}$.
-$\gamma$ is set to 1.5.
-This initialization leads to a conservative MPC policy, preventing the arm from approaching the target at the beginning.
-VI-C2 Human Correction Interface
-The motion of the robot arm is rendered in the free-view mode in MuJoCo visualization (Fig.
-12), and users can drag to select his/her comfortable view.
-The slot and the random reaching target are visible to the users.
-Human users apply their corrections to the robot arm via a keyboard.
-Users are informed with the following key mapping between the key press and the corresponding correction in the game.
-The safety constraint (38) in Franka arm’s safe MPC policy is updated immediately once a user correction is detected.
-When user press "ENTER", the arm is reset to the initial position and the reaching target will also change.
-TABLE II: keyboard commands in robot arm user study
-Key Correction Interpretation
-"left" $[0,0,0,0,0,1]^{\mkern-1.5mu\mathsf{T}}$ Clock-wise rotation of the gripper
-"right" $[0,0,0,0,0,-1]^{\mkern-1.5mu\mathsf{T}}$ Anti-clock-wise rotation of the gripper
-"shift" $[0,0,0,-1,0,0]^{\mkern-1.5mu\mathsf{T}}$ Go up in world frame
-"ctrl" $[0,0,0,1,0,0]^{\mkern-1.5mu\mathsf{T}}$ Go down in world frame
-"Enter" N/A Reset robot position and switch target.
+(b) One example of successful reaching, and long green arrow denotes the motion direction.
+Users can see the motion of the robot arm by adjusting the view.
+The slot, the robot arm, and current motion target (cyan ball) are all visible to users.
 (a) After 2 corrections
@@
 (c) After 12 corrections
-Figure 13: Illustrations of the drone MPC rollout trajectory after $k=2,6,12$ corrections.
-The red stars denote the target positions and the blue curves denote the trajectories.
-The gate is drew in black lines and collision is denoted by green crosses.
-User’s corrections typically lower the flying drone to make it pass through the gate.
-VI-D Participants, Procedures, and Metrics
-VI-D1 Participants
-We recruited 15 participants from the student population of Arizona State University, covering undergraduate, master’s, and PhD students with different backgrounds.
+Figure 17: Illustrations of the drone MPC rollout trajectory after $2,6,12$ corrections.
+The red stars denote the targets and blue curves denote the trajectories.
+The gate is drawn as black frame and the collision is highlighted in green crosses.
+User corrections typically lower the drone to make it pass through the gate.
+VII-B Participants, Procedures, and Metrics
+VII-B1 Participants
+We recruited 15 participants from Arizona State University, including undergraduates, masters, and PhD students with different backgrounds.
 Among all participants, 13 are male and 2 female.
-All participants are novices to the user study.
+All participants are new to the user study.
 This user study had been reviewed and approved by the Institutional Review Board (IRB) of Arizona State University, and all participants had signed the consent forms.
-VI-D2 Procedure
+VII-B2 Procedure
 Each participant was instructed to play the previously developed two simulation games. In each game, each participant plays 10 trials.
-In the drone navigation game, a trial is deemed successful when the drone navigates through the gate without collision and successfully reaches different targets 6 consecutive times with the learned safe MPC policy.
-A trial is deemed failed when there are too many corrections, causing the crash of the MVE solver, or the user believes the robot cannot be corrected anymore and gives up.
-In the Franka reaching game, a trial is deemed successful when the robot arm reaches different targets without collision with the slot frame with the learned safe MPC controller.
-The failure of one trial is determined when there are too many corrections or the user believes the robot cannot be corrected anymore and gives up.
+In the drone navigation game, a trial is deemed successful when the drone navigates through the gate without collision and reaches different targets 6 consecutive times with the learned safe MPC policy.
+A trial is deemed failed when there are too many corrections (exceeding 25), or the user believes the robot cannot be corrected anymore and gives up.
+In the Franka reaching game, a trial is deemed successful when the robot arm reaches different targets with the end-effector having no collision with the slot.
+The failure of one trial is determined when there are too many corrections (exceeding 25) or the user believes the robot cannot be corrected anymore and gives up.
 The result of every trial is recorded.
-Before formally starting the study, every user is instructed about the settings of the game and is given 2 chances for hands-on trying of the environments to get familiar with the keys and commands.
-VI-D3 Metrics for Evaluation
-Two major metrics are used to evaluate the performance of the proposed safe MPC alignment algorithm in our user study:
+Before formally starting the study, every user is instructed about the settings of the game and is given 2 chances to interact with the environment hands-on to get familiar with the keys.
+VII-B3 Metrics for Evaluation
+Two major metrics are used to evaluate the performance of the proposed algorithm:
   • (I)
@@
   • (II)
-    Numbers of corrections needed for a successful trial: The averaged number of corrections performed by the users in successful trials is a measure for the efficiency of the proposed method.
-Fewer corrections in successful trials suggest better efficiency in teaching the robot.
-VI-E Result and Analysis
-VI-E1 User Study Metrics Report
-TABLE III: Result of Simulated User Study
+    Numbers of corrections needed for a successful trial: The averaged number of corrections performed by users in successful trials is a measure for the efficiency of the proposed method.
+Fewer corrections in successful trials suggest better human data efficiency.
+VII-C Result and Analysis
+VII-C1 User Study Metrics Report
+TABLE II: Result of user study for simulation games
 Environment Success Rate Correction count
@@
 Franka arm reaching $89.3\%\pm 11.2\%$ $16.5\pm 6.1$
-For the user study, we report the success rate and correction count in a successful game trial in Table III.
-Both metrics are averaged across individual users.
-The success rate is $89.3\%\pm 7.7\%$ for the drone navigation game and $89.3\%\pm 11.2\%$ for the Franka arm constrained reaching game.
-The number of corrections for a successful trial is $15.7\pm 10.6$ for the drone navigation game and $16.5\pm 6.1$ Franka arm reaching games.
-The above result illustrates that the proposed method is effective (i.e., with a high success rate) in enabling a robot to learn a safety constraint for varying tasks (i.e.
-reaching targets are changing).
-It also shows the efficiency of the proposed method — successful learning of a safety constraint only requires tens of human corrections.
-Since this method is online and the user can perform multiple corrections in one robot rollout trajectory, the constraint can be learned to guide the robot to complete safety-critical tasks in a few runs.
-In Table III, one can note that the drone navigation task has a higher variance in the correction count than the Franka arm reaching task.
-This could be attributed to the larger motion range and longer task rollout horizon of drone navigation.
-As a result, in the drone game, the human correction strategies, including the correction position and correction timing, can vary significantly.
-Such diverse correction strategies can lead to a larger variance of total correction count.
-VI-E2 Analysis for Learned Behavior and Learned Constraint
-To demonstrate the learning progress, in Fig.
+For the user study, we report the success rate and correction count in a successful game trial in Table II.
+Both metrics are averaged across users.
+The success rate is $89.3\%\pm 7.7\%$ for the drone navigation game and $89.3\%\pm 11.2\%$ for the Franka arm reaching game.
+The number of corrections for a successful trial is $15.7\pm 10.6$ for the drone navigation game and $16.5\pm 6.1$ Franka arm reaching game.
+In both games, each learning update of $\boldsymbol{\theta}$ takes approximately 0.15 seconds, during which the simulation briefly pauses.
+This brief pause is imperceptible to the user, especially since human corrections occur infrequently.
+The above result illustrates that the proposed method is effective (i.e., with a high success rate) in learning a safety constraint for varying tasks (recall targets are changing across trials).
+It also shows the efficiency of the proposed method — successful learning only requires tens of human corrections.
+Since this method is online and the user can perform multiple corrections in one robot rollout trajectory, the constraint can be learned in a few runs.
+In Table II, one can note that the drone navigation task has a higher variance in the correction count than the Franka arm reaching task.
+This could be because the drone game has a larger motion range and longer task rollout horizon.
+As a result, human correction strategies, such as correction position and timing, can vary significantly.
+This correction diversity leads to a larger variance of correction count.
+VII-C2 Analysis for Learned Constraint and Robot Motion
+To illustrate learning progress, in Fig.
 17 and Fig.
-14, we illustratively show the intermediate robot motion at different numbers of human corrections.
+18, we show the intermediate robot motion at different numbers of human corrections.
 Fig.
-13 shows the flight trajectory of the drone after different numbers of human corrections.
+17 shows the trajectory of the drone after different numbers of human corrections.
 Fig.
-14 shows the position and altitude of the robot arm’s end effector when passing the slot after the different number of human corrections.
-From the illustration examples, we can conclude that the learned constraint generally guides the robot to complete the task.
-Figure 14: Illustrations of the robot arm pose when passing the $yz$-plane of the bars after $k=4,8,17$ corrections.
-(a): The motion after 4 corrections with three different targets (cyan ball).
-(b): The motion after 8 corrections.
-(c): The motion after 17 corrections.
-The ends of the gripper are denoted with red balls.
-At the beginning, the gripper cannot pass the slot.
-But after several corrections, it succeeds in passing the slot for all target positions.
-To better understand the learned safety constraint function ${g}_{\boldsymbol{\theta}}(\boldsymbol{\xi})$, in Fig.
-15, we draw the heat map of the value of ${g}_{\boldsymbol{\theta}_{H}}(\boldsymbol{\xi})$ with $\boldsymbol{\theta}_{H}$ the averaged value of weights learned in all successful trials.
-In drawing such heat map, since ${g}_{\boldsymbol{\theta}_{H}}$ is defined by RBFs spreading only on $yz$-plane, we fix $r_{x,t}=10$ (recall $x_{\text{gate}}=9.8$) and $r_{y,t},r_{z,t}$ are generated by the grid in $[0,10]\times[0,10]$ for all states $\boldsymbol{x}_{t}$ in $\boldsymbol{\xi}$.
-In the heat map, we draw the 00-level contour of ${g}_{\boldsymbol{\theta}_{H}}$ in the dashed line, which is the learned safety boundary.
+18 shows the pose of the robot gripper when passing the slot after a different number of human corrections.
+(a) After 4 Corrections
+(b) After 17 Corrections
+Figure 18: Screenshots of the robot gripper when passing through the slot after $k=4,17$ corrections, given three different targets (cyan ball).
+(a): After 4 corrections, the gripper collides with the slot.
+(b): After 17 corrections, the gripper succeeds in passing the slot for all targets.
+To better understand the learned safety constraint ${g}_{\boldsymbol{\theta}}(\boldsymbol{\xi})$, in Fig.
+19, we draw the heat map of the learned ${g}_{\boldsymbol{\theta}^{*}}(\boldsymbol{\xi})$ for the drone navigation game, with $\boldsymbol{\theta}^{*}$ the averaged of weights learned.
+Recall in drone navigation game, ${g}_{\boldsymbol{\theta}}$ is defined on y-z plane at $x=10$ (true gate position is $x_{\text{gate}}=9.8$).
+In the heat map, we draw 00-level contour of ${g}_{\boldsymbol{\theta}^{*}}$ in dashed lines, which is the learned safety boundary.
 Also, we plot the $yz$-plane projection of the gate, which is the black box.
-In Fig.
-15, we can note the learned safety boundary (corresponding to ${g}_{\boldsymbol{\theta}_{H}}(\boldsymbol{\xi})\leq 0$) is only similar to the top edge of the gate.
+Fig.
+19 shows that the learned safety boundary (corresponding to ${g}_{\boldsymbol{\theta}^{*}}(\boldsymbol{\xi})\leq 0$) is only similar to the top edge of the gate.
 This is explained by the following reasons.
 As shown in Fig.
-11, the target positions are set higher than the gate, and thus during the learning process, the motion of the drone is always going up, as shown in Fig.
-13.
+15, the targets are set higher than the gate, and thus during the learning process, the motion of the drone is always going up, as shown in Fig.
+17.
+As a result, the majority of human corrections happen when the drone approaches the top edge of the gate, leaving the learned safety boundary accurate near the top edge of the gate and bottom edge untouched.
+It is expected that to learn a safety constraint that fully represents the gate, lower target positions should be added so that the drone will receive human corrections near the bottom edge of the gate.
+Figure 19: Heatmap for the learned constraint in the drone game.
+The black solid line denotes the actual position of the gate in $yz$-plane.
+The black dashed curve denotes the boundary of the learned safety zone.
+Since the targets are higher than the gate (see Fig.
+15), during the learning process, the motion of the drone is always going up, as shown in Fig.
+17.
 As a result, the majority of human directional corrections happen when the drone approaches the top edge of the gate, leaving the learned safety boundary accurate near the top edge of the gate and bottom boundary untouched.
-It is expected that in order to learn a safety constraint that fully represents the gate frame, lower target positions should be added so that the drone will receive human corrections near the bottom of the gate.
-Figure 15: Heatmap for the learned constraint function in the drone user study.
-The black solid line denotes the actual position of the gate in $yz$-frame.
-The black dashed curve denotes the boundary of the safety zone.
-Since the target positions are higher than the gate (see Fig.
-11), thus during the learning process, the motion of the drone is always going up, as shown in Fig.
-13.
-As a result, the majority of human directional corrections happen when the drone approaches the top edge of the gate, leaving the learned safety boundary accurate near the top edge of the gate and bottom boundary untouched.
-VII User Study in Real-world Robot Experiment
-In this section, we conduct a user study to evaluate the performance of our safe MPC alignment algorithm using a hardware Franka Research 3 robot experiment.
-We focus on the task of mobile liquid pouring, i.e., the robot performs liquid pouring while moving from one location to another.
+VIII User Study in Real-world Experiment
+In this section, we conduct a user study to evaluate the performance of our safe MPC alignment algorithm in a hardware Franka Research 3 experiment.
+We focus on the task of mobile liquid pouring, i.e., the robot performs liquid pouring while moving from one location to targets.
 In certain scenarios, this kind of task can be safety-critical: for example, if the liquid is corrosive/chemical/toxic, inappropriate mobile pouring will cause safety concerns.
 The task is challenging because it is hard to manually define a safety constraint for the robot to accomplish liquid pouring without spilling.
-The Franka robot arm will learn such a safe MPC controller from human physical corrections.
-VII-A Hardware Experiment Setup
+The Franka robot arm will learn a safe MPC controller from human physical corrections.
+VIII-A Hardware Setup
 […]
-We specifically focus on mobile water pouring by the Franka robot arm.
-A task setup is shown in Fig.
-16.
+The setup of mobile water pouring task is shown in Fig.
+20.
 A water pot is gripped by the robot end-effector.
-The robot needs to move from the starting point $[0.22,-0.5,0.48]^{\mkern-1.5mu\mathsf{T}}$ (with orientation quaternion $[0,-0.707,0,0.707]^{\mkern-1.5mu\mathsf{T}}$) to target point $[-0.65,-0.5,0.5]^{\mkern-1.5mu\mathsf{T}}$ (with orientation quaternion $[0,-0.707,0,0.707]^{\mkern-1.5mu\mathsf{T}}$).
+The robot moves from the starting point $[0.22,-0.5,0.48]^{\mkern-1.5mu\mathsf{T}}$ (with quaternion $[0,-0.707,0,0.707]^{\mkern-1.5mu\mathsf{T}}$) to target point $[-0.65,-0.5,0.5]^{\mkern-1.5mu\mathsf{T}}$ (with quaternion $[0,-0.707,0,0.707]^{\mkern-1.5mu\mathsf{T}}$).
 During the motion, the robot needs to pour more than 50 grams of water into a bar sink.
-The bar sink is approximately placed in the middle of the robot’s moving path, but the precise size and location of the bar sink are unknown.
-Figure 17: Components of the hardware Franka arm learning platform
+The bar sink is placed in the middle of the robot path, but the precise size and location of the bar sink are unknown.
+Figure 21: System diagram for the hardware implementation. Different blocks indicate separate processes connected via arrows that indicate communication via ROS.
 As shown Fig.
-17, the hardware Franka arm platform has three main components: the parameterized safe MPC controller (high-level controller), the low-level controller, and a learning module.
-The high-level safe MPC controller follows the safe MPC formulation of Franka arm reaching game in Section VI-C (recall that the unknown constraint is parameterized by RBFs).
-The other control parameters of this Safe MPC controller are listed in Table IV.
-The high-level safe MPC controller receives the end-effector’s current pose $\boldsymbol{x}$ from the low-level controller and solves for the current desired velocity $\boldsymbol{u}$ (defined in (36)) of the end-effector.
-The low-level controller implements the operation space velocity control to track the desired velocity $\boldsymbol{u}$ given by the high-level controller.
-The learning module updates the parameterize $\boldsymbol{\theta}$ of the high-level safe MPC controller.
-The update only happens when the learning module detects a human correction $\boldsymbol{a}$, which is always accompanied by an end-effector pose $\boldsymbol{x}_{\boldsymbol{a}}$ where such human correction $\boldsymbol{a}$ is applied to.
-The learning module implements the proposed safe MPC alignment algorithm.
-TABLE IV: parameter list in real-world user study
+21, the hardware Franka arm platform has three main components: the parameterized safe MPC controller (high-level controller), the low-level controller, and safe MPC alignment module.
+The safe MPC controller follows the safe MPC formulation of Franka arm reaching game in Appendix B-D (recall that the learnable safety constraint is parameterized by RBFs).
+The parameters of this Safe MPC controller are listed in Table III.
+The safe MPC controller receives the end-effector’s current pose $\boldsymbol{x}$ and generate the action $\boldsymbol{u}$ (see (58) in Appendix), which is the desired velocity of the end-effector.
+The low-level controller implements the operational space control to track the $\boldsymbol{u}$ given by the safe MPC controller.
+The safe MPC alignment module updates the parameterize $\boldsymbol{\theta}$ of the safe MPC controller.
+The update only happens when the alignment module detects a human correction $\boldsymbol{a}$.
+The module implements our safe MPC alignment algorithm.
+Note that the alignment module and controller modules runs different process, thus the learning update does not affect the real-time MPC control of the robot arm.
+ROS is used for communication between different processes.
+TABLE III: parameter list in real-world user study
 Parameter Value
@@
 $\mu_{q}$ 80
-In the high-level MPC policy, the cost of the MPC controller is set to penalize the current and target gripper position in Cartesian space, as in Section VI-C.
-Since it is difficult to explicitly define a parameterized constraint using the state of pouring water, we instead define the constraints of safe water pouring using the robot state.
-Specifically, we define the same RBF (40) parameterized constraint in Section VI-C, constraining the tilt angle and height of the robot gripper for pouring the water into the sink.
-Figure 18: Illustrations of human correction in the real-world user study.
-(a) A human user applies force to the robot.
-(b) The human moves the robot to the desired configuration to pour the water into the bar sink.
-The delta position of the end effector is used as human directional correction, as in (42).
+In the safe MPC controller, the cost of the MPC controller is set to penalize the current and target, as in Section B-D.
+Since it is difficult to explicitly define a parameterized safety constraint using the state of pouring water, we instead define the constraints only using the robot state.
+Specifically, we use the weighted RBFs to parameterize the safety constraints constraint, similar to Appendix B-D.
 As shown in Fig.
-18, to give a physical correction to the Franka robot arm, a human user can physically touch/hold any link of the robot arm, and the applied force will immediately activate the zero-torque control mode.
+20, to give a physical correction to the Franka robot arm, a human user need physically touch/hold any link of the robot arm, and the applied force will immediately activate the zero-torque control mode.
 Suppose the joint positions of the robot arm before and after human physical correction are $\mathbf{q}_{\text{before}}$ and ${\mathbf{q}_{\text{after}}}$, respectively, we calculate the corresponding raw human correction $\boldsymbol{a}_{r}$ using
 $\boldsymbol{a}_{r}=\mathrm{J}(\mathbf{q}_{\text{before}})({\mathbf{q}}_{\text{after}}-\mathbf{q}_{\text{before}}),$ (35)
-with $\mathrm{J}$ the Jacobian of the end-effector.
-The above calculated $\boldsymbol{a}_{r}\in\mathbb{R}^{6}$ is considered as the directional displacement of the end-effector pose, intended by the human user.
-In the computed $\boldsymbol{a}_{r}\in\mathbb{R}^{6}$, we mask the entries in the other dimensions with 00s and only keep the entries corresponding to the tilt angle along $x$-axis and the height in $z$-axis, since they are the only dimensions relevant to the given task.
-The masked vector are is as the actual human correction $\boldsymbol{a}$ fed to the learning module.
+with $\mathrm{J}$ the Jacobian of the end-effector. The calculated $\boldsymbol{a}_{r}\in\mathbb{R}^{6}$ is considered as the displacement of the end-effector pose, intended by the human user.
 (a) After 2 Corrections
@@
 (c) After 5 Corrections
-Figure 19: Illustrations of one trial of the real-world study.
+Figure 22: Illustrations of one trial of the real-world experiment.
 The trajectories of the watering pot is denoted by blue dashed lines.
@@
 (c) After 5 corrections, the arm learns to tilt only along the direction of the $x-$axis and successfully aligns with human intent.
-VII-B Participants, Procedures, and Metrics
-All participants recruited for the two simulation games in Section VI-D also take part in the hardware experiment user study.
+VIII-B Participants, Procedures, and Metrics
+All participants recruited for the previous two simulation games in Section VII-B also participate the hardware experiment.
 They are told of the objective before starting trials.
-Similar to part VI-C, one episode is defined as one robot motion from the starting gripper position to the target position.
-One trial means one run of the control with multiple episodes (with a maximum of 15 corrections).
-During corrections, the nozzle of the water pot is sealed, and we switch to the unsealed nozzle and measure the amount of water poured into the flower pot after the user is satisfied with the motion of the robot motion.
-The success trial is defined when the amount of poured water is over 50g.
-Each user performs 5 trials, and the same metrics in Section VI-D are recorded.
-VII-C Result and Analysis
-In real-world robot experiments, the success rate is $84.0\%\pm 16.7\%$ and the average number of corrections in successful trials is $6.88\pm 3.26$.
-The results show the efficacy and efficiency of the proposed method in real-world robot learning applications.
+One episode is defined as one robot motion from the starting position to the targets, during which the human can apply corrections at any time instances.
+One learning trial may include multiple episodes (with a maximum of 15 corrections).
+During learning, the nozzle of the water pot is sealed, and we switch to the unsealed nozzle and measure the amount of water poured into the flower pot after the user is satisfied with the motion of the robot motion.
+The success learning trial is defined when the amount of poured water is over 50g.
+Each user performs 5 trials, and the same metrics in Section VII-B are recorded.
+VIII-C Result and Analysis
+In the real-world robot experiment, the success rate is $84.0\%\pm 16.7\%$ and the average number of corrections needed for successful trials is $6.88\pm 3.26$.
+The results show the efficacy and efficiency of our method in real-world robot learning applications.
 To illustrate the robot learning progress, in Fig.
-19, we use one trial of the user study to show the robot behavior at different numbers of human corrections.
-Since the task of this real-world Franka arm experiment is similar to the simulated Franka arm reaching game in Section VI-C, we compare the user study results for both and have the following conclusion and analysis.
-(I) The real-world experiment has a slightly lower success rate: While the difference between simulations and real-world experiments is small in success rate, we have identified some factors during our experiment that may contribute to such lower success rate, which includes both the human factors and hardware factors.
-On human factors, most of the participants have no experience or knowledge in operating robot hardware, and thus when directly interacting with hardware, their corrections can be conservative or ineffective due to safety considerations.
-On the hardware side, some error sources, such as noisy correction signals, and the possible timing lag between the recorded correction-happening robot state and the human intended correction-happening state, can contribute to the failure of the learning.
+22, we use one trial to show the robot behavior at different numbers of human corrections.
+Since this real-world Franka arm experiment is similar to the simulated Franka arm reaching game in Appendix B-D, we compare the user study results for both and have the following conclusion and analysis.
+(I) The real-world experiment has a slightly lower success rate: While the difference between simulations and real-world experiments is small in success rate, we have identified some factors during our experiment that may contribute to such lower success rate, including both the human aspects and hardware aspects.
+On human aspects, most of the participants have no experience or knowledge in operating robot hardware, and thus when directly interacting with hardware, their corrections can be conservative or ineffective due to safety considerations.
+On hardware side, noisy correction signals and the possible timing delay between the recorded human correction and robot state can contribute to the failure of the learning.
 (II) The real-world experiment has a less number of human corrections: This could be attributed to two reasons.
-First, the constraint needed for accomplishing the water pouring task in the real-world experiment could be more relaxed than the constraint for accomplishing the narrow slot passing task in the simulation game in Section VI-C.
+First, the constraint needed for accomplishing the water pouring task could be more flexible than that for reaching task in the simulation game Appendix B-D.
 This makes the safety constraints easy to learn for the real-world task.
-Second, different from simulation games, where human users can only perceive the motion of the robot arm from the screen, a human user has a better perception of the robot’s motion in a real-world experiment.
-A good perception of the robot’s motion could help humans give more effective feedback each time, reducing the total number of human interactions for successful robot learning.
-VII-D Discussion
-Based on the observations in both the hardware and simulation user studies, we discuss some design choices and potential limitations of the proposed safe MPC alignment algorithm.
-VII-D1 Learning Task-Independent Safety Constraints
-In the real-world and simulation experiments, due to the insufficiency of task sampling, the learned safety constraint can be task-dependent, instead of task-independent.
-For instance, in the drone navigation game in Section VI-B, since the three task targets are placed above the gate, the learned constraint shown in Fig.
-15 can only reflect the upper edge of the gate constraints.
-One limitation of this is that the learned task constraints have poor generalization to new tasks, which are far from previously encountered.
+Second, different from simulation games, where human users can only perceive the robot arm from the screen, a human user has a better perception of the robot’s motion in real world.
+A good perception of the robot’s state could help humans give more effective feedback each time, reducing the total number of human interactions for successful robot learning.
+VIII-D Discussion
+Based on the observations in both the hardware and simulation user studies, we discuss some design choices and potential limitations of the proposed method.
+VIII-D1 Learning Task-Independent Safety Constraints
+In the real-world and simulation experiments, due to the insufficiency of task sampling, the learned safety constraint could be task-specific, instead of task-independent.
+For instance, in the drone navigation game, since the three task targets are placed above the gate, the learned constraint shown in Fig.
+19 can only reflect the upper edge of the unknown gate.
+One limitation of this is that the learned task constraints have poor generalization to new tasks.
 For example, in drone navigation, if a new target is placed below the gate, the drone with the learned safety constraint Fig.
@@
 To improve the generalization of the learned safety constraints or learning task-independent safety constraints, one solution is to increase the diversity of task samples.
-By doing so, it is expected that the robot can receive human correction on the entire boundary of the safety region, making the learned safety constraints task irrelevant.
-VII-D2 Robustness Against Human Irrationality
-Our proposed safe MPC alignment method is based on the cutting hypothesis.
-In our formulation, we define a hypothesis cut (7) based on an average of multiple human corrections, such an averaging treatment is to mitigate human mistakes for one cut.
-But still, in practice, human irrationality can lead to the wrong cut of the hypothesis space, which constitutes the main reason for failure in user studies.
-For example, in the drone navigation game, we found that failure game trials usually happen when a human gives too many wrong directional corrections at the beginning of the trial.
-A failure case is shown in Fig.
-20.
-Figure 20: One illustration of the failure case in the drone user study. A human user makes a false "up" correction at the beginning and fails to correct the motion of the drone.
-To further improve the robustness of the method against human mistakes, two extensions of the current method can be used on top of our proposed algorithm.
+By doing so, it is expected that the robot can receive human correction on the entire boundary of the safety region, making the learned safety constraints reusable in new tasks.
+VIII-D2 Robustness Against Human Mistakes
+The proposed safe MPC alignment is based on the hypothesis cutting.
+In our formulation, a hypothesis cut (7) is based on an average of multiple human corrections, such an averaging treatment can mitigate human mistakes for one cut.
+But still, in practice, human irrationality can lead to the wrong cut of the hypothesis space, which is the main reason for failure in user studies.
+For example, in the drone navigation game, we found that failure game trials usually happen when a human gives too many wrong directional corrections.
+A failure case is in Fig.
+23.
+Figure 23: One illustration of the failure case in the drone game user study. A human user makes a false "up" correction at the beginning and fails to correct the motion of the drone.
+To further improve the robustness of the method against human mistakes, two extensions of the current method can be explored.
 The first is to add a relaxation factor $\epsilon\geq 0$ in our cutting hypothesis.
@@
 $\forall\boldsymbol{\theta}_{H}\in\bar{\Theta}_{H},\ \operatorname{\mathbb{E}}_{\mathbf{a}}\left\langle\mathbf{a},-\nabla B(\boldsymbol{\xi}_{\boldsymbol{\theta}}^{\gamma},\boldsymbol{\theta}_{H})\right\rangle+\epsilon\geq 0.$ (36)
-Another extension is to add a selection mechanism on human correction.
-For example, one can use some statistical method to distinguish and rule out mistake corrections before performing a hypothesis cut.
+Another is to add a selection mechanism on human correction.
+For example, one can use some statistical method to rule out mistake corrections before performing a hypothesis cut.
 We will leave those possible extensions in our future work.
-VIII Conclusion
-This paper proposes safe MPC alignment method to online learn a safety constraint in robot safe MPC controllers from human directional feedback.
+IX Conclusion
+This paper proposes safe MPC alignment method to online learn a safety constraint in safe MPC controllers from human directional feedback.
 The method, named hypothesis space cutting, has geometric interpretation: it exponentially cuts the hypothesis space based on each received human directional feedback.
-The method has a provable guarantee for the upper bound of the total number of human feedback, and also can declare hypothesis misspecification.
-A numeric example, two user studies on simulation games, and a user study on the real-world robot arm, covering varying robot tasks, were used to validate the effectiveness and efficiency of the proposed approach.
+The method has a provable guarantee for the total number of human feedback needed for learning, and also can declare hypothesis misspecification.
+Numeric examples, user studies on simulation games, and a user study on the real-world robot arm, covering varying robot tasks, were used to validate the effectiveness of the proposed approach.
 The results show that the proposed method can learn safety constraints effectively and efficiently.
@@
 Proof.
-For any $i=1,2,3,...$, the plan $\boldsymbol{\xi}_{\boldsymbol{\theta}_{i}}^{\gamma}=\{\boldsymbol{u}^{\boldsymbol{\theta}_{i},\gamma}_{0:T-1},\boldsymbol{x}^{\boldsymbol{\theta}_{i},\gamma}_{0:T}\}$ is a solution to the optimization (3), with $\boldsymbol{\theta}=\boldsymbol{\theta}_{i}$.
+(14) can be directly obtained from (12), (2) and (5).
+Based on the definition of the penalty objective function (2), $\mathbf{\bar{a}}_{i},\nabla B(\boldsymbol{\xi}_{\boldsymbol{\theta}_{i}}^{\gamma},\boldsymbol{\theta})\right>\leq 0$ is a linear fractional inequality
+$\mathbf{\bar{a}}_{i},\nabla J(\boldsymbol{\xi}_{\boldsymbol{\theta}_{i}}^{\gamma})\right>+\gamma\frac{1}{-g_{\boldsymbol{\theta}}(\boldsymbol{\xi}_{\boldsymbol{\theta}_{i}}^{\gamma})}\left<\mathbf{\bar{a}}_{i},\nabla g_{\boldsymbol{\theta}}(\boldsymbol{\xi}_{\boldsymbol{\theta}_{i}}^{\gamma})\right>\leq 0,$ (37)
+with a positive denominator $-g_{\boldsymbol{\theta}}(\boldsymbol{\xi}_{\boldsymbol{\theta}_{i}}^{\gamma})>0$ due to (12).
+Multiplying the denominator on both side and substituting the weight-feature expression of $g_{\boldsymbol{\theta}}$ in (5) yields $\boldsymbol{\theta}^{\mkern-1.5mu\mathsf{T}}\boldsymbol{h}_{i}\leq b_{i}$ in Lemma 3.
+The condition $-g_{\boldsymbol{\theta}}(\boldsymbol{\xi}_{\boldsymbol{\theta}_{i}}^{\gamma})>0$ itself yields $\ {\boldsymbol{\theta}}^{\mkern-1.5mu\mathsf{T}}\boldsymbol{\phi}(\boldsymbol{\xi}_{\boldsymbol{\theta}_{i}}^{\gamma})<-\phi_{0}(\boldsymbol{\xi}_{\boldsymbol{\theta}_{i}}^{\gamma})$ in Lemma 3.
+This completes the proof. ∎
+A-C Proof of Lemma 4
+Proof.
+For any $i=1,2,3,...$, the plan $\boldsymbol{\xi}_{\boldsymbol{\theta}_{i}}^{\gamma}$ is a solution to the optimization (3), with $\boldsymbol{\theta}=\boldsymbol{\theta}_{i}$.
 Recall Lemma 1, $g_{\boldsymbol{\theta}_{i}}(\boldsymbol{\xi}_{\boldsymbol{\theta}_{i}}^{\gamma})<0$.
@@
 ∎
-A-C Proof of Lemma 5
+A-D Proof of Lemma 5
 Proof.
@@
 based on (12), (13), one can get
-$\frac{\mathbf{Vol}(\Theta_{i})}{\mathbf{Vol}(\Theta_{i-1})}\leq 1-\frac{1}{r}.$ (49)
+$\frac{\mathbf{Vol}(\Theta_{i})}{\mathbf{Vol}(\Theta_{i-1})}\leq 1-\frac{1}{r},$ (43)
 which concludes the proof. ∎
-A-D Proof of Theorem 1
+A-E Proof of Theorem 1
 Proof.
@@
 ∎
-A-E Proof of Theorem 2
+A-F Proof of Theorem 2
 Proof.
@@
 which contradicts $\mathbf{Vol}(\Theta^{*})=0$ from (46). Thus, $\Theta^{*}\cap\mathbf{int}\ \Theta_{0}=\emptyset$ must hold. This completes the proof. ∎
+Appendix B Settings of Experiments and User Studies
+B-A Inverted Pendulum
+In the dynamics of an inverted pendulum (26), the physical parameters is set in Table IV.
+TABLE IV: Physics Parameters in Pendulum Numerical Study
+Parameter Value Interpretation
+$m$ 1kg Weight of the pendulum
+$l$ 1m Length of the pendulum
+$\mathrm{g}$ $10\mathrm{m/s}^{2}$ Gravitational Acceleration
+$d$ 0.1 Dampening Coefficient
+When the system state violates ${g}_{\boldsymbol{\theta}_{H}}(\boldsymbol{\xi}_{\boldsymbol{\theta}}^{\gamma})\leq 0$ or it is near the target, the pendulum state is uniformly randomly reset between $[0,0]^{\mkern-1.5mu\mathsf{T}}$ and $[\frac{2\pi}{3},3]^{\mkern-1.5mu\mathsf{T}}$.
+Simulated human correction will continue to be applied until the learned parameters $\boldsymbol{\theta}_{i}$ fall in $\bar{\Theta}_{H}$, which terminates one trial.
+B-B Two-link Robot Arm
+We assume each links has unit length and mass, and the mass is evenly distributed along the link. The corresponding inertia matrix and Coriolis term can then be calculated.
+B-C Drone Navigation Game
+B-C1 Environment
+The drone starts from $[0.1,y_{\text{init}},0.5]^{\mkern-1.5mu\mathsf{T}}$ with $y_{\text{init}}$ uniformly randomly in an interval $[4,6]$.
+The starting quaternion $\boldsymbol{q}_{\text{init}}=[1,0,0,0]^{\mkern-1.5mu\mathsf{T}}$.
+We set three different targets at positions $[19,1,9]^{\mkern-1.5mu\mathsf{T}}$, $[19,5,9]^{\mkern-1.5mu\mathsf{T}}$ and $[19,9,9]^{\mkern-1.5mu\mathsf{T}}$, with quaternion $[1,0,0,0]^{\mkern-1.5mu\mathsf{T}}$.
+For the narrow gate (yellow rectangle) in Fig.
+15, both its size and position are unknown, and one only roughly knows its x position $x_{\text{gate}}\approx 9.8$.
+Since the location of the gate is unknown, the drone needs to learn a safety constraint in its safe MPC policy with humans guidance in order to successfully fly through the gate.
+In the drone MPC policy (1), the dynamics and its parameters are set following ([20]); and we generically set the constraint function ${g}_{\boldsymbol{\theta}}(\boldsymbol{\xi})$ using a set of $N$ Gaussian radial basis functions (RBFs)
+${g}_{\boldsymbol{\theta}}(\boldsymbol{\xi})=\phi_{0}+\sum\nolimits_{i=1}^{N}\theta_{i}\phi_{i}(\boldsymbol{\xi}).$ (54)
+As discussed in Section III-C4, we fix $\phi_{0}=-1$ to avoid learning ambiguity, and use $N=20$ temporally accumulated RBFs defined as
+$\phi_{i}(\boldsymbol{\xi})=\sum\nolimits_{t=0}^{T}\beta_{t}\mathrm{RBF}_{i}(\boldsymbol{x}_{t}),$ (55)
+where $\beta_{t}$ is a weighting factor to balance the immediate and future constraint satisfaction.
+If one emphasizes instant safety, more weights will be given to the initial time step.
+In our implementation, we choose an exponentially decaying $\beta_{t}$ starting at $t=5$, i.e.
+$\beta_{t}=0,\ t<5$ and $\beta_{t}=0.9^{t-5},\ t\geq 5$ to achieve better predictive safety performance.
+In (55), each RBF is defined as
+$\mathrm{RBF}_{i}(\boldsymbol{x}_{t})=\begin{cases}\exp(-\epsilon^{2}\left\lVert r_{xy,t}-c_{xy,i}\right\rVert^{2}),\,\,i=1,...,10\\ \exp(-\epsilon^{2}\left\lVert r_{xz,t}-c_{xz,i}\right\rVert^{2}),\,\,i=11,...,20\end{cases}.$ (56)
+Here, the variables $r_{xy,t}=[r_{x,t},r_{y,t}]^{\mkern-1.5mu\mathsf{T}}$ and $r_{xz,t}=[r_{x,t},r_{z,t}]^{\mkern-1.5mu\mathsf{T}}$ with $r_{x},r_{y},r_{z}$ from $\boldsymbol{x}_{t}$.
+The $i$th center $c_{xy,i}=[c_{x,i},c_{y,i}]^{\mkern-1.5mu\mathsf{T}}$ or $c_{xz,i}=[c_{x,i},c_{z,i}]^{\mkern-1.5mu\mathsf{T}}$ is defined with $c_{x,i}={x_{\text{gate}}}=9.8$, $c_{y,i}$ and $c_{z,i}$ are evenly chosen on interval $[0,10]$.
+We set $\epsilon=0.45$.
+The initial hypothesis space $\Theta_{0}$ for possible $\boldsymbol{\theta}$ in (54) is set as $\boldsymbol{c}_{l}=[-80,\dots,-80]^{\mkern-1.5mu\mathsf{T}}$ and $\boldsymbol{c}_{u}=[100,\dots,100]^{\mkern-1.5mu\mathsf{T}}$.
+$\gamma$ is 50.
+This initialization blocks the drone from advancing to the gate at the beginning of the game, leading to conservative behavior in the learning process.
+The drone will proceed toward the gate after few corrections.
+B-C2 Human Correction Interface
+The motion of the drone is rendered in two camera views in MuJoCo visualization (Fig.
+15) for better human perception: one is the first-person view tracking the drone, the other a third-person fixed view.
+The gate, the current MPC planning solution, and the target position are visible to the users.
+Human users apply corrections to the drone via a keyboard during the drone flying.
+The user is informed of the mappings between the key press and the corresponding correction in the game, listed in Table V.
+The safety constraint (54) in the drone’s MPC policy is immediately updated once a user correction is detected.
+When the user presses "ENTER", the drone is reset to the initial position and target will randmize.
+TABLE V: keyboard commands in uav user study
+Key Correction Interpretation
+"up" $[-1,0,1,0]^{\mkern-1.5mu\mathsf{T}}$ Go forward (adding positive y-axis torque)
+"down" $[1,0,-1,0]^{\mkern-1.5mu\mathsf{T}}$ Go backward (adding negative y-axis torque)
+"left" $[0,1,0,-1]^{\mkern-1.5mu\mathsf{T}}$ Go left (adding negative x-axis torque)
+"right" $[0,-1,0,1]^{\mkern-1.5mu\mathsf{T}}$ Go right (adding positive x-axis torque)
+"shift" $[1,1,1,1]^{\mkern-1.5mu\mathsf{T}}$ Go up (increasing all thrusts)
+"ctrl" $-[1,1,1,1]^{\mkern-1.5mu\mathsf{T}}$ Go down (decreasing all thrusts)
+"enter" N/A Reset robot position and switch target.
+B-D Franka Constrained Reaching Game
+B-D1 Environment
+The arm starts at a configuration in which the end-effector position $\boldsymbol{r}_{\text{init}}=[0.22,-0.5,0.48]^{\mkern-1.5mu\mathsf{T}}$ and altitude $\boldsymbol{q}_{\text{init}}=[0,-0.707,0,0.707]^{\mkern-1.5mu\mathsf{T}}$.
+The reaching target position $\boldsymbol{r}^{*}$ will be randomly chosen from three options $[-0.6,-0.5,0.4]^{\mkern-1.5mu\mathsf{T}}$, $[-0.6,-0.5,0.5]^{\mkern-1.5mu\mathsf{T}}$, $[-0.6,-0.5,0.6]^{\mkern-1.5mu\mathsf{T}}$.
+The reaching target attitude is $\boldsymbol{q}^{*}=[0,-0.707,0,0.707]^{\mkern-1.5mu\mathsf{T}}$.
+The slot size and position are unknown except its rough x-position which is $x_{\text{slot}}\approx-0.15$.
+We adopt a two-level control scheme, where a safe MPC policy (high-level) is defined for the end effector, which generates the desired velocity for a given low-level operational space controller to track.
+In the safe MPC policy, we model the end-effector dynamics as
+$\begin{gathered}\dot{\boldsymbol{r}}=\boldsymbol{v}^{d},\\ \dot{\boldsymbol{q}}=\frac{1}{2}\Omega(R^{\mkern-1.5mu\mathsf{T}}(\boldsymbol{q})\boldsymbol{w}^{d})\boldsymbol{q},\end{gathered}$ (57)
+where $\boldsymbol{r}$ and $\boldsymbol{q}$ is the position and the quaternion of the arm end-effector in the world frame.
+$R^{\mkern-1.5mu\mathsf{T}}(\boldsymbol{q})$ is the rotation matrix from the end-effector frame to the body frame.
+$\Omega(\cdot)$ is the matrix operator of the angular velocity for quaternion multiplication.
+Define the robot state and control input
+$\boldsymbol{x}_{t}=[\boldsymbol{r}_{t},\ \boldsymbol{q}_{t}]\in\mathbb{R}^{7},\quad\boldsymbol{u}_{t}=[\boldsymbol{v}^{d}_{t},\ \boldsymbol{w}^{d}_{t}]\in\mathbb{R}^{6},$ (58)
+respectively.
+The above end-effector dynamics is discretized with $\triangle t=0.1\mathrm{s}$.
+In (1), the MPC horizon is $T=20$.
+Let $\boldsymbol{x}^{*}=[\boldsymbol{r}^{*}\ \boldsymbol{q}^{*}]$ denote the target end-effector pose.
+The step cost $c(\boldsymbol{x}_{t},\boldsymbol{u}_{t})$ and terminal cost $h(\boldsymbol{x}_{T})$ in (1) is set to
+$c(\boldsymbol{x}_{t},\boldsymbol{u}_{t})=$ $(\boldsymbol{r}_{t}-\boldsymbol{r}^{*})^{\mkern-1.5mu\mathsf{T}}Q_{r}(\boldsymbol{r}_{t}-\boldsymbol{r}^{*})+k_{q}\mathbf{diff}(\boldsymbol{q}_{t},\boldsymbol{q}^{*})$ (59)
+$\quad\quad+\boldsymbol{v}_{I,t}^{\mkern-1.5mu\mathsf{T}}R_{v}\boldsymbol{v}_{I,t}+\boldsymbol{w}_{I,t}^{\mkern-1.5mu\mathsf{T}}R_{w}\boldsymbol{w}_{I,t},$
+$h(\boldsymbol{x}_{T})=$ $(\boldsymbol{r}_{T}-\boldsymbol{r}^{*})^{\mkern-1.5mu\mathsf{T}}S_{r}(\boldsymbol{r}_{T}-\boldsymbol{r}^{*})+\mu_{q}\mathbf{diff}(\boldsymbol{q}_{T},\boldsymbol{q}^{*}).$
+In this game, the cost parameters are set as $Q=\mathrm{diag}(2,2,2)$, $k_{q}=1$, $R_{v}=\mathrm{diag}(30,30,1)$, $R_{w}=\mathrm{diag}(0.85,0.85,0.85)$, $S_{r}=\mathrm{diag}(80,80,80)$, and $\mu_{q}=60$.
+The output of the MPC controller is sent to the low-level controller to track.
+Similar to Appendix B-C, to constrain $r_{z}$ and $\boldsymbol{q}$ for proper height and tilt in order to pass through the slot.
+We use a set of $N=20$ Gaussian radial basis function (RBFs) to parameter the safety constraint:
+${g}_{\boldsymbol{\theta}}(\boldsymbol{\xi})=\phi_{0}+\sum\nolimits_{i=1}^{N}\theta_{i}\phi_{i}(\boldsymbol{\xi}),$ (60)
+with
+$\phi_{i}(\boldsymbol{\xi})=\sum\nolimits_{t=0}^{T}\beta_{t}\mathrm{RBF}_{i}(\boldsymbol{x}_{t}),\quad\text{and}$ (61)
+$\mathrm{RBF}_{i}(\boldsymbol{x}_{t}){=}\begin{cases}\sigma(r_{x,t})k_{z}\exp{(-\epsilon_{z}^{2}(r_{z,t}-c_{z,i})^{2})},\,\,i{=}1,...,10\\ \sigma(r_{x,t})\exp{(-\epsilon_{q}^{2}\mathbf{diff}(\boldsymbol{q}_{t},\boldsymbol{q}_{c,i}))},\,\,i{=}11,...,20\end{cases}.$ (62)
+Here, $k_{z}$ is used to balance the scaling between position and orientation $\mathrm{RBF}$s, $\sigma(r_{x,t})$ is defined as
+$\sigma(r_{x,t})=\mathrm{sigmoid}(50\big{(}(r_{x,t}-x_{\text{slot}})^{2}-(\frac{w}{2})^{2}\big{)}),$ (63)
+which is used to constrain the $\mathrm{RBF}_{i}$ only taking effect around $x$ range $[x_{\text{slot}}-\frac{w}{2},x_{\text{slot}}+\frac{w}{2}]$ (recall slot’s $x$ location, slot, is a approximate value).
+In the above definition, $w=0.3$, $k_{z}=0.05$, $\epsilon_{z}=12$ and $\epsilon_{q}=1.8$.
+The center of positional $\mathrm{RBF}$, ${c_{z,i}}$, is evenly chosen on interval $[0,1]$, and the center of orientation $\mathrm{RBF}$, ${\boldsymbol{q}_{c,i}}$, is generated by $(\cos{\frac{\alpha}{2}},\sin{\frac{\alpha}{2}},0,0)*\boldsymbol{q}_{\text{init}}$ with $\alpha$ evenly choosen from $[-\pi/4,3\pi/4]$ to make sure ${\boldsymbol{q}_{c,i}}$ covers all possible orientations.
+$\beta_{t}=0.9^{t}$ if $t<5$ and 00 otherwise.
+The bounds of the initial hypothesis space $\Theta_{0}$ is set to $\boldsymbol{c}_{l}=[-3,\dots,-3]^{\mkern-1.5mu\mathsf{T}}$ and $\boldsymbol{c}_{u}=[5,\dots,5]^{\mkern-1.5mu\mathsf{T}}$.
+$\gamma$ is set to 1.5.
+This initialization leads to a conservative MPC policy, preventing the arm from approaching the target at the beginning.
+B-D2 Human Correction Interface
+The motion of the robot arm is rendered in the free-view mode in MuJoCo visualization (Fig.
+16), and users can drag to select his/her comfortable view.
+The slot and the random reaching target are visible to the users.
+Human users apply their corrections to the robot arm via a keyboard.
+Users are informed with the following key mapping between the key press and the corresponding correction in the game.
+The safety constraint (60) in Franka arm’s safe MPC policy is updated immediately once a user correction is detected.
+When user press "ENTER", the arm is reset to the initial position and the reaching target will also change.
+TABLE VI: keyboard commands in robot arm user study
+Key Correction Interpretation
+"left" $[0,0,0,0,0,1]^{\mkern-1.5mu\mathsf{T}}$ Clock-wise rotation of the gripper
+"right" $[0,0,0,0,0,-1]^{\mkern-1.5mu\mathsf{T}}$ Anti-clock-wise rotation of the gripper
+"shift" $[0,0,0,-1,0,0]^{\mkern-1.5mu\mathsf{T}}$ Go up in world frame
+"ctrl" $[0,0,0,1,0,0]^{\mkern-1.5mu\mathsf{T}}$ Go down in world frame
+"Enter" N/A Reset robot position and switch target.

### v3

@@
 In safety-critical robot planning or control, manually specifying safety constraints or learning them from demonstrations can be challenging.
-In this article, we propose a certifiable alignment method for a robot to learn a safety constraint in its model predictive control (MPC) policy with human online directional feedback.
+In this article, we propose a certifiable alignment method for a robot to learn a safety constraint in its model predictive control (MPC) policy from human online directional feedback.
 To our knowledge, it is the first method to learn safety constraints from human feedback.
@@
 The method only requires the direction of human feedback to update the learning hypothesis space.
-It is certifiable, providing an upper bound on the total number of human feedback in the case of successful learning, or declaring the hypothesis misspecification, i.e., the true implicit safety constraint cannot be found within the specified hypothesis space.
-We evaluated the proposed method using numerical examples and user studies in two simulation games.
-Additionally, we implemented and tested the proposed method on a real-world Franka robot arm performing mobile water-pouring tasks.
+It is certifiable, providing an upper bound on the total number of human feedback in the case of successful learning, or declaring the hypothesis misspecification, i.e., the true safety constraint cannot be found within the specified hypothesis space.
+We evaluated the proposed method in numerical examples and user studies with two simulation games.
+Additionally, we tested the proposed method on a real-world Franka robot arm performing mobile water-pouring tasks.
 The results demonstrate the efficacy and efficiency of our method, showing that it enables a robot to successfully learn safety constraints with a small handful (tens) of human directional corrections.
-Index Terms:
-Learning from Human Feedback, Constraint Inference, Model Predictive Control, Human-Robot Interaction, Cutting-Plane Method.
 I Introduction
-In robot policy design, safety is always the top priority, especially in safety-critical applications, such as autonomous vehicles or robots operating near humans, where safety violation can lead to irreversible loss or damage.
-While extensive work in safe control [1] and reinforcement learning [2] focuses on designing safe policy given robot motion constraints, less attention has been paid to how to specify the safety constraints in the first place.
+In robot policy design, safety is always the top priority, especially in safety-critical applications, such as robots operating near humans, where safety violations can lead to irreversible losses.
+While extensive work in safe control [wabersich2023data] and reinforcement learning [gu2022review] focuses on deducing robot policy given safety constraints, less attention has been paid to how to specify the safety constraints in the first place.
 In many applications, specifying proper safety constraints can be challenging.
-First, explicit safety boundary for robot behavior can be hard to model.
-For example, when delivering food in a crowded restaurant, a service robot must balance the tray to avoid spills; this requires constraints on robot speed, food weight distributions, and interaction with its surroundings, and all those constraint aspects can be complex to model.
-Second, safety perception can vary among users.
-In the same example of the delivery robot: some diner may feel secure with the robot staying within a foot or two when delivering items, while others may feel uncomfortable with close proximity and prefer the robot to stay at a slight distance.
-Figure 1: Overall illustration of the proposed safe MPC alignment.
-Our method online updates the safety constraint in robot safe MPC policy based on human directional feedback, such that the robot eventually aligns the safe MPC policy with human intent for accomplishing safety-critical tasks.
-The challenges in safety specifications lead to two questions, which this paper seeks to answer.
-Q1: can a robot learn and align its safety constraints with human intent in an interactive manner?
-Q2: can such learning process to be user-data efficient such that a robot can quickly adapt to different users?
-Recent work [3, 4, 5] has explored learning safety constraints from human demonstrations.
-However, acquiring safety-informative demonstrations could be challenging.
-An alternative is learning from human preference feedback [6, 7, 8, 9, 10, 11, 12], where users express preferences between trajectories to guide robot learning.
-Despite the success in reward alignment [6], the query trajectories can be difficult for humans to judge when they are equally suboptimal, especially in early learning stages.
-To address such issue, [13, 14, 15, 16, 17, 18, 19] proposed learning from human physical corrections.
-While promising, the recent work [20] suggests those methods are potentially sensitive to correction magnitude: improper human corrections can lead to overshooting and inefficiency of learning.
-Importantly, all those feedback learning methods primarily focus on learning the reward function and do not address the challenges of learning safety constraints from human feedback.
-This paper aims to answer the previous two questions by proposing a new paradigm for a robot to efficiently learn safety constraints in its model predictive control (MPC) policy with human feedback.
+First, explicit boundary for safe robot behavior can be hard to model.
+For example, when delivering food in a crowded restaurant, a service robot must balance a tray to avoid spills; this requires constraints on robot speed, food weight distributions, and interaction with its surroundings, all of which can be complex to model.
+Second, user’s perception of safety may vary.
+In the same example, some diner may feel secure with the robot staying within a foot when delivering, while others may feel uncomfortable with such proximity.
+Figure 1: Illustration of the proposed safe MPC alignment.
+Our method online updates the safety constraint using human directional feedback, such that the robot eventually aligns the safe MPC policy with human intent for the safety-critical task.
+The challenges in safety specification lead to two questions.
+Q1: can a robot learn safety constraints to align with human intent in an interactive manner?
+Q2: can such a learning process be user-data efficient?
+Recent work [chou2020learning_1, lindner2023learning, scobee2019maximum] explored learning safety constraints from human demonstrations.
+However, acquiring safety-informative demonstrations could be difficult.
+Learning from human preference [christiano2017deep, Sadigh2017ActivePL, ibarz2018reward, pmlr-v139-lee21i, hejna2023few, 10160439, pmlr-v164-myers22a] offers a new paradigm, where a user’s preference of one trajectory over the other is used for robot learning.
+The preference, however, can be difficult for a human to make, especially in early learning stages.
+[bajcsy2017learning, losey2018including, bajcsy2018learning, zhang2019learning, li2021learning, losey2022physical, mehta2023unified] proposed learning from human physical corrections.
+While promising, [jin2022learning] suggests those methods are potentially sensitive to correction magnitude: corrections can lead to overshooting and inefficiency of learning.
+All those feedback-driven learning methods primarily focus on learning rewards instead of safety constraints.
+This paper aims to answer the previous two questions by proposing a new paradigm for a robot to efficiently learn safety constraints in its model predictive control (MPC) policy from human feedback.
 We name the method safe MPC alignment.
-As shown in Fig.
-1, a human user, observing the robot’s potentially unsafe behavior, provides feedback to the robot.
-Upon receiving human feedback, the robot immediately updates the safety constraints in its MPC policy and continues with the updated policy.
-Different from existing learning from feedback methods, our method only requires human directional corrections [20], which can be sparse in time.
-For example, when a human user corrects an autonomous car, he/she only needs to give the command of ‘turn left‘ or ‘turn right‘, without dictating how much the car needs to turn.
-The safe MPC alignment only utilizes the direction of human correction, regardless of the magnitude, for the safety constraints update.
-As shown in our previous work [20] or later, the type of directional feedback provides a consistently large and flexible correction space, effectively overcoming the over-shooting correction, compared to existing magnitude feedback.
-Importantly, the proposed alignment is able to guarantee the exponential convergence with finite rounds of human corrections, or output failure with certified hypothesis misspecification.
+As in Fig.
+1, a user, observing the robot’s potentially unsafe behavior, provides online corrections to the robot.
+Upon receiving human corrections, the robot immediately updates the safety constraints in its MPC policy and continues with the updated policy.
+Our method only requires/utilizes direction of a human correction [jin2022learning].
+For example, when a human user corrects an autonomous car, he/she only needs to give the command of ‘turn left’ or ‘turn right’, without dictating how much the car needs to turn.
+Directional corrections permit a consistently large and flexible correction space.
+Importantly, the proposed method guarantees the exponential convergence with finite rounds of human corrections, or outputs failure with certified hypothesis misspecification.
 Specifically, we highlight our main contributions below.
-(I) We propose the method of safe MPC alignment with human directional feedback.
-The method enables a robot to online learn a safe MPC policy from human directional corrections.
+(I) We propose the method of safe MPC alignment from human directional correction.
+It enables a robot to learn a safe MPC policy from online human directional corrections.
 To our knowledge, it is the first work for interactive learning of safety constraints from human feedback.
-(II) We theoretically show the human-data efficiency of the safe MPC alignment: the proposed method is guaranteed to have exponential convergence with finite human corrections.
-Furthermore, we prove that the proposed method enables the certification of hypothesis space misspecification failure.
-(III) We conduct extensive user studies to validate the proposed method, including two MuJoCo-based simulation games and a real-world experiment of teaching a Franka robot arm for mobile liquid pouring.
+(II) We show that safe MPC alignment is certifiable: it guarantees exponential convergence with a finite number of human corrections and enables verification when the true safety constraint falls outside the specified hypothesis space.
+(III) Extensive user studies have been conducted to validate the proposed method, including a real-world experiment on the Franka robot arm.
 The results show the effectiveness and efficiency of our method.
@@
 The main method is presented in Section IV.
-Section V gives the implementation details and presents the main results.
-Sections VI and VII provide the experiments and user studies.
-Conclusions are in Section VIII.
+Section V gives the implementation details and the main results.
+Sections VI-VIII provide the experiments and user studies.
+Discussions and Conclusions are in Section IX.
 II Related Work
 II-A Learning from Demonstrations
-Under the general theme of learning from human data, the proposed method is loosely connected to the topic of learning from demonstrations, where cost/reward functions are learned from expert (human) data using inverse reinforcement learning (IRL) and inverse optimal control (IOC) methods [21, 22].
-IRL/IOC methods typically run offline: human demonstration data is first obtained, and then reward/cost learning is performed [22, 23, 24, 25, 26, 27].
-As pointed in [28], acquiring optimal demonstration data can be difficult in practice.
-The challenge can be either attributed to the higher degrees of freedom of the robot, or the lack of prior knowledge about the desired robot behaviour to perform demonstrations with high-quality [29].
+Under the general theme of learning from human inputs, the proposed method falls into the category of learning from demonstrations, where costs/rewards are learned from expert data using inverse reinforcement learning (IRL) [ng2000algorithms].
+IRL methods typically run offline: human demonstrations are first collected, and then reward/cost is learned [ng2000algorithms, abbeel2004apprenticeship, ratliff2006maximum, ziebart2008maximum, mombaur2010human, englert2017inverse].
+As pointed by [ravichandar2020recent], acquiring demonstration data can be difficult in practice, due to either the high degrees of freedom of a robot, or the lack of knowledge about the desired robot behavior [jain2015learning].
 In this work, we focus on interactive learning of safety constraints.
-Different from learning from demonstrations that needs the exhibition of desired robot motion, the proposed method only requires a human user to incrementally improve the current robot motion by providing directional feedback, thus more accessible to human users.
+Different from learning from demonstrations, the proposed method only requires a user to incrementally improve the current robot motion by providing directional feedback.
 II-B Learning from Human Feedback
-The line of this work typically focuses on learning rewards/costs from human feedback, which can be given in preference and corrections.
-In learning from human preferences, human preference is typically given as user’s ranking of two or multiple robot candidate trajectories [6, 7, 8, 9, 10, 11, 12], and Bradley-Terry model [30] is used to update the rewards/costs.
-The recent variant seeks to directly learn a policy from human preference [31].
-Human feedback can also come as corrections, where human users may intervene and correct the robot at runtime.
-An intended robot reward/cost function can be inferred from human corrections.
-Specific human corrections include language [32, 33, 34, 35, 36] and physical corrections [13, 15, 18].
-Since human corrections can be sparse in time, many methods [13, 14, 15, 16, 17, 18, 19] use trajectory deformation methods to interpolate sparse corrections as a continuous trajectory.
-The rewards/costs are updated to explain the deformed trajectories.
-Some strategies are further proposed to deal with the uncertainty and improve the robustness of the learning process, like updating one parameter dimension at a time [15] or formulating the reward/cost learning in the framework of Kalman Filter [14].
+This line typically focuses on learning rewards/costs from human feedback in forms of preferences and corrections.
+Human preference is typically given as user’s ranking of two or multiple robot trajectories [christiano2017deep, Sadigh2017ActivePL, ibarz2018reward, pmlr-v139-lee21i, hejna2023few, 10160439, pmlr-v164-myers22a], which is modeled using the Bradley-Terry formulation [bradley1952rank].
+Human corrections require a user to intervene and correct the robot at runtime, from which a intended reward/cost can be inferred.
+Human corrections can include language command [karamcheti2022shared, cui2023no, srivastava2023generating, zha23droc, shi2024yell] and physical corrections [bajcsy2017learning, bajcsy2018learning, losey2022physical].
+Since human corrections can be sparse in time, many methods [bajcsy2017learning, losey2018including, bajcsy2018learning, zhang2019learning, li2021learning, losey2022physical, mehta2023unified] use trajectory deformation methods to interpolate sparse corrections as a continuous trajectory.
+Some strategies are further proposed to deal with learning robustness and uncertainty, such as updating one feature at a time [bajcsy2018learning] or formulating the problem in the Kalman Filter framework [losey2018including].
 Our work belongs to the category of learning from human corrections.
 But instead of learning rewards/costs, we focus on learning safety constraints from human corrections, which has not been addressed before.
-Methodologically, we propose a certifiable method, which guarantees the exponential learning convergence with finite human corrections, and certifies the hypothesis misspecification if the learning hypothesis space does not contain the true human intent.
-Those theoretical results are missing in existing methods.
+Methodologically, we propose a certifiable method, which guarantees the learning convergence with finite human corrections, and certifies the hypothesis misspecification.
+Those results are missing in existing methods.
 II-C Learning Constraints from Demonstrations
 An increasing number of work focuses on learning safety constraints from demonstrations.
-One vanilla idea is to formulate constraint learning as a binary classification.
-However, since a robot usually runs within a safe region, negative demonstration samples that violate safety are challenging to collect.
-To address such, [3, 37] consider the trajectories of lower cost as the unsafe samples.
-[38] proposes adjusting the constraints toward decreasing the safety margin of feasible policies.
-Safety constraints can also be learned via mixed integer programming [3, 37], pessimistic adjustment [38], and no-regret algorithm [39].
-Another line of work attempts to learn safety constraints only from safe demonstration data.
-For example, some work exploits the optimality of demonstration with unknown constraints [40, 41, 42].
-Different probabilistic formulations are also proposed to learn constraints from demonstration data, such as maximum likelihood [5, 43, 44], maximum causal entropy [45, 46], and Bayesian inference [47, 48].
-In the above constraint learning, safety constraints can be represented as a set of safe states [40], the belief of potential constraints [41], Gaussian process [42], or directly as a convex hull in policy space [4].
-The vast majority of the above methods formulate the constraint learning problem in the setting of reinforcement learning, and thus the learning process usually involves repetitive policy optimization.
-Those methods cannot be deployed in an online setting, where a robot is expected to on-the-fly update its safe policy and respond to human feedback.
-To fill this gap, we will develop the first method, to our best knowledge, to enable online learning of safe constraint from human feedback.
+One straightforward idea is to cast the problem as a binary classification.
+[chou2020learning_1, chou2020learning_2] consider the trajectories of lower cost as the unsafe samples.
+[gaurav2023learning] proposes adjusting the constraints toward decreasing the safety margin of feasible policies.
+Safety constraints can also be learned via mixed integer programming [chou2020learning_1, chou2020learning_2], pessimistic adjustment [gaurav2023learning], and no-regret algorithm [kim2024learning].
+Those methods typically require collecting negative (unsafe) samples, which can be challenging in real robot systems.
+Another line of work attempts to learn safety constraints only from safe demonstration.
+For example, [chou2020learning_3, chou2021uncertainty, chou2022gaussian] infer the safety constraints by exploiting the optimality conditions of demonstrations.
+Different probabilistic formulations are also used, such as maximum likelihood [scobee2019maximum, stocking2022maximum, malik2021inverse], causal entropy [mcpherson2021maximum, baert2023maximum], and Bayesian inference [park2020inferring, papadimitriou2022bayesian].
+Among them, safety constraints can be represented as a set of safe states [chou2020learning_3], the belief of constraints [chou2021uncertainty], Gaussian process [chou2022gaussian], or a convex hull in policy space [lindner2023learning].
+Using demonstrations, the vast majority of the above methods formulate the constraint learning in offline settings, and the learning process usually involves repetitive policy optimization.
+Those methods can be difficult to be deployed online, where a robot is expected to quickly update its safety estimate and respond to human feedback.
+To fill this gap, we will develop the first method, to our best knowledge, to enable online learning of safe constraints from human feedback.
 III Preliminary and Problem Formulation
 III-A Safe Model Predictive Control
-We consider the following safe model predictive control (Safe MPC) for a robotic system, where the safety constraint is learnable and parameterized by $\boldsymbol{\theta}\in\mathbb{R}^{r}$.
-$\min_{\boldsymbol{u}_{0:T-1}}&\quad J(\boldsymbol{u}_{0:T-1};\boldsymbol{x}_{0})=\sum_{t=0}^{T-1}c(\boldsymbol{x}_{t},\boldsymbol{u}_{t})+h(\boldsymbol{x}_{T})\\ \text{subject to}&\quad\boldsymbol{x}_{t+1}=\boldsymbol{f}(\boldsymbol{x}_{t},\boldsymbol{u}_{t})\\ &\quad{g}_{\boldsymbol{\theta}}(\boldsymbol{x}_{0:T},\boldsymbol{u}_{0:T-1})\leq 0\end{aligned}.$ (1)
-Here, $\boldsymbol{x}_{t}\in\mathbb{R}^{n}$ and $\boldsymbol{u}_{t}\in\mathbb{R}^{m}$ are robot state and action, respectively; $t=0,1,2,...,T$ is the time step with prediction horizon $T$; $\boldsymbol{x}_{0:T}=(\boldsymbol{x}_{0},\boldsymbol{x}_{1},...,\boldsymbol{x}_{T})$ and $\boldsymbol{u}_{0:T-1}=(\boldsymbol{u}_{0},\boldsymbol{u}_{1},...,\boldsymbol{u}_{T-1})$; $J(\cdot)$ is the given cost function including the running cost ${c}(\cdot)$ and final cost $h(\cdot)$, representing the given task; $\boldsymbol{f}(\cdot,\cdot)$ is the system dynamics; ${g}_{\boldsymbol{\theta}}(\cdot)$ is the safety constraint, which is parameterized by $\boldsymbol{\theta}\in\mathbb{R}^{r}$.
-Given starting state $\boldsymbol{x}_{0}$ and constraint parameter $\boldsymbol{\theta}$, the trajectory solution to (1), denoted as $\boldsymbol{\xi}_{\boldsymbol{\theta}}=\{\boldsymbol{u}^{\boldsymbol{\theta}}_{0:T-1},\boldsymbol{x}^{\boldsymbol{\theta}}_{0:T}\}$, is an implicit function of $\boldsymbol{\theta}$.
-In a receding horizon implementation, $\boldsymbol{x}_{0}$ is set as the robot’s current state (estimation) $\boldsymbol{{x}}_{k}$, where $k=0,1,2,...,$ is the policy rollout time step; and only the first action $\boldsymbol{u}^{\boldsymbol{\theta}}_{0}$ in $\boldsymbol{{u}}_{0:T}^{\boldsymbol{\theta}}$ is applied to the robot to move the robot to the next state $\boldsymbol{x}_{k+1}$.
-This leads to a closed-loop Safe MPC policy for the robot, denoted as $\boldsymbol{\pi}_{\boldsymbol{\theta}}$, mapping from robot current state $\boldsymbol{x}_{k}$ to its action $\boldsymbol{u}^{\boldsymbol{\theta}}_{0}$.
+Consider a robot equipped with a safe model predictive control (Safe MPC) policy defined below, where the safety constraint is learnable and parameterized by $\boldsymbol{\theta}\in\mathbb{R}^{r}$.
+$\begin{aligned} \min_{\boldsymbol{u}_{0:T-1}}&\quad J(\boldsymbol{u}_{0:T-1};\boldsymbol{x}_{0})=\sum\nolimits_{t=0}^{T-1}c(\boldsymbol{x}_{t},\boldsymbol{u}_{t})+h(\boldsymbol{x}_{T})\\ \text{subject to}&\quad\boldsymbol{x}_{t+1}=\boldsymbol{f}(\boldsymbol{x}_{t},\boldsymbol{u}_{t})\quad\text{given}\quad\boldsymbol{x}_{0}\\ &\quad{g}_{\boldsymbol{\theta}}(\boldsymbol{x}_{0:T},\boldsymbol{u}_{0:T-1})\leq 0\end{aligned}.$ (1)
+Here, $\boldsymbol{x}_{t}\in\mathbb{R}^{n}$ and $\boldsymbol{u}_{t}\in\mathbb{R}^{m}$ are robot state and action, respectively; t=0,1,2,,Tt=0,1,2,...,T is the time step over a short prediction horizon $T$, starting from a given $\boldsymbol{x}_{0}$.
+Define $\boldsymbol{x}_{0:T}=(\boldsymbol{x}_{0},\boldsymbol{x}_{1},...,\boldsymbol{x}_{T})$ and $\boldsymbol{u}_{0:T-1}=(\boldsymbol{u}_{0},\boldsymbol{u}_{1},...,\boldsymbol{u}_{T-1})$; $J(\cdot)$ is the given cost function including the running cost ${c}(\cdot)$ and final cost $h(\cdot)$; $\boldsymbol{f}(\cdot,\cdot)$ is the system dynamics; ${g}_{\boldsymbol{\theta}}(\cdot)$ is the safety constraint parameterized by $\boldsymbol{\theta}\in\mathbb{R}^{r}$.
+Given starting state $\boldsymbol{x}_{0}$ and parameter $\boldsymbol{\theta}$, the optimal plan (solution) to (1), denoted as $\boldsymbol{\xi}_{\boldsymbol{\theta}}=\{\boldsymbol{u}^{\boldsymbol{\theta}}_{0:T-1},\boldsymbol{x}^{\boldsymbol{\theta}}_{0:T}\}$, is an implicit function of $\boldsymbol{\theta}$.
+In a receding horizon implementation, $\boldsymbol{x}_{0}$ is set as the real robot’s current state (estimation) $\boldsymbol{{x}}_{k}^{\text{real}}$, where k=0,1,2,,k=0,1,2,..., is the global rollout time step of the Safe MPC policy.
+Only the first action $\boldsymbol{u}^{\boldsymbol{\theta}}_{0}$ in each $\boldsymbol{\xi}_{\boldsymbol{\theta}}$ is applied to the robot to move the robot to the next state $\boldsymbol{x}^{\text{real}}_{k+1}$.
+This process repeats at each rollout time step $k=0,1,$....
+Such a receding horizon implementation leads to a closed-loop policy, denoted as $\boldsymbol{\pi}_{\boldsymbol{\theta}}$, mapping from robot current state $\boldsymbol{x}_{k}^{\text{real}}$ to its action $\boldsymbol{u}^{\boldsymbol{\theta}}_{0}$.
 III-B Safe Asymptotic Approximation of Safe MPC
@@
 $B(\boldsymbol{\xi},\boldsymbol{\theta})=J(\boldsymbol{\xi})-\gamma\ln(-g_{\boldsymbol{\theta}}(\boldsymbol{\xi})).$ (2)
-Here, the above defined penalty $B(\boldsymbol{\xi},\boldsymbol{\theta})$ includes the original control cost in (1) and safety constraint $g_{\boldsymbol{\theta}}(\boldsymbol{\xi})$ via logarithmic barrier transformation, $\gamma>0$ is named the barrier parameter, controlling the weight of safety penalty $-\ln(-g_{\boldsymbol{\theta}}(\boldsymbol{\xi}))$ relative the control cost.
-With $B(\boldsymbol{\xi},\boldsymbol{\theta})$, one can establish the following penalty-based MPC formulation
+Here, $B(\boldsymbol{\xi},\boldsymbol{\theta})$ includes the cost $J(\boldsymbol{\xi})$ and safety constraint $g_{\boldsymbol{\theta}}(\boldsymbol{\xi})$ in a logarithmic barrier transformation.
+$\gamma>0$ is the barrier parameter, controlling the weight of safety penalty $\ln(-g_{\boldsymbol{\theta}}(\boldsymbol{\xi}))$ relative the cost.
+With $B(\boldsymbol{\xi},\boldsymbol{\theta})$, one can establish the following penalty-based MPC
 $\min_{\boldsymbol{u}_{0:T-1}}\quad B(\boldsymbol{\xi},\boldsymbol{\theta})\qquad\text{subject to}\quad\boldsymbol{x}_{t+1}=\boldsymbol{f}(\boldsymbol{x}_{t},\boldsymbol{u}_{t}).$ (3)
-We denote the solution to the above penalty-based MPC (3) as $\boldsymbol{\xi}_{\boldsymbol{\theta}}^{\gamma}=\{\boldsymbol{u}_{0:T-1}^{\boldsymbol{\theta},\gamma},\boldsymbol{x}_{0:T-1}^{\boldsymbol{\theta},\gamma}\}$.
-The following result is from our prior work [49], stating that the penalty-based MPC (3) is an asymptotic and safe approximation to the Safe MPC (1).
+We denote the optimal plan to (3) as $\boldsymbol{\xi}_{\boldsymbol{\theta}}^{\gamma}=\{\boldsymbol{u}_{0:T-1}^{\boldsymbol{\theta},\gamma},\boldsymbol{x}_{0:T-1}^{\boldsymbol{\theta},\gamma}\}$.
+The following result from [jin2021safe] states the penalty-based MPC (3) is an asymptotic, safe approximation to Safe MPC (1).
 Lemma 1.
-(Theorem 2 in [49]) If the original Safe MPC (1) satisfies certain mild conditions (i.e., local second-order condition and strict complementarity), then for any small $\gamma>0$, the penalty-based MPC (3) has a (local) solution $\boldsymbol{\xi}_{\boldsymbol{\theta}}^{\gamma}$, and $\boldsymbol{\xi}_{\boldsymbol{\theta}}^{\gamma}$ is safe with respect to the original constraint, i.e., $g_{\boldsymbol{\theta}}(\boldsymbol{\xi}_{\boldsymbol{\theta}}^{\gamma})<0$, and $\boldsymbol{\xi}_{\boldsymbol{\theta}}^{\gamma}$ satisifies the first-order condition of (3).
+(Theorem 2 in [jin2021safe]) If the Safe MPC (1) satisfies certain mild conditions (i.e., local second-order condition and strict complementarity), then for any small $\gamma>0$, the penalty-based MPC (3) has a (local) solution $\boldsymbol{\xi}_{\boldsymbol{\theta}}^{\gamma}$, and $\boldsymbol{\xi}_{\boldsymbol{\theta}}^{\gamma}$ is safe with respect to the original constraint, i.e., $g_{\boldsymbol{\theta}}(\boldsymbol{\xi}_{\boldsymbol{\theta}}^{\gamma})<0$, and $\boldsymbol{\xi}_{\boldsymbol{\theta}}^{\gamma}$ satisifies the first-order condition of (3).
 Furthermore, $\boldsymbol{\xi}_{\boldsymbol{\theta}}^{\gamma}$ is asymptotically approximate $\xi_{\boldsymbol{\theta}}$: $\boldsymbol{\xi}_{\boldsymbol{\theta}}^{\gamma}\rightarrow\boldsymbol{\xi}_{\boldsymbol{\theta}}$ as $\gamma\rightarrow 0$.
 The above lemma asserts that solving the Safe MPC (1) can be replaced by solving penalty-based MPC (3).
-The solution approximation is controlled by $\gamma$ for arbitrary accuracy, as $\boldsymbol{\xi}_{\boldsymbol{\theta}}^{\gamma}\rightarrow\boldsymbol{\xi}_{\boldsymbol{\theta}}$ as $\gamma\rightarrow 0$.
-A good approximation is achieved by smaller $\gamma$.
-on the contrary, if $\gamma$ is larger, penalty-based MPC (3) will lead to more safety-conservative solution $\boldsymbol{\xi}_{\boldsymbol{\theta}}^{\gamma}$.
+The solution approximation is controlled by $\gamma$ for arbitrary accuracy: $\boldsymbol{\xi}_{\boldsymbol{\theta}}^{\gamma}\rightarrow\boldsymbol{\xi}_{\boldsymbol{\theta}}$ as $\gamma\rightarrow 0$.
+A good approximation is achieved by a smaller $\gamma$.
+On the contrary, if $\gamma$ is larger, penalty-based MPC will lead to more safety-conservative solution $\boldsymbol{\xi}_{\boldsymbol{\theta}}^{\gamma}$.
 In the development below, we fix the parameter $\gamma$ in (3).
@@
 III-C1 Representation of Human Intended Safety
-Consider a human-robot coexistence task, where a human is supervising the robot motion.
-The user has his/her intent of robot motion safety.
-In Safe MPC (1), we hypothesize user-intended safety can be represented by a subset of the trajectory space:
+Consider a human is supervising the robot’s behavior. In Safe MPC (1), we hypothesize that user-intended safety can be represented by a subset of the trajectory space:
 $\begin{split}\mathcal{S}(\boldsymbol{\theta}_{H})=\{(\boldsymbol{x}_{0:T},\boldsymbol{u}_{0:T-1})\,\,|\,\,\ {g}_{\boldsymbol{\theta}_{H}}(\boldsymbol{x}_{0:T},\boldsymbol{u}_{0:T-1})\leq 0\}.\end{split}$ (4)
-Here, $\boldsymbol{\theta}_{H}$ is intent parameter from an implicit set $\boldsymbol{\theta}_{H}\in\Theta_{H}\subset\mathbb{R}^{r}$.
-$\Theta_{H}$ is implicit to human and unknown to the robot.
-For the robot policy $\boldsymbol{\pi}_{\boldsymbol{\theta}}^{\gamma}$, if $\boldsymbol{\theta}\in\Theta_{H}$, the user considers the robot motion to be safe; otherwise if $\boldsymbol{\theta}\notin\Theta_{H}$, the robot motion by $\boldsymbol{\pi}_{\boldsymbol{\theta}}^{\gamma}$ might not be safe at certain time in the view of the user, and thus the user provides corrections to improve its safety.
-We use an implicit set $\Theta_{H}\in\mathbb{R}^{r}$, instead of single point $\boldsymbol{\theta}_{H}$, to capture the potential variation of user intents.
-We suppose that a volume measure of $\Theta_{H}$, $\operatorname{{Vol}}(\Theta_{H})$, is non-zero, and a user intent $\boldsymbol{\theta}_{H}\in\Theta_{H}$ can vary at different steps and context, but always within $\Theta_{H}$.
+Here, $\boldsymbol{\theta}_{H}$ is from a set $\boldsymbol{\theta}_{H}\in\Theta_{H}$, encoding all possible user intended safety.
+$\Theta_{H}$ is implicit and unknown to the robot.
+For the robot policy $\boldsymbol{\pi}_{\boldsymbol{\theta}}^{\gamma}$, if $\boldsymbol{\theta}\in\Theta_{H}$, the user considers the planned robot motion to be safe; otherwise if $\boldsymbol{\theta}\notin\Theta_{H}$, the robot motion by $\boldsymbol{\pi}_{\boldsymbol{\theta}}^{\gamma}$ might not be safe at certain time.
+We use an implicit set $\Theta_{H}\in\mathbb{R}^{r}$, instead of a single point $\boldsymbol{\theta}_{H}$, to capture the potential variance of user intents.
+We suppose that a volume measure of $\Theta_{H}$, $\operatorname{{Vol}}(\Theta_{H})$, is non-zero, and a user intent $\boldsymbol{\theta}_{H}\in\Theta_{H}$ can vary at different steps and contexts, but always within $\Theta_{H}$.
 III-C2 Human Directional Correction
-We consider human gives correction to a robot in an online fashion while the robot is executing its most updated $\boldsymbol{\pi}_{\boldsymbol{\theta}}^{\gamma}$.
+As shown in Fig.
+2, we consider human correcting a robot online while the robot is executing $\boldsymbol{\pi}_{\boldsymbol{\theta}}^{\gamma}$.
 Specifically, let $i$ be the index of learning update, and suppose robot current policy is $\boldsymbol{\pi}_{\boldsymbol{\theta}_{i}}^{\gamma}$.
-During rolling out $\boldsymbol{\pi}_{\boldsymbol{\theta}_{i}}^{\gamma}$, the robot receives a sparse human directional correction $\boldsymbol{a}_{k_{i}}\in\mathbb{R}^{m}$ at the rollout time step $k_{i}$.
-Here, we do not assume precise timing $k_{i}$ or duration of the human correction due to the sparsity or intermittency of human feedback.
-In other words, human can apply one or multiple corrections at any time in robot rollout time steps for any duration as he/she finds necessary.
-But we here assume human directional corrections $\boldsymbol{a}_{k_{i}}$ are in robot action space.
-For example, in autonomous driving where the robot action space is steering angle and acceleration, human can directly apply feedback with steering wheel and pedals.
-In some other applications, the action space correction can be achieved via specific mapping or interface.
-For example, to correct a torque-controlled robot arm, human’s correction on robot end-effector can be mapped into joint torque via Jacobian transpose matrices.
-The reason why we only consider the action space correction is because first, corrections in input spaces may be easier to implement, as shown in our later experiments, and second, corrections in state spaces can be infeasible for the under-actuated robot system [20].
-After receiving each $\boldsymbol{a}_{k_{i}}$, the robot updates the safety parameter $\boldsymbol{\theta}_{i}\rightarrow\boldsymbol{\theta}_{i+1}$ in (1) for its MPC policy.
-Then, the robot continues with $\boldsymbol{\pi}_{\boldsymbol{\theta}_{i+1}}^{\gamma}$ until the next possible user corrections.
-This process repeats until the convergence of the policy update or the user is satisfied with the robot motion.
-An illustration of one update is shown in Fig.
-2.
+During rollout of $\boldsymbol{\pi}_{\boldsymbol{\theta}_{i}}^{\gamma}$, the robot receives a human correction $\boldsymbol{a}_{k_{i}}\in\mathbb{R}^{m}$ at the rollout time step $k_{i}$.
+Here, we do not pose any assumption on correction time $k_{i}$ and its duration due to the sparsity or intermittency of human feedback.
+In other words, the human can apply one or multiple corrections at any time as necessary.
+But we assume human correction $\boldsymbol{a}_{k_{i}}$ is in the robot action space.
+For example, in autonomous driving where the robot action space is steering angle and acceleration, human can directly apply feedback with the steering wheel and pedals.
+In some other applications, the action space correction can be achieved via a specific user interface.
+For example, to correct a torque-controlled robot arm, human’s correction on the end-effector can be mapped into joint torque via the Jacobian transpose.
+The reason why we only consider the action space correction is because first, corrections in input spaces may be easier to implement, as shown in our later experiments, and second, corrections in state spaces can be infeasible for the under-actuated robot system.
+After receiving a human correction $\boldsymbol{a}_{k_{i}}$, the robot updates the safety parameter $\boldsymbol{\theta}_{i}\rightarrow\boldsymbol{\theta}_{i+1}$ in (1) for its Safe MPC.
+Then, the robot continues with $\boldsymbol{\pi}_{\boldsymbol{\theta}_{i+1}}^{\gamma}$ until the next possible user correction.
+This process repeats until the convergence of the policy update or the user is satisfied with the robot’s motion.
 Figure 2: Update of Safe MPC policy with human directional corrections.
-In the above learning process, we consider that a human user only provides directional correction $\boldsymbol{a}$, i.e., the magnitude $\left\lVert\boldsymbol{a}\right\rVert$ does not matter (theoretical justification will be given later).
-For example, when correcting a mobile robot, the user’s correction can be "left" or "right", without specifying how much the robot should go.
-As shown in our previous work [20] or later, the type of directional feedback provides a consistently large and flexible correction space, effectively overcoming the over-shooting correction, compared to existing magnitude feedback.
+While the correction $\boldsymbol{a}$ given by human is real-valued, only its direction is used in the learning process.
+This means that we will use the sign (for scalar signal) or its normalized unit vector (for vector-valued correction signal) for learning update.
+The magnitude of the correction is NOT taken into account in the learning process.
+As we will show later, directional feedback provides a consistently large and flexible correction space, effectively overcoming the over-shooting correction, compared to magnitude feedback.
 III-C3 Robot Operation Safety and Emergency Stop
-In the aforementioned learning procedure, the robot will respond to human corrections by executing the updated safe MPC policy $\pi_{\boldsymbol{\theta}}^{\gamma}$.
-However, it is possible that the robot may still exit the user-intended safety region.
-In such cases, a human user may trigger an emergency stop mechanism to stop the robot motion and potentially restart it from a desired safety region.
-In this sense, all trajectories before the human triggers the emergency stop are considered safe in human intent.
-In other words, until an emergency stop is triggered, human corrections typically occur while the robot is still within the intended safety zone; otherwise, the corrections would be useless.
+In the aforementioned learning procedure, the robot will respond to human correction by executing the updated safe MPC policy $\pi_{\boldsymbol{\theta}}^{\gamma}$.
+However, it is possible that the robot may still exit the user-intended safety.
+In such cases, the human may trigger an emergency stop to stop the robot and potentially restart it from a desired safety region.
+In such case, the executed motion before the emergency stop is assumed safe.
+In other words, until an emergency stop is triggered, human correction typically occur when the robot is still in the intended safety set; otherwise, the correction would be useless.
 For instance, correcting a vehicle after it is crashed is futile.
 III-C4 Parametric Safety Constraints
-In the safe MPC (4), the safety set is defined as a zero sublevel set of a safety value function ${g}_{\boldsymbol{\theta}}$.
-We note that if ${g}_{\boldsymbol{\theta}}$ has sufficient representation power, the safety set can represent complex (non-convex) safety region, which can be alternatively represented by the intersection of multiple constraints.
-In the following, we mainly focus on a weighted-features parameterization of ${g}_{\theta}$, which has been commonly used by the prior work on learning cost/reward [4, 20, 50, 51, 52].
-Specifically, we parameterize $\boldsymbol{g}_{\theta}$
+In (4), the safety set is defined as a zero sublevel set of ${g}_{\boldsymbol{\theta}}$.
+If ${g}_{\boldsymbol{\theta}}$ has sufficient representation power, the safety set can represent a complex safety region.
+In the following, we focus on a weighted-features parameterization of ${g}_{\boldsymbol{\theta}}$, which has been commonly used by the prior work on learning cost/reward [lindner2023learning, jin2022learning, zhang2023dual, biyik2022aprel, bobu2020quantifying].
+Specifically,
 ${g}_{\boldsymbol{\theta}}(\boldsymbol{\xi})=\phi_{0}(\boldsymbol{\xi})+\theta_{1}\phi_{1}(\boldsymbol{\xi})+\theta_{2}\phi_{2}(\boldsymbol{\xi})+\cdots+\theta_{r}\phi_{r}(\boldsymbol{\xi}),$ (5)
-with $\boldsymbol{\theta}=[\theta_{1},\theta_{2},...\theta_{r}]^{\mkern-1.5mu\mathsf{T}}\in\mathbb{R}^{r}$ and $\boldsymbol{\phi}=[\phi_{1},\phi_{2},\dots,\phi_{r}]^{\mkern-1.5mu\mathsf{T}}$ the weight and feature vectors, respectively.
-We fix the weight of $\phi_{0}$ as unit, because otherwise, $\boldsymbol{\theta}$ has scaling ambiguity, i.e., any scaled $c\cdot\boldsymbol{\theta}$ with $c>0$ lead to the same MPC solutions.
-The dimension (number) $r$ of features can be chosen such that $\boldsymbol{\phi}$ has enough expressiveness to represent the human-intended safety constraint.
-We do not impose any constraints on the choice of feature vector as long as it is differentiable, for example, $\boldsymbol{\phi}$ can be neural features.
-In practice the feature vector can be chosen by trial and error, i.e., starting from a few features and gradually adding more until the learned constraint is satisfactory to the human user.
-Later, we will discuss our proposed method for generic parameterization $\boldsymbol{g}_{\boldsymbol{\theta}}$.
+with $\boldsymbol{\theta}=[\theta_{1},\theta_{2},...\theta_{r}]^{\mkern-1.5mu\mathsf{T}}\in\mathbb{R}^{r}$ and $\boldsymbol{\phi}=[\phi_{1},\phi_{2},\dots,\phi_{r}]^{\mkern-1.5mu\mathsf{T}}$ the learnable weight and feature vectors, respectively.
+We fix the weight of $\phi_{0}$ as unit, because otherwise, $\boldsymbol{\theta}$ has scaling ambiguity, i.e., any scaled $c\cdot\boldsymbol{\theta}$ with $c>0$ lead to the same constraint.
+In this work, we assume that the features in $\boldsymbol{\phi}$ are given as a priori.
+These features, defined in system states (e.g., position, velocity) and actions, can be either manually defined or obtained from pre-trained neural representations, leveraging prior knowledge of the environment and tasks, as demonstrated in our experiments.
+The feature selection itself is beyond the scope of this paper and is left for future work.
+A more detailed discussion of this assumption and its limitations is in Section IX-A3.
 III-C5 Problem of Safe MPC Alignment
-Following the procedure of human correction and robot policy update in Subsection III-C and given a weighted-feature safety constraint (5), we aim to develop a policy update rule: $\boldsymbol{\theta}_{i}\rightarrow\boldsymbol{\theta}_{i+1}$ for each received human directional correction, such that $\boldsymbol{\theta}_{i}\rightarrow\boldsymbol{\theta}_{H}\in\Theta_{H}$, as $i=1,2,3,\dots,$.
+Following the procedure of human correction and robot policy update in Section III-C, we aim to develop a policy update rule: $\boldsymbol{\theta}_{i}\rightarrow\boldsymbol{\theta}_{i+1}$ for each received human correction, such that $\boldsymbol{\theta}_{i}\rightarrow\boldsymbol{\theta}_{H}\in\Theta_{H}$, as $i=1,2,3,\dots,$.
 i.e., safety constraints in the robot’s safe MPC policy will be aligned with human intent.
-IV Overview of Safe MPC Alignment Algorithm
+IV Overview of Safe MPC Alignment
 IV-A Hypothesis on Human Directional Corrections
 Informally, we have the following intuitions on human feedback.
-First, human corrections, on average, improve the robot safety.
-Second, human correction on robot safety is typically proactive and happens before the robot leaves the human-intended safety set.
+First, human corrections, on average, improve the robot’s safety.
+Second, human correction is typically proactive and happens before the robot leaves the human-intended safety set.
 In other words, human correction is made when the robot is still in $\mathcal{S}(\boldsymbol{\theta}_{H})$ in (4).
 For example, it is useless to correct a vehicle after its crash.
-Third, human feedback is typically safety-margin aware: as a robot approaches the safety boundary, human users tend to prioritize safety over the task objective.
+Third, human feedback is typically safety-margin aware: as a robot approaches the safety boundary, human users tend to prioritize safety over the task cost.
 For instance, when a vehicle moves closer to obstacles, humans may instinctively shift their focus to avoiding collisions rather than pursuing the original target.
@@
 Without loss of generality, suppose the motion plan $\boldsymbol{\xi}_{\boldsymbol{\theta}}^{\gamma}$ is a local solution to the penalty MPC (3), and when the robot execute the first action $\boldsymbol{u}_{0}^{\boldsymbol{\theta},\gamma}$ (receding horizon implementation), a human applies a directional correction $\boldsymbol{a}$ in robot action space.
+In this subsection, the subscription $k_{i}$ is omitted for notation simplicity.
 Define
 $\mathbf{a}=(\boldsymbol{a}^{\mkern-1.5mu\mathsf{T}},0,0,...)\in\mathbb{R}^{mT}.$ (6)
-We assume there is a nonzero volume set $\bar{\Theta}_{H}$ (containing a ball) in user implicit set $\Theta_{H}$, i.e., $\bar{\Theta}_{H}\subseteq\Theta_{H}$, such that for any $\boldsymbol{\theta}_{H}\in\bar{\Theta}_{H}$, the human correction $\boldsymbol{a}$, on average, indicates a direction of lowering the penalty value $B(\boldsymbol{\xi}_{\boldsymbol{\theta}}^{\gamma},\boldsymbol{\theta}_{H})$ for the robot motion plan $\boldsymbol{\xi}_{\boldsymbol{\theta}}^{\gamma}$.
+We assume there is a nonzero volume set $\bar{\Theta}_{H}$ (containing a ball) in user implicit set $\Theta_{H}$, i.e., $\bar{\Theta}_{H}\subseteq\Theta_{H}$, such that for any $\boldsymbol{\theta}_{H}\in\bar{\Theta}_{H}$, the human correction $\mathbf{a}$, on average, indicates a direction of lowering the penalty value $B(\boldsymbol{\xi}_{\boldsymbol{\theta}}^{\gamma},\boldsymbol{\theta}_{H})$.
 Concretely, it means $\mathbf{a}$ on averages points in the direction of the negative gradient of $B(\boldsymbol{\xi},\boldsymbol{\theta}_{H})$ at $\boldsymbol{\xi}_{\boldsymbol{\theta}}^{\gamma}$, i.e.,
 $\forall\boldsymbol{\theta}_{H}\in\bar{\Theta}_{H},\ \operatorname{\mathbb{E}}_{\mathbf{a}}\left\langle\mathbf{a},-\nabla B(\boldsymbol{\xi}_{\boldsymbol{\theta}}^{\gamma},\boldsymbol{\theta}_{H})\right\rangle=\\ \left\langle\mathbf{\bar{a}},-\nabla B(\boldsymbol{\xi}_{\boldsymbol{\theta}}^{\gamma},\boldsymbol{\theta}_{H})\right\rangle\geq 0.$ (7)
-[…]
-(7) indicates that the magnitude of $\mathbf{a}$ does not affect the inequality.
-This brings one advantage of directional correction compared to the magnitude-based correction: regardless of robot plan $\boldsymbol{\xi}_{\boldsymbol{\theta}}^{\gamma}$, the allowable human corrections satisfying (7) always account for half of the action space, making directional corrections more accessible for human users, as shown in Fig.
+Here, $\left\langle\cdot,\cdot\right\rangle$ denotes the inner product operation, and ${\operatorname{\mathbb{E}}}_{\mathbf{a}}(\mathbf{a})=\mathbf{\bar{a}}$ denotes the expectation value of $\mathbf{a}$.
+$\nabla B(\boldsymbol{\xi},\boldsymbol{\theta})$ denotes the full gradient of $B$ with respect to $\boldsymbol{u}_{0:T-1}$ evaluated at trajectory $\boldsymbol{\xi}=\{\boldsymbol{u}_{0:T-1},\boldsymbol{x}_{0:T}\}$.
+Similar notations apply to $J$ and $g$ below.
+Since $\nabla B(\boldsymbol{\xi},\boldsymbol{\theta})$ is not dependent on $\mathbf{a}$, the expectation can be moved inside the inner product.
+Because state sequence $\boldsymbol{x}_{0:T}$ is the rollout of $\boldsymbol{u}_{0:T-1}$ on dynamics, the differentiation goes through the dynamics equation.
+Equation (7) shows that although the human correction $\boldsymbol{a}$ is provided as a real-valued vector, only its direction—not its magnitude—affects the inequality.
+This brings one advantage of directional correction compared to the magnitude-based correction: the allowable human correction satisfying (7) always account for half of the action space, making directional corrections more accessible for human users, as shown in Fig.
 3.
@@
 $\forall\boldsymbol{\theta}_{H}\in\bar{\Theta}_{H},\left<\mathbf{\bar{a}},-\nabla J(\boldsymbol{\xi}_{\boldsymbol{\theta}}^{\gamma})\right>+\\ \gamma\frac{1}{-g_{\boldsymbol{\theta}_{H}}(\boldsymbol{\xi}_{\boldsymbol{\theta}}^{\gamma})}\left<\mathbf{\bar{a}},-\nabla g_{\boldsymbol{\theta}_{H}}(\boldsymbol{\xi}_{\boldsymbol{\theta}}^{\gamma})\right>\geq 0.$ (8)
-Figure 3: Illustration of directional corrections in the space of robot motion plan $\boldsymbol{\xi}$s.
+Figure 3: Illustration of directional correction in the space of robot plan $\boldsymbol{\xi}$s.
 The grey circles represent the level contours of $B(\boldsymbol{\xi},\boldsymbol{\theta}_{H})$.
-The black dot stands for current motion plan $\boldsymbol{\xi}_{\boldsymbol{\theta}}^{\gamma}$ and the red point stands for the plan $\boldsymbol{\xi}_{\boldsymbol{\theta}_{H}}^{\gamma}$ with human intended safety.
-The blue region stands for the half-space of all possible human corrections under our assumption.
+The black dot stands for the current motion plan $\boldsymbol{\xi}_{\boldsymbol{\theta}}^{\gamma}$ and the red point stands for optimal plan $\boldsymbol{\xi}_{\boldsymbol{\theta}_{H}}^{\gamma}$ with $\theta_{H}$.
+The blue region stands for the half-space of all possible human correction under our assumption (7).
 Some samples of human corrections are shown by blue arrows.
 For the above equation, we have the following interpretations.
-(I) The first and second terms on the left side of (8) correspond to human correction toward lowering the cost and improving the safety of $\boldsymbol{\xi}_{\boldsymbol{\theta}}^{\gamma}$, respectively.
-$\gamma$ serves as a weight between two improvements.
-(II) If we interpret the value of $-{g_{\boldsymbol{\theta}_{H}}(\boldsymbol{\xi}_{\boldsymbol{\theta}}^{\gamma})}$ as the safety margin of $\boldsymbol{\xi}_{\boldsymbol{\theta}}^{\gamma}$, (8) suggests that the human correction is safety-margin aware: the closer the robot motion get to the unsafe region, the more will human user emphasize on improving the robot’s safety than minimizing the objective.
+(I) The first and second terms of (8) correspond to human correction toward lowering the cost and improving the safety of $\boldsymbol{\xi}_{\boldsymbol{\theta}}^{\gamma}$, respectively.
+$\gamma$ weighs between two improvements.
+(II) If we interpret the value of $g_{\boldsymbol{\theta}_{H}}(\boldsymbol{\xi}_{\boldsymbol{\theta}}^{\gamma})$} as the safety margin of the plan $\boldsymbol{\xi}_{\boldsymbol{\theta}}^{\gamma}$, (8) suggests that the human correction is safety-margin aware: the closer the robot motion get to the unsafe region, the more will human improve the robot safety than minimizing the cost.
 This property is consistent with our previous intuition.
@@
 $\forall\boldsymbol{\theta}_{H}\in\bar{\Theta}_{H},\ \quad g_{\boldsymbol{\theta}_{H}}(\boldsymbol{\xi}_{\boldsymbol{\theta}}^{\gamma})<0.$ (9)
-This corresponds to the third intuition: a user typically gives correction only when the robot is in user intended safety region, i.e., $\boldsymbol{\xi}_{\boldsymbol{\theta}}^{\gamma}\in\mathcal{S}(\boldsymbol{\theta}_{H})$.
-In sum, we have used (7) and (9) to formalize the intuitions we have observed at the beginning of this session.
-Those two constraints on $\boldsymbol{\theta}_{H}$ will be used in the robot learning next.
-IV-B Alignment Algorithm: Learning by Hypothesis Space Cutting
-To find $\boldsymbol{\theta}_{H}\in\bar{\Theta}_{H}\subseteq\Theta$ for the safe MPC (1), our idea is that for each human directional correction received $i=0,1,2,....$, the robot performs one learning iteration, by updating a hypothesis space $\Theta_{i}\in\mathbb{R}^{r}$, a set containing all possible $\bar{\Theta}_{H}$.
-At initial, the robot may choose a large enough hypothesis space $\Theta_{0}$.
-Specifically, one can choose $\Theta_{0}$ as a box space
+This corresponds to the third intuition: a user typically gives correction only when the robot is in user intended safety region.
+In sum, the two constraints(7) and (9) formalize the intuitions we have observed at the beginning of this subsection. Those constraints will be used in the robot learning next.
+IV-B Main Algorithm: Learning by Hypothesis Space Cutting
+To find $\boldsymbol{\theta}_{H}\in\bar{\Theta}_{H}\subseteq\Theta$ for the Safe MPC, our idea is that for each human correction $i=0,1,2,.$..., the robot performs one learning iteration, by updating a hypothesis space $\Theta_{i}\in\mathbb{R}^{r}$, a set containing all possible $\bar{\Theta}_{H}$.
+Initially, the robot may choose a large box hypothesis space $\Theta_{0}$:
 $\Theta_{0}=\{\boldsymbol{\theta}\ |\ \boldsymbol{c}_{l}\leq\boldsymbol{\theta}\leq\boldsymbol{c}_{u}\}\subset\mathbb{R}^{r},$ (10)
-in which $\boldsymbol{c}_{l},\boldsymbol{c}_{u}$ are the vectors of lower and upper bounds of the box space, resepctively, and the inequalities applied element-wise.
+in which $\boldsymbol{c}_{l},\boldsymbol{c}_{u}$ are the vectors of lower and upper bounds of the box space, respectively, and the inequalities are applied element-wise.
 By saying a $\Theta_{0}$ large enough, we assume
 $\bar{\Theta}_{H}\subset\Theta_{0}.$ (11)
-Later, we will show that our method can self-access the learning failure in case of misspecification, where (11) is not met.
-Recall the procedure of human correction in Section III-C and Fig.
-2.
-At $i$-th learning iteration, $i=1,2,3...$, the robot executes $\boldsymbol{\pi}_{\boldsymbol{\theta}_{i}}^{\gamma}$ with $\boldsymbol{\theta}_{i}\in\Theta_{i-1}$.
-The user gives a directional correction $\boldsymbol{a}_{k_{i}}$ at rollout step ${k_{i}}$.
-By letting $\mathbf{\bar{a}}_{i}=(\boldsymbol{a}_{k_{i}}^{T},0,0,...)$ as in (6), we can establish the following two constraints on $\bar{\Theta}_{H}$ from the previous hypotheses (7) and (9):
-$\bar{\Theta}_{H}\subset\mathcal{C}_{i}=\left\{\boldsymbol{\theta}|\left<\mathbf{\bar{a}}_{i},\nabla B(\boldsymbol{\xi}_{\boldsymbol{\theta}_{i}}^{\gamma},\boldsymbol{\theta})\right>\leq 0,\,\,\,g_{\boldsymbol{\theta}}(\boldsymbol{\xi}_{\boldsymbol{\theta}_{i}}^{\gamma})<0\right\},$ (12)
-which we refer to as the cutting set.
-Note that the first inequality $\mathbf{\bar{a}}_{i},\nabla B(\boldsymbol{\xi}_{\boldsymbol{\theta}_{i}}^{\gamma},\boldsymbol{\theta})\right>\leq 0$ is deduced by multiplying -1 on both sides of (7).
-Using $\mathcal{C}_{i}$, one can update the hypothesis space
+Later, we will show our method can self-access the misspecification, where (11) is not met.
+[…]
+$\mathcal{C}_{i}=\left\{\boldsymbol{\theta}|\left<\mathbf{\bar{a}}_{i},\nabla B(\boldsymbol{\xi}_{\boldsymbol{\theta}_{i}}^{\gamma},\boldsymbol{\theta})\right>\leq 0,\,\,\,g_{\boldsymbol{\theta}}(\boldsymbol{\xi}_{\boldsymbol{\theta}_{i}}^{\gamma})<0\right\},$ (12)
+which we refer to as the cutting set. Using $\mathcal{C}_{i}$, one can update the hypothesis space
 $\Theta_{i}=\Theta_{i-1}\cap\mathcal{C}_{i}.$ (13)
-The above update can be viewed as a step of using (12) to cut (intersect with) the current hypothesis space $\Theta_{i-1}$ to remove the portion where $\bar{\Theta}_{H}$ is impossible.
-The remaining hypothesis space $\Theta_{i}$ will be new hypothesis used for the next iteration.
+The above update can be viewed as using (12) to cut the hypothesis space $\Theta_{i-1}$ to remove portions where $\bar{\Theta}_{H}$ is impossible.
+The new hypothesis space $\Theta_{i}$ will be used for the next iteration.
 With the above update, we have the following result.
@@
 See the proof in Appendix A-A. ∎
-As shown in Fig.
-4, the procedure of robot learning is to cut and move hypothesis space based on human directional correction.
-Therefore, we name our alignment method Hypothesis Space Cutting.
+As in Fig.
+4, the procedure of robot learning is to cut and move the hypothesis space based on human directional correction.
+Therefore, we name our learning method Safe MPC Alignment by Hypothesis Space Cutting.
 As a summary, we present the three main steps of our algorithm below.
-Figure 4: The update of constraint parameter (2D examples).
+Figure 4: The update of hypothesis space (2D example).
 Two cutting lines are used to cut $\Theta_{i-1}$ to $\Theta_{i}$.
-The user implicit intent set $\bar{\Theta}_{H}$ is contained in $\Theta_{i}$.
-New $\boldsymbol{\theta}_{i+1}$ is chosen as the parameter for the next safe MPC policy.
+The user intent set $\bar{\Theta}_{H}$ is contained in $\Theta_{i}$.
+New $\boldsymbol{\theta}_{i+1}$ is chosen from $\Theta_{i}$ for the updated Safe MPC policy.
 Safe MPC Alignment by Hypothesis Space Cutting Initialize $\Theta_{0}$ that contains $\bar{\Theta}_{H}$.
-Repeat the following three steps for $i=1,2,...,K$ until $\boldsymbol{\theta}_{i}\in\bar{\Theta}_{H}$.
-Step 1: Robot chooses parameter $\boldsymbol{\theta}_{i}\in\Theta_{i-1}$ for its current safe MPC policy $\boldsymbol{\pi}_{\boldsymbol{\theta}_{i}}^{\gamma}$.
+Repeat the following three steps for i=1,2,,Ki=1,2,...,K until $\boldsymbol{\theta}_{i}\in\bar{\Theta}_{H}$.
+Step 1: Robot chooses parameter $\boldsymbol{\theta}_{i}\in\Theta_{i-1}$ for its current Safe MPC $\boldsymbol{\pi}_{\boldsymbol{\theta}_{i}}^{\gamma}$.
 Step 2: Human user applies a directional correction $\boldsymbol{a}_{k_{i}}$ while robot executes $\boldsymbol{\pi}_{\boldsymbol{\theta}_{i}}^{\gamma}$ Step 3: Perform hypothesis space cut using (13) to get new hypothesis space $\Theta_{i}$.
-At $i$th learning iteration, the robot chooses $\boldsymbol{\theta}_{i}\in\Theta_{i-1}$ from the most recent hypothesis space $\Theta_{i-1}$ for its current MPC $\boldsymbol{\pi}_{\boldsymbol{\theta}_{i}}^{\gamma}$.
-While executing $\boldsymbol{\pi}_{\boldsymbol{\theta}_{i}}^{\gamma}$, the robot receives human directional correction $\boldsymbol{a}_{k_{i}}$ at the rollout time step $k_{i}$.
-This correction leads to a cutting set $\mathcal{C}_{i}$ (12), which will be used to update $\Theta_{i-1}$ to $\Theta_{i}$ in (13).
+At $i$th learning update, the robot chooses $\boldsymbol{\theta}_{i}\in\Theta_{i-1}$ for its current Safe MPC $\boldsymbol{\pi}_{\boldsymbol{\theta}_{i}}^{\gamma}$.
+While executing $\boldsymbol{\pi}_{\boldsymbol{\theta}_{i}}^{\gamma}$, the robot receives human correction $\boldsymbol{a}_{k_{i}}$ at the rollout time step $k_{i}$.
+This correction leads to a cutting set $\mathcal{C}_{i}$ (12), which will update $\Theta_{i-1}$ to $\Theta_{i}$ in (13).
 The above iteration repeats $i=1,2,3,$....
@@
 Given a volume measure, it follows $\operatorname{{Vol}}\Theta_{0}\geq\operatorname{{Vol}}\Theta_{1}\geq...,\geq\operatorname{{Vol}}{\bar{\Theta}_{H}}$.
-Thus, the volume of the hypothesis space is at least not increasing after each user correction.
-It is hoped that the volume of the hypothesis space decreases quickly, such that new $\boldsymbol{\theta}_{i+1}$ increases its chance of falling inside $\bar{\Theta}_{H}$ — a successful safety alignment between the robot and human.
-For the above Safe MPC Alignment method, we will later (in Section V) show that within finite number $K$ of corrections, the method will either
+Thus, the volume of the hypothesis space is at least not increasing after each correction.
+It is hoped that the volume of the hypothesis space decreases quickly, such that new $\boldsymbol{\theta}_{i+1}$ increases its chance of choosing from $\bar{\Theta}_{H}$ — meaning a successful safety alignment between the robot and human.
+For the Safe MPC Alignment, we will later (Section V) show within finite number $K$ of correction, the method will either
   • (I)
@@
 IV-C Properties and Geometric Interpretations
-In this session, we present some properties and the geometric interpretations for the proposed safe MPC alignment.
-Those properties will be used to establish the convergence of the algorithm in the next section.
-IV-C1 Piecewise Linearity of Cuts
+We next show properties and geometric interpretations for the proposed safe MPC alignment process. Those properties will be used to establish the convergence of the algorithm.
+IV-C1 Piecewise Linear Cut
 By recalling the weight-feature parameterization of the safety constraint (5), one can examine that the inequities in (12) are both linear in $\theta$.
-This means that the cutting set $\mathcal{C}_{i}$ in (12) is a piecewise linear cut, as stated formally in the following lemma.
+This means that the cutting set $\mathcal{C}_{i}$ is a piecewise linear cut, as stated below.
 Lemma 3.
@@
 $\small\boldsymbol{h}_{i}=-\left<\mathbf{\bar{a}}_{i},\nabla J(\boldsymbol{\xi}_{\boldsymbol{\theta}_{i}}^{\gamma})\right>\boldsymbol{\phi}(\boldsymbol{\xi}_{\boldsymbol{\theta}_{i}}^{\gamma})+\gamma\frac{\partial\boldsymbol{\phi}}{\partial\boldsymbol{u}}(\boldsymbol{\xi}_{\boldsymbol{\theta}_{i}}^{\gamma})\mathbf{\bar{a}}_{i},$ (15)
-[…]
+$b_{i}=\left<\mathbf{\bar{a}}_{i},\nabla J(\boldsymbol{\xi}_{\boldsymbol{\theta}_{i}}^{\gamma})\right>\phi_{0}(\boldsymbol{\xi}_{\boldsymbol{\theta}_{i}}^{\gamma})-\gamma\left<\mathbf{\bar{a}}_{i},\nabla\phi_{0}(\boldsymbol{\xi}_{\boldsymbol{\theta}_{i}}^{\gamma})\right$>.
+(16)
 Proof.
 See Appendix A-B. ∎
-According to the above lemma, $\Theta_{i-1}\cap\mathcal{C}_{i}$ can be viewed as cutting $\Theta_{i-1}$ with the boundary of $\mathcal{C}_{i}$ which is two hyperplanes:
+According to the above lemma, $\Theta_{i-1}\cap\mathcal{C}_{i}$ can be viewed as cutting $\Theta_{i-1}$ with two hyperplanes (Fig. 4):
 $\begin{split}\mathbf{bd}\ \mathcal{C}_{i}=\{\boldsymbol{\theta}\ |\ \boldsymbol{\theta}^{\mkern-1.5mu\mathsf{T}}\boldsymbol{h}_{i}=b_{i},{\boldsymbol{\theta}}^{\mkern-1.5mu\mathsf{T}}\boldsymbol{\phi}(\boldsymbol{\xi}_{\boldsymbol{\theta}_{i}}^{\gamma})\leq-\phi_{0}(\boldsymbol{\xi}_{\boldsymbol{\theta}_{i}}^{\gamma})\}\ \cup\\ \{\boldsymbol{\theta}\ |\ \boldsymbol{\theta}^{\mkern-1.5mu\mathsf{T}}\boldsymbol{h}_{i}\leq b_{i},{\boldsymbol{\theta}}^{\mkern-1.5mu\mathsf{T}}\boldsymbol{\phi}(\boldsymbol{\xi}_{\boldsymbol{\theta}_{i}}^{\gamma})=-\phi_{0}(\boldsymbol{\xi}_{\boldsymbol{\theta}_{i}}^{\gamma})\},\end{split}$ (17)
-shown in Fig. 4.
-As the initial hypothesis space $\Theta_{0}\in\mathbb{R}^{r}$ in (10) is a convex polytope, iteratively cutting with hyperplane pairs yields a series of convex polytopes in $\mathbb{R}^{r}$.
-Consequently, all $\Theta_{i}$s, $i=0,1,2,...$, are convex polytopes.
+As the initial hypothesis space $\Theta_{0}\subset\mathbb{R}^{r}$ in (10) is a convex polytope, iteratively cutting with hyperplanes yields a series of convex polytopes $\Theta_{i}$s, $i=0,1,2,$....
 IV-C2 Relationship between $\boldsymbol{\theta}_{i}$ and $\mathbf{bd}\ \mathcal{C}_{i}$
-According to (14), the cutting set $\mathcal{C}_{i}$ depends on plan $\boldsymbol{\xi}_{\boldsymbol{\theta}_{i}}^{\gamma}$ which is a solution to penalty-based MPC (3) with $\boldsymbol{\theta}_{i}\in\Theta_{i-1}$.
 The following assertion states a relationship between $\boldsymbol{\theta}_{i}$ and $\mathbf{bd}\ \mathcal{C}_{i}$.
@@
 We briefly discuss the implication of different choices of $\boldsymbol{\theta}_{i}\in\Theta_{i-1}$.
-When choosing $\boldsymbol{\theta}_{i}$ near the boundary of $\Theta_{i-1}$, as shown in Fig.
+When choosing $\boldsymbol{\theta}_{i}$ near the boundary of $\Theta_{i-1}$, as in Fig.
 5 and Fig.
 5, the removal of the hypothesis space depends on the specific direction of cut $\mathbf{bd}\,\mathcal{C}_{i}$.
-Since the direction of the cut $\mathbf{bd}\,\mathcal{C}_{i}$ will be eventually determined by the direction of user correction according to (17), the removal volume of the hypothesis space can be relatively small, as in Fig.
+Since the direction of $\mathcal{C}_{i}$ will be eventually determined by the direction of user correction according to (17), the removal of the hypothesis space can be relatively small, as in Fig.
 5, or large, as in Fig.
 5, depending on the specific direction of the correction.
-As the specific direction of user correction is hard to assume, one can intuitively see that a good choice of $\boldsymbol{\theta}_{i}\in\Theta_{i-1}$ will be near the center of $\Theta_{i-1}$.
-As such, it can lead to an averaged large cut of the hypothesis space regardless of the direction of $\mathbf{bd}\ \mathcal{C}_{i}$, as shown in Fig.
+As the specific direction of user correction is hard to assume, a good choice of $\boldsymbol{\theta}_{i}\in\Theta_{i-1}$ will be near the center of $\Theta_{i-1}$.
+As such, it can lead to an averaged large cut of the hypothesis space regardless of the direction of $\mathcal{C}_{i}$, as in Fig.
 5.
-We will rigorously discuss the centering choice of $\boldsymbol{\theta}_{i}$ and the guarantee of volume reduction rate of the hypothesis space in the next section.
+We will rigorously discuss the centering choice of $\boldsymbol{\theta}_{i}$ and volume reduction rate of the hypothesis space in the next section.
 Figure 5: Illustration of the different choices of $\boldsymbol{\theta}_{i}$.
-(a) and (b): When $\boldsymbol{\theta}_{i}$ is chosen near the set boundary, the removal volume depends on the specific direction of $\mathbf{bd}\ \mathcal{C}_{i}$.
+(a) and (b): When $\boldsymbol{\theta}_{i}$ is chosen near the set boundary, the removed volume depends on the specific direction of $\mathbf{bd}\ \mathcal{C}_{i}$.
 (c): When $\boldsymbol{\theta}_{i}$ is near the center, the cut volume is large on average regardless of the specific user directional correction.
-V Main Results and Implementation Details
-In this section, we will present the details for how to select $\boldsymbol{\theta}_{i}\in\Theta_{i-1}$ at each learning iteration $i$ in our Safe MPC Alignment algorithm.
-Based on the previous geometric intuition, our choice of $\boldsymbol{\theta}_{i}$ is to achieve a large averaged reduction rate in the hypothesis space volume for all possible human directional corrections.
-The reduction rate enables us to establish an exponential convergence of the learning process.
-As a result, we can find an upper bound of total human corrections for successful learning, i.e., $\boldsymbol{\theta}_{i}\in\bar{\Theta}_{H}$, or identify the misspecification of hypothesis space.
+V Implementation Details and Main Results
 V-A The Choice of $\boldsymbol{\theta}_{i}$
-As discussed previously, $\boldsymbol{\theta}_{i}$ needs to be picked from the "center" of $\Theta_{i-1}$ in each learning iteration.
+At each learning iteration $i$, based on the previous geometric intuition, our choice of $\boldsymbol{\theta}_{i}$ is to achieve a large reduction of the hypothesis space for all possible human directional corrections.
+Thus, $\boldsymbol{\theta}_{i}$ needs to be picked from the "center" of $\Theta_{i-1}$.
 One option is to choose the mass center of $\Theta_{i-1}$.
-However, computing the centroid of a polytope can be #P-Hard [53].
-While methods like the Hit-And-Run algorithm [54] estimate the mass center by sampling, they can suffer from high sample complexity.
-Another possible choice is to use the center of the maximum inscribed ball of a polytope [55], but it lacks of the guarantee in volume reduction rate.
-For computational efficiency and guaranteed convergence, we will choose $\boldsymbol{\theta}_{i}$ as the center of the Maximum Volume Ellipsoid (MVE) inscribed in the convex set $\Theta_{i-1}$, shown in Fig.
-6.
-Figure 6: $\boldsymbol{\theta}_{i}$ is chosen as the center of MVE of $\Theta_{i-1}$ (2D example).
-Definition 1 (Maximum Volume inscribed Ellipsoid (MVE)[56]).
+However, computing the centroid of a polytope can be #P-Hard [rademacher2007approximating].
+For computational efficiency, we choose $\boldsymbol{\theta}_{i}$ as the center of the Maximum Volume Ellipsoid (MVE) inscribed in $\Theta_{i-1}$, and will later compare its performance with other choices of $\theta_{i}$ in our experiments.
+Definition 1 (Maximum Volume inscribed Ellipsoid (MVE)[boyd2004convex]).
 The MVE of a given convex set $\Theta\subset\mathbb{R}^{r}$ is defined as the image of the unit ball under an affine transformation
-=$\mathcal{E}=\{\bar{H}\boldsymbol{\theta}+\bar{\boldsymbol{d}}\ |\ \|\boldsymbol{\theta}\|_{2}\leq 1\},$ (19)
+$\mathcal{E}=\{\bar{H}\boldsymbol{\theta}+\bar{\boldsymbol{d}}\ |\ \|\boldsymbol{\theta}\|_{2}\leq 1\},$ (19)
 where the affine transformation parameter $\bar{H}\in\mathbb{S}^{r}_{++}$ (a $r\times r$ symmetric positive definite matrix), and $\bar{\boldsymbol{d}}\in\mathbb{R}^{r}$ is called the center of $\mathcal{E}$.
@@
 $\begin{split}\bar{H},\bar{\boldsymbol{d}}=\underset{H\in\mathbb{S}^{r}_{++},\boldsymbol{d}}{\arg\max}\quad&\log\det H\\ \rm{s.t.}\quad&\sup\nolimits_{\|\boldsymbol{\theta}\|_{2}\leq 1}\mathbbm{1}_{\Theta}(H\boldsymbol{\theta}+\boldsymbol{d})\leq 0,\end{split}$ (20)
-where the indicator function $\mathbbm{1}_{\Theta}(\boldsymbol{\theta})=0$ for $\boldsymbol{\theta}\in\Theta$ and otherwise $\mathbbm{1}_{\Theta}(\boldsymbol{\theta})=+\infty$.
-Specifically, when $\Theta$ is a polytope $\{\boldsymbol{\theta}^{\mkern-1.5mu\mathsf{T}}\boldsymbol{n}_{i}\leq c_{i},\ i=1,\dots,m\}$, the problem can be converted to a convex optimization with $m$ constraints, see chapter 8.4.2 in [56]:
+where the indicator $\mathbbm{1}_{\Theta}(\boldsymbol{\theta})=0$ for $\boldsymbol{\theta}\in\Theta$ and otherwise $\mathbbm{1}_{\Theta}(\boldsymbol{\theta})=+\infty$.
+Specifically, when $\Theta$ is a polytope $\{\boldsymbol{\theta}^{\mkern-1.5mu\mathsf{T}}\boldsymbol{n}_{i}\leq c_{i},\ i=1,\dots,m\}$, the problem can be converted to a convex optimization with $m$ constraints, see chapter 8.4.2 in [boyd2004convex]:
 $\begin{split}\bar{H},\bar{\boldsymbol{d}}=\underset{H\in\mathbb{S}^{r}_{++},\boldsymbol{d}}{\arg\min}\quad&\log\det H^{-1}\\ \rm{s.t.}\quad&\left\lVert H\boldsymbol{n}_{i}\right\rVert_{2}+\boldsymbol{d}^{\mkern-1.5mu\mathsf{T}}\boldsymbol{n}_{i}\leq c_{i},\ i=1,...,m.\end{split}$ (21)
-Recall that in our Safe MPC Alignment the learning update (13) is to iteratively cut the hypothesis space by a pair of hyperplanes (17), and that the initial hypothesis space (11) is a box.
+Recall that in Safe MPC Alignment the learning update (13) is to iteratively cut the hypothesis space by hyperplanes (17), and that the initial hypothesis space (11) is a box.
 By induction, it leads to that any updated hypothesis space $\Theta_{i}$ is a convex polytope, $i=0,1,2,$....
@@
 $\boldsymbol{\theta}_{i}=\bar{\boldsymbol{d}}_{i-1}.$ (22)
-The convex optimization (20) can be solved efficiently [57].
+The convex optimization (20) can be solved efficiently [cvx].
 V-B Result 1: Finite Iteration Convergence
@@
 Lemma 5.
-In the Safe MPC Alignment algorithm, when choosing $\boldsymbol{\theta}_{i}\in\mathbb{R}^{r}$ as the center of MVE of $\Theta_{i-1}$, one has
+In the Safe MPC Alignment, when choosing $\boldsymbol{\theta}_{i}\in\mathbb{R}^{r}$ as the center of MVE of $\Theta_{i-1}$, one has
 $\frac{\mathbf{Vol}(\Theta_{i})}{\mathbf{Vol}(\Theta_{i-1})}\leq 1-\frac{1}{r},$ (23)
@@
 See Appendix A-D. ∎
-The lemma above asserts an upper bound of the volume of the updated hypothesis space $\mathbf{Vol}(\Theta_{i})$ relative to the previous volume.
+The lemma above asserts an upper bound of the volume of the hypothesis space $\mathbf{Vol}(\Theta_{i})$ relative to the previous volume.
 With this, we can now establish the finite iteration convergence of the Safe MPC Alignment, stated below.
@@
 See Appendix A-E. ∎
-The above main result asserts a maximum iteration count for the proposed algorithm to find $\boldsymbol{\theta}_{i}\in\Theta_{H}$.
-In other words, within $K$ rounds of human corrections, the robot can successfully learn a safety constraint intended by the user.
+The above result asserts the maximum learning iteration to find $\boldsymbol{\theta}_{i}\in\Theta_{H}$.
+In other words, within $K$ rounds of human correction, the robot can successfully learn a safety constraint.
 The above result requires a pre-set of $\rho_{H}^{r}$, which is the radius of a ball contained in $\bar{\Theta}_{H}$.
-Since the size of $\bar{\Theta}_{H}$ is unknown in practice, $\rho_{H}>0$ can be interpreted as an underestimated size of the unknown user intent set $\bar{\Theta}_{H}$.
-It also serves as a termination threshold for learning convergence: a smaller $\rho_{H}$ can lead to a more concentrated of $\boldsymbol{\theta}_{H}$ learned at convergence.
-However, the finite iteration convergence in Theorem 1 comes with a prerequisite: $\bar{\Theta}_{H}\in\Theta_{0}$ has to be satisfied in the first place in the Safe MPC Alignment algorithm.
-Although such a condition can be achieved by empirically choosing $\Theta_{0}$ large enough to contain all possible $\bar{\Theta}_{H}$ (10), it is also desired to have a guaranteed method to detect the hypothesis misspecification, where such condition is not met.
-Next, we will discuss the declaration of hypothesis misspecification.
-V-C Result 2: Hypothesis Misspecification Declaration
-We consider the case of hypothesis space misspecification, where $\bar{\Theta}_{H}\cap\Theta_{0}=\emptyset$.
+Since the size of $\bar{\Theta}_{H}$ is unknown in reality, $\rho_{H}>0$ can be interpreted as an estimated size of the unknown user intent set $\bar{\Theta}_{H}$.
+It also serves as a termination threshold for learning convergence: a smaller $\rho_{H}$ can lead to a more concentrated $\boldsymbol{\theta}_{H}$ at convergence.
+The finite iteration convergence in Theorem 1 comes with a prerequisite: $\bar{\Theta}_{H}\in\Theta_{0}$ has to be satisfied in the first place for the Safe MPC Alignment.
+Although such a condition can be achieved by empirically choosing $\Theta_{0}$ large enough to contain all possible $\bar{\Theta}_{H}$ (10), it is also desired to have a verification for the hypothesis misspecification, where such condition is not met.
+We will provide such result in the next subsection.
+V-C Result 2: Hypothesis Misspecification Certification
+Consider hypothesis space misspecification, where $\bar{\Theta}_{H}\cap\Theta_{0}=\emptyset$.
 In this case, no $\boldsymbol{\theta}\in\Theta_{0}$ can account for the user intended safety (4).
-As shown in [52, 58], the hypothesis misspecification can lead to the undesired behavior of the robot.
-The following result states the proposed Safe MPC Alignment algorithm can automatically certify the misspecification of the initial hypothesis space.
+The hypothesis misspecification can lead to the undesired behavior of the robot [bobu2020quantifying, ghosal2023effect].
+The following result states the Safe MPC Alignment can automatically certify the misspecification of $\Theta_{0}$.
 Theorem 2 (Certifying misspecification).
-In the Safe MPC Alignment algorithm, $\boldsymbol{\theta}_{i}$ is selected as the center of MVE of $\Theta_{i-1}$, $i=1,2,3,...$.
+In the Safe MPC Alignment, $\boldsymbol{\theta}_{i}$ is selected as the center of MVE of $\Theta_{i-1}$, $i=1,2,3,$....
 If $\bar{\Theta}_{H}\cap\Theta_{0}=\emptyset$, the algorithm converges to a set $\Theta^{*}$ as $i\rightarrow\infty$, such that $\mathbf{Vol}(\Theta^{*})=0$ and $\Theta^{*}\cap\mathbf{int}\ \Theta_{0}=\emptyset$.
@@
 See Appendix A-F. ∎
-The above result states that when the initial hypothesis space is misspecified, $\bar{\Theta}_{H}\cap\Theta_{0}=\emptyset$, the Safe MPC Alignment algorithm will finally converge to a set of zero volume on the boundary of the initial hypothesis space $\Theta_{0}$.
+The above result states when the initial hypothesis space is misspecified, $\bar{\Theta}_{H}\cap\Theta_{0}=\emptyset$, the Safe MPC Alignment algorithm will finally converge to a set of zero volume on the boundary of the initial hypothesis space $\Theta_{0}$.
 The above theorem inspires us to empirically identify the misspecification by monitoring the position of the MVE center of $\Theta_{i}$ during the learning update.
 As shown in Fig.
-7, if after the maximum iteration $K-1$ in (24), the MVE center $\boldsymbol{\theta}_{K-1}$ is very close to the boundary of the initial hypothesis set $\Theta_{0}$, there is likely hypothesis space misspecification, and the algorithm terminates with misspecification failure.
-Figure 7: When there is misspecification, the Safe MPC Alignment algorithm will converge to some point on the boundary of $\Theta_{0}$.
+6, if after the maximum iteration K1K-1 in (24), the MVE center $\boldsymbol{\theta}_{K-1}$ is very close to the boundary of the initial hypothesis set $\Theta_{0}$, there is likely hypothesis space misspecification, and the algorithm terminates with misspecification failure.
+Figure 6: When there is misspecification, the Safe MPC Alignment will converge to some point on the boundary of $\Theta_{0}$.
 So we can use the distance of $\boldsymbol{\theta}_{K-1}$ to the boundary as an indicator for misspecification.
-Specifically, recall that $\Theta_{0}$ in (10) is a box with bounds $\boldsymbol{c}_{l},\boldsymbol{c}_{h}$, we can pre-define a small threshold $\epsilon$ to examine if $\boldsymbol{\theta}_{K-1}$ is on the boundary of $\Theta_{0}$.
+Specifically, if $\Theta_{0}$ is a box as defined in (10) with bounds $\boldsymbol{c}_{l},\boldsymbol{c}_{h}$, we can pre-define a small threshold $\epsilon$ to examine if $\boldsymbol{\theta}_{K-1}$ is on the boundary of $\Theta_{0}$.
 If $\boldsymbol{\theta}_{K-1}$ satisfies
@@
 V-D Detailed Algorithm of Safe MPC Alignment
-With the main results on convergence and misspecification, we can establish the full safe MPC alignment algorithm in Algorithm 1.
-With initializing $\Theta_{0}$ and pre-define a termination threshold $\rho_{H}$, three steps in the hypothesis space cutting are executed until the user satisfies robot motion or the iteration count reaches the maximum count $K$ in (24).
-Note that during the execution of the MPC policy, $\boldsymbol{\pi}_{\boldsymbol{\theta}_{i}}^{\gamma}$, the human can perform emergency stopping and reset the robot’s state, which corresponds to Section III-C3.
-After $K-1$ iterations, the algorithm checks the misspecification by (25).
-Initialization: Parameterized constraint $\boldsymbol{g}_{\theta}$, initial hypothesis space $\Theta_{0}$, termination threshold $\rho_{H}$ (i.e., approximate radius of $\bar{\Theta}_{H}$), $i=1$
-Compute maximum correction count $K$ from $\rho_{H}$ in (24);
+With the main results on convergence and misspecification, we present the detailed algorithm in Algorithm 1.
+With $\Theta_{0}$ and a termination threshold $\rho_{H}$ pre-set, three steps in the hypothesis space cutting are executed until the user satisfies robot motion or the iteration count reaches the maximum count $K$ in (24).
+Note that during the execution of the MPC policy $\boldsymbol{\pi}_{\boldsymbol{\theta}_{i}}^{\gamma}$, the human can perform emergency stopping and reset the robot’s state, which corresponds to Section III-C3.
+After K1K-1 iterations, the algorithm checks the misspecification by (25).
+Initialization: Initial hypothesis space $\Theta_{0}$, termination threshold $\rho_{H}$, $i=1$
+Compute maximum iteration count $K$ from $\rho_{H}$ in (24);
 while $i<K$ do
@@
  else if Human triggers emergency stop then
-             Human restarts robot from safe region.;
+     Human restarts robot from safe region;
  else if Human user is satisfied then
@@
 Algorithm 1 Safe MPC Alignment Algorithm
-VI Numerical Examples and Analysis
-In this section, we will evaluate the proposed Safe MPC Alignment using two numerical examples: inverted pendulum and two-link torque-controlled robot arm moving in constrained environment.
-Simulated human corrections knowing ground-truth safety constraints will be used.
-We will compare the proposed method with two baselines, and conduct ablation studies on the hyperparameters of the algorithm.
-All numerical experiments are implemented in Python, running on a Dell desktop with Intel i9-13900K CPU.
-All our experiment codes are open-sourced at at https://github.com/asu-iris/Safe-MPC-Alignment.
+VI Numerical Examples
+In this section, we will evaluate the Safe MPC Alignment using two numerical examples: A pendulum task, where we will illustrate all theoretical aspects of the proposed method, and a thrust-controlled quadrotor navigation task, where we will demonstrate the capability of the method to learn safety constraints in complex 3D environments.
+Simulated human correction knowing ground-truth safety constraints will be used.
+We will compare the proposed method with baselines, and ablate the hyperparameters of the algorithm.
+All numerical experiments are implemented in Python, running on a desktop with Intel i9-13900K CPU.
+All our experiment codes are open-sourced at https://github.com/asu-iris/Safe-MPC-Alignment.
+(a) Initial trajectory (with $\boldsymbol{\theta}_{1}$)
+(b) After 3 corrections (with $\boldsymbol{\theta}_{4}$)
+(c) After 6 corrections (with $\boldsymbol{\theta}_{7}$)
+(d) After 14 corrections (with $\boldsymbol{\theta}_{15}$)
+Figure 7: The Safe MPC rollout trajectory after different learning interaction.
+The learned constraint with $\boldsymbol{\theta}_{1}$ is omitted since it covers the whole first quadrant.
+As learning proceeds, the constraint boundary (orange line) converges to the ground-truth (red line).
+When the learning process terminates, the learned and ground-truth constraint boundaries almost overlap, suggesting the success of constraint learning from correction.
 VI-A Inverted Pendulum
@@
 with $\alpha$ being the angle between the pendulum and gravity, $\mathrm{g}$ the gravity constant, $u$ input torque, $m$, $l$ and $d$ the length, inertia, and damping ratio.
-We set the system state as $\boldsymbol{x}=[\alpha,\dot{\alpha}]^{\mkern-1.5mu\mathsf{T}}$ and discretize (26) by the Euler method with $\triangle t=0.02\mathrm{s}$.
-A zero-mean Gaussian noise with covariance $\Sigma=\mathbf{diag}(1,4)\times 10^{-5}$ is added to $\boldsymbol{x}$ at each control step to simulate the state noise.
-The task is to control the pendulum to the target $\boldsymbol{x}^{*}=[\pi,0]^{\mkern-1.5mu\mathsf{T}}$.
-We additionally pose an unknown global state constraint on pendulum motion:
+The system state is $\boldsymbol{x}=[\alpha,\dot{\alpha}]^{\mkern-1.5mu\mathsf{T}}$, (26) discretized by the Euler method with $\triangle t=0.02\mathrm{s}$.
+A zero-mean Gaussian noise with covariance $\Sigma=\mathbf{diag}(1,4)\times 10^{-5}$ is added to $\boldsymbol{x}$ at each MPC rollout to simulate the state noise.
+The task is to control the pendulum to a target $\boldsymbol{x}^{*}=[\pi,0]^{\mkern-1.5mu\mathsf{T}}$.
+We pose an unknown global state constraint on pendulum motion:
 $0.6\alpha+\dot{\alpha}=\boldsymbol{\theta}^{H}\boldsymbol{x}\leq 3,$ (27)
@@
 respectively, with $R=0.1$ and $Q=\mathrm{diag}(25,10)$.
-To ensure the safe MPC policy to satisfy a global constraint, it is sufficient to learn a safety constraint defined on the closest controllable state.
-Thus, in (1), we define
+To ensure the safe MPC to satisfy a global constraint, it is sufficient to learn a safety constraint defined on the closest controllable state.
+Thus, we define
 ${g}_{\boldsymbol{\theta}}(\boldsymbol{\xi})=-3+\boldsymbol{\theta}^{\mkern-1.5mu\mathsf{T}}\boldsymbol{x}_{1}.$ (29)
@@
 Suppose the user intent set $\bar{\Theta}_{H}=\{\boldsymbol{\theta}|\left\lVert\boldsymbol{\theta}-\boldsymbol{\theta}_{H}\right\rVert\leq 0.02\}$.
-The simulated human corrections $\mathbf{a}$ are generated as follows: when the system state is near the safety boundary, i.e., ${g}_{\boldsymbol{\theta}_{H}}(\boldsymbol{\xi}_{\boldsymbol{\theta}}^{\gamma})>-\epsilon_{g}$ ($\epsilon_{g}=0.25$ here), a directional correction $\mathbf{a}$ is be generated in probability $p=0.3$ using $\mathrm{sign}(-\nabla B(\boldsymbol{\xi}_{\boldsymbol{\theta}}^{\gamma},\boldsymbol{\theta}_{H}))$ (from which only the first entry is taken considering (6)).
+The simulated (synthetic) correction $\mathbf{a}$ are generated as follows: when the system state is near the safety boundary, i.e., ${g}_{\boldsymbol{\theta}_{H}}(\boldsymbol{\xi}_{\boldsymbol{\theta}}^{\gamma})>-\epsilon_{g}$ ($\epsilon_{g}=0.25$ here), a directional correction $\mathbf{a}$ is be generated in probability $p=0.3$ using $\mathrm{sign}(-\nabla B(\boldsymbol{\xi}_{\boldsymbol{\theta}}^{\gamma},\boldsymbol{\theta}_{H}))$ (from which only the first step is taken as in (6)).
 The correction is probabilistic to simulate the randomness of human providing corrections.
@@
 VI-A2 Convergence Analysis
+(a) Logarithm MVE Volume
+(b) Parameter Estimation Error
+Figure 8: (a) The plot of $\log\det\bar{H}$ with respect to correction numbers.
+In all algorithm runs (different colors), the plot shows the near-linear decrease of the logarithm volume, suggesting the exponential decrease of the volume of MVE of the hypothesis spaces, as asserted in Theorem 1.
+(b) The plot of the parameter estimation error $\left\lVert\boldsymbol{\theta}_{i}-\boldsymbol{\theta}_{H}\right\rVert_{2}$ versus correction numbers.
+In all runs, the parameter estimation error can converge to a value lower than 0.02 within 17 corrections, consistent with the theoretical bound in Theorem 1.
 We first evaluate the convergence of our algorithm without hypothesis misspecification.
@@
 The initial $\boldsymbol{\theta}_{1}=[-2,-2]^{\mkern-1.5mu\mathsf{T}}$ (i.e., the center of $\Theta_{0}$).
-With each simulated human correction, we track $\boldsymbol{\theta}_{i}$ in each learning update.
+With each simulated correction, we track $\boldsymbol{\theta}_{i}$ in each learning update.
 Each update of ${\Theta}_{i}$ takes less than 30ms.
@@
 The MPC rollout trajectory (phase plot) with each $\boldsymbol{\theta}_{i}$ is shown in Fig.
+7.
+The ground-truth constraint (27) and the learned constraints $\boldsymbol{g}_{\boldsymbol{\theta}_{i}}$ are also plotted.
+Fig.
+7 shows that the true safety constraint can be successfully learned from simulated correction.
+To analyze the convergence, we run the algorithm for 10 times.
+In each run, we keep track of two metrics over corrections (iterations) and show them in Fig.
 8.
-The ground-truth constraint (27) and the learned constraints $\boldsymbol{g}_{\boldsymbol{\theta}_{i}}$ are also plotted in Fig.
-8.
-From Fig.
-8, we can conclude that the ground-truth constraint can be successfully learned from simulated corrections.
-(a) Initial trajectory (with $\boldsymbol{\theta}_{1}$)
-(b) After 3 corrections (with $\boldsymbol{\theta}_{4}$)
-(c) After 6 corrections (with $\boldsymbol{\theta}_{7}$)
-(d) After 14 corrections (with $\boldsymbol{\theta}_{15}$)
-Figure 8: The pendulum rollout trajectory at different correction interactions: (a) trajectory with $\boldsymbol{\theta}_{1}$, (b) with $\boldsymbol{\theta}_{4}$, (c) with $\boldsymbol{\theta}_{7}$, (d) with $\boldsymbol{\theta}_{15}$.
-The learned constraint with $\boldsymbol{\theta}_{1}$ is omitted since it covers the whole first quadrant.
-As the learning proceeds, the learned constraint boundary (orange line) converges to the ground-truth constraint boundary (red line).
-When the learning process terminates, the learned and ground-truth constraint boundaries almost overlap, suggesting the success of constraint learning from corrections.
-To further analyze the learning convergence, we run the algorithm for 10 times.
-In each run, we keep track of two metrics over correction (iteration) count and show them in Fig.
-9.
-The first metric (left panel) is the logarithm volume measure of the solved MVE $\log\det\bar{H}$ in (20); this provides a lower bound of $\Theta_{i}$ to show the exponential convergence property.
+The first metric (left panel) is the logarithm volume measure of the solved MVE $\log\det\bar{H}$ in (20); this provides a lower bound of the volume of $\Theta_{i}$ to show the exponential convergence.
 The second (right panel) is the parameter error $\left\lVert\boldsymbol{\theta}_{i}-\boldsymbol{\theta}_{H}\right\rVert_{2}$.
 From Fig.
-9, we can conclude the proposed method is
+8, we can conclude the proposed method is
   • (I)
-    Fast-converging. The logarithm volume $\log\det\bar{H}$ has near-linear decreasing speed. This is consistent with the exponential convergence in theory stated in Lemma 5.
+    Fast-converging. The logarithm volume $\log\det\bar{H}$ has a near-linear decreasing speed. This is consistent with the theory in Lemma 5.
   • (II)
-    Correction-efficient.
+    Data-efficient.
 $\left\lVert\boldsymbol{\theta}_{i}-\boldsymbol{\theta}_{H}\right\rVert_{2}$ converges to the value lower than 0.02 within 17 corrections.
-This is consistent with the theoretical upper bound of correction count from given by Theorem 1 :
+This is consistent with the theoretical upper bound of correction count in Theorem 1:
     ${\ln\frac{\tau_{r}\rho_{H}^{r}}{\mathbf{Vol}(\Theta_{0})}}/{\ln(1-\frac{1}{r})}=17.28.$ (30)
-(a) Logarithm MVE Volume
-(b) Parameter Estimation Error
-Figure 9: (a) The plot of $\log\det\bar{H}$ with respect to correction numbers.
-In all algorithm runs, the plot shows the near-linear decrease of the logarithm volume, suggesting the exponential decrease of the volume of MVE of the hypothesis spaces, as asserted in Theorem 1.
-(b) The plot of the parameter estimation error $\left\lVert\boldsymbol{\theta}_{i}-\boldsymbol{\theta}_{H}\right\rVert_{2}$ versus correction numbers.
-In all runs, the parameter estimation error can converge to a value lower than 0.02 within 17 corrections.
-VI-A3 Analysis for Misspecification Case
+VI-A3 Misspecification Analysis
+Figure 9: Experiment of misspecification case.
+The blue dashed lines are the boundary of $\Theta_{0}$.
+Path of $\theta$ convergence, $\boldsymbol{\theta}_{i},\ i=1,\dots,7$, is shown in orange dots with arrows showing the update direction.
+The ground-truth intent set is marked as the red star.
+Since misspecification happens for this case, $\boldsymbol{\theta}$ will converge toward the boundary of $\Theta_{0}$ (the green triangle).
+This is consistent with the statement in Theorem 2.
 To analyze the performance of our algorithm to certify misspecification cases, we set the initial hypothesis space $\Theta_{0}$ as ${\boldsymbol{c}_{l}}=[-1,-1]^{\mkern-1.5mu\mathsf{T}}$ and ${\boldsymbol{c}_{u}}=[0.8,0.8]^{\mkern-1.5mu\mathsf{T}}$ which does not contain $\bar{\Theta}_{H}$.
@@
 The path of $\boldsymbol{\theta}_{i},\ i=1,\dots,8$ is shown in Fig.
+9.
+$\boldsymbol{\theta}_{i}$ is continuously moving towards the boundary of $\Theta_{0}$.
+This fact verifies Theorem 2 and justifies the proposed method is capable of misspecification detection.
+VI-B Quadrotor Navigation in Complex Environment
+Figure 10: Quadrotor navigation in an 3D zigzag tube.
+VI-B1 Setup
+We consider a thrust-controlled quadrotor navigating in an 3D zigzag tube, shown in Fig.
 10.
-In Fig.
-10, $\boldsymbol{\theta}_{i}$ is continuously moving towards the boundary of $\Theta_{0}$.
-This fact verifies Theorem 2 and justifies the proposed method is capable of misspecification detection.
-Figure 10: Illustration of misspecification case.
-The blue dashed lines are the boundary of $\Theta_{0}$.
-$\boldsymbol{\theta}_{i},\ i=1,\dots,7$ is shown in he orange dots in one run.
-Black arrows show the update directions of $\boldsymbol{\theta}_{i}$s.
-The center of the ground-truth intent set is marked as the red star, and the $\boldsymbol{\theta}_{8}$ at the termination of learning is marked as the green triangle.
-When misspecification happens, $\boldsymbol{\theta}$ will converge toward the boundary of $\Theta_{0}$, as stated by Theorem 2.
-VI-B Two-link Robot Arm Reaching with Complex Constraint
-VI-B1 Setup
-The dynamics if torque controlled two-link robot arm moving in the horizontal plane is $M(\boldsymbol{q})\ddot{\boldsymbol{q}}+\boldsymbol{c}(\boldsymbol{q},\dot{\boldsymbol{q}})=\boldsymbol{\tau}$, where $M(\boldsymbol{q})$ is the inertia matrix; $\boldsymbol{c}(\boldsymbol{q},\dot{\boldsymbol{q}})$ is the Coriolis term.
-$\boldsymbol{q}=[q_{1},q_{2}]^{\mkern-1.5mu\mathsf{T}}$ defines the two joint angles of the robot.
-The system state is defined as $\boldsymbol{x}=[\boldsymbol{q},\dot{\boldsymbol{q}}]$.
-The control input $\boldsymbol{u}$ is the joint torque $\boldsymbol{\tau}=[\tau_{1},\tau_{2}]^{\mkern-1.5mu\mathsf{T}}$, i.e.
-$\boldsymbol{u}=\boldsymbol{\tau}$.
-The other detailed physical parameters are given in Appendix.
-As shown in Fig.
-12, the robot starts at $\boldsymbol{x}_{\mathrm{init}}=[-\frac{\pi}{2},0,0,0]^{\mkern-1.5mu\mathsf{T}}$ and the task is to reach a target $\boldsymbol{x}^{*}=[\frac{\pi}{2},0,0,0]^{\mkern-1.5mu\mathsf{T}}$.
-In Safe MPC (1), we set $T=30$, the step cost functions
-$\begin{gathered}c(\boldsymbol{x}_{t},\boldsymbol{u}_{t})=\boldsymbol{x}_{t}^{\mkern-1.5mu\mathsf{T}}P\boldsymbol{x}_{t}+\boldsymbol{u}_{t}^{\mkern-1.5mu\mathsf{T}}R\boldsymbol{u}_{t},\\ h(\boldsymbol{x}_{T})=(\boldsymbol{x}_{T}-\boldsymbol{x}^{*})^{T}Q(\boldsymbol{x}_{T}-\boldsymbol{x}^{*}),\end{gathered}$ (31)
-with $P=\mathrm{diag}(1,1,2,2)$, $R=\mathrm{diag}(0.1,0.1)$ and $Q=\mathrm{diag}(5,5,10,10)$.
-As shown by the black lines (obstacles) in Fig.
-12, we define a complex reaching environment with non-convex obstacles and narrow corridor.
-The ground-truth global safety constraint corresponding to Fig.
-12 is given
-${g}_{\boldsymbol{\theta}_{H}}(\boldsymbol{\xi})=-1+\boldsymbol{\theta}_{H}^{\mkern-1.5mu\mathsf{T}}\boldsymbol{\phi}$ (32)
-with $\boldsymbol{\phi}{=}[\mathrm{tanh}(x_{e}y_{e}),\mathrm{tanh}(x_{e}^{2},\mathrm{tanh}(y_{e}^{2}),\mathrm{tanh}(x_{e}),\mathrm{tanh}(y_{e})]^{\mkern-1.5mu\mathsf{T}}$, $\boldsymbol{\theta}_{H}=[0.78,0.91,-0.52,0.52,-1.04]^{\mkern-1.5mu\mathsf{T}}$, and $[x_{e},y_{e}]^{\mkern-1.5mu\mathsf{T}}$ is the end-effector position of the robot arm.
-The human intent set is $\bar{\Theta}_{H}{=}\{\boldsymbol{\theta}|\left\lVert\boldsymbol{\theta}-\boldsymbol{\theta}_{H}\right\rVert\leq 0.2\}$.
-We set the initial hypothesis $\Theta_{0}$ with box boundaries $\boldsymbol{c}_{l}=[-3,\dots,-3]^{\mkern-1.5mu\mathsf{T}}$ and $\boldsymbol{c}_{u}=[3,\dots,3]^{\mkern-1.5mu\mathsf{T}}$.
-The simulated human corrections are generated in the similar way to that in Section VI-A.
-Figure 11: The heatmap of a learned constraint $g$. The learned and ground-truth constraints are almost identical, showing the success of our method for learning complex constraint from corrections.
+The simulation environment is based on MuJoCo [todorov2012mujoco], and the quadrotor dynamics follows [jin2022learning].
+Initially, the quadrotor has no knowledge of the safety constraint.
+When it nears the tube wall, a simulated (synthetic) correction is generated to push it away from the tube wall.
+Details of the setting are provided in Appendix B-B.
+To learn positional safety constraints, we experiment with three feature vectors $\boldsymbol{\phi}(\boldsymbol{\xi})$ in (5):
+Polynomial feature vector
+A vector of polynomial basis functions up to 3rd order to properly capture the zig-zag pattern of the tube) is defined in the quadrotors’ 3D position $[x,y,z]$ at the 5th timestep in motion plan $\boldsymbol{\xi}$:
+$\boldsymbol{\phi}(\boldsymbol{\xi}){=}[x_{5}^{3},y_{5}^{3},z_{5}^{3},x_{5}^{2},y_{5}^{2},\\ z_{5}^{2},x_{5},y_{5},z_{5},x_{5}y_{5},y_{5}z_{5},x_{5}z_{5}]^{\top}.$ (31)
+The features are defined using the quadrotor’s 3D position at the 5th prediction step.
+The hypothesis space is initialized as a box space with $\boldsymbol{c}_{l}=[-80,\dots,-80]^{\mkern-1.5mu\mathsf{T}}$ and $\boldsymbol{c}_{u}=[200,\dots,200]^{\mkern-1.5mu\mathsf{T}}$.
+We select $\phi_{0}=1$ and $\gamma=60$.
+Pre-trained neural feature vector
+We obtain the neural spacial features from a pre-trained Signed Distance Function (SDF) [park2019deepsdf].
+Specifically, we pre-train an SDF neural network using the point cloud data samples near the tube surface.
+This simulates the application where some geometric information of the environment is available.
+The neural SDF is pre-trained in a regression-like manner, with distance, normal, and Eikonal losses (see details in Appendix B-B).
+With the pretrained SDF function, we take the 16-dimensional output of the second-to-last layer of the SDF network as our feature vector $\boldsymbol{\phi}(\boldsymbol{\xi})$.
+The input to $\boldsymbol{\phi}(\boldsymbol{\xi})$ is the quadrotor’s 3D position $[x,y,z]$ at the 5th perdition step.
+Such a choice of feature vector $\boldsymbol{\phi}(\boldsymbol{\xi})$ means that we will use the proposed Safe MPC Alignment to fine tune the last layer of the learned SDF, $g(\boldsymbol{\xi})=\boldsymbol{\theta}^{\mkern-1.5mu\mathsf{T}}\boldsymbol{\phi}(\boldsymbol{\xi})$, for its safe navigation control.
+To facilitate the real-time Safe MPC with the neural safety constraint $g(\boldsymbol{\xi})=\boldsymbol{\theta}^{\mkern-1.5mu\mathsf{T}}\boldsymbol{\phi}(\boldsymbol{\xi})$, the optimization (1) uses the linearized $g(\boldsymbol{\xi})$ for calculation, which is linearized at the real robot state $\boldsymbol{x}^{\text{real}}_{k}$.
+The initial bounds of the hypothesis space are set to $\boldsymbol{c}_{l}=[-0.2,\dots,-0.2]^{\mkern-1.5mu\mathsf{T}}$, $\boldsymbol{c}_{u}=[0.1,\dots,0.1]^{\mkern-1.5mu\mathsf{T}}$.
+We select $\phi_{0}=0.1$ and $\gamma=50$.
+To ensure the feasibility of the learned constraint, in the above feature vectors, two additional cuts are added in each MVE update to enforce that the start and goal remain within the intended safe region.
+The task is successful if the drone reaches the goal without triggering this stop.
+Polynomial features with velocity features
+We consider the safety constraints are additionally defined in the velocity space.
+After learning the positional constraint with the polynomial feature vector, a new set of velocity-based features over all prediction steps, $\boldsymbol{\phi}(\boldsymbol{\xi})=[\|\boldsymbol{v}_{1}\|^{2},\dots,\|\boldsymbol{v}_{10}\|^{2}]^{\top}$, is added.
+In this case, we will freeze the weights for polynomial features, and only learn the weights for velocity features.
+Our objective is to impose an additional constraint that caps the quadrotor’s velocity magnitude below 0.45 [m/s], while ensuring that the previously learned positional constraint remains satisfied for collision avoidance.
+When the quadrotor’s velocity exceeds a correction triggering threshold (0.4), a simulated correction is applied to push the quadrotor in the direction of negative velocity.
+The hypothesis space for the velocity features is initialized with bounds $\boldsymbol{c}_{l}=[0,\dots,0]^{\top}$ and $\boldsymbol{c}_{u}=[0.16,\dots,0.16]^{\top}$.
+We set the bias feature $\phi_{0}=1$ and select $\gamma=100$.
+Figure 11: The Safe MPC rollout trajectories (in green) for learning safety constraints with polynomial features after 26, 38, 45 corrections (different columns).
+The first row and second row show the top and the side views of the navigation environment.
+Figure 12: The Safe MPC rollout trajectories (in green) for learning safety constraints with neural features after 6, 9, 11 corrections (different columns).
+The first row show the top view, and second row shows the side view of the navigation environment.
 VI-B2 Result
-Our method can successfully learn the safety constraint within 25-30 simulated corrections.
-One instance of the learned constraint is shown in Fig.
-11.
-The comparison of the robot reaching motion without and with the learned safety constraint is shown in Fig.
-12.
-The robot is capable of operating in the intended safe region after the algorithm converges.
-Those results show that the efficacy of the proposed method for solving system with complex safety constraints.
-(a) Trajectory Before Corrections
-(b) Trajectory After Corrections
-Figure 12: The trajectory of the two-link robot arm (a) without and (b) with learned safety constraint. The robot successfully achieves the task with the learned safety constraint.
+Fig.
+11 and Fig.
+12 show the rollout trajectories of the learned Safe MPC policy using polynomial and neural features, respectively.
+With polynomial features, the task is completed in 45 corrections, while the neural features require only 11.
+This is because the neural features, constructed from the second-to-last layer of pre-trained neural SDF, have already encoded the partial knowledge of spatial constraints.
+Figure 13: Velocity profile during the first 100 timesteps (acceleration phase) of the drone trajectory.
+After 12 corrections, the quadrotor successfully learns to stay below the intended velocity cap of 0.45 (dashed line).
+For the learning of velocity constraint, Fig.
+13 shows the velocity profile of the quadrotor during the Safe MPC rollout in the 100 time steps (acceleration phase) after different number of simulated corrections (different colors).
+After 12 corrections, the quadrotor successfully learns to stay below the intended velocity cap of 0.45 (dashed line).
+Throughout this process, the positional safety constraint remains satisfied, with the quadrotor continuing to navigate within the tube to the target.
+This result demonstrates the flexibility of the proposed method in scenarios involving both positional and velocity constraints.
 VI-C Baseline Comparison
-In this session, we compare our method with two baseline methods using magnitude corrections in the pendulum environment (the task and model settings are identical to Section VI-A).
-As reviewed previously, the direct literature learning safety constraints from human feedback is sparse to the authors’ best knowledge.
-Thus, the chosen baselines are an adaptation from the most relevant methods.
-Baseline 1: gradient matching.
-This baseline is inspired by [59], where raw human correction (with magnitude) is directly used as the gradient direction.
-In a similar spirit, when a human feedback $\mathbf{a}$ is made on robot trajectory $\xi_{\boldsymbol{\theta}}^{\gamma}$, we directly use the human magnitude correction as the true gradient label to learn the constraint $\boldsymbol{\theta}$.
+In this session, we compare our method with two baseline methods using magnitude corrections in the pendulum environment.
+Since the literature on directly learning safety constraints from human feedback is sparse, the chosen baselines are an adaptation from the most relevant methods.
+VI-C1 Baseline 1: Gradient Matching
+This baseline is inspired by [losey2019learning], where raw human correction (with magnitude) is directly used as the gradient direction.
+In a similar spirit, when a human correction $\mathbf{a}$ is made on robot motion $\xi_{\boldsymbol{\theta}}^{\gamma}$, we directly use the human correction as the true gradient label to learn the constraint $\boldsymbol{\theta}$.
 This leads to a learning loss
-$\mathcal{L}=\left\lVert\mathbf{a}+\nabla B(\boldsymbol{\xi}_{\boldsymbol{\theta}}^{\gamma})\right\rVert^{2}$ (33)
-Baseline 2: maximum likelihood learning.
-This baseline follows [60], where it views $e^{{g}_{\boldsymbol{\theta}}(\boldsymbol{\xi})}$ as a probability of human applying the "stop-feedback" to robot motion $\boldsymbol{\xi}$, when the robot reaches in a $\epsilon_{g}$-proximity of the safety boundary, i.e., when ${g}_{\boldsymbol{\theta}_{H}}(\boldsymbol{\xi}_{\boldsymbol{\theta}}^{\gamma})>-\epsilon_{g}$ Thus, a dataset of trajectory-correction paris, i.e., $\mathcal{D}=\{\mathbf{a}_{i},\boldsymbol{\xi}_{i}\}_{i=0}^{N}$, can be used as the training data to learn $\boldsymbol{\theta}$ using maximum likelihood, while also penalizing positivity of ${g}_{\boldsymbol{\theta}}(\boldsymbol{\xi}_{i})$ as a regulizer term.
+$\mathcal{L}(\boldsymbol{\theta})=\left\lVert\mathbf{a}+\nabla B(\boldsymbol{\xi}_{\boldsymbol{\theta}}^{\gamma},\boldsymbol{\theta})\right\rVert^{2}.$ (32)
+VI-C2 Baseline 2: Maximum Likelihood
+.
+This baseline follows [poletti2023learning], where it views $e^{{g}_{\boldsymbol{\theta}}(\boldsymbol{\xi})}$ as a probability of human applying the "stop-feedback" to robot motion $\boldsymbol{\xi}$, when the robot reaches in a $\epsilon_{g}$-proximity of the safety boundary, i.e., when ${g}_{\boldsymbol{\theta}_{H}}(\boldsymbol{\xi}_{\boldsymbol{\theta}}^{\gamma})>-\epsilon_{g}$ .
+Thus, a dataset of trajectory-correction paris, i.e., $\mathcal{D}=\{\mathbf{a}_{i},\boldsymbol{\xi}_{i}\}_{i=0}^{N}$, can be used as the training data to learn $\boldsymbol{\theta}$ using maximum likelihood, while also penalizing positivity of ${g}_{\boldsymbol{\theta}}(\boldsymbol{\xi}_{i})$ as a regulizer term.
 This leads to the following loss
-$\mathcal{L}=\sum\nolimits_{i=1}^{N}\big{(}-{g}_{\boldsymbol{\theta}}(\boldsymbol{\xi}_{i})+\alpha\cdot\mathrm{softplus}(g_{\boldsymbol{\theta}}(\boldsymbol{\xi}_{i}))\big{)}$ (34)
-In our comparison, we choose $N=15$ (similar to correction number needed for our method to converge) and $\alpha=5$ for its best performance.
-The comparison between our method and Baseline 1 focuses on data efficiency.
+$\mathcal{L}(\boldsymbol{\theta})=\sum\nolimits_{i=1}^{N}\big(-{g}_{\boldsymbol{\theta}}(\boldsymbol{\xi}_{i})+\alpha\cdot\mathrm{softplus}(g_{\boldsymbol{\theta}}(\boldsymbol{\xi}_{i}))\big).$ (33)
+We choose $N=15$ (similar to the correction number needed for our method to converge) and $\alpha=5$ for its best performance.
+The comparison between our method and the gradient matching baseline focuses on data efficiency.
 The convergence criteria of $\left\lVert\boldsymbol{\theta}-\boldsymbol{\theta}_{H}\right\rVert\leq 0.02$ is set for both methods.
-50 independent runs are performed and the number of correction needed for convergence are counted.
-We use Adam optimizer [61] with learning rate 0.02.
+50 independent runs are performed and the total number of correction at convergence are counted.
+We use Adam optimizer [diederik2014adam] with a learning rate of 0.02.
 We summarize the results in Table I.
-The parameter estimation error $\left\lVert\boldsymbol{\theta}-\boldsymbol{\theta}_{H}\right\rVert$ versus learning iteration (correction count) for a best run (i.e., best performance for the baseline) is plotted in Fig.
-13(a).
-The comparison results demonstrate the significant data efficiency of the proposed method compared to gradient matching baseline, even though the latter uses both magnitude and direction information of corrections to learn.
+The parameter error $\left\lVert\boldsymbol{\theta}-\boldsymbol{\theta}_{H}\right\rVert$ versus learning iteration for a best run (for the baseline) is plotted in Fig.
+14(a).
+The comparison results demonstrate data efficiency of the proposed method compared to the gradient matching baseline, even though the latter uses both magnitude and direction information of corrections to learn.
 TABLE I: Comparison of Safe MPC Alignment and Baseline I
@@
 Gradient Matching $240.38\pm 120.70$
-(a) Comparison with Baseline 1
-(b) Comparison with Baseline 2
-Figure 13: (a) Comparison with gradient matching baseline: the parameter estimation error $\left\lVert\boldsymbol{\theta}-\boldsymbol{\theta}_{H}\right\rVert$ versus learning iteration (correction count).
-This is best run (among 50 runs) with the least corrections for the baseline.
-Our method converges after 12 corrections, while the baseline needs more than 160 corrections to converge.
-(b) Comparison with the maximum likelihood baseline under the different $\epsilon_{g}$.
-In the plot, the y-axis is $\left\lVert\boldsymbol{\theta}-\boldsymbol{\theta}_{H}\right\rVert$ at convergence and the error bar denotes the standard deviation.
-The maximum likelihood baseline fails to learn the correct weights when $\epsilon_{g}$ is large, while our method consistently shows high learning accuracy.
-The comparison between our method and maximum likelihood baseline focuses on the varying positions of simulated correction, i.e., varying $\epsilon_{g}$.
-Recall that the simulated corrections will be generated only when ${g}_{\boldsymbol{\theta}_{H}}(\boldsymbol{\xi}_{\boldsymbol{\theta}}^{\gamma})>-\epsilon_{g}$.
-The value of $\epsilon_{g}$ determines the position of the correction applied.
+(a) Comparison: gradient matching
+(b) Comparison: maximum likelihood
+Figure 14: (a) Comparison with gradient matching baseline: the parameter error $\left\lVert\boldsymbol{\theta}-\boldsymbol{\theta}_{H}\right\rVert$ versus learning iteration.
+This is best run with the fewest corrections for the baseline.
+Our method converges after 12 corrections, while the baseline needs more than 160 corrections.
+(b) Comparison with the maximum likelihood baseline under the different $\epsilon_{g}$: the y-axis is $\left\lVert\boldsymbol{\theta}-\boldsymbol{\theta}_{H}\right\rVert$ at convergence with the error bar denoting the standard deviation.
+The maximum likelihood baseline fails to learn the correct weights when $\epsilon_{g}$ is large.
+The comparison with maximum likelihood baseline uses the varying positions of simulated correction, i.e., varying $\epsilon_{g}$.
+Recall that the simulated corrections are generated only when ${g}_{\boldsymbol{\theta}_{H}}(\boldsymbol{\xi}_{\boldsymbol{\theta}}^{\gamma})>-\epsilon_{g}$.
+The value of $\epsilon_{g}$ determines the position of correction.
 In our comparison, we vary $\epsilon_{g}$ from 0.1 to 0.9 and track the converged parameter error for both methods.
 As shown in Fig.
-13(b), safe MPC alignment consistently converges to the ground-truth $\boldsymbol{\theta}_{H}$ regardless of $\epsilon_{g}$, while the maximum likelihood baseline fails when $|\epsilon_{g}|$ is large.
-The results suggest the robustness of the proposed method against different correcting conditions compared to maximum likelihood baseline method.
+14(b), Safe MPC alignment consistently converges to the true $\boldsymbol{\theta}_{H}$ regardless of $\epsilon_{g}$, while the maximum likelihood baseline fails when $|\epsilon_{g}|$ is large.
+The results suggest the robustness of the proposed method against different correcting conditions compared to the maximum likelihood baseline.
 VI-D Ablation Study
-In this session, we conduct two ablation studies to show the performance of our algorithm given different design parameters/aspects.
-The task is still based on pendulum environment with settings identical to Section VI-A).
+We conduct two ablation studies to show the performance of our algorithm given different design aspects. The task is still based on the pendulum task in Section VI-A.
 VI-D1 Choosing $\boldsymbol{\theta}_{i}$: MVE centering versus random sampling
-In the first study, we compare two strategies in selecting $\boldsymbol{\theta}_{i}$ from the hypothesis space $\Theta_{i-1}$ in our algorithm: MVE centering (used in our method) and random sampling.
-For the latter, $\boldsymbol{\theta}_{i}$ is chosen by sampling from the uniform distribution supported by $\Theta_{i-1}$.
-We compare the total number of corrections needed at learning convergence under the two strategies.
-With 50 independent runs, we observe that the MVE centering requires $11.76\pm 2.14$ corrections, while random sampling requires $11.88\pm 3.50$ corrections.
+We compare two strategies in selecting $\boldsymbol{\theta}_{i}$ from the hypothesis space $\Theta_{i-1}$: MVE centering (used in our method) and random sampling.
+For the latter, $\boldsymbol{\theta}_{i}$ is sampled from the uniform distribution supported by $\Theta_{i-1}$.
+We compare the total number of corrections at learning convergence under the two strategies.
+With 50 independent runs, we observe the MVE centering requires $11.76\pm 2.14$ corrections, while random sampling $11.88\pm 3.50$ corrections.
 While the mean value of the correction count looks similar, the random sampling shows a slightly larger variance.
-This is understandable given the sampling randomness.
-We also observed that sampling process introduced large computational overhead, which is around 25% time slower than MVE centering.
-In contrast, the MVE centering used in our algorithm shows low variance and computational efficiency in finding $\boldsymbol{\theta}_{i}$.
-Because selecting $\boldsymbol{\theta}_{i}$ is casted as a convex programming for MVE centering, such advantage would be more significant for high-dimensional parameter space.
+This is understandable given sampling randomness.
+We also observed that the sampling process introduced a large computational overhead, which is around 25% time slower than MVE centering.
+In contrast, the MVE centering used in our algorithm shows low variance and computational efficiency.
+Because selecting $\boldsymbol{\theta}_{i}$ is cast as a convex programming for MVE centering, such an advantage would be more significant for a high-dimensional parameter space.
 VI-D2 Choosing $\gamma$
-In the second study, we investigate how the setting of $\gamma$ affects the learning performance.
-We vary $\gamma$ from 0.01 to 10.
-Fig.
-14 shows the total number of corrections needed at convergence versus different $\gamma$s.
-The result in Fig.
-14 shows that the learning performance remains the same given different settings of $\gamma$.
-Since $\gamma$ is weights parameter to control the safety conservativeness in the penalty-based MPC policy in (3), it means our method can successfully learn a safety constraint when the robot executes with safe MPC policy of different conservativeness levels.
+[…]
 Figure 15: The total number of corrections needed for convergence under different $\gamma$, the x-axis is of log scale.
-Safe MPC alignment can converge with a similar number of corrections (iterations) under different levels of $\gamma$.
 VII User Study in Simulation Games
-In this section, we test the Safe MPC Alignment method in two real-human user studies using two simulation games developed based on MuJoCo [62].
+We conduct two human user studies to evaluate the Safe MPC Alignment.
 One is the drone navigation game and the other the Franka robot arm reaching game.
-In both games, the robot is controlled by an MPC controller and human participants provide real-time corrections through a keyboard.
-The objective of each game is to teach a robot to successfully learn a safe MPC to move through environments with unknown obstacles.
-In following, Subsection VII-A describes the basic setting of the two games.
-Subsection VII-B details the user study procedure and the recruitment of human participants.
-The results and analysis are given in VII-C.
-VII-A Setting of Two Simulation Games
-In both user studies, human participants are asked to give correction to the following two physics-based simulation games developed based on MuJoCo.
+In both games, the robot is controlled by an MPC policy and human participants provide real-time corrections through a keyboard.
+The objective of each game is to teach a robot to successfully learn Safe MPC to move through environments with obstacles.
+Section VII-A describes the settings.
+Section VII-B details the user study procedure.
+The results are given in Section VII-C.
+VII-A Two Simulation Games
+In both user studies, human participants are asked to give correction to the following two physics-based simulation games developed based on MuJoCo [todorov2012mujoco].
 (1) Drone Navigation Game.
 As shown in Fig.
-15, the goal of this game is to let a drone to successfully navigate through a narrow gate (yellow rectangle) whose position is unknown, reaching different targets (shown in yellow points behind the gate).
+16(a), the goal of this game is to let a drone to successfully navigate through a narrow gate (yellow rectangle), reaching different targets (shown in yellow points behind the gate).
 The safety constraint $\boldsymbol{\theta}^{\mkern-1.5mu\mathsf{T}}\boldsymbol{\phi}$ is parameterized using a linear combination of spacial radial basis functions (RBFs), defined on y-z plane at $x=10$.
-The drone’s control inputs are the thrusts of four rotors and human correction is applied via keyboard, with specific keys mapping to specific correction commands to the drone’s control space.
-See the detailed task setup and keyboard interface in Appendix B-C.
+The drone’s control inputs are the thrusts of four rotors and human correction is applied via keyboard, with specific keys mapping to the drone’s control space.
+See the details in Appendix B-C.
 (2) Franka arm reaching game.
-As in Fig.
-16, the goal of the Franka arm reaching game is to teach a Franka arm such that its gripper hand successfully reaches different target positions (cyan ball) through an unknown narrow slot (two yellow bars).
-Note that the safety constraint is only applied on the end effector and the collision between the link and the slot is ignored.
+As shown in Fig.
+16(b), the goal of the Franka arm reaching game is to teach a Franka arm such that its gripper hand successfully reaches different target positions (cyan ball) through an narrow slot (two yellow bars).
+Note that the safety constraint is only applied to the end effector and the collision between the link and the slot is ignored.
 The safety constraint function $\boldsymbol{\theta}^{\mkern-1.5mu\mathsf{T}}\boldsymbol{\phi}$ is parameterized as a weighted combination of RBFs defined on the $z$ axis and rotations around the $x$ axis.
-Human corrections are applied via keyboard, with specific keys mapping to specific correction commands applied to the robot control space $\mathbb{R}^{6}$ (we use operational space control).
-See the detailed task setup and keyboard interface in Appendix B-D.
-In both games, the robots are controlled with MPC controller (3) but fail to maintain safety with the initial safety constraints.
+Human corrections are applied via keyboard, with specific keys mapping to the robot control space $\mathbb{R}^{6}$ (we use operational space control).
+See the details in Appendix B-D.
+In both games, the robots are controlled with Safe MPC policy (3) but fail to maintain safety with the initial safety constraints.
 To simulate an emergency stop, the “Enter” key is used as a reset command, allowing users to reset the environment to its initial state whenever the participant deems the robot violating the safety constraint.
-(a) Tracking View
-(b) Fixed View
-Figure 15: Drone navigation game.
-(a) Tracking view.
-(b) Fixed view.
-Users can see the motion of the drone with the help of these two views.
-The $xyz$-axis (red, green, and blue axes), the gate (yellow bars), the drone, the local plan solved from the current MPC (red curve) and the motion target (yellow balls) are visible to users.
-In each game reset, the user will see one target randomly.
-(a) Starting Configuration
-(b) Motion Illustration
-Figure 16: Franka arm reaching game.
-(a) The starting pose of the arm.
-RGB arrows denote the world $xyz$-frame.
-(b) One example of successful reaching, and long green arrow denotes the motion direction.
-Users can see the motion of the robot arm by adjusting the view.
-The slot, the robot arm, and current motion target (cyan ball) are all visible to users.
+(a) Drone Navigation game
+(b) Franka arm reaching game
+Figure 16: Two simulated games in user study. (a) Drone Navigation game. (b) Franka arm reaching game.
 (a) After 2 corrections
@@
 All participants are new to the user study.
-This user study had been reviewed and approved by the Institutional Review Board (IRB) of Arizona State University, and all participants had signed the consent forms.
+This user study had been reviewed and approved by the Institutional Review Board (IRB), and all participants had signed the consent forms.
 VII-B2 Procedure
-Each participant was instructed to play the previously developed two simulation games. In each game, each participant plays 10 trials.
-In the drone navigation game, a trial is deemed successful when the drone navigates through the gate without collision and reaches different targets 6 consecutive times with the learned safe MPC policy.
+Each participant was asked to play the two simulation games.
+In each game, each participant plays 10 trials.
+In the drone navigation game, a trial is successful when the drone navigates through the gate without collision and reaches different targets 6 consecutive times with the learned safe MPC policy.
 A trial is deemed failed when there are too many corrections (exceeding 25), or the user believes the robot cannot be corrected anymore and gives up.
@@
 VII-B3 Metrics for Evaluation
-Two major metrics are used to evaluate the performance of the proposed algorithm:
-  • (I)
-    Success rate: The overall and individual success rate suggest the effectiveness of the proposed method.
-  • (II)
-    Numbers of corrections needed for a successful trial: The averaged number of corrections performed by users in successful trials is a measure for the efficiency of the proposed method.
+Two metrics are used to evaluate the outcomes: (I) Success rate: The overall success rate suggest the effectiveness of the proposed method.
+(II) The number of corrections needed for a successful trial: The averaged number of user corrections for successful trials is a measure for the efficiency of the proposed method.
 Fewer corrections in successful trials suggest better human data efficiency.
 VII-C Result and Analysis
-VII-C1 User Study Metrics Report
+VII-C1 Results
 TABLE II: Result of user study for simulation games
@@
 Franka arm reaching $89.3\%\pm 11.2\%$ $16.5\pm 6.1$
-For the user study, we report the success rate and correction count in a successful game trial in Table II.
+We report the success rate and correction count in a successful game trial in Table II.
 Both metrics are averaged across users.
 The success rate is $89.3\%\pm 7.7\%$ for the drone navigation game and $89.3\%\pm 11.2\%$ for the Franka arm reaching game.
-The number of corrections for a successful trial is $15.7\pm 10.6$ for the drone navigation game and $16.5\pm 6.1$ Franka arm reaching game.
+The number of corrections for a successful trial is $15.7\pm 10.6$ for the drone navigation game and $16.5\pm 6.1$ for Franka arm reaching game.
 In both games, each learning update of $\boldsymbol{\theta}$ takes approximately 0.15 seconds, during which the simulation briefly pauses.
 This brief pause is imperceptible to the user, especially since human corrections occur infrequently.
-The above result illustrates that the proposed method is effective (i.e., with a high success rate) in learning a safety constraint for varying tasks (recall targets are changing across trials).
-It also shows the efficiency of the proposed method — successful learning only requires tens of human corrections.
+The above result shows that the proposed method is effective (i.e., with a high success rate) in learning a safety constraint for varying tasks (recall targets are changing across trials).
+It also shows the efficiency of the proposed method: successful learning only requires tens of human corrections.
 Since this method is online and the user can perform multiple corrections in one robot rollout trajectory, the constraint can be learned in a few runs.
 In Table II, one can note that the drone navigation task has a higher variance in the correction count than the Franka arm reaching task.
-This could be because the drone game has a larger motion range and longer task rollout horizon.
-As a result, human correction strategies, such as correction position and timing, can vary significantly.
-This correction diversity leads to a larger variance of correction count.
-VII-C2 Analysis for Learned Constraint and Robot Motion
+This could be because the drone game has a larger motion range and longer rollout horizon.
+As a result, human correction strategies, such as correction position and timing, vary more.
+VII-C2 Analysis
 To illustrate learning progress, in Fig.
 17 and Fig.
-18, we show the intermediate robot motion at different numbers of human corrections.
+18, we show the intermediate robot motion after different numbers of human corrections.
 Fig.
-17 shows the trajectory of the drone after different numbers of human corrections.
+17 shows the safe MPC rollout trajectory of the drone after different numbers of human corrections.
 Fig.
-18 shows the pose of the robot gripper when passing the slot after a different number of human corrections.
+18 shows the keyframes of the robot gripper when passing the slot after a different human corrections.
 (a) After 4 Corrections
 (b) After 17 Corrections
-Figure 18: Screenshots of the robot gripper when passing through the slot after $k=4,17$ corrections, given three different targets (cyan ball).
+Figure 18: Screenshots of the robot gripper when passing through the slot after $k=4,17$ corrections, given three different targets (cyan ball in different columns).
 (a): After 4 corrections, the gripper collides with the slot.
 (b): After 17 corrections, the gripper succeeds in passing the slot for all targets.
-To better understand the learned safety constraint ${g}_{\boldsymbol{\theta}}(\boldsymbol{\xi})$, in Fig.
+To understand the learned safety constraint ${g}_{\boldsymbol{\theta}}(\boldsymbol{\xi})$, in Fig.
 19, we draw the heat map of the learned ${g}_{\boldsymbol{\theta}^{*}}(\boldsymbol{\xi})$ for the drone navigation game, with $\boldsymbol{\theta}^{*}$ the averaged of weights learned.
-Recall in drone navigation game, ${g}_{\boldsymbol{\theta}}$ is defined on y-z plane at $x=10$ (true gate position is $x_{\text{gate}}=9.8$).
-In the heat map, we draw 00-level contour of ${g}_{\boldsymbol{\theta}^{*}}$ in dashed lines, which is the learned safety boundary.
-Also, we plot the $yz$-plane projection of the gate, which is the black box.
+Recall in drone navigation game, ${g}_{\boldsymbol{\theta}}$ is defined on y-z plane at $x=10$.
+In the heat map, we draw 0-level contour of ${g}_{\boldsymbol{\theta}^{*}}$ in dashed lines, which is the learned safety boundary.
+Also, we plot y-z plane projection of the gate, which is the black box.
 Fig.
 19 shows that the learned safety boundary (corresponding to ${g}_{\boldsymbol{\theta}^{*}}(\boldsymbol{\xi})\leq 0$) is only similar to the top edge of the gate.
-This is explained by the following reasons.
+This is explained below.
 As shown in Fig.
-15, the targets are set higher than the gate, and thus during the learning process, the motion of the drone is always going up, as shown in Fig.
+16(a), the targets are set higher than the gate, and thus during the learning process, the motion of the drone is always going up, as shown in Fig.
 17.
@@
 Figure 19: Heatmap for the learned constraint in the drone game.
-The black solid line denotes the actual position of the gate in $yz$-plane.
+The black solid line denotes the actual position of the gate in y-z plane.
 The black dashed curve denotes the boundary of the learned safety zone.
-Since the targets are higher than the gate (see Fig.
-15), during the learning process, the motion of the drone is always going up, as shown in Fig.
-17.
-As a result, the majority of human directional corrections happen when the drone approaches the top edge of the gate, leaving the learned safety boundary accurate near the top edge of the gate and bottom boundary untouched.
-VIII User Study in Real-world Experiment
-In this section, we conduct a user study to evaluate the performance of our safe MPC alignment algorithm in a hardware Franka Research 3 experiment.
-We focus on the task of mobile liquid pouring, i.e., the robot performs liquid pouring while moving from one location to targets.
-In certain scenarios, this kind of task can be safety-critical: for example, if the liquid is corrosive/chemical/toxic, inappropriate mobile pouring will cause safety concerns.
-The task is challenging because it is hard to manually define a safety constraint for the robot to accomplish liquid pouring without spilling.
-The Franka robot arm will learn a safe MPC controller from human physical corrections.
+Since the majority of human directional corrections happen when the drone approaches the top edge of the gate, the learned safety boundary is accurate near the top edge of the gate and the bottom boundary is untouched.
+VIII Real-world Experiment
+We conduct a user study to evaluate the performance of the Safe MPC Alignment in a real-world experiment.
+We focus on the task of mobile liquid pouring: the Franka robot arm performs liquid pouring while moving from one location to targets.
+The task is challenging because it is hard to manually define a safety constraint to accomplish liquid pouring without spilling.
+The robot arm will learn a safe MPC controller from human physical corrections.
 VIII-A Hardware Setup
 […]
-The setup of mobile water pouring task is shown in Fig.
+The setup of the mobile water pouring task is shown in Fig.
 20.
-A water pot is gripped by the robot end-effector.
-The robot moves from the starting point $[0.22,-0.5,0.48]^{\mkern-1.5mu\mathsf{T}}$ (with quaternion $[0,-0.707,0,0.707]^{\mkern-1.5mu\mathsf{T}}$) to target point $[-0.65,-0.5,0.5]^{\mkern-1.5mu\mathsf{T}}$ (with quaternion $[0,-0.707,0,0.707]^{\mkern-1.5mu\mathsf{T}}$).
+A water pot is gripped by the robot’s end-effector.
+The robot moves from the starting point $0.22,-0.5,0.48]^{\mkern-1.5mu\mathsf{T}}$ (with quaternion $0,-0.707,0,0.707]^{\mkern-1.5mu\mathsf{T}}$) to target point $0.65,-0.5,0.5]^{\mkern-1.5mu\mathsf{T}}$ (with quaternion $0,-0.707,0,0.707]^{\mkern-1.5mu\mathsf{T}}$).
 During the motion, the robot needs to pour more than 50 grams of water into a bar sink.
-The bar sink is placed in the middle of the robot path, but the precise size and location of the bar sink are unknown.
-Figure 21: System diagram for the hardware implementation. Different blocks indicate separate processes connected via arrows that indicate communication via ROS.
-As shown Fig.
-21, the hardware Franka arm platform has three main components: the parameterized safe MPC controller (high-level controller), the low-level controller, and safe MPC alignment module.
+The bar sink is placed in the middle of the robot’s path, but the precise size and location of the bar sink are unknown.
+Figure 21: System diagram for the hardware implementation. Different blocks indicate separate processes communicated via ROS.
+Fig.
+21 shows three main components of the system: the parameterized Safe MPC controller (high-level controller), the low-level controller, and the Safe MPC alignment module.
 The Safe MPC controller follows the safe MPC formulation of Franka arm reaching game in Appendix B-D (recall that the learnable safety constraint is parameterized by RBFs).
 The parameters of this Safe MPC controller are listed in Table III.
-The safe MPC controller receives the end-effector’s current pose $\boldsymbol{x}$ and generate the action $\boldsymbol{u}$ (see (58) in Appendix), which is the desired velocity of the end-effector.
+The Safe MPC controller receives the end-effector’s current pose $\boldsymbol{x}$ and generates the action $\boldsymbol{u}$ (see (59) in Appendix), which is the desired velocity of the end-effector.
 The low-level controller implements the operational space control to track the $\boldsymbol{u}$ given by the safe MPC controller.
-The safe MPC alignment module updates the parameterize $\boldsymbol{\theta}$ of the safe MPC controller.
+The Safe MPC Alignment module updates the parameter $\boldsymbol{\theta}$ of the Safe MPC controller.
 The update only happens when the alignment module detects a human correction $\boldsymbol{a}$.
 The module implements our Safe MPC Alignment algorithm.
-Note that the alignment module and controller modules runs different process, thus the learning update does not affect the real-time MPC control of the robot arm.
+Note that the alignment module and controller modules runs in different processes, thus the learning update does not affect the real-time MPC control of the robot arm.
 ROS is used for communication between different processes.
@@
 $\mu_{q}$ 80
-In the safe MPC controller, the cost of the MPC controller is set to penalize the current and target, as in Section B-D.
-Since it is difficult to explicitly define a parameterized safety constraint using the state of pouring water, we instead define the constraints only using the robot state.
-Specifically, we use the weighted RBFs to parameterize the safety constraints constraint, similar to Appendix B-D.
 As shown in Fig.
@@
 (a) After 2 corrections, the arm begins tilting the pot but spills out the water.
-(b) After 3 corrections, the arm can pour water into the sink but the human is not satisfied with its height and altitude of pouring.
+(b) After 3 corrections, the arm can pour water into the sink but the human is not satisfied with its height and the altitude of pouring.
 (c) After 5 corrections, the arm learns to tilt only along the direction of the $x$-axis and successfully aligns with human intent.
 VIII-B Participants, Procedures, and Metrics
-All participants recruited for the previous two simulation games in Section VII-B also participate the hardware experiment.
+All participants recruited in Section VII-B also participate in the hardware experiment.
 They are told of the objective before starting trials.
@@
 One learning trial may include multiple episodes (with a maximum of 15 corrections).
-During learning, the nozzle of the water pot is sealed, and we switch to the unsealed nozzle and measure the amount of water poured into the flower pot after the user is satisfied with the motion of the robot motion.
-The success learning trial is defined when the amount of poured water is over 50g.
+During learning, the nozzle of the water pot is sealed, and we switch to the unsealed nozzle and measure the amount of water poured into the flower pot after the user is satisfied with the motion of the robot.
+The successful trial is defined when the amount of poured water is over 50g.
 Each user performs 5 trials, and the same metrics in Section VII-B are recorded.
 VIII-C Result and Analysis
-In the real-world robot experiment, the success rate is $84.0\%\pm 16.7\%$ and the average number of corrections needed for successful trials is $6.88\pm 3.26$.
-The results show the efficacy and efficiency of our method in real-world robot learning applications.
-To illustrate the robot learning progress, in Fig.
-22, we use one trial to show the robot behavior at different numbers of human corrections.
+In the real-world experiment, the success rate is $84.0\%\pm 16.7\%$ and the average number of corrections needed for successful trials is $6.88\pm 3.26$.
+The results show the efficacy and efficiency of our method in real-world applications.
+To illustrate the learning progress, Fig.
+22 shows the robot behavior after different numbers of corrections in one trial.
 Since this real-world Franka arm experiment is similar to the simulated Franka arm reaching game in Appendix B-D, we compare the user study results for both and have the following conclusion and analysis.
-(I) The real-world experiment has a slightly lower success rate: While the difference between simulations and real-world experiments is small in success rate, we have identified some factors during our experiment that may contribute to such lower success rate, including both the human aspects and hardware aspects.
+(I) Real-world experiment has slightly lower success rate: While the difference between simulation and real-world studies is small in success rate, we have identified some factors for such a difference.
 On human aspects, most of the participants have no experience or knowledge in operating robot hardware, and thus when directly interacting with hardware, their corrections can be conservative or ineffective due to safety considerations.
-On hardware side, noisy correction signals and the possible timing delay between the recorded human correction and robot state can contribute to the failure of the learning.
-(II) The real-world experiment has a less number of human corrections: This could be attributed to two reasons.
-First, the constraint needed for accomplishing the water pouring task could be more flexible than that for reaching task in the simulation game Appendix B-D.
+On hardware side, noisy correction signals and the possible timing delay between the recorded correction and the robot state can contribute to failure.
+(II) Real-world experiment needs less human corrections: This could be attributed to two reasons.
+First, the constraint needed for accomplishing the water pouring task could be more flexible than that for reaching task in the simulation game.
 This makes the safety constraints easy to learn for the real-world task.
-Second, different from simulation games, where human users can only perceive the robot arm from the screen, a human user has a better perception of the robot’s motion in real world.
+Second, different from simulation games, where human users can only perceive the robot arm from the screen, a human user has a better perception of the robot’s motion in the real world.
 A good perception of the robot’s state could help humans give more effective feedback each time, reducing the total number of human interactions for successful robot learning.
-VIII-D Discussion
-Based on the observations in both the hardware and simulation user studies, we discuss some design choices and potential limitations of the proposed method.
-VIII-D1 Learning Task-Independent Safety Constraints
-In the real-world and simulation experiments, due to the insufficiency of task sampling, the learned safety constraint could be task-specific, instead of task-independent.
+IX Discussion and Conclusion
+IX-A Discussion
+Based on the above hardware and simulation user studies, we discuss some design choices and potential limitations of the proposed method.
+IX-A1 Learning Task-Independent Safety Constraints
+In the real-world and simulation experiments, the learned safety constraint could be task-specific, instead of task-independent.
 For instance, in the drone navigation game, since the three task targets are placed above the gate, the learned constraint shown in Fig.
 19 can only reflect the upper edge of the unknown gate.
-One limitation of this is that the learned task constraints have poor generalization to new tasks.
+One limitation is that the learned constraints have poor generalization to new tasks.
 For example, in drone navigation, if a new target is placed below the gate, the drone with the learned safety constraint Fig.
@@
 To improve the generalization of the learned safety constraints or learning task-independent safety constraints, one solution is to increase the diversity of task samples.
-By doing so, it is expected that the robot can receive human correction on the entire boundary of the safety region, making the learned safety constraints reusable in new tasks.
-VIII-D2 Robustness Against Human Mistakes
+By doing so, it is expected that the robot can receive human correction on the entire boundary of the safety region.
+IX-A2 Robustness Against Human Mistakes
 The proposed Safe MPC Alignment is based on the hypothesis cutting.
@@
 23.
-Figure 23: One illustration of the failure case in the drone game user study. A human user makes a false "up" correction at the beginning and fails to correct the motion of the drone.
+Figure 23: A failure case in the drone game user study. A user makes a wrong "up" correction at the beginning and fails to correct the motion of the drone.
 To further improve the robustness of the method against human mistakes, two extensions of the current method can be explored.
@@
 $\forall\boldsymbol{\theta}_{H}\in\bar{\Theta}_{H},\ \operatorname{\mathbb{E}}_{\mathbf{a}}\left\langle\mathbf{a},-\nabla B(\boldsymbol{\xi}_{\boldsymbol{\theta}}^{\gamma},\boldsymbol{\theta}_{H})\right\rangle+\epsilon\geq 0.$ (35)
-Another is to add a selection mechanism on human correction.
-For example, one can use some statistical method to rule out mistake corrections before performing a hypothesis cut.
-We will leave those possible extensions in our future work.
-IX Conclusion
-This paper proposes safe MPC alignment method to online learn a safety constraint in safe MPC controllers from human directional feedback.
-The method, named hypothesis space cutting, has geometric interpretation: it exponentially cuts the hypothesis space based on each received human directional feedback.
+Another is to add a selection mechanism for human correction.
+For example, one can use a voting system to preserve more fraction of the hypothesis space and avoid cutting out the human safety set, as shown in [xie2025robust].
+IX-A3 Limitation of Weight-Feature Constraints
+In this work, we assume the set of task-relevant features in $\boldsymbol{\phi}(\boldsymbol{\xi})$ is selected as a priori.
+These features define the space in which the safety constraints are represented.
+However, in practice, identifying appropriate features can be challenging.
+Poorly chosen features may fail to capture the intended safety constraints, leading to undesired learning outcomes.
+Furthermore, feature representations are often task-dependent—features that are effective in one task may not generalize to another.
+This task-specific dependence inherently limits the development of a unified framework for constraint learning across diverse tasks, making the lack of task-generalizable feature selection a key limitation of the proposed method.
+At the same time, learning unstructured constraints entirely from scratch would demand extensive human interaction and involvement, rendering such approaches impractical for most real-world applications.
+This challenge likely explains why recent works on learning from human feedback [bajcsy2017learning, jin2022learning, losey2022physical] commonly assume task-specific feature structures or prior knowledge of the environment/tasks.
+We recognize that identifying or learning a universal feature representation that can flexibly describe a wide variety of safety constraints remains an important and open research challenge.
+We consider this an exciting direction for future research, as it could significantly increase the applicability and effectiveness of safety constraint learning in real world.
+IX-B Conclusion
+This paper proposes a Safe MPC Alignment method to online learn a safety constraint in Safe MPC controllers from human directional feedback.
+The method, named hypothesis space cutting, has a geometric interpretation: it exponentially cuts the hypothesis space based on each received human directional feedback.
+To our knowledge, this is the first certifiable method to enable online learning of safety constraints with human feedback.
 The method has a provable guarantee for the total number of human feedback needed for learning, and also can declare hypothesis misspecification.
@@
 The results show that the proposed method can learn safety constraints effectively and efficiently.
-To our knowledge, this is the first certifiable method to enable online learning of safety constraints with human feedback.
 Appendix A Proof of lemmas and theorems
@@
 Suppose $\bar{\Theta}_{H}\subset\Theta_{i-1}$, $i=1,2,3,$....
-Due to the condition $\bar{\Theta}_{H}\subset\mathcal{C}_{i}$ in (12), following the update (13), we have $\bar{\Theta}_{H}\subset\Theta_{i}=\Theta_{i-1}\cap\mathcal{C}_{i}$, which completes the proof.
+From the previous assumptions (7) and (9), $\bar{\Theta}_{H}\subset\mathcal{C}_{i}$.
+Following the update (13), we have $\bar{\Theta}_{H}\subset\Theta_{i}=\Theta_{i-1}\cap\mathcal{C}_{i}$, which completes the proof.
 ∎
@@
 (14) can be directly obtained from (12), (2) and (5).
-Based on the definition of the penalty objective function (2), $\mathbf{\bar{a}}_{i},\nabla B(\boldsymbol{\xi}_{\boldsymbol{\theta}_{i}}^{\gamma},\boldsymbol{\theta})\right>\leq 0$ is a linear fractional inequality
-$\mathbf{\bar{a}}_{i},\nabla J(\boldsymbol{\xi}_{\boldsymbol{\theta}_{i}}^{\gamma})\right>+\gamma\frac{1}{-g_{\boldsymbol{\theta}}(\boldsymbol{\xi}_{\boldsymbol{\theta}_{i}}^{\gamma})}\left<\mathbf{\bar{a}}_{i},\nabla g_{\boldsymbol{\theta}}(\boldsymbol{\xi}_{\boldsymbol{\theta}_{i}}^{\gamma})\right>\leq 0,$ (37)
-with a positive denominator $-g_{\boldsymbol{\theta}}(\boldsymbol{\xi}_{\boldsymbol{\theta}_{i}}^{\gamma})>0$ due to (12).
+Based on the definition of the penalty objective function (2), $\left<\mathbf{\bar{a}}_{i},\nabla B(\boldsymbol{\xi}_{\boldsymbol{\theta}_{i}}^{\gamma},\boldsymbol{\theta})\right>\leq 0$ is a linear fractional inequality
+$\left<\mathbf{\bar{a}}_{i},\nabla J(\boldsymbol{\xi}_{\boldsymbol{\theta}_{i}}^{\gamma})\right>+\gamma\frac{1}{-g_{\boldsymbol{\theta}}(\boldsymbol{\xi}_{\boldsymbol{\theta}_{i}}^{\gamma})}\left<\mathbf{\bar{a}}_{i},\nabla g_{\boldsymbol{\theta}}(\boldsymbol{\xi}_{\boldsymbol{\theta}_{i}}^{\gamma})\right>\leq 0,$ (36)
+with a positive denominator $g_{\boldsymbol{\theta}}(\boldsymbol{\xi}_{\boldsymbol{\theta}_{i}}^{\gamma})>0$ due to (12).
 Multiplying the denominator on both side and substituting the weight-feature expression of $g_{\boldsymbol{\theta}}$ in (5) yields $\boldsymbol{\theta}^{\mkern-1.5mu\mathsf{T}}\boldsymbol{h}_{i}\leq b_{i}$ in Lemma 3.
-The condition $-g_{\boldsymbol{\theta}}(\boldsymbol{\xi}_{\boldsymbol{\theta}_{i}}^{\gamma})>0$ itself yields $\ {\boldsymbol{\theta}}^{\mkern-1.5mu\mathsf{T}}\boldsymbol{\phi}(\boldsymbol{\xi}_{\boldsymbol{\theta}_{i}}^{\gamma})<-\phi_{0}(\boldsymbol{\xi}_{\boldsymbol{\theta}_{i}}^{\gamma})$ in Lemma 3.
+The condition $g_{\boldsymbol{\theta}}(\boldsymbol{\xi}_{\boldsymbol{\theta}_{i}}^{\gamma})>0$ itself yields ${\boldsymbol{\theta}}^{\mkern-1.5mu\mathsf{T}}\boldsymbol{\phi}(\boldsymbol{\xi}_{\boldsymbol{\theta}_{i}}^{\gamma})<-\phi_{0}(\boldsymbol{\xi}_{\boldsymbol{\theta}_{i}}^{\gamma})$ in Lemma 3.
 This completes the proof. ∎
@@
 With (37),
-$\mathbf{a},\nabla B(\boldsymbol{\xi}_{\boldsymbol{\theta}_{i}}^{\gamma},\boldsymbol{\theta}_{i})\right>=0,$ (39)
+$\left<\mathbf{a},\nabla B(\boldsymbol{\xi}_{\boldsymbol{\theta}_{i}}^{\gamma},\boldsymbol{\theta}_{i})\right>=0,$ (38)
 always holds. Extending the above (38) leads to
-$\mathbf{a},\nabla J(\boldsymbol{\xi}_{\boldsymbol{\theta}_{i}}^{\gamma})\right>-\gamma\frac{1}{g_{\boldsymbol{\theta}_{i}}(\boldsymbol{\xi}_{\boldsymbol{\theta}_{i}}^{\gamma})}\left<\mathbf{a},\nabla g_{\boldsymbol{\theta}_{i}}(\boldsymbol{\xi}_{\boldsymbol{\theta}_{i}}^{\gamma})\right>=0.$ (40)
-Due to $-g_{\boldsymbol{\theta}_{i}}(\boldsymbol{\xi}_{\boldsymbol{\theta}_{i}}^{\gamma})>0$, one can multiple $-g_{\boldsymbol{\theta}_{i}}(\boldsymbol{\xi}_{\boldsymbol{\theta}_{i}}^{\gamma})$ on both sides of the above equality, which leads to $\boldsymbol{\theta}_{i}^{\mkern-1.5mu\mathsf{T}}\boldsymbol{h}_{i}=b_{i}$ based on the definition (15).
+$\left<\mathbf{a},\nabla J(\boldsymbol{\xi}_{\boldsymbol{\theta}_{i}}^{\gamma})\right>-\gamma\frac{1}{g_{\boldsymbol{\theta}_{i}}(\boldsymbol{\xi}_{\boldsymbol{\theta}_{i}}^{\gamma})}\left<\mathbf{a},\nabla g_{\boldsymbol{\theta}_{i}}(\boldsymbol{\xi}_{\boldsymbol{\theta}_{i}}^{\gamma})\right>=0.$ (39)
+Due to $g_{\boldsymbol{\theta}_{i}}(\boldsymbol{\xi}_{\boldsymbol{\theta}_{i}}^{\gamma})>0$, one can multiply $g_{\boldsymbol{\theta}_{i}}(\boldsymbol{\xi}_{\boldsymbol{\theta}_{i}}^{\gamma})$ on both sides of the above equality, which leads to $\boldsymbol{\theta}_{i}^{\mkern-1.5mu\mathsf{T}}\boldsymbol{h}_{i}=b_{i}$ based on the definition (15).
 The second term in (18) directly follows $g_{\boldsymbol{\theta}_{i}}(\boldsymbol{\xi}_{\boldsymbol{\theta}_{i}}^{\gamma})<0$.
@@
 By Lemma 4, the MVE center $\boldsymbol{\theta}_{i}=\bar{\boldsymbol{d}}_{i-1}$ is on the cutting hyperplane $\boldsymbol{\theta}^{\mkern-1.5mu\mathsf{T}}\boldsymbol{h}_{i}=b_{i}$.
-Directly from [63], one has
+Directly from [tarasov1988method], one has
 $\frac{\mathbf{Vol}(\Theta_{i-1}\cap\{\boldsymbol{\theta}|\boldsymbol{\theta}^{\mkern-1.5mu\mathsf{T}}\boldsymbol{h}_{i}\leq b_{i}\})}{\mathbf{Vol}(\Theta_{i-1})}\leq 1-\frac{1}{r}.$ (40)
@@
 We complete the proof in two parts.
-First, we prove the number of learning iterations is less than $K$, then we use this to prove there will be a $\boldsymbol{\theta}_{k}\in\bar{\Theta}_{H}$, $k=0,1,2...,K-1.$
+First, we prove the number of learning iterations is less than $K$, then we use this to prove there will be a $\boldsymbol{\theta}_{k}\in\bar{\Theta}_{H}$, k=0,1,2,K1.k=0,1,2...,K-1.
 We first prove the total number of iterations must be less than $K$ by contradiction. Suppose the learning iteration number $i\geq K$. With recursively applying lemma 5, one can obtain
@@
 We first prove the algorithm converges to $\Theta^{*}$ which has $\mathbf{Vol}(\Theta^{*})=0$.
-Given the condition $\Theta_{0}\cap\bar{\Theta}_{H}=\emptyset$, the algorithm will never terminate because $\forall i=1,2,3,...\infty$, $\mathcal{C}_{i}\supset\bar{\Theta}_{H}$ due to Lemma 2.
+Given the condition $\Theta_{0}\cap\bar{\Theta}_{H}=\emptyset$, the algorithm will never terminate because $\forall i=1,2,3,$$\infty$, $\mathcal{C}_{i}\supset\bar{\Theta}_{H}$ due to Lemma 2.
 Since the volume of the initial hypothesis space is finite, by applying Lemma 5 we have
@@
 $d$ 0.1 Dampening Coefficient
-When the system state violates ${g}_{\boldsymbol{\theta}_{H}}(\boldsymbol{\xi}_{\boldsymbol{\theta}}^{\gamma})\leq 0$ or it is near the target, the pendulum state is uniformly randomly reset between $[0,0]^{\mkern-1.5mu\mathsf{T}}$ and $[\frac{2\pi}{3},3]^{\mkern-1.5mu\mathsf{T}}$.
+When the system state violates ${g}_{\boldsymbol{\theta}_{H}}(\boldsymbol{\xi}_{\boldsymbol{\theta}}^{\gamma})\leq 0$ or it is near the target, the pendulum state is uniformly randomly reset between $0,0]^{\mkern-1.5mu\mathsf{T}}$ and $\frac{2\pi}{3},3]^{\mkern-1.5mu\mathsf{T}}$.
 Simulated human correction will continue to be applied until the learned parameters $\boldsymbol{\theta}_{i}$ fall in $\bar{\Theta}_{H}$, which terminates one trial.
-B-B Two-link Robot Arm
-We assume each links has unit length and mass, and the mass is evenly distributed along the link. The corresponding inertia matrix and Coriolis term can then be calculated.
-B-C Drone Navigation Game
+B-B Simulated Quadrotor Navigation
+B-B1 Environment
+The quadrotor starts from $0.0,0.0,5.0]^{\mkern-1.5mu\mathsf{T}} [$m].
+The starting quaternion $\boldsymbol{q}_{\text{init}}=[1,0,0,0]^{\mkern-1.5mu\mathsf{T}}$.
+The target is to navigate to the target position $15.0,0.0,10.0]^{\mkern-1.5mu\mathsf{T}}[$m] within a zigzag tube.
+The average radius of the tube section is set to be 2.5m.
+B-B2 Learning Neural Features
+We construct the neural feature $\boldsymbol{\phi}$ using a neural Signed Distance Function (SDF) model trained on synthetic 3D point samples around the tube surface.
+Each input coordinate (scaled for stability) is first encoded using sinusoidal positional embeddings, then passed through a 5-layer MLP (width 128, ReLU activations).
+The positional embeddings are skip-connected to the input of the 5th hidden layer.
+The penultimate layer outputs a 16D feature vector used as $\boldsymbol{\phi}$; the final layer predicts a scalar signed distance supervised during training.
+Training data is generated by randomly perturbing surface points on the tube.
+Signed distance labels are computed as the offset from the nearest point on the tube surface to the sample point, and normals are estimated from the same displacement vector pointing from the nearest surface point to the sample point.
+The model is trained with three loss terms: $L_{2}$ distance loss, cosine similarity loss on normals (via auto-diff), and Eikonal loss to regularize gradient norm, which is similar to the work [ortiz2022isdf].
+Optimization is done using Adam for 250 epochs with a batch size of 1024.
+B-B3 Simulated Correction and Stop Behavior
+Correction signals $\boldsymbol{a}_{x}=[-1,0,1,0]^{\mkern-1.5mu\mathsf{T}},\boldsymbol{a}_{y}=[0,1,0,-1]^{\mkern-1.5mu\mathsf{T}},\boldsymbol{a}_{z}=[1,1,1,1]^{\mkern-1.5mu\mathsf{T}}$ in drone control space is used throughout the experiments, corresponding to different motion directions in cartesian space.
+During the process of learning positional constraints with polynomial features or neural features, if the distance between the quadrotor and the nearest point on the surface of the tube is less than 0.6, a simulated correction $\boldsymbol{a}$ is applied in the action space of the drone, based on the vector $\boldsymbol{d}=(d_{x},d_{y},d_{z})$ pointing from the closest tube surface point to the drone position:
+$\boldsymbol{a}=d_{y}\boldsymbol{a}_{y}+d_{z}\boldsymbol{a}_{z}.$ (53)
+Intuitively, the simulated correction "pulls" the drone to the center of the tube (designated safe region).
+After each correction, the navigation restarts, emulating an emergency stop.
+During the process of learning velocity constraint, Simulated human corrections against the velocity direction of form:
+$\boldsymbol{a}=-v_{x}\boldsymbol{a}_{x}-v_{y}\boldsymbol{a}_{y}-v_{z}\boldsymbol{a}_{z}$ (54)
+are applied whenever the velocity $\boldsymbol{v}=(v_{x},v_{y},v_{z})$ exceeds 0.4, and an emergency stop is triggered if it surpasses 0.45.
+B-C Quadrotor Navigation Game with Real Human Feedback
 B-C1 Environment
-The drone starts from $[0.1,y_{\text{init}},0.5]^{\mkern-1.5mu\mathsf{T}}$ with $y_{\text{init}}$ uniformly randomly in an interval $[4,6]$.
+The drone starts from $0.1,y_{\text{init}},0.5]^{\mkern-1.5mu\mathsf{T}}$ with $y_{\text{init}}$ uniformly randomly in an interval $[4,6]$.
 The starting quaternion $\boldsymbol{q}_{\text{init}}=[1,0,0,0]^{\mkern-1.5mu\mathsf{T}}$.
-We set three different targets at positions $[19,1,9]^{\mkern-1.5mu\mathsf{T}}$, $[19,5,9]^{\mkern-1.5mu\mathsf{T}}$ and $[19,9,9]^{\mkern-1.5mu\mathsf{T}}$, with quaternion $[1,0,0,0]^{\mkern-1.5mu\mathsf{T}}$.
+We set three different targets at positions $19,1,9]^{\mkern-1.5mu\mathsf{T}}$, $19,5,9]^{\mkern-1.5mu\mathsf{T}}$ and $19,9,9]^{\mkern-1.5mu\mathsf{T}}$, with quaternion $1,0,0,0]^{\mkern-1.5mu\mathsf{T}}$.
 For the narrow gate (yellow rectangle) in Fig.
-15, both its size and position are unknown, and one only roughly knows its x position $x_{\text{gate}}\approx 9.8$.
+16(a), both its size and position are unknown, and one only roughly knows its x position $x_{\text{gate}}\approx 9.8$.
 Since the location of the gate is unknown, the drone needs to learn a safety constraint in its safe MPC policy with humans guidance in order to successfully fly through the gate.
-In the drone MPC policy (1), the dynamics and its parameters are set following ([20]); and we generically set the constraint function ${g}_{\boldsymbol{\theta}}(\boldsymbol{\xi})$ using a set of $N$ Gaussian radial basis functions (RBFs)
+In the drone MPC policy (1), the dynamics and its parameters are set following [jin2022learning]; and we generically set the constraint function ${g}_{\boldsymbol{\theta}}(\boldsymbol{\xi})$ using a set of $N$ Gaussian radial basis functions (RBFs)
 ${g}_{\boldsymbol{\theta}}(\boldsymbol{\xi})=\phi_{0}+\sum\nolimits_{i=1}^{N}\theta_{i}\phi_{i}(\boldsymbol{\xi}).$ (55)
@@
 B-C2 Human Correction Interface
-The motion of the drone is rendered in two camera views in MuJoCo visualization (Fig.
-15) for better human perception: one is the first-person view tracking the drone, the other a third-person fixed view.
-The gate, the current MPC planning solution, and the target position are visible to the users.
-Human users apply corrections to the drone via a keyboard during the drone flying.
-The user is informed of the mappings between the key press and the corresponding correction in the game, listed in Table V.
-The safety constraint (54) in the drone’s MPC policy is immediately updated once a user correction is detected.
-When the user presses "ENTER", the drone is reset to the initial position and target will randmize.
+The drone’s motion is visualized in MuJoCo using two complementary camera views: a first-person tracking view and a fixed third-person view.
+The gate, the current MPC plan, and the target position are always shown to the user.
+Human users provide corrections through keyboard inputs during flight.
+Key–command mappings are given in Table V.
+Whenever a correction is detected, the safety constraint (55) in the MPC policy is updated instantly.
+Pressing ENTER resets the drone to its initial state and reset the target.
 TABLE V: keyboard commands in uav user study
 Key Correction Interpretation
-"up" $[-1,0,1,0]^{\mkern-1.5mu\mathsf{T}}$ Go forward (adding positive y-axis torque)
-"down" $[1,0,-1,0]^{\mkern-1.5mu\mathsf{T}}$ Go backward (adding negative y-axis torque)
-"left" $[0,1,0,-1]^{\mkern-1.5mu\mathsf{T}}$ Go left (adding negative x-axis torque)
-"right" $[0,-1,0,1]^{\mkern-1.5mu\mathsf{T}}$ Go right (adding positive x-axis torque)
-"shift" $[1,1,1,1]^{\mkern-1.5mu\mathsf{T}}$ Go up (increasing all thrusts)
-"ctrl" $-[1,1,1,1]^{\mkern-1.5mu\mathsf{T}}$ Go down (decreasing all thrusts)
+"up" $1,0,1,0]^{\mkern-1.5mu\mathsf{T}}$ Go forward (adding positive y-axis torque)
+"down" $1,0,-1,0]^{\mkern-1.5mu\mathsf{T}}$ Go backward (adding negative y-axis torque)
+"left" $0,1,0,-1]^{\mkern-1.5mu\mathsf{T}}$ Go left (adding negative x-axis torque)
+"right" $0,-1,0,1]^{\mkern-1.5mu\mathsf{T}}$ Go right (adding positive x-axis torque)
+"shift" $1,1,1,1]^{\mkern-1.5mu\mathsf{T}}$ Go up (increasing all thrusts)
+"ctrl" $1,1,1,1]^{\mkern-1.5mu\mathsf{T}}$ Go down (decreasing all thrusts)
 "enter" N/A Reset robot position and switch target.
@@
 The arm starts at a configuration in which the end-effector position $\boldsymbol{r}_{\text{init}}=[0.22,-0.5,0.48]^{\mkern-1.5mu\mathsf{T}}$ and altitude $\boldsymbol{q}_{\text{init}}=[0,-0.707,0,0.707]^{\mkern-1.5mu\mathsf{T}}$.
-The reaching target position $\boldsymbol{r}^{*}$ will be randomly chosen from three options $[-0.6,-0.5,0.4]^{\mkern-1.5mu\mathsf{T}}$, $[-0.6,-0.5,0.5]^{\mkern-1.5mu\mathsf{T}}$, $[-0.6,-0.5,0.6]^{\mkern-1.5mu\mathsf{T}}$.
+The reaching target position $\boldsymbol{r}^{*}$ will be randomly chosen from three options $0.6,-0.5,0.4]^{\mkern-1.5mu\mathsf{T}}$, $0.6,-0.5,0.5]^{\mkern-1.5mu\mathsf{T}}$, $0.6,-0.5,0.6]^{\mkern-1.5mu\mathsf{T}}$.
 The reaching target attitude is $\boldsymbol{q}^{*}=[0,-0.707,0,0.707]^{\mkern-1.5mu\mathsf{T}}$.
@@
 The output of the MPC controller is sent to the low-level controller to track.
+TABLE VI: keyboard commands in robot arm user study
+Key Correction Interpretation
+"left" $0,0,0,0,0,1]^{\mkern-1.5mu\mathsf{T}}$ Clock-wise rotation of the gripper
+"right" $0,0,0,0,0,-1]^{\mkern-1.5mu\mathsf{T}}$ Anti-clock-wise rotation of the gripper
+"shift" $0,0,0,-1,0,0]^{\mkern-1.5mu\mathsf{T}}$ Go up in world frame
+"ctrl" $0,0,0,1,0,0]^{\mkern-1.5mu\mathsf{T}}$ Go down in world frame
+"Enter" N/A Reset robot position and switch target.
 Similar to Appendix B-C, to constrain $r_{z}$ and $\boldsymbol{q}$ for proper height and tilt in order to pass through the slot.
@@
 $\sigma(r_{x,t})=\mathrm{sigmoid}(50\big((r_{x,t}-x_{\text{slot}})^{2}-(\frac{w}{2})^{2}\big)),$ (64)
-which is used to constrain the $\mathrm{RBF}_{i}$ only taking effect around $x$ range $[x_{\text{slot}}-\frac{w}{2},x_{\text{slot}}+\frac{w}{2}]$ (recall slot’s $x$ location, slot, is a approximate value).
+which is used to constrain the $\mathrm{RBF}_{i}$ only taking effect around $x$ range $x_{\text{slot}}-\frac{w}{2},x_{\text{slot}}+\frac{w}{2}]$ (recall slot’s $x$ location, slot, is a approximate value).
 In the above definition, $w=0.3$, $k_{z}=0.05$, $\epsilon_{z}=12$ and $\epsilon_{q}=1.8$.
-The center of positional $\mathrm{RBF}$, ${c_{z,i}}$, is evenly chosen on interval $[0,1]$, and the center of orientation $\mathrm{RBF}$, ${\boldsymbol{q}_{c,i}}$, is generated by $(\cos{\frac{\alpha}{2}},\sin{\frac{\alpha}{2}},0,0)*\boldsymbol{q}_{\text{init}}$ with $\alpha$ evenly choosen from $[-\pi/4,3\pi/4]$ to make sure ${\boldsymbol{q}_{c,i}}$ covers all possible orientations.
-$\beta_{t}=0.9^{t}$ if $t<5$ and 00 otherwise.
+The center of positional $\mathrm{RBF}$, $c_{z,i}$}, is evenly chosen on interval $[0,1]$, and the center of orientation $\mathrm{RBF}$, $\boldsymbol{q}_{c,i}$}, is generated by $(\cos{\frac{\alpha}{2}},\sin{\frac{\alpha}{2}},0,0)*\boldsymbol{q}_{\text{init}}$ with $\alpha$ evenly choosen from $\pi/4,3\pi/4]$ to make sure $\boldsymbol{q}_{c,i}$} covers all possible orientations.
+$\beta_{t}=0.9^{t}$ if $t<5$ and 0 otherwise.
 The bounds of the initial hypothesis space $\Theta_{0}$ is set to $\boldsymbol{c}_{l}=[-3,\dots,-3]^{\mkern-1.5mu\mathsf{T}}$ and $\boldsymbol{c}_{u}=[5,\dots,5]^{\mkern-1.5mu\mathsf{T}}$.
@@
 B-D2 Human Correction Interface
-The motion of the robot arm is rendered in the free-view mode in MuJoCo visualization (Fig.
-16), and users can drag to select his/her comfortable view.
-The slot and the random reaching target are visible to the users.
-Human users apply their corrections to the robot arm via a keyboard.
-Users are informed with the following key mapping between the key press and the corresponding correction in the game.
-The safety constraint (60) in Franka arm’s safe MPC policy is updated immediately once a user correction is detected.
-When user press "ENTER", the arm is reset to the initial position and the reaching target will also change.
-TABLE VI: keyboard commands in robot arm user study
-Key Correction Interpretation
-"left" $[0,0,0,0,0,1]^{\mkern-1.5mu\mathsf{T}}$ Clock-wise rotation of the gripper
-"right" $[0,0,0,0,0,-1]^{\mkern-1.5mu\mathsf{T}}$ Anti-clock-wise rotation of the gripper
-"shift" $[0,0,0,-1,0,0]^{\mkern-1.5mu\mathsf{T}}$ Go up in world frame
-"ctrl" $[0,0,0,1,0,0]^{\mkern-1.5mu\mathsf{T}}$ Go down in world frame
-"Enter" N/A Reset robot position and switch target.
+The robot arm’s motion is visualized in MuJoCo using a free-view mode (Fig.
+16(b)), allowing users to adjust the camera to a preferred angle.
+The slot and the randomized reaching target remain visible throughout the task.
+Users provide corrections to the robot arm via keyboard input, following the specified key–command mappings.
+Whenever a correction is detected, the safety constraint (61) in the Franka arm’s safe MPC policy is updated immediately.
+Pressing ENTER resets the arm to its initial position and randomizes the reaching target.
+[Uncaptioned image] Zhixian Xie Zhixian Xie is currently a Ph.D.
+student in Arizona State University, Tempe, AZ, USA.
+In 2022, he received B.
+Eng degree in automation at Shanghai Jiaotong University, Shanghai, China.
+His research focuses mainly on controls, robotic learning from human and robot dexterous manipulation.
+[Uncaptioned image] Wenlong Zhang Wenlong Zhang is an associate professor in the School of Manufacturing Systems and Networks at Arizona State University, where he serves as the Research Director.
+He is a core faculty member in the Robotics and Autonomous Systems and Systems Engineering PhD programs.
+He joined ASU faculty in 2015 after earning his doctoral degree in mechanical engineering from the University of California, Berkeley.
+At ASU, he leads the Robotics and Intelligent Systems Laboratory (RISE Lab), and teaches undergraduate- and graduate-level classes on dynamics, control systems, optimization, and robotics.
+[Uncaptioned image] Yi Ren Yi (Max) Ren is an associate professor of aerospace and mechanical engineering in the School for Engineering of Matter, Transport and Energy at Arizona State University.
+he directs the Design Informatics Laboratory at ASU.
+His group studies computational and data-driven methods that will augment or transform the activities of engineering and industrial design.
+Ren’s current research interests include optimization, product/configuration design, human-computer interaction and machine learning.
+[Uncaptioned image] Zhaoran Wang Zhaoran Wang is an associate professor with Northwestern University, working at the interface of machine learning, statistics, and optimization.
+He is the recipient of the AISTATS (Artificial Intelligence and Statistics Conference) notable paper award, ASA (American Statistical Association) best student paper in statistical learning and data mining, INFORMS (Institute for Operations Research and the Management Sciences) best student paper finalist in data mining, Microsoft PhD Fellowship, SimonsBerkeley/J.P.
+Morgan AI Research Fellowship, Amazon Machine Learning Research Award, and NSF CAREER Award.
+[Uncaptioned image] George J.
+Pappas George J.
+Pappas is the UPS Foundation Professor at the Department of Electrical and Systems Engineering at the University of Pennsylvania.
+He also holds a secondary appointment in the Departments of Computer and Information Sciences, and Mechanical Engineering and Applied Mechanics.
+Previously, he served as the Deputy Dean for Research in the School of Engineering and Applied Science.
+Pappas’s research interest focus on control systems, robotics and autonomous systems, formal methods, machine learning for safe and secure cyber-physical systems.
+He has received numerous awards including the NSF PECASE, the Antonio Ruberti Young Researcher Prize, the George S.
+Axelby Award, the O.
+Hugo Schuck Best Paper Award, and the George H.
+Heilmeier Faculty Excellence Award.
+Pappas has mentored more than fifty students and postdocs, now faculty in leading universities worldwide.
+He is a Fellow of IEEE, IFAC, and was elected to the National Academy of Engineering in 2024.
+[Uncaptioned image] Wanxin Jin Wanxin Jin is an assistant professor the School for Engineering of Matter, Transport and Energy at Arizona State University (ASU).
+Previously, he was a postdoctoral fellow at the GRASP Laboratory at the University of Pennsylvania.
+He obtained his doctorate from Purdue University in 2021.
+His research includes robotics, control and machine learning, with a focus on the autonomy of robots interacting with humans and objects.